A Complete Bibliography of Publications in the Journal of Mathematical Physics: 2020–2024

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

13 September 2023 Version 1.08

Title word cross-reference

#P [67]. #P-complete [67].

(2 < p < 4) [203]. (k, a) [1484]. (n + 1) [870]. (p, k) [971]. (p, q) [1233]. (t_{2}, t_{3}) [1170]. (U_{q}(∫⊓(∞, ∞)), ≀_{⨿^{∞/∈}}(∈\)) [95]. 1 [276, 619, 838, 726, 82, 307, 122, 671, 884, 775]. 1 + 1 [255]. 1/N [687]. 1/r^{2} [897]. 1 < p < 2 < q < 6 [996]. 2 [355, 321, 575, 229, 43, 236, 743, 832, 160, 615, 302, 63, 677, 975, 839]. 2 × 2 [188]. 3 [604, 459, 837, 743, 366, 61, 21, 871, 354]. 6j [1372, 1388]. ^{∗} [241]. ^{−1} [731]. ^{2} [748, 280]. ^{3} [353]. _{2} [1398, 1265]. _{IV} [639]. _{n} [1084]. _{V} [639]. A[(λx) + i tanh(λx)] [1505]. A_{n} [781, 1149]. α [219, 486]. α − z [325]. α → 1 [746]. B [1239, 1479]. B_{n} [1447]. ¯∂ [765]. β [687, 1277, 922]. C^{∗} [1526, 880]. C^{5} ⊗ C^{5} [505]. C^{∞} [634]. C_{n}(1) [1498]. N = 2 [510]. PT [1505]. Cur(sl_{2}(C)) [1314]. D [225, 328, 656, 667, 258, 710, 789]. ẍ + f(x)ẋ^{2} + g(x) = 0 [115, 114, 8, 5, 6]. Ḃ_{∞,∞}^{−1} [731]. E^{2} [1322]. E_{8} [1108]. E_{τ,η}gl_{3} [151]. F (r) [1228]. g [1306, 303, 1009]. gl(n) ⊗ gl(n) [646]. GL_{q}(2) [686]. H^{2} [1047]. H^{s} [654]. K [1450, 1390, 789, 1386, 1416]. k + 1 [576]. κ [247]. K′_{4} [1148]. L^{1} [1427]. L^{2} [208, 726, 120]. L^{∞} [574]. L^{p} [660, 167, 285]. L_{∞} [371, 592, 999]. λ [953]. M [542, 1358, 1380]. M_{2} [1158]. C [1153]. P^{1} [30]. R^{+} × R^{3} [1165]. R^{3} [1146, 935]. R^{6} [150]. R^{d}, d = 1, 2, 3 [282]. R^{N} [814, 922]. R^{n+1} [1084]. R^{n+2} [1084]. Z^{d} [1329, 1236]. Z_{2}^{2} [568]. Z_{2}^{n} [134]. Z_{2} × Z_{2} [28, 650]. D^{⌊}(X) [169]. N = ∈ [1151]. ON [1177]. S(N) [113]. W_{⨿,⊔}(∫↕(∈|∞)) [585]. }↕(∞) [1277]. ≀∫√(∞|∈\) [1033]. ≀∫√(∞|∈\) ⊃ }↕(\) [1033]. ∫↕_{∈} [157, 914].

∫↕_{∈}[ltimes]L(△) [914]. ∫↕_{\} [1162]. SU(2) [213]. μ [1382]. N [995, 1458, 669, 1098, 1127, 1457, 199, 1448, 1245, 1523, 245, 262, 294, 452]. n^{2} + 3 [1223]. ♮_{α} [1082, 1487]. O [1497, 389]. O(3) [1104]. osp(1|2) [568]. P [984, 1452, 579, 653, 116, 1264, 1516, 1253, 658, 935, 857, 1081, 1025, 841, 471]. p(x) [20]. p· [814]. ψ [1190, 703]. q [1433, 440, 830, 223, 1407, 526, 95, 186]. Q(1) [1500]. q_{xx} [1283]. R [878, 953, 1312, 1177, 791, 646, 1149]. R^{1+1} [574]. ρ [695]. S [667, 462]. S^{2} [1047]. SDiff(S^{2}) [35]. σ [529]. sl(7) [667]. SL_{2} [662]. SL_{q}^{∗}(2) [188]. SO(3)_{p} [1081]. SU(2) [1395]. SU(N) [493, 1372]. τ [1333, 443]. S̃L(2, R) [1398, 1265]. U(1) [1399]. U(1)^{N} [510]. U(1)_{B−L} [430]. U(h) [1032]. U_{q}(D_{4}^{(3)}) [1255]. U_{q}(sl_{2}^{∗}) [188]. V_{D}(x) = min[(x + d)^{2}, (x − d)^{2}] [1353]. V_{S}(x) = max[(x + d)^{2}, (x − d)^{2}] [1353]. φ^{2k} [286]. φ^{4} [556]. φ_{4}^{4} [368, 1165]. ∨ [469]. VOA[M_{4}] [36]. W [585, 1313, 997]. W_{1+∞} [643]. W_{3} [694]. X [1360, 824]. Z [553, 905]. Z_{2} [1037]. Z_{2}^{2} [1151]. Z_{n} [1149]. ζ [1278].

-abc [199]. -adic [1452, 1081]. -algebras [241, 371, 592, 1313, 1526]. -Araki [830]. -based [660]. -body [1271, 294]. -bridge [832]. -calculus [526]. -Camassa [1382]. -coherent [526]. -component [462]. -contact [1416]. -convex [1380]. -critical [120]. -D [276]. -deformation [1433]. -deformations [999]. -Deformed [247, 526, 186, 1277]. -dependent [824]. -difference [1407]. -dimensional [667, 82, 1245, 316, 416, 452, 328]. -dimensions [258, 710]. -dressing [765]. -evolution [922]. -exact-solvability [667]. -extension [430]. -fidelity [1264]. -form [529]. -free [1032]. -functions [443, 1278]. -generalized [1484]. -graded [134, 1151, 568, 1153, 28, 650, 553, 1149]. -Harish-Chandra [914]. -Hilfer [1190, 703]. -Jacobsthal [1233]. -Kropina [1358]. -Laplace [814]. -Laplacian [653, 1306, 579, 116, 20, 1516, 1253, 658, 935, 1025, 841, 471]. -level [669]. -Lie [695]. -manifolds [852]. -matrices [646]. -matrix [1312, 1149, 878, 542]. -matrix-Nijenhuis [1177]. -minimal [1360]. -modules [1033, 1498]. -Navier [303, 486, 1009]. -operators [1497, 389]. -order [870]. -oscillators [95]. -particle [262]. -plane [223]. -point [1457]. -polynomials [526]. -Rényi [1082, 1487]. -semitoric [1479]. -soliton [995]. -solutions [656]. -spaces [285]. -spacetime [1119]. -spin [791, 857]. -stable [219]. -state [1037]. -states [667]. -structures [1177]. -superalgebra [585]. -superalgebras [997]. -supercritical [208, 120]. -systems [469]. -tensor [241, 880]. -term [1283]. -th [1386, 1448]. -theory [1450, 1390, 789]. -topological [1037]. -torsion [789]. -type [764, 1239]. -valued [646]. -weak [905]. -wise [984]. -Zakharov [1170].

1 [510]. 1/ [280]. 1/2 [645, 885, 510]. 1/2-BPS [510]. 1D [1341, 1455].

2-BPS [510]. 20 [479]. 2D [1095, 1099, 1362, 992, 1298, 1333, 1297, 1494, 1444, 1188, 1209, 1467, 1044, 1175, 1009, 1489, 1086].

3 [695]. 3- [695]. 3D [1095, 1060, 1444, 1207, 1210, 1343, 1422, 1425, 1340, 1381, 1478].

4-regular [1056].

53 [384]. 54 [115, 114, 8, 6]. 55 [8]. 56 [447]. 59 [808, 835].

60 [606, 54, 80]. 61 [195, 651, 448, 196, 1203, 269, 1336]. 62 [652, 653, 970, 1095, 865, 607]. 63 [1337, 1374, 1293, 932, 1133, 1487, 988].

ABC [764, 199]. abelian [1401, 1198, 1472, 69, 1225, 536]. Ablowitz [961]. Absence [1436, 224, 860, 1367, 846]. Absolute [48]. absolutely [741]. absorbed [959]. absorbing [918]. absorption [285]. abundance [1403]. acceleration [1048]. acceptor [212]. Accessible [742]. accidental [589]. according [581]. acknowledgment [15]. acoustic [26]. acoustic-type [26]. Across [216]. acting [669]. action [1408]. action-dependent [1408]. ad [317]. adaptation [617]. additive [1514, 138]. additivity [1397]. adeles [1259]. adelic [806]. Adiabatic [879, 974, 312, 284, 214, 1243, 279, 1197]. adic [1452, 1081]. adjoint [147, 25, 448, 148, 820]. adjointness [426]. admitted [154]. AdS [1398, 1265, 1228]. advected [142]. advection [44]. affects [1053]. Affine [1328, 997, 618, 1513, 1359, 38, 628, 1277]. affine-null [628]. age [956]. age-structured [956]. aggregation [331]. Aharonov [497]. Airy [569]. Algebra [334, 1151, 188, 1035, 1110, 878, 1106, 113, 29, 287, 667, 494, 11, 197, 792, 1312, 1073, 328, 256, 511, 186, 342, 478, 129, 1485, 1289, 694, 915]. Algebraic [661, 883, 1002, 886, 1179, 783, 713, 1334, 1291, 820, 639, 1146, 455]. algebraically [333]. Algebras [1109, 931, 134, 845, 191, 695, 543, 495, 1153, 239, 440, 830, 1069, 371, 736, 999, 27, 367, 95, 140, 592, 241, 979, 715, 62, 650, 1313, 1386, 1257, 129, 389, 422, 1526, 268, 968, 424, 361, 1032, 260]. algebroid [592, 50]. algebroids [507, 410, 998]. algorithm [1407, 34, 1197]. algorithmic [467, 1322]. algorithms [466]. aligned [722]. alignment [1341]. all-real [190]. Allen [911]. Almost [920, 540, 210, 560, 82, 458, 1014, 1451]. almost-periodic [82]. almost-periodically [458]. along [1169, 771]. also [1292]. Ambarzumyan [896]. Ambarzumyan-type [896]. ambient [1084]. Ambrosetti [1137]. among [1498]. amplification [659]. amplitude [1224, 1231, 214, 1004]. amplitudes [557]. an-harmonic [477]. Analog [1509]. Analogies [1074]. analogs [1198, 1241]. Analogues [1305]. analogy [1260]. Analysis [1018, 1323, 1335, 872, 800, 1484, 991, 1413, 670, 45, 1068, 123, 801, 508, 617, 151, 44, 1341, 621, 351, 192, 2, 1247, 913, 1031]. Analytic [1045, 981, 347, 25, 448, 313, 1353, 1458]. Analytical [409, 838, 1172, 801, 33, 1125]. Analyticity [146, 1446, 902, 958]. Anderson [472, 671, 884, 1189]. Ando [1041]. angle [724]. Angular [833]. Angular-radial [833]. anharmonic [945, 1061, 1212, 787]. Anisotropic [1149, 1402, 1438, 655, 1142, 232, 1175]. annihilation [144, 296]. annular [161]. annulus [1001]. anomalies [1391]. Anomalous [861, 636]. Anomaly [1400]. ansatz [713]. anti [1252, 186, 847, 1208, 982]. anti-Chaplygin [1252]. anti-de [982, 186, 1208]. anti-Zeno [847]. antiferromagnetic [541, 105]. antivortices [166, 805]. any [1137, 347]. anyonic [497]. anyons [1472, 210]. aperiodic [411, 1390]. appearance [1038]. Application [675, 1055, 102, 1182, 34, 524, 985, 215, 1526, 266]. Applications [411, 173, 955, 1153, 1484, 1111, 1267, 1129, 574, 979, 148, 465, 650, 707, 789, 108]. applied [129]. Applying [722]. approach [1194, 1371, 1242, 1039, 1513, 1110, 858, 1413, 1346, 661, 897, 378, 995, 443, 864, 1520, 820, 476, 657, 1168, 1404, 1322, 132, 300, 966, 1285, 193, 1098, 1081, 1237, 1469, 1448, 281]. approaches [1219]. Approximate [757, 1125, 880]. approximating [466]. approximation [1117, 1096, 1282, 284, 210, 383, 518, 1432]. approximations [1147, 1079, 633, 750, 632, 689]. Araki [830]. arbitrary [795, 1417, 1167, 245, 857]. Archimedean [49]. area [974]. area-constrained [974]. Arf [791]. Arf-invariant [791]. arising [166, 1059, 122, 571]. arrays [774]. Artificial [1260]. ASD [439]. Askey [95]. aspects [438, 144, 447, 7, 512, 139, 50, 410]. assisted [822, 1461, 947]. associated [783, 1069, 878, 29, 964, 340, 1065, 1211, 465, 545, 171, 1447, 649, 1196, 151, 965, 1521]. associative [1194, 845, 188, 999]. asymmetric [1389, 1319]. Asymptotic [800, 890, 950, 473, 1342, 1298, 92, 1297, 809, 1065, 218, 434, 414, 1307, 1007, 307, 1341, 222, 141, 158, 636, 491, 986, 267, 762, 195, 81, 87, 196, 1460, 1213, 1185, 131, 1278, 21, 1240]. asymptotic-preserving [1460]. asymptotical [104]. Asymptotically [817, 1009, 1008, 1268, 329, 1511, 270, 824, 982]. Asymptotics [719, 637, 1058, 971, 836, 234, 1523, 1352, 1283, 913]. Atiyah [1453]. atmosphere [279]. atmospheric [1442]. atom [496, 1040]. atomic [33]. atoms [408, 947, 176, 1318]. attraction [816, 818, 732, 315, 532]. attraction-repulsion [816, 818, 732, 532]. attractive [549, 492, 63]. attractor [303, 291, 922, 1087]. attractors [474, 954, 1188, 956, 374, 1253, 677, 815, 1361, 632, 1441, 471, 873]. automata [743, 1095, 290, 1053]. automorphism [1158]. autonomous [1187, 333, 473, 291, 374, 1322, 1516, 222, 221, 1087, 1009, 632, 689]. Average [210, 1234, 889]. Averaging [1350]. Avoided [111]. Aw [1252, 1432]. axial [1402]. axially [1166, 1067]. axiom [1356]. axiomatic [1219]. axis [854]. axisymmetric [246, 837, 1422]. Aztec [1365].

b [388, 452]. b-family [388]. Babylonian [1518]. back [1102]. background [604, 69, 663, 481, 1049, 56, 347]. backgrounds [1474, 1199]. Backward [613]. Bakry [1360]. Bakry-Émery [1360]. balance [122]. balancing [827]. ball [910]. Ballistic [605]. balls [727]. band [712]. Bands [741]. bandwidth [712]. Bannai [239]. Bardeen [846]. Bargmann [213]. barotropic [123, 1274, 480]. based [188, 154, 634, 660, 401, 1444, 322, 1404, 788, 1375, 1448]. Bases [1033, 1002, 171, 1470]. basis [1262]. bath [619, 1220]. Baxter [736, 1150, 999, 1314, 715, 343, 968, 915, 1092]. Bayesian [1004]. BCD [643]. beam [156, 458, 1308, 1031]. BEC [587]. behavior [491, 473, 762, 809, 81, 947, 576, 530, 162, 1063, 58, 307, 732, 167, 222, 94, 141, 21, 1324, 418, 354, 1240]. behaviors [702, 853, 1010, 103, 1318, 1286]. behind [635]. Bell [740, 505]. Bell-test [740]. below [236, 941, 796, 201]. Bénard [1209]. bending [630]. Benson [586]. Benson-Ratcliff [586]. Berezin [269, 189]. Berezinian [1107]. Berger [1087]. Bernouilli [522]. Bernoulli [193, 671, 884, 1237, 1189]. Berry [661]. Bessel [1415, 603, 1132]. beta [985]. Bethe [713]. between [1507, 682, 1389, 322, 876, 903, 641]. Beyond [1504, 1071, 98, 1246, 1264]. BGK [977, 1345]. bi [684, 305]. bi-drifting [305]. bi-Hamiltonian [684]. bialgebras [1451]. Bianchi [1301]. biconfluent [808, 10]. bidiagonal [749]. Bifurcation [1469, 142, 1231, 261, 1477, 1464]. bifurcations [598, 400]. bihamiltonian [1411]. bilayer [724, 1377]. Bilinear [443, 308, 499, 1413]. billiards [293]. bilocal [1473]. bimodules [241]. bipartite [214]. birational [292]. Birkhoff [886]. birth [1054]. Bistritzer [1377]. biwave [906]. BKP [1015, 443, 462]. black [1474, 673, 515, 398, 14, 606, 329, 1055, 691, 874, 1228, 594]. black-hole [1055]. blackhole [1260]. Bloch [780, 408, 1147, 995, 307]. Block [1435, 1093, 712]. Block-circulant [1435]. block-diagonalization [1093]. blocks [561]. Blow [388, 575, 816, 272, 725, 1296, 1496, 55, 612, 537, 534, 155]. Blow-up [388, 816, 725, 1296, 1496, 55, 612, 537, 534, 155]. blowing [910]. blowing-up [910]. blowup [862, 842]. BMO [731]. bodies [1043]. body [1271, 1124, 667, 1021, 1072, 433, 294]. Bogoliubov [320, 664]. Bogoliubov-Fröhlich [320]. Bogoyavlensky [1413]. Bohm [497, 849]. Boltzmann [1330, 481, 934, 1301, 58, 907, 1100]. Boltzmann-scalar [934, 1301]. bond [1154, 1389]. bonding [1326]. Bonnet [370]. Bonus [350]. Born [47]. Bose [1071, 1020, 1113, 492, 170, 63, 1021, 1072, 380, 1278, 152, 315]. boson [549, 1077]. Bosonic [1338, 11, 197, 526, 210, 696, 894, 323]. bosons [543, 990, 1070, 334, 718]. both [199]. Bound [175, 652, 774, 651, 1117, 1472, 862, 1071, 386, 1003, 170, 622]. boundaries [330, 872, 1172, 759, 313]. Boundary [557, 1471, 982, 1008, 91, 470, 1062, 1362, 459, 1524, 1254, 1446, 969, 1266, 535, 310, 612, 1144, 1141, 47, 963, 876, 453, 1428, 690, 332, 1102, 540, 755, 200, 1067, 1303, 1423, 1086]. Bounded [1024, 1376, 488, 21, 769]. Boundedness [1442, 816, 1462]. bounds [822, 784, 1329, 574, 1236, 688, 1417, 394, 74, 671, 747, 621]. Bourgain [989].

Boussinesq [80, 1099, 1175, 1241, 684, 416, 12, 975, 1283, 1465, 1422, 540, 839]. box [970, 786, 492]. box-shaped [492]. BPS [900, 510]. brackets [1440, 301]. Bragg [1390, 1318]. braid [126]. braided [126, 880]. braiding [673]. branching [1033]. branes [555]. breakdown [284]. breaking [1299, 206, 1246, 1169, 224, 1192, 100, 1064]. breather [864]. breathers [1123]. Bresse [1103]. bridge [832]. Brillouin [796]. broadcast [101, 623]. broken [547]. Brown [1125]. BTZ [794]. buckling [836]. building [1282]. Bulk [1351, 1258, 933, 1051]. bump [1216]. bundle [295, 525, 1284, 1255, 885, 889]. bundles [763, 644, 30, 804, 1256]. Burgers [613]. BV [1163].

C [241]. Caffarelli [772]. Cahn [1214, 1381, 911]. Calabi [139]. Calculation [639]. calculations [33]. calculi [686]. calculus [526, 494, 83, 1203, 1200]. Callias [1501]. Calogero [609, 107]. Camassa [273, 1299, 1382, 870, 1246, 340, 576, 55, 972, 728, 108, 165, 199]. cancellation [1400]. cancer [1136]. Canonical [1410, 508, 256, 627]. canonoid [1410]. Cantor [1392]. Capacity [101]. capillarity [352]. Caputo [545]. Carleman [617]. Carroll [1334]. Carroll/ [1334]. Carrollian [257]. carrying [69]. Cartan [371, 461, 1036]. cartoon [394]. case [208, 883, 861, 379, 157, 209, 796, 1208, 1059, 280, 1380]. cases [1430]. Casimir [113, 715, 422, 782]. categories [126, 241, 880]. Cattaneo [835, 9]. Cauchy [992, 490, 682, 1330, 1209, 654, 302, 844, 1476, 1029, 531, 939, 871, 23]. causal [254, 367]. causality [516, 72, 1321]. cavitation [520, 1468, 770]. cavity [947, 417, 1067]. Cayley [1333, 979, 300]. cell [1136, 1460, 1323]. cell-to-cell [1323]. cellular [743, 1095, 290, 1053]. center [41, 1324, 161]. central [827, 1271, 1043, 1040, 892, 1508, 433, 1235]. centrally [662]. CFT [1097]. chain [541, 565, 1162, 784, 687, 799, 45, 338, 645, 300, 375, 1462]. chains [683, 1258, 1178, 348, 1389, 1317, 1455, 1366]. Chandra [914]. Change [802]. changes [856]. changing [118, 1023, 488, 730, 16, 314, 926, 729, 1173]. channel [850, 101]. channels [177, 1161, 1083, 620, 623, 253, 1221, 250]. Chaos [1086]. Chaotic [332]. Chaplygin [530, 1063, 1252, 936, 1248, 94, 1432]. characteristic [345, 1407]. characteristics [747]. Characterization [716, 463, 237]. Characterizations [1042]. characterized [595]. characters [348, 1239]. Charge [1346, 1383, 69, 663, 1074, 353]. Charged [1232, 1474, 663, 404, 1051]. charged-particle [663]. charges [753, 1007]. Chebyshev [1073]. Chemical [1331]. chemically [991]. chemicals [816]. chemo [537]. chemo-repulsion [537]. chemotaxis [57, 816, 818, 533, 732, 1143, 939, 1024]. chemotaxis-Navier [939]. chemotaxis-Stokes [1024]. Chern [489, 73, 370, 574, 996]. chevron [162]. Chiral [287, 53, 1350, 1258, 1502, 724]. chirality [1346, 260]. Choi [1159]. choice [947]. Choquard [653, 579, 1244, 386, 733, 17, 153, 1105].

Choquard-type [1244, 17]. choreographic [294]. CHSH [740]. Chua [1477]. circle [1111, 402]. circuit [561, 1477]. circuits [949, 401]. circulant [1435, 572]. Circular [712]. CKP [718]. class [330, 1452, 118, 597, 259, 1363, 910, 1244, 927, 229, 834, 969, 488, 1125, 232, 17, 730, 151, 184, 1384, 266, 646, 436, 33, 252, 843, 869, 1340, 1429, 1296, 278, 873, 590]. classes [1499, 1450, 875]. Classical [766, 369, 1000, 1440, 1519, 1196, 563, 1038, 722, 177, 1222, 344, 98, 1036, 687, 559, 1065, 1083, 128, 243, 524, 1313, 227, 955, 1367, 646, 1149, 968, 361, 1394, 1395, 375, 353, 247, 281, 390]. Classification [1148, 115, 489, 1414, 140, 429, 114, 650, 1158, 8, 5, 6, 568, 1512, 696, 416, 505]. Clifford [1095, 134, 297, 67, 949, 743, 1178]. climate [994]. close [409, 1510]. close-packed [409]. closed [395, 1358, 1068]. cloud [299]. Cluster [1088, 1329, 1327]. CMA [439]. co [1410, 169, 125]. co-tropical [169]. coalescing [959]. coarse [281]. coarse-grained [281]. cocycles [672]. code [1434]. codes [101, 429, 1160]. codimension [1267]. COE [640]. coefficient [1109]. coefficients [168, 355, 1074, 802, 824]. Coexisting [805]. Coherence [551]. Coherent [244, 78, 591, 648, 1289, 526, 548, 321, 668, 1396, 97, 323, 334, 550, 738, 1319]. cohomogeneity [256]. cohomogeneity-one [256]. Cohomologies [1497]. Cohomology [1150, 845, 139, 363]. cohomotopy [555]. coincide [1219]. collection [1528]. Collective [944]. collisions [903]. coloring [683, 948]. Combinatorial [1498, 1261]. Combined [172, 996, 484]. combustion [1344]. Comment [652, 115, 195, 651, 447, 808, 970, 114, 196, 384, 8, 351, 1336]. common [177, 465]. communication [98]. communications [890]. Commutative [789, 624, 780, 292, 324, 1503]. commutativity [1034]. commutator [1257]. Commutators [1238]. compact [1104, 1250, 373, 1466, 1525]. compactification [625, 1118]. compactness [1306]. comparison [798, 1263]. compatibility [1039, 554]. competing [840, 877, 300, 1384, 571]. competitive [1335]. Complementarity [641]. Complete [331, 256, 404, 853, 67, 139, 1473, 305]. Complex [918, 1006, 190, 1483, 1052, 1435, 473, 954, 358, 1456, 826, 865, 293, 617, 1451, 632, 1505, 913]. Complexity [1437, 971, 1168, 857, 889]. component [1299, 1066, 777, 462, 55, 764, 765, 152, 534, 165, 1247]. components [1425]. composite [734]. Compositional [1368]. compound [101]. Compressible [608, 417, 991, 679, 992, 838, 1172, 575, 577, 1254, 812, 1206, 1063, 352, 1274, 1493, 938, 60, 75, 22, 453, 167, 1341, 61, 1210, 1343, 89, 21, 480, 1340, 23]. computation [44, 672]. computer [822]. computer-assisted [822]. computers [34]. Computing [1312, 381, 1404]. concave [996]. concentrated [1276, 601]. Concentration [1182, 1468, 857, 520, 877, 164, 770, 418, 1105]. conceptual [13, 54]. concise [621, 1294]. Condensate [1504]. condensates [881, 1115, 737, 492, 152, 315]. condensation [543, 1113, 380]. condensed [1391, 449]. condition [741, 47, 690, 332, 1441, 1303, 1086]. conditional [43, 1003, 1080, 802]. conditions [91, 1363, 1062, 1254, 1446, 1137, 941, 1266, 535, 52, 932, 888, 1144, 232, 393, 963, 1005, 1457, 325, 1423, 278, 982]. conductance [319]. conducting [991, 608, 753, 183, 1102, 354]. conduction [23]. conductive [600, 938, 167]. conductivity [475, 420, 1051, 942, 460, 1380]. confidential [101]. configuration [136]. configurations [1271, 1043, 711, 1453, 433, 1235]. confluent [1415, 1116, 1058, 1521]. Conformal [1314, 263, 1177, 845, 1148, 1153, 1512, 1110, 1069, 887, 356, 1320, 218, 150, 511, 1032, 71, 793]. conformally [527, 1164]. conformation [832]. conicity [147]. conjecture [1453]. conjugacy [1499]. conjugation [1346]. Conley [1518]. connecting [93]. connection [708, 507, 644, 293]. connections [570, 1068]. Connes [102]. consequences [1412]. Conservation [130, 403, 1383, 355, 233, 720, 485, 599, 316, 769]. conservative [912, 777, 524, 728, 221]. conserved [570, 708]. conserving [1460]. consistency [1000]. consistent [898]. constant [722, 1229, 951, 1101, 372, 564]. constants [397, 1510, 562]. Constrained [855, 1447, 1138, 1221, 974]. constraint [503]. constraints [506, 607, 337]. constructing [154]. Construction [1311, 900, 867, 1261, 962, 1450, 794, 625, 1118]. Constructions [1479]. Constructive [567]. consumption [868]. contact [1410, 1094, 1013, 42, 1416, 1186, 855]. Contextuality [1075]. continuation [981]. Continuity [1010, 253, 411, 48, 1329, 1174, 303, 1003, 747, 815]. Continuous [1169, 511, 1383, 900, 741, 373, 648, 327, 149, 796, 1492, 778, 1288]. continuum [774, 501, 1133, 1076, 1217]. contraction [772, 511]. contractions [465]. Contractivity [830]. contributions [557]. control [858, 136, 1190, 566]. Controllability [308, 499, 1230, 704]. controls [1224]. convection [1364, 74]. Convergence [1300, 640, 1489, 1224, 737, 633, 796, 253, 621]. converging [409]. Converting [848, 666]. convex [933, 669, 996, 1380]. Convexity [1067]. Cooper [846]. cooperating [623]. coordinate [848, 641]. coordinates [1157, 854, 628, 666]. Coprimeness [795]. Coprimeness-preserving [795]. copy [1401]. coral [939]. corank [178]. correction [174, 1229]. corrections [719]. correlated [405, 1080]. Correlation [177, 1365, 45, 958, 717]. correlations [137, 1012, 920]. Correlators [1162]. correspondence [1351, 1097, 642, 591, 96]. cosine [759]. cosmological [528, 722, 951, 1405, 631]. cotangent [295, 525]. Couette [1099]. Coulomb [479, 1415, 1062, 833, 1, 1019, 428, 1012, 49, 899]. Coulomb-like [833]. counterpart [1121]. Counting [67]. coupled [431, 1348, 1291, 864, 451, 693, 1044, 4, 384, 202, 776, 1304, 416, 1504, 1428, 765, 1240]. Coupling [377, 1130, 427]. course [1519]. Covariant [64, 686, 494, 1164, 1152]. CP [489]. CPn [513]. cradle [1123]. Creation [296]. criteria [552, 798, 837, 1066, 105, 55]. criterion [575, 842, 1422, 1425, 779]. Critical [1304, 828, 579, 653, 822, 840, 1244, 413, 490, 575, 116, 386, 610, 814, 1236, 1171, 414, 861, 957, 160, 813, 1347, 1007, 1208, 658, 201, 817, 167, 16, 314, 750, 976, 1087, 120, 1204, 418, 1105, 909, 1173, 1430]. criticality [174]. cross [84]. cross-Kerr [84]. Crossing [1111]. crossings [111]. crystal [992, 1066, 1139, 711, 1343, 1294, 23]. crystalline [265]. crystals [73, 339, 1255]. Cubic [1322, 1052, 409, 431, 356, 549, 660, 1247]. Cucker [702, 853, 1121, 890, 451]. curl [1492]. current [792]. currents [283, 680, 1480, 1486]. curvature [1211, 372, 1360, 102]. curved [592, 966, 102, 1049]. curves [213]. cut [1052, 1267]. cutoffs [556]. cuts [744]. Cwikel [987]. cycles [34, 522]. cycloacene [589]. cylinder [356]. Cylindrical [1456, 1357, 1402, 1438, 1118].

d [639, 276, 619, 604, 355, 321, 459, 838, 575, 229, 43, 726, 255, 236, 837, 743, 160, 615, 302, 63, 307, 122, 671, 884, 677, 975, 366, 61, 21, 871, 839, 775, 354, 639]. d-P [639, 639]. Daele [1450]. damped [1010, 1342, 419, 633, 1429]. damping [474, 1172, 156, 59, 537, 1245, 237, 184, 1280, 306, 141, 158, 922, 1465, 200, 1308, 452]. Darboux [1195, 93, 972, 266, 79, 716]. Darcy [1044]. dark [675, 581, 536]. Darken [681]. Darmois [66]. data [1008, 900, 762, 660, 1427, 1376, 941, 737, 1209, 1213, 313, 1256, 122, 1005, 839, 1340, 325, 937, 988]. data-processing [1005]. Davey [842]. death [1054]. Decay [517, 600, 1139, 160, 275, 237, 1080, 89, 719, 59, 19, 612, 1467, 615, 977, 1405, 616, 1379, 1146, 1012, 1210, 1465, 540, 871, 200, 614]. decaying [919, 1249, 464]. decimation [1014]. Decimations [1288, 967]. decoders [623]. Decomposable [745]. Decomposition [215, 1036, 424, 717]. Decompositions [96, 757]. decoupling [1224]. deep [1059]. defect [1060]. defects [287, 194, 1051]. deficiency [470]. defined [148, 1179, 424]. definite [191, 1285, 550]. Deformation [763, 525, 260, 1433, 1522]. deformations [396, 1497, 1150, 999, 1287, 626, 157]. Deformed [756, 247, 526, 585, 186, 342, 280, 1149, 1277, 778, 529, 738]. deformed- [778]. degeneracy [701, 589, 1434]. Degenerate [1234, 1279, 1337, 234, 1414, 212, 61, 578, 942, 460]. Degenerations [62]. degree [822, 576]. degrees [1320]. delay [474, 156, 908, 1241, 202, 456, 1324, 704, 1325]. delay-difference [1241]. delay-differential [1241]. delayed [40, 947, 1215]. delayed-choice [947]. delays [374, 1270, 1335, 1190, 677, 163, 471]. Delta [1248, 872, 1302]. Demazure [781]. dense [1194]. densities [902]. Density [438, 569, 420, 838, 1446, 663, 1139, 615, 891, 985, 383, 1072, 453, 671, 884, 550, 942]. density-dependent [838, 1139, 615, 891, 942]. density-depending [420]. density-of-states [663]. dependence [330, 1350, 1051, 266, 839]. Dependent [1112, 244, 168, 477, 960, 147, 92, 838, 1172, 1139, 387, 97, 826, 865, 615, 1044, 891, 1245, 4, 323, 384, 1143, 1408, 79, 716, 849, 942, 1029, 824, 460, 532, 247, 738, 1319]. depending [420]. depolarizing [850]. depth [797]. dequantizer [322]. Derivation [1315, 454, 549, 1386, 880, 1179]. derivative [1482, 345, 545, 963, 1190, 703, 690, 622, 1445]. Desargues [292]. descent [858]. description [412, 548, 1328, 1108]. detect [516]. detection [636]. detectors [1038]. Determinant [1333, 1407, 1523]. determinants [1350, 1065]. determining [666]. Detonation [1344]. developed [1309]. development [1135, 1031]. deviation [707]. deviations [597, 1040, 1280, 1366]. diagonal [745]. diagonalizability [31]. diagonalization [1093, 647]. diamond [1365, 223]. Dickson [979]. dielectric [1476]. diffeomorphisms [359]. difference [1230, 1407, 1389, 1241, 965]. differences [1074]. different [808, 10, 1235]. differentiable [754, 258]. Differential [1091, 333, 686, 751, 927, 494, 112, 434, 1109, 896, 119, 157, 132, 1241, 965, 621, 457, 450, 704, 807]. differential-difference [965]. differential-subtraction [621]. diffracted [1318]. diffuse [1524]. Diffusion [773, 378, 91, 168, 946, 474, 234, 1297, 90, 387, 801, 710, 1495, 1467, 1018, 47, 44, 1259, 806, 1361, 1024, 636, 566, 873, 1027, 1331]. diffusive [1269, 1325]. diffusivity [905]. Digital [349]. Dijkgraaf [1111]. dilation [253]. Dilations [364]. Dilute [1021, 1088, 170]. dimension [220, 286, 1154, 696, 1275, 1417, 546, 554, 96]. Dimensional [924, 213, 676, 860, 1033, 946, 1227, 1078, 912, 887, 182, 526, 1288, 967, 1136, 1411, 667, 1291, 1126, 255, 429, 883, 328, 629, 493, 795, 124, 105, 104, 508, 1002, 82, 458, 1245, 1493, 598, 170, 194, 432, 1292, 400, 316, 1228, 22, 1167, 416, 365, 819, 1295, 1420, 1012, 1102, 460, 869, 480, 939, 435, 315, 1024, 452, 1303, 614, 23, 1356]. dimensions [1371, 993, 1223, 317, 950, 1440, 567, 655, 1266, 258, 1168, 710, 454, 313, 393, 1316, 1419, 1205, 380, 1514, 519]. Dirac [448, 1293, 56, 251, 917, 665, 415, 24, 872, 1398, 895, 1062, 1346, 1351, 1355, 413, 833, 25, 289, 691, 1266, 1017, 1226, 1313, 1250, 186, 342, 442, 441, 86]. direct [1122, 584, 664, 735]. direct-sum [584]. Directional [1193, 649]. Dirichlet [91, 1001, 48, 109, 1427, 1085, 583, 535, 310, 1303]. Disappearance [757]. discontinuity [253]. discontinuous [355, 658]. Discrete [720, 792, 444, 759, 627, 426, 1449, 853, 1415, 1433, 783, 1481, 1363, 344, 346, 954, 1288, 341, 156, 1073, 1126, 378, 795, 1226, 1293, 929, 884, 1054, 141, 632, 1423]. discrete-time [1126]. discretization [961]. discretizations [233]. discretized [739]. discussion [373]. disformal [1320]. disk [1339, 792, 64]. dislocations [112]. disordered [687, 799, 224, 671, 1502]. dispersal [908, 163]. Dispersion [9, 835, 659, 341, 158, 235, 199]. Dispersive [810, 1240]. dissipation [1251, 1188, 1142, 612, 1175, 1204, 199]. dissipative [388, 421, 588, 436, 855]. distance [1268, 892, 1321, 1181]. distortion [635]. distributed [163]. distribution [242, 637, 1059, 245, 262, 787]. distributional [1201]. distributions [1232, 1289, 1458, 625, 1118]. disturbance [1302]. disturbances [207]. disturbed [1197]. divergence [746, 1443, 1082, 1487, 1005, 1296]. divergence-free [1443]. divisibility [552]. divisors [375]. DKP [1015]. Do [516, 1219]. domain [1298, 1524, 809, 275, 21, 769, 775, 1478, 161]. domains [872, 420, 473, 539, 1376, 488, 374, 406, 1516, 1492, 222, 750, 1343, 460, 590]. Domar [617]. dominates [506, 607]. donor [212]. donor-acceptor [212]. Double [1418, 187, 592, 963, 1514, 29, 1401, 1137, 1427, 277, 50, 410, 1070, 1353, 662, 158, 976, 1058]. double-confluent [1058]. Double-graded [187]. double-phase [1137]. Doubled [50, 512, 410]. doubles [128]. doubling [822]. doubly [868, 1309]. dressing [765]. drift [168]. drifting [305]. driven [1439, 973, 823, 37, 1128, 1215, 1027]. drop [1248, 1469]. drums [1048]. dS [1084, 150]. dual [709, 1006, 489, 95, 657, 129, 1353, 1466, 622]. dualities [240]. duality [1334, 880, 589, 125, 366, 1091]. dually [805]. Dubrovin [1183, 1447]. due [475, 332]. Duffing [1185]. Dullin [206]. Dunkl [966]. dusty [520]. Dynamic [144, 1503, 856, 645, 994]. Dynamical [1472, 826, 865, 546, 889, 1215, 1224, 1010, 336, 1040, 45, 1083, 104, 1322, 400, 1190, 703, 221]. Dynamics [1187, 142, 549, 692, 1044, 699, 1516, 582, 532, 619, 1096, 1279, 1337, 1088, 1020, 1000, 41, 890, 1362, 1424, 675, 548, 1496, 156, 69, 951, 600, 1251, 559, 40, 596, 1517, 530, 1079, 768, 242, 426, 1515, 933, 1419, 1112, 1053, 427, 1377, 1009, 755, 1327, 775, 1429, 855]. Dyson [127, 687, 799, 861, 1419, 1528].

E6 [508]. earth [279, 1181]. eddy [1442]. Edge [112, 183, 709, 178, 1351, 1258, 1455]. edged [1471]. Effect [797, 860, 273, 681, 537, 847, 818, 760, 203, 484, 841]. Effective [1481, 1096, 523, 518, 1429, 542, 1144, 1400, 247, 738]. effects [504, 680, 996, 614]. Efficient [466]. Efimov [860]. eigenfunctions [822, 727]. eigenmodes [312, 509]. eigenstates [1133, 1076, 1526, 547]. Eigenvalue [282, 1059, 960, 637, 572, 1458, 779, 226]. Eigenvalues [53, 1250, 1249, 110, 147, 1193, 1085, 124, 987, 1133, 1076, 1179, 1058, 305]. eight [365]. eight-dimensional [365]. Einstein [356, 371, 1129, 934, 1301, 1113, 492, 700, 628, 1466, 380, 152, 315, 793]. Ekman [1442, 74]. elastic [59, 306, 456]. elastodynamics [1028]. Electric [1051, 1078, 73, 447, 401, 408, 7, 663]. electrical [475, 1481]. Electrodynamics [519, 678]. electromagnetic [403, 722, 604, 73, 69, 931]. electromagnetism [924]. electron [475, 1154, 1226, 1293, 212]. electron-phonon [475]. electronic [902]. Electrostatic [402, 145, 602]. electroweak [1400]. Elementary [1055, 1394]. elements [422]. elephant [892, 1286]. ellipses [753]. Elliptic [609, 1371, 1421, 118, 204, 1244, 20, 343, 1026, 617, 17, 658, 270, 1428, 976]. embedded [1084, 1167]. Embedding [257, 446, 794]. embeddings [1305, 794]. emergence [251]. Emergent [702, 853, 103, 40, 596]. Émery [1360]. emission [944]. encodes [1321]. end [1268, 26]. endpoint [209]. energies [173, 26, 884, 857]. Energy [43, 485, 612, 769, 264, 350, 1071, 1020, 990, 1023, 723, 1349, 1460, 414, 488, 932, 888, 957, 1491, 546, 170, 730, 648, 1072, 664, 130, 266, 79, 716, 1221, 1179, 556, 16]. energy-conserving [1460]. energy-constrained [1221]. energy-dependence [266]. energy-dependent [79, 716]. engineered [1004]. Enhancement [683]. ensemble [1371, 1052, 40, 742, 637, 1521]. ensembles [1036, 687, 682, 767, 1017, 985, 1012, 529]. entails [31]. Entangled [178, 947, 1318]. Entanglement [98, 784, 100, 980, 591]. entity [740]. entries [572]. entropies [802, 639, 744, 1184, 325]. Entropy [220, 1057, 1050, 597, 784, 379, 1263, 1003, 933, 453, 1366, 566]. enveloping [845]. environment [707, 1269]. epidemic [163, 1269]. Equality [325, 979, 294, 1005]. equation [1198, 491, 392, 1421, 24, 773, 499, 1166, 579, 653, 912, 1030, 115, 356, 273, 166, 539, 1342, 358, 762, 549, 1114, 1382, 1145, 916, 413, 490, 1496, 660, 292, 206, 198, 808, 111, 870, 1330, 1246, 121, 921, 9, 835, 81, 1314, 386, 655, 18, 777, 1266, 811, 501, 867, 114, 387, 576, 535, 973, 310, 795, 10, 414, 83, 1203, 861, 801, 1125, 160, 481, 1468, 303, 55, 291, 454, 612, 1144, 972, 1495, 58, 82, 654, 458, 813, 1213, 977, 1307, 1431, 728, 47, 393, 119, 164, 404, 844, 1405, 721, 633, 316, 8, 1208, 963, 907, 307, 903, 130]. equation [193, 1488, 153, 1253, 1167, 935, 1140, 268, 968, 961, 1476, 181, 770, 1201, 690, 5, 6, 787, 1091, 158, 16, 360, 518, 631, 815, 1283, 922, 1465, 531, 538, 824, 1087, 1009, 235, 120, 165, 725, 200, 1025, 347, 199, 418, 1105, 913, 1448, 1303, 1423, 1429, 778, 859, 1086, 1064, 1331, 1430]. equation-Wigner [787]. equations [1482, 1300, 395, 604, 118, 368, 1165, 1311, 719, 1000, 333, 1348, 1363, 1413, 1010, 1362, 1424, 376, 182, 473, 474, 1176, 1230, 954, 1298, 1407, 204, 233, 1346, 659, 234, 910, 93, 92, 751, 810, 584, 734, 1214, 927, 388, 838, 1172, 1297, 720, 1475, 1401, 833, 635, 1315, 809, 575, 482, 208, 229, 398, 687, 116, 1406, 1147, 112, 378, 1376, 236, 610, 837, 558, 733, 1188, 814, 758, 434, 812, 574, 715, 123, 1174, 1460, 357, 451, 1142, 657, 1206, 1439, 162, 485, 117, 1273, 487, 1207, 944, 20, 737, 1063, 1468, 710, 343, 374, 352, 905, 1305, 232, 1467, 275, 823, 896, 1274]. equations [891, 1245, 302, 237, 60, 1185, 1175, 911, 1170, 274, 26, 1241, 875, 184, 707, 419, 1384, 931, 486, 1248, 1280, 22, 1205, 453, 628, 1304, 266, 79, 817, 1341, 222, 1186, 573, 416, 262, 94, 677, 961, 469, 484, 975, 108, 1146, 272, 843, 522, 571, 141, 332, 61, 765, 578, 731, 750, 942, 976, 1029, 1422, 564, 450, 869, 540, 1361, 480, 632, 871, 1204, 1324, 704, 839, 1058, 1340, 1521, 1445, 1308, 769, 775, 536, 909, 1092, 793, 614, 1381, 1296, 354, 807, 471, 873]. equilibria [827, 402, 1049, 1492]. Equilibrium [1047, 858, 1191, 313, 435]. Equivalence [73, 941, 876]. equivalent [67]. ergodic [1508]. Ergodicity [1419]. Ericksen [1385]. Erratum [1337, 385, 197, 835, 1095, 54, 479, 1374, 1203, 1133, 1487, 80, 6, 8]. error [109]. Essential [1034]. estimate [1307]. Estimates [305, 1268, 1182, 810, 987, 441, 1525, 871, 1375, 1380, 778]. Estimating [1258]. Estimation [1210, 219, 1004, 622]. Euclidean [1008, 395, 722, 685, 186, 180, 1167, 1492, 1394, 625, 1118]. Euclidean-signature [722]. Euler [80, 93, 734, 1172, 1494, 642, 123, 1063, 1468, 1305, 1275, 1245, 1248, 22, 453, 1492, 94, 12, 1100, 869, 1426, 769]. Eulerian [143]. Evaluation [1116]. even [887, 1389]. even-dimensional [887]. Evolution [1174, 1157, 84, 1078, 1363, 601, 561, 299, 1439, 627, 263, 374, 823, 1202, 875, 922, 281]. evolutionary [1324]. Exact [604, 1078, 638, 1191, 328, 393, 1480, 75, 748, 1458, 1308, 1319, 1311, 1402, 1438, 667, 416]. Exactly [1054, 1353, 1505]. examples [46, 1305]. Exceptional [701, 590, 1499, 1196]. exchange [771, 621]. exchange-integration [621]. excitations [1504]. exclusion [1126, 215, 193]. exist [980]. Existence [1060, 395, 579, 653, 991, 166, 474, 1176, 413, 208, 723, 1378, 921, 610, 864, 52, 104, 492, 1216, 1026, 164, 1443, 573, 724, 418, 491, 420, 1299, 906, 1104, 752, 726, 576, 451, 82, 458, 615, 1275, 977, 938, 1138, 100, 1143, 380, 1146, 61, 815, 1465, 21, 200, 1308, 354, 161]. exothermically [1494]. exotic [790]. exotics [493]. expanded [904]. expansion [986, 267, 168, 1088, 225, 459, 567, 179, 663, 476, 796, 556, 869]. expansions [1339, 1291, 443, 1278]. experiment [740, 947]. Explicit [46, 1405, 1457, 1448]. Exploring [1318, 180]. exponent [386, 813, 419, 1087]. Exponential [751, 1012, 579, 653, 1244, 615, 1128, 876, 817, 621, 1379, 314, 540, 1441]. Exponentially [3, 385, 864, 464]. exponents [204, 116, 688, 1253, 750, 976]. Expressions [1524, 1458]. Extended [42, 469, 808, 1243, 10, 662]. extension [191, 568, 430, 1163, 983, 793]. extensions [134, 1305, 1511, 129]. extensivity [802]. exterior [539, 515, 809, 275, 1343]. external [1279, 1337, 604, 810, 931, 955, 538, 873, 562, 923]. extremal [1193, 329, 140, 594]. extremals [570].

Faber [428]. facilitated [1126]. factor [541, 1437]. Factorization [966, 528, 249, 1263]. factorized [737]. factors [446, 1077]. falling [261]. families [190]. family [959, 29, 388, 338, 875, 1434, 199]. Fano [774, 169]. fashion [666]. fast [858, 409, 184, 1280, 1249, 518, 1046, 1429]. feature [1395]. features [717]. Fermi [56, 1048, 929, 1029]. fermion [254, 367, 1317]. fermionic [179, 1284, 443, 1243, 97]. fermions [1096, 946, 1520, 334]. ferrite [768]. ferromagnetic [940]. ferromagnetism [1471]. fertilization [939]. Feynman [1259, 806]. fiber [644]. Fibonacci [1122]. fibrations [407]. Fick [681]. fidelity [1264]. field [264, 1326, 722, 1279, 1337, 1398, 708, 1348, 887, 172, 295, 69, 567, 1315, 11, 197, 205, 447, 408, 7, 951, 626, 1378, 246, 496, 210, 592, 1265, 952, 1332, 150, 224, 1164, 1101, 379, 934, 1301, 509, 894, 404, 1120, 50, 410, 936, 194, 931, 1248, 176, 1416, 335, 791, 1051, 1492, 1098, 1504, 1127, 1184, 1469, 86, 536, 793, 1356, 390]. fields [396, 1006, 604, 1078, 359, 634, 3, 385, 397, 1510, 691, 1163, 1456, 1007, 523, 955, 1443]. Filippov [62]. filling [818]. finding [34]. Finite [946, 571, 1131, 213, 1117, 1148, 676, 569, 1224, 706, 1494, 1225, 1163, 1258, 207, 465, 77, 1523, 176, 1259, 926, 899]. finite-dimensional [213]. Finite-part [571]. finite-time [207]. finitely [258, 1080]. Finsler [952]. First [399, 1187, 333, 431, 1011, 339, 494, 103, 1456, 1322, 1443, 556, 221, 724, 228]. first-order [333, 431, 1011, 103, 1456]. Fisher [578]. Fisher-type [578]. fission [1469]. FitzHugh [1215, 689]. Five [433, 558, 508, 1235]. Five-body [433]. five-manifolds [558]. fixed [822, 1406, 864]. fixing [254]. FKG [224, 983]. FKMM [1412]. flat [1008, 1298, 329, 397, 647, 934, 279, 26, 432, 1292, 439, 519]. flatness [754, 950]. flavor [547]. Flett [965]. flights [638, 93]. flip [144]. flip-annihilation [144]. Floating [711]. flock [890]. flocking [890]. flocks [702, 1326]. Floquet [1518, 381]. flow [1153, 368, 1165, 459, 331, 263, 1057, 52, 797, 1063, 1250, 1044, 313, 1138, 391, 231, 417, 1102, 974, 1067, 1432]. flows [1099, 679, 992, 1494, 1442, 1066, 370, 230, 1493, 75, 44, 1492, 167, 1343, 23, 705, 1309]. FLRW [951, 481, 934, 1464]. Fluctuations [1016, 572, 406, 960, 638, 957, 1070, 828]. fluid [991, 514, 1254, 1057, 19, 1063, 1305, 75, 933, 456, 871, 925]. fluids [415, 1357, 1402, 1438, 1135, 600, 352, 615, 938, 21, 755]. flux [3, 385, 663, 319, 936, 377, 771, 1432]. flux-limited [771]. fluxes [355, 1309]. Fock [1242, 1114, 1459, 240, 931, 697]. focusing [549, 1496, 1213, 153, 120]. Fokas [1448, 1423]. Fokker [1101, 1141, 486, 1379, 168]. fold [1231]. fold-Hopf [1231]. folded [683]. foldings [1387]. foliating [1492]. folklore [411]. food [1462]. force [1326, 827, 1341]. forced [182, 458]. forces [873, 562, 923]. Forchheimer [679, 52, 230]. forcing [82, 1303]. forest [1091]. form [541, 1358, 256, 749, 1077, 1221, 859, 529]. Formal [1069, 41]. formalism [106, 527, 524, 322, 301, 365]. formality [146]. Formation [207, 1302]. forms [854, 634, 644, 401, 1084, 148, 1116]. formula [1089, 1333, 109, 642, 781, 739, 131]. formulas [113, 642, 1405, 621]. formulation [438, 790, 230, 546, 953, 628, 807]. Formulations [1331]. Fornberg [534]. FORQ [1064]. Foundations [1004, 952]. Four [667, 598, 442, 1453, 931, 227]. Four-body [667]. four-dimensional [598]. four-parameter [931]. four-sphere [442]. four-wave [227]. Fourier [783, 649, 829]. fourth [754, 490, 654, 1307]. fourth-order [754, 490, 654, 1307]. Fractional [330, 964, 813, 779, 1300, 304, 1306, 579, 653, 474, 358, 204, 1145, 751, 500, 692, 208, 1378, 9, 835, 81, 386, 610, 387, 487, 20, 710, 545, 1491, 911, 730, 334, 153, 658, 1190, 222, 416, 703, 922, 978, 120, 1204, 1105, 1027]. fractons [790]. fracture [580]. frame [143, 1404]. frames [1225]. framework [213, 665, 543, 34, 167]. Free [264, 957, 1202, 946, 1172, 331, 1520, 170, 1443, 540, 1032, 718, 1067]. freedom [1320, 760]. Freeman [1528]. Freezing [379, 709, 1050]. frequencies [181]. frequency [1484]. friction [916, 1252]. Friedland [772]. Friedmann [272]. Friedrichs [498]. Frobenius [46, 1183, 1447]. Fröhlich [320]. front [578]. fronts [911]. frustration [596]. full [575, 458, 1275, 1295]. full-dimensional [458]. Fully [1357, 1309, 57]. function [1415, 541, 754, 223, 706, 882, 45, 1236, 1, 479, 767, 343, 1202, 863, 1132, 32]. functional [179, 434, 1491, 664, 846]. functionals [1174, 1366]. functions [1117, 1415, 1370, 1365, 267, 118, 3, 385, 1333, 964, 1459, 1055, 443, 919, 159, 1332, 551, 462, 1004, 645, 983, 322, 466, 151, 211, 698, 958, 262, 441, 787, 1091, 571, 618, 1278, 899]. fundamental [1075]. Further [373]. Future [951, 793]. Fuzzy [1119, 930, 186, 342, 442].

galaxies [675]. Galilean [257, 694, 1032]. Galilei [1334]. Galois [132]. game [1222]. games [708, 503]. gap [346, 1399, 879, 1390, 978]. gapped [1506, 1093, 319, 1156, 880]. gaps [276, 1085]. gas [1071, 1088, 1020, 1446, 1330, 299, 530, 1063, 313, 170, 1021, 1072, 22, 231, 167, 94, 1012, 1278, 460, 869, 775, 1489, 1432]. gasdynamics [520]. gaseous [123]. gases [230, 1019, 63, 1367, 49, 899]. gaskets [901]. Gaudin [348, 1149]. Gaudin-type [1149]. Gauge [1391, 663, 714, 760, 295, 254, 510, 30, 463, 591, 821, 50, 536]. Gauge-invariant [663]. gauge/gravity [591]. Gauged [186, 1104, 805, 574]. Gauss [370, 958, 1366]. Gaussian [984, 1242, 597, 742, 465, 379, 1050, 823, 1523, 697, 1233, 1458, 1521]. gCH [725]. Gelfand [739, 1201]. Gelfand-Yaglom [739]. General [1407, 94, 1482, 912, 661, 1254, 741, 90, 742, 28, 59, 288, 454, 1493, 122, 280, 1392, 968, 1462, 859, 1086, 1309]. general-relativistic [288]. Generalization [131, 524, 186, 609]. Generalizations [804, 24, 1373]. generalize [214]. Generalized [593, 548, 668, 1129, 159, 801, 1077, 1257, 129, 1485, 316, 782, 130, 506, 607, 266, 739, 1195, 1483, 1484, 719, 1078, 1396, 759, 842, 1211, 1332, 139, 576, 657, 52, 230, 1170, 77, 1470, 211, 334, 794, 1286, 252, 522, 158, 505, 1465, 1425, 531, 1131, 996, 1092, 1432, 1064]. generated [408, 805, 290, 77, 1080, 226]. generates [439]. Generating [51, 620]. generation [177]. generators [806]. Generic [437, 341, 717, 766]. genuinely [712]. genus [986, 46, 1131]. Geodesic [398]. geodesics [1509]. Geometric [213, 1450, 580, 923, 504, 378, 1328, 370, 1211, 886, 1197, 1272]. geometrical [507, 457]. geometrically [1160]. geometries [507, 1287, 595, 1360, 257]. Geometry [137, 1433, 430, 1359, 1218, 624, 1272, 1281, 318, 1094, 218, 512, 169, 852, 521, 1155, 684]. geophysical [755]. geostropic [1142]. Gerdjikov [963]. Geroch [700]. gFQXL [725]. gFQXL/gCH [725]. ghost [1163]. Gibbs [1180]. Gibbsian [766]. Ginibre [1371, 1193]. Ginsparg [186, 342]. Ginzburg [473, 954, 358, 169, 973, 846, 222, 538, 632]. Girsanov [1282]. given [898]. GJS [241]. glass [264, 984]. Glauber [1337, 1279]. Global [528, 91, 420, 1299, 906, 1114, 1136, 752, 660, 777, 512, 576, 451, 162, 905, 1209, 1467, 57, 615, 1275, 977, 908, 1493, 938, 533, 1143, 794, 1379, 1146, 61, 578, 868, 942, 1465, 1343, 460, 21, 1462, 871, 200, 1340, 1025, 452, 155, 199, 1310, 1464, 841, 354, 161, 1371, 491, 726, 104, 509, 1443, 975, 939, 1308]. Global-in-time [1114]. Globally [728, 516, 1164, 894]. Glogower [668]. go [1181]. Godunov [933]. going [60]. Golden [560, 1041]. Good [171, 1283]. Gordon [931, 729, 1474, 604, 236, 328, 844, 316, 518, 582]. Gottwald [206]. governed [123]. GP [587]. GP-limit [587]. GR [1357, 1402, 1438]. graded [134, 1151, 568, 1153, 187, 28, 650, 553, 1149]. Gradient [1380, 858, 1066]. grained [281]. graph [1222, 67, 1333]. graphene [321, 1471, 724, 1377]. graphite [800]. Graphs [899, 308, 1056, 223, 214, 1158, 553, 926]. Grassmann [407, 119]. Grassmannian [584]. Grassmannian-parameterized [584]. Grassmannians [1370, 461, 1515]. gravitating [1476]. Gravitational [1229, 257, 1104, 1403, 796, 71]. gravity [673, 1513, 371, 792, 1164, 1226, 1293, 508, 591, 38, 748, 1228, 228]. Green [479, 1459, 1]. Gregory [515]. Gregory-Laflamme [515]. grids [440]. Groeneveld [476]. Gross [1071, 990, 454]. Ground [311, 1145, 496, 357, 1171, 105, 693, 1072, 1434, 152, 271, 1173, 1060, 423, 1071, 236, 941, 657, 492, 164, 880, 664, 556, 314, 483, 909]. group [251, 188, 719, 634, 525, 1475, 743, 1095, 468, 1178, 463, 821, 700, 1212, 1454, 1092]. group-quasialgebras [1092]. groups [1499, 217, 154, 946, 126, 1225, 349, 1158, 1404, 931, 1238]. growing [1496]. Growth [1128, 304, 579, 653, 1244, 1378, 1137, 610, 814, 1171, 232, 394, 817, 843, 16, 314, 418, 1105, 1441, 1173, 278]. GUE [1058]. Guiding [41]. Gurevich [581].

Haag [880]. Hadamard [1435, 894]. Hahn [149]. hair [14, 606]. half [1507, 1165, 811, 973, 690, 235, 1489]. half-infinite [1507]. half-line [811, 973, 235]. Hall [1332, 485, 319, 1503, 88, 978, 89, 614]. Hall-magnetohydrodynamic [89]. Hamilton [106, 1166, 527, 42, 524, 404, 931, 1186, 923]. Hamiltonian [865, 1133, 1410, 665, 248, 1474, 1522, 761, 412, 661, 1414, 1244, 498, 34, 246, 1013, 826, 320, 17, 1076, 553, 227, 684, 1234]. Hamiltonians [259, 1093, 148, 820, 611, 381, 324, 133, 547, 590, 1220]. hand [333, 431, 1011]. Hankel [1034, 1065, 1523, 1527, 226]. haptotaxis [1136]. hard [1088, 58]. Hardy [386]. Harish [914]. Harmonic [1212, 110, 477, 84, 667, 559, 805, 883, 258, 37, 4, 384, 1503, 796, 365, 849]. harmonics [470]. Harris [1375]. Harris-type [1375]. Hartree [1346, 1114, 737]. Hausdorff [437]. Hawking [1260]. Hayman [772]. Heat [391, 619, 719, 991, 420, 600, 121, 9, 835, 608, 1144, 938, 74, 377, 167, 1102, 942, 460, 1445, 354, 23]. heat-bath [619]. heat-conducting [991, 608]. heat-conductive [938, 167]. heat-conductivity [420, 942]. heated [1364]. Hecke [1106]. Heegaard [366]. Hegerfeldt [351, 2]. Heineman [904]. Heisenberg [541, 297, 878, 545, 129, 1485, 1289, 1212, 662, 778]. Heisenberg-type [297]. helicity [760, 1140, 925]. helicity-A [925]. helium [377]. Helmholtz [797, 1201]. Hénon [1347]. Herglotz [330]. Hermite [526, 697, 334, 529]. Hermite-like [697]. Hermitian [970, 898, 248, 65, 225, 1506, 53, 1016, 596, 786, 133, 547, 1285, 1356, 590]. Hermiticity [1260, 641, 31]. Hessian [521]. Heun [808, 392, 239, 713, 1055, 10, 266, 1058, 1521]. Heun-type [392]. hexagonal [409]. Hiai [1041]. Hidden [685, 589, 667]. hierarchical [223, 1236]. hierarchies [462, 764, 1170]. hierarchy [412, 962, 1015, 159, 454, 1473, 1447, 684, 609, 1131, 718, 1239, 107]. Higgs [763, 1399, 1349, 30, 804, 1256, 1466]. High [45, 22, 986, 1446, 1168, 1495, 238, 181, 1423]. high-genus [986]. high-low [238]. high-order [1423]. High-temperature [45]. Higher [79, 570, 1371, 950, 1440, 182, 1407, 1382, 1475, 784, 995, 655, 678, 88, 1234, 235, 689, 859]. Higher-order [79, 1407, 1382, 1475, 995, 655, 678, 859]. Hilbert [917, 81, 995, 979, 171, 1285, 1448, 1423]. Hilfer [1190, 703]. Hill [752]. Hilliard [1214, 1381]. Hirota [765]. history [491]. Hitchin [1525]. Hitchin-Thorpe [1525]. HIV [1323]. Hochschild [845]. Hohenberg [291, 815]. hole [1474, 673, 1055, 691, 594, 449]. holes [14, 606, 329, 874, 1228]. Holevo [622]. Holm [273, 1299, 1382, 206, 870, 1246, 340, 576, 55, 972, 728, 108, 165, 199]. Holm-type [340]. Holomorphic [88]. holonomic [599, 399]. holonomy [286]. Hom [695, 389, 998]. hom-algebroids [998]. hom-Lie [389]. homogeneous [431, 1011, 1414, 1124, 481, 293, 274, 748, 907]. Homogenization [415, 44]. Homological [513]. Homology [926, 72]. Homology-changing [926]. homomorphism [1222]. Homotopy [1401, 999, 513]. honeycomb [759]. honor [1528]. Hookean [1028]. Hopf [1231, 494, 1477, 598, 125, 1092]. Hopfield [405]. horizon [1360]. horizontal [1142, 1467]. Horowitz [372]. Hubbard [380]. human [1090, 1374]. Hurwitz [51, 46, 1333, 1277]. hybrid [1287, 34]. Hydrodynamic [961, 415, 275, 1295, 1302]. hydrodynamical [531]. hydrodynamics [255]. hydrogen [288]. hydrogenic [496, 288]. hydromagnetic [1057]. hydrostatic [416]. hyper [365]. hyper-Kähler [365]. hyperbolic [516, 993, 1268, 1459, 207, 1164, 1113, 894, 1511, 1515, 628, 1179, 238, 76, 1086]. hyperbolicity [528]. hyperboloid [702, 1326, 216, 103]. hyperchaotic [598]. hypercontractivity [850]. hypercube [1237]. hypergeometric [1415, 571]. Hypergraph [744]. hypergraphs [948]. hypersurface [1512]. hypersurfaces [1360]. hypocoercivity [1375]. Hypoellipticity [807]. hysteresis [1013].

i.i.d [1016]. ice [800, 1424]. ideal [1446, 485, 938, 1278]. ideas [617]. identifying [1369]. identity [1349]. II [1038, 719, 1402, 1288, 1281, 759, 170, 748, 366, 228]. III [396, 1198, 1058]. Ill [201, 993, 1376, 778]. Ill-posedness [201, 993, 778]. ill-prepared [1376]. illposedness [121]. Images [753, 1403]. imaginary [1449]. immersed [580]. impedance [523]. implies [555, 252]. Improved [850, 63, 1525]. impulse [75]. impulsive [376, 751, 434, 1270]. incidence [163, 1269]. inclusion [38, 436]. inclusions [29, 1380]. incompatibility [1355]. Incomplete [77, 1502]. Incompressible [1376, 1028, 1100, 485, 1207, 1467, 1044, 1274, 75, 416, 1102, 452, 769]. increasing [138]. indefinite [1145, 814, 693, 484, 1464]. independent [572, 666]. index [1037, 856, 819, 1420, 1455, 778]. indexed [444, 149, 1449]. indices [1421, 1518, 423, 1258]. indirect [1143]. induced [565, 675, 551, 138]. inelastic [907]. inequalities [1211, 1264, 428]. inequality [740, 619, 772, 850, 1041, 983, 1525, 1005, 325]. inertia [893]. infectious [1325]. inference [1004]. Infinite [500, 1415, 244, 308, 1507, 1364, 1370, 1033, 1062, 642, 59, 1335, 183, 1419, 1205, 353, 1327, 199, 1309]. Infinitely [487, 538, 980]. infinitesimal [806]. infinity [395, 992, 723, 1510, 218, 1120, 1007, 935, 793]. inflation [1473]. inflaton [106]. influences [1495]. information [177, 1182, 742, 521, 215, 623, 1080, 550, 788]. information-theoretic [623]. infrared [1315]. Inhomogeneous [776, 1496, 1068, 1207, 1274, 1304, 769, 566, 1296]. initial [1011, 762, 660, 969, 941, 207, 85, 737, 1209, 1213, 1141, 953, 122, 839, 1340]. initial-value [1011, 953]. insertions [319]. Instability [1231, 515, 208, 797, 202]. instanton [342]. insulated [1380]. insulators [1383, 1338, 171, 1420]. integer [1388]. integer-valued [1388]. Integrability [854, 833, 42, 256, 804, 445]. Integrable [1178, 1241, 875, 1117, 412, 229, 1191, 1192, 227, 337, 1239]. Integral [322, 719, 964, 1055, 249, 1284, 189, 269, 1259, 450, 64]. integral-partial [450]. integrals [68, 179, 1322, 132, 931, 1443, 334, 399, 221, 806]. Integrated [569, 1357, 671, 884]. Integration [1483, 621, 571]. integro [434]. integro-differential [434]. integrodifferential [118]. interacting [1096, 1020, 447, 7, 319, 60, 176, 1420]. interaction [1507, 110, 549, 692, 105, 383, 1021, 1053, 1049]. interactions [984, 872, 84, 954, 40, 866, 282, 428, 300, 1072, 907]. intercritical [1296]. interest [33]. interesting [1233]. interface [91]. Interior [1266, 1357, 1402, 1438, 932, 888, 1428]. Interior-boundary [1266]. interlacing [1418]. intermediate [779]. Intermittency [754]. Internal [1225, 1480]. internally [1364]. Interpolating [334]. interpretation [783, 145, 249]. intersection [672, 1445]. intersections [147, 139]. interstellar [299]. intertwining [783]. interval [1433]. intricacy [889]. intrinsic [218, 770]. Introduction [989, 955]. invariance [1391]. Invariant [570, 259, 233, 1227, 683, 586, 1412, 429, 663, 767, 1439, 362, 508, 458, 823, 791, 1492, 1186, 1394, 1395, 625, 1118, 815, 582, 775]. invariants [213, 1512, 1056, 1387, 1291, 1156, 826, 865, 595, 1197]. invasion [1136]. Inverse [937, 988, 941, 387]. inverted [1289]. invertible [696]. Investigation [1052, 421, 1031]. invicid [938]. invisibility [393]. involving [1306, 579, 653, 118, 355, 575, 1066, 814, 20, 119, 184, 658, 1425, 779]. ion [602, 614]. ion-slip [614]. ions [288]. Irreducible [1151, 568, 1148, 1081]. isentropic [520, 577, 1376, 812, 530, 230, 22, 231, 94, 770, 775]. Ising [541, 1288, 967, 224, 105, 300, 335]. isochrone [825]. isolated [363, 982]. isometric [1305]. isometries [1492]. Isomonodromic [157, 1522]. isospectral [583]. isothermal [992, 838, 1381, 23]. Isotropic [461, 183]. isotropy [1404]. Issue [989]. issues [388]. iteration [195, 87, 196]. iterative [1093]. Ito [239]. IV [1438, 529]. Ivanov [963]. Izergin [151].

J [652, 1337, 653, 115, 385, 197, 195, 651, 447, 808, 970, 606, 448, 835, 1095, 54, 114, 196, 479, 1374, 1203, 1293, 865, 932, 1133, 384, 8, 607, 269, 1487, 80, 351, 6, 1336, 988]. Jacobi [106, 1166, 544, 527, 682, 834, 42, 444, 402, 404, 931, 1186, 923]. Jacobsthal [1233]. jammer [177]. Jean [989]. jellium [711]. Jordan [1199, 401, 561].

Kac [962, 29, 339, 1259, 806]. Kadomtsev [1170, 903, 735]. Kagome [1130]. Kähler [1517, 365]. Kaluza [558, 631]. KAM [182, 1457]. Kantorovich [1181]. Kaup [859]. Kawai [1401]. KdV [795, 117, 1241, 235]. Keller [841].

Kelvin [635, 797]. Kepler [1002]. kernel [168, 719]. kernels [191, 1242]. Kerr [1293, 932, 84, 1226, 888, 86]. Kertész [1329]. Keyfitz [866]. Khovanov [72]. Killing [396, 256, 439]. kind [729]. kinds [980]. kinetic [1121, 673, 451, 338, 944, 925]. kinklike [1273]. Kirchhoff [539, 1023, 877, 1378, 357, 813, 1216, 1491, 1431, 730, 1488, 935, 484, 16, 483, 418, 1105, 909, 1173, 1027]. Kirchhoff-like [1431]. Kirchhoff-type [1023, 877, 813, 1216, 1488, 484, 16, 1173, 1027]. Kirillov [586]. Kirkpatrick [264]. Kirkwood [1355]. Kitaev [1192, 242]. Klein [1474, 604, 236, 558, 328, 844, 931, 316, 631, 729]. klystrons [1045]. knots [1387, 832]. Knotted [1056, 337]. known [392, 1019]. Kontsevich [1277]. Korepin [151]. Korteweg [811, 1206, 352, 769]. Korteweg-de [811]. KP [962, 136, 443, 462, 764, 1447, 1241, 609, 107]. KPZ [1520]. Krahn [428]. Kramers [1101]. Kranzer [866]. Krawtchouk [440, 1196]. Krein [820]. Kretschmann [1221]. Kronecker [343]. Kropina [1358]. Kruskal [768]. KTAP [893]. Kuiper [1305]. Kundt [1227]. Kuper [340]. Kuznetsov [160].

labeling [1390]. Lack [798, 1306]. Ladik [961]. Laflamme [515]. Lagrange [376, 524, 1185]. Lagrangian [1451, 855]. Lagrangians [1408, 228]. Laguerre [1065, 444, 637, 226]. lake [1030]. lambda [136]. Landau [473, 954, 358, 1351, 624, 1147, 169, 973, 307, 88, 846, 1140, 222, 538, 632, 1310]. landscape [1168]. Langevin [184, 707, 1280]. Laplace [1509, 814, 1084, 965, 1027]. Laplacian [653, 1306, 579, 346, 109, 116, 20, 603, 1516, 1253, 658, 935, 1025, 305, 841, 471]. Large [597, 1040, 1155, 707, 732, 1366, 225, 601, 53, 549, 1209, 1523, 1059, 1508, 506, 607, 1098, 167, 1210, 1343, 564, 354, 226]. Large-scale [1155]. large-solutions [1210]. large-time [167]. larger [873]. largest [637]. lateral [771]. Latin [1354]. lattice [1187, 1198, 1370, 368, 409, 229, 741, 1399, 1243, 759, 795, 319, 696, 1156, 749, 1270, 1400, 985, 1131, 435, 1215, 1441]. lattices [292, 429, 443, 864, 362, 553, 582]. Law [1508, 403, 1300, 355, 233, 681, 712, 19, 1493, 1417, 478]. law-preserving [233]. laws [740, 1364, 1069, 720, 599, 316, 122, 869]. Lax [708, 758, 553, 185]. layer [459, 1339, 1442, 680]. layers [1102]. leading [685, 846]. learning [1461]. learning-assisted [1461]. Least [1023, 488, 730, 16, 1372, 723, 1408]. Lebesgue [1034]. Leblond [24]. lecture [904]. Lee [498, 958]. Left [686, 494, 1227]. Left-covariant [686, 494]. Legendre [267, 964, 1013]. Leibniz [507]. Lemaître [272]. lemma [586]. Lenells [1448, 1423]. length [280]. lengths [33]. lenses [1403]. Lenz [1509]. Leslie [1385]. Leur [962]. Level [883, 111, 408, 133, 1498, 176, 669, 1331]. levels [88]. Levinson [617]. Levitan [1201]. Lévy [24, 219]. Lewellen [1401]. Lewis [1291]. Lichnerowicz [921]. Lie [115, 114, 8, 6, 695, 495, 713, 1069, 1497, 736, 525, 1334, 1291, 28, 715, 650, 1386, 389, 1451, 1107, 782, 968, 5, 260]. Liénard [115, 114, 404, 8, 5, 6]. Lifespan [1144, 906]. Lifshitz [1147, 307, 1228, 1140]. Lifshitz-like [1228]. lift [558]. Light [1166, 1094, 1412, 1524, 398, 630, 944]. like [1509, 756, 440, 833, 214, 1431, 748, 1228, 697, 365]. likelihood [1004].

Limit [709, 336, 1478, 543, 1365, 773, 746, 1276, 172, 959, 339, 1040, 1376, 879, 842, 1206, 1063, 1468, 352, 313, 892, 936, 1508, 22, 122, 189, 269, 522, 1028, 1127, 1278, 1100, 390, 587]. limitations [364]. limited [771]. Limiting [1286, 285, 1324, 530, 94]. Limits [1432, 1121, 369, 1342, 734, 60, 231, 464]. Lindblad [669]. line [1506, 1169, 1284, 1329, 811, 973, 263, 185, 1201, 690, 735, 235]. line-gapped [1506]. Linear [1424, 1230, 862, 440, 960, 552, 1407, 355, 233, 234, 799, 503, 236, 28, 627, 896, 588, 1185, 572, 776, 424, 484, 1527, 824, 1445, 1086]. linearities [870]. linearity [685]. linearizability [445]. linearizable [432, 1292]. Linearized [123, 117, 58]. linearly [693]. lines [1372]. link [516]. linked [1500]. Liouville [1421, 887, 910, 482, 882, 1085, 1274, 891, 77, 656, 1049, 262]. Liouville-type [1274, 891, 1049, 262]. Lippmann [393]. Lipschitz [441]. liquid [992, 339, 261, 1066, 1139, 231, 1343, 1469, 23]. liquid-gas [231]. Littlewood [386, 1332]. Local [1365, 577, 367, 812, 1156, 303, 55, 994, 1384, 1397, 1140, 505, 480, 1371, 98, 1180, 20, 291, 1317, 383, 1341, 1445, 793, 899, 1093]. Locality [68, 1053]. localization [287, 949, 727, 472, 1155, 1316, 1501, 1502, 351, 2, 901]. Localized [116, 733, 1369, 3, 385, 1251, 864]. localizer [1506]. locally [766, 527]. locus [1267, 1386]. Loewner [263]. Log [669, 899]. Log-convex [669]. log-Coulomb [899]. logarithm [1300, 1307]. logarithm-law [1300]. Logarithmic [1205, 345, 619, 579, 653, 719, 1176, 1468, 1417, 573, 770, 622, 1025]. logistic [1018, 533, 868, 1462]. loglog [617]. Lohe [331, 40, 596]. Long [762, 234, 1496, 156, 1251, 848, 1283, 1240, 1115, 954, 741, 1236, 472, 576, 688, 105, 1417, 907, 33, 913]. Long-range [848, 954, 1236, 472, 688, 907, 33]. Long-time [762, 234, 156, 1251, 1283, 1240, 913]. longitudinal [541]. loop [557, 41, 126, 687, 1284]. Lorentz [969, 129]. Lorentzian [1042]. lossless [401]. Lotka [1335]. Low [993, 1020, 1411, 350, 990, 567, 104, 379, 546, 1072, 238, 1379, 1240, 975]. Low-dimensional [1411, 104]. Low-energy [1020, 546]. Lower [862, 357]. Lowest [493, 568]. Lowest-dimensional [493]. Lucas [1233, 958]. LUE [1527]. lumped [659]. lumps [903]. lunar [1518, 752]. lying [660].

M5 [555]. M5-branes [555]. MacDonald [1377]. Mach [1206, 22]. Machine [1461]. MacMahon [915]. Magic [1354, 364, 1219, 724]. Magnetic [309, 497, 1486, 438, 1279, 1337, 174, 854, 1078, 1145, 3, 385, 73, 780, 1297, 11, 197, 447, 798, 7, 1378, 246, 496, 210, 610, 663, 867, 1101, 1456, 1200, 1467, 1232, 936, 1248, 375, 538, 1027]. magnetized [321]. magneto [205, 275, 21, 871]. magneto-hydrodynamic [275]. magneto-micropolar [21, 871]. magnetogasdynamics [936]. magnetohydrodynamic [809, 482, 1376, 302, 1210, 942, 89, 480, 1478]. magnetohydrodynamics [1339, 19, 656]. magnetostatics [760]. Maillet [268]. Majorana [213]. managed [659]. Manifold [821, 41, 198, 1068, 580, 550]. manifolds [853, 395, 1410, 761, 219, 1042, 46, 527, 42, 558, 1211, 169, 1439, 852, 1250, 823, 1183, 1447, 26, 1324]. Manin [78]. manipulation [98]. many [757, 214, 980, 487, 538]. Mapping [641, 392]. maps [461, 822, 552, 906, 292, 805, 468, 373, 745, 138, 1046]. Marchenko [1482, 1201]. Marginal [719, 740, 1181]. Markov [1282, 1409, 300, 184, 1366]. Markovian [552, 1280, 427]. martingales [613]. Maryland [647]. Maryland-type [647]. maser [1040]. Mass [1399, 56, 244, 216, 1062, 1268, 490, 397, 630, 370, 1511, 1488, 1127, 247, 738, 1319]. mass-critical [490]. masses [827, 1235, 294]. Massive [150, 541, 885, 228]. massless [951, 150, 645, 481, 347]. massless-scalar [951]. master [378, 83, 1203, 1331]. Math [652, 1337, 653, 115, 385, 197, 195, 651, 447, 808, 970, 606, 448, 835, 1095, 54, 114, 196, 479, 1374, 1203, 1293, 865, 932, 1133, 384, 8, 607, 269, 1487, 80, 351, 6, 1336, 988]. Mathematical [952, 1090, 1374, 438, 904, 455]. Mathematics [243, 298]. Mathieu [560, 1014, 633]. Mathieu-type [1014]. matrices [188, 1052, 640, 960, 1435, 53, 544, 1016, 561, 1350, 712, 749, 1159, 572, 985, 1238, 1059, 646, 717, 550, 1527, 226, 96]. matrix [709, 766, 878, 225, 1036, 898, 542, 799, 682, 1154, 1312, 40, 1168, 1177, 394, 1485, 609, 1149, 803, 779, 590]. matter [174, 1111, 359, 1391, 581, 449, 1220]. maximal [1085]. Maximally [432]. Maximum [622]. Maxwell [835, 995, 1466, 172, 1104, 1315, 408, 9, 1460, 487, 1120, 209, 729]. Maxwell-scalar [1120]. mCH [725]. MDS [1160]. mean [708, 172, 560, 379, 985, 1480, 1504, 1127, 1514, 390]. mean-field [379, 1504, 390]. Measure [907, 882, 896, 1237, 305]. measurement [83, 1203, 252]. measurements [1524, 425, 554, 327]. measures [437, 43, 99, 551, 265, 1270, 625, 1118, 815, 582, 788, 1478]. mechanic [739]. mechanical [666, 1394, 1395, 562]. mechanics [1194, 134, 1452, 563, 1000, 187, 500, 1075, 1272, 1281, 642, 1519, 981, 599, 1526, 1449, 825, 391, 337, 788, 828, 247, 738]. mechanism [635, 1143, 452]. Meckler [1073]. media [679, 1094, 52, 944, 580, 306, 1476]. medium [93, 408, 1024]. Mehler [1459]. Meixner [149]. meltings [1294]. memories [1194]. memory [1342, 336, 989, 1090, 1374, 59, 1361]. Mermin [1073]. Merola [1131]. Meromorphy [927]. mesoscopic [1365]. Message [177]. messages [101]. Metastability [1279, 1337, 565]. metastable [1013]. method [1195, 1482, 722, 154, 195, 808, 87, 196, 1460, 918, 10, 35, 794, 238, 621, 765, 455]. methods [409, 825]. Metric [507, 1513, 1094, 97, 521, 373, 1285, 439, 365, 305, 594]. metric-affine [1513]. Metric-connection [507]. Metrics [372, 1008, 898, 1129, 138]. metrizability [1358]. metrology [425]. MHD [420, 1298, 459, 1297, 577, 837, 1188, 1385, 485, 1207, 1467, 1492, 677, 731, 1425, 839, 614]. micropolar [615, 21, 871]. MICZ [1002]. Mills [1349, 570, 1006, 1134, 1068, 594]. min [744]. min-cuts [744]. Mini [1519]. Mini-course [1519]. minimal [52, 650, 1511, 1360, 280]. Minimally [1292]. minimizers [1491]. Minkowski [367, 969, 1510, 1290]. mirror [513]. Mixed [96, 830, 762, 680, 230, 1175, 1428]. Mixing [949, 1061, 676, 373, 227, 636]. mixture [737]. mixtures [991, 1251]. mKdV [1213, 913]. mobile [1051]. Möbius [589]. MOCH [728]. Mode [111, 465]. Model [1031, 264, 405, 853, 1121, 971, 430, 1279, 1337, 91, 1089, 991, 1229, 489, 1104, 1136, 897, 205, 286, 408, 329, 284, 1399, 331, 103, 596, 663, 1329, 956, 1192, 451, 338, 105, 379, 463, 945, 957, 348, 1252, 1018, 908, 1323, 242, 394, 994, 1373, 300, 966, 1419, 580, 953, 609, 732, 1143, 122, 1049, 56, 671, 884, 581, 1458, 201, 1217, 231, 1466, 380, 770, 1345, 556, 1277, 868, 1302, 724, 1189, 163, 1269, 89, 1469, 1462, 1024, 452, 1489, 1432, 1325, 1108]. Modeling [706, 1461, 1330, 939]. models [1418, 106, 709, 528, 722, 1199, 1513, 893, 225, 1344, 1288, 967, 507, 1282, 567, 90, 1520, 224, 1090, 1374, 104, 1050, 1077, 1485, 748, 335, 1149, 857, 366, 828]. Moderate [1280]. modes [600, 874, 313, 755]. modification [675]. Modified [1513, 1015, 655, 619, 713, 273, 962, 790, 1211, 530, 1063, 291, 972, 728, 94, 1283]. modularity [287]. modules [568, 1148, 495, 1033, 140, 171, 914, 1500, 1498, 1032]. moduli [763, 1386]. moisture [856]. molecular [683, 1282, 1331]. molecule [1507]. molecules [902, 955]. moment [447, 7]. Moments [345, 640, 682, 1417]. Momentum [897, 424]. monodromy [955]. monoids [220]. monopole [536]. monopoles [900, 1232, 594]. monotone [138]. Monotonicity [536, 1041, 1329]. Mori [807]. Morse [1421, 756, 513]. Morse-like [756]. Moser [609, 107]. most [1394, 1395]. motility [1143, 532]. motion [91, 41, 358, 312, 663, 1232, 931, 562]. motions [1305, 931]. Mountain [1466]. Moutard [296]. Moutard-type [296]. movement [581]. mover [1181]. moves [1056]. moving [1362, 447, 7]. Moyal [787]. Multi [1348, 526, 313, 154, 225, 92, 964, 1461, 883, 444, 387, 764, 1216, 1493, 1384, 149, 1449, 538, 735, 939, 582]. multi-bump [1216]. multi-component [764]. Multi-dimensional [526, 883, 1493, 939]. multi-dimensions [313]. multi-indexed [444, 149, 1449]. multi-integral [964]. multi-matrix [225]. multi-parameter [154]. Multi-peak [1348, 1384]. Multi-scale [313, 1461]. multi-soliton [92]. multi-stochastic [582]. multi-term [387]. multi-vortex [538]. multicavity [1045]. multiform [412]. Multigraded [127]. multimonopole [1311]. Multiparameter [425, 1239]. multipartite [747]. Multiplane [1403]. Multiple [386, 1428, 824, 263, 313, 1190, 859]. multiplication [188, 478]. multiplicative [973, 1263, 291, 815]. Multiplicity [20, 277, 1105, 395, 579, 653, 840, 877, 610, 573]. multipliers [1518]. Multipole [14, 606]. multiscale [219]. multisymplectic [852]. Multivariate [1264]. mutations [445]. mutual [215, 1080]. Myers [372].

n [983, 345, 317]. Nagumo [1215, 689]. Nahm [1311, 900]. Nambu [1440]. nanosystems [1051]. Nash [1305]. Navier [1426, 1362, 1214, 838, 1401, 575, 261, 812, 1206, 303, 352, 1044, 1274, 891, 237, 60, 486, 1341, 1379, 1146, 779, 1009, 939, 1204, 1340, 1310, 1381, 532, 354, 161]. near [1008, 1099, 1120, 1360]. nearly [761]. Necessary [552]. negative [1279, 1337, 1427, 124, 372, 987]. Nehari [311]. Nehari-Pohozaev [311]. Nematic [339, 992, 1343, 23]. Nested [1492]. network [1130, 1481, 40, 663]. networks [1270, 1059]. Neumann [830, 48, 811, 603, 1486, 1464]. neural [1270, 1059]. neutral [1348, 1363, 751, 434, 721, 1318, 704]. Newell [859]. Newman [1293, 397, 1226]. Newton [526, 1123, 314]. Newtonian [827, 675, 1254, 1327]. next [846]. next-to-leading [846]. Nijenhuis [1177]. Nikiforov [808, 10]. Nil [396]. Nil-Killing [396]. nine [1002]. nine-dimensional [1002]. Nishimori [1169]. NLS [549, 692, 726, 941, 238]. NLSE [1060]. NLTS [1261]. no [439, 1181]. no-go [1181]. Nodal [470, 723, 116, 733, 406]. Noether [1156, 399]. Noether’s [1336, 382]. Noether’s-type [1336, 382]. Noise [504, 1444, 638, 219, 273, 1174, 1439, 535, 973, 291, 141, 815, 138, 582, 1215]. noises [823, 193, 1237]. Non [110, 910, 734, 292, 148, 1347, 242, 1317, 26, 796, 380, 839, 49, 1187, 1499, 993, 917, 190, 248, 65, 754, 1506, 53, 172, 473, 992, 234, 1016, 624, 1414, 780, 69, 1401, 1254, 970, 1446, 1191, 870, 25, 448, 1427, 685, 236, 866, 629, 820, 530, 786, 524, 20, 291, 374, 823, 1077, 1183, 313, 599, 1404, 1133, 1076, 847, 1503, 133, 547, 1285, 1516, 22, 1488, 776, 337, 222, 416, 94, 1422, 1087, 564, 1009, 435, 138, 632, 1445, 775, 536, 1423, 689, 1381, 23, 278, 590, 1198, 630]. non-abelian [1401, 69, 536, 1198]. Non-Archimedean [49]. non-autonomous [1187, 473, 291, 374, 1516, 222, 1087, 1009, 632, 689]. Non-commutative [292, 624, 780, 1503]. non-conservative [524]. non-constant [564]. Non-convergence [796]. non-degenerate [234]. non-differentiable [754]. non-equilibrium [1191, 313, 435]. non-Euclidean [685]. non-exceptional [1499]. Non-existence [380]. non-Gaussian [823]. non-Hermitian [970, 248, 65, 1506, 53, 1016, 786, 133, 547, 1285, 590]. non-holonomic [599]. non-homogeneous [1414]. non-hydrostatic [416]. non-ideal [1446]. non-increasing [138]. non-integrable [337]. non-isentropic [530, 22, 94, 775]. non-isothermal [992, 1381, 23]. non-linear [234, 236, 776, 1445]. non-linearities [870]. Non-local [1317, 20]. non-negative [1427]. non-Newtonian [1254]. non-normalizable [1077]. non-parity-time-symmetric [190]. non-PDE [630]. Non-perturbative [242]. Non-power [1347]. Non-real [110]. Non-relativistic [734, 172]. Non-scattering [26]. Non-self-adjoint [148, 25, 448, 820]. non-semisimple [1183]. non-separable [917]. non-standard [278]. non-strictly [993]. Non-symmetric [910, 866]. non-trapping [1488]. non-trivial [1404, 1133, 1076]. Non-uniform [839]. non-universality [847]. non-vacuum [629]. non-vanishing [1423]. non-zero [1422]. Nonautonomous [1416, 1188]. nonclassical [1361]. nonclassicality [1355]. Noncommutative [1373, 430, 987, 102, 391]. noncommutativity [1355]. Noncontextual [948]. noncritical [578]. nondegeneracy [1457]. Nonequilibrium [1013, 289]. Nonexistence [657, 200]. nonholonomic [1197]. nonhomogeneous [485, 1209, 615, 302, 871, 354]. nonisentropic [1028]. Nonlinear [1194, 548, 631, 1482, 491, 1300, 719, 143, 1011, 204, 233, 659, 1145, 92, 810, 490, 1496, 660, 156, 1406, 995, 655, 758, 867, 357, 207, 1439, 310, 414, 1273, 59, 19, 454, 612, 232, 82, 57, 458, 1216, 313, 419, 963, 306, 130, 732, 416, 961, 1428, 690, 141, 332, 360, 163, 1269, 582, 1215, 200, 1308, 155, 271, 841, 278]. nonlinearities [1421, 576, 996]. nonlinearity [579, 653, 1176, 1496, 726, 454, 658, 360, 578, 1247, 1025, 909, 1296]. nonlinearly [864]. Nonlocal [972, 439, 1421, 118, 1363, 182, 1273, 612, 908, 164, 301, 815, 163, 120, 271, 1325]. nonlocality [801]. nonlocally [154]. nonnegative [96]. nonperiodic [1171]. Nonrelativistic [1121]. nonsimple [517]. nonstandard [232, 1441]. nonzero [290, 963]. Nordström [691, 594]. norm [1039, 551, 621]. Normal [761, 634, 755, 309, 854, 749]. normalizable [1077]. Normalized [1488, 120, 483, 909, 208, 1026]. Norman [249]. norms [1436]. Note [653, 448, 1293, 865, 932, 607, 269, 988, 1315]. notion [1526]. Novel [1273, 1415]. Novikov [777]. nozzle [1494, 417]. NP [1510]. NP-constants [1510]. NS [736]. NSR [1022]. null [213, 1267, 1358, 218, 1007, 628, 1321, 793]. null-infinity [218]. Number [290, 1364, 827, 1350, 406, 22, 1393]. numbers [51, 73, 1333, 706, 600, 1508, 1233, 672]. Numerical [421, 517, 801, 1197]. NUT [365]. NUT-like [365].

observability [1230, 876]. observables [318, 821, 245, 366]. observations [1518]. Observing [947]. obstacle [152]. obstruction [499]. obstructs [1155]. ocean [680]. Octions [1108]. octonionic [979]. octonions [943]. Odd [343, 110, 1389, 1395]. ODE’s [293, 431]. Okubo [1041]. Oldroyd [452]. Oldroyd-B [452]. one [928, 860, 912, 437, 178, 46, 1288, 1287, 1126, 1066, 1358, 1084, 256, 696, 105, 944, 454, 1275, 546, 119, 400, 22, 1049, 1295, 1420, 1012, 460, 435, 1303, 96]. one- [860, 1288]. one-dimensional [912, 1126, 105, 22, 1420, 1012, 460, 435]. one-form [1358]. one-forms [1084]. one-packing [437]. one-photon [944]. one-way [1049]. onto [407]. Open [563, 289, 1258, 1409, 1021, 262]. operational [1219]. operations [98, 1221, 1220]. Operator [394, 438, 713, 1001, 569, 25, 448, 1436, 140, 99, 1262, 1084, 715, 727, 1101, 560, 1250, 186, 342, 442, 242, 301, 1249, 1393, 192, 803, 1027]. operators [1452, 1157, 276, 783, 1481, 872, 118, 113, 1062, 147, 1497, 1351, 497, 736, 1150, 999, 1034, 1414, 341, 670, 964, 834, 882, 741, 1085, 282, 1109, 472, 688, 647, 605, 124, 1200, 929, 1417, 428, 322, 987, 389, 1014, 422, 782, 666, 965, 185, 1501, 441, 424, 457]. Optical [698, 1094, 192]. optics [457]. Optimal [1085, 1221, 1426, 136, 1202, 566]. optimization [895]. optimized [1057]. orbit [35]. Orbital [1382]. orbits [1509, 404]. order [1187, 1071, 754, 1115, 333, 431, 1011, 1407, 204, 1382, 490, 706, 1475, 339, 494, 870, 995, 103, 655, 678, 1456, 654, 896, 1307, 119, 1317, 846, 79, 1082, 1487, 556, 228, 1448, 1441, 1423, 859]. ordered [1436, 523]. ordering [1290]. ordinary [333, 927, 112, 119]. organization [337]. orientation [1130]. oriented [1387]. origin [959, 925]. Ornstein [830, 1018, 1323, 1335, 1128]. Orthogonal [1059, 709, 1106, 344, 1065, 338, 77, 129, 1523, 1451, 149, 1449, 1238, 1256]. orthogonality [444, 1449, 948]. orthosymplectic [1157]. oscillation [1525, 518, 1429]. oscillations [279, 332, 1343]. oscillator [110, 477, 1078, 667, 11, 197, 559, 883, 328, 258, 945, 37, 966, 1503, 1289, 849, 1179, 738, 1319]. oscillators [84, 526, 95, 1291, 4, 384, 1212, 787]. Otelbaev [882]. other [790]. out-going [60].

P [639]. Pachner [1056]. packed [409]. packing [437]. Painlevé [1198, 1097, 927, 1058, 1521, 529]. Painlevé/ [1097]. pair [95, 1504]. pairs [1349, 553, 185]. Palatini [371]. para [440, 1035, 119]. para-Grassmann [119]. para-Krawtchouk [440]. para-Racah [1035]. parabolic [24, 862, 969, 1002, 804, 232, 57, 1307, 1516, 1256, 1025]. paradox [351, 2]. parallel [459, 644]. parameter [175, 652, 330, 154, 219, 147, 651, 727, 931, 1361]. parameter-dependent [147]. parameterization [1437]. parameterized [584]. parameters [715, 1317]. parametric [1078, 1350]. parametrization [465]. parity [190]. Parseval [979]. part [571, 779]. partial [178, 980, 1188, 1175, 1317, 1425, 450]. Partially [1515, 1204, 1409]. Particle [1199, 449, 1509, 1038, 172, 69, 1191, 103, 663, 947, 1460, 885, 1232, 404, 1318, 581, 262, 1127]. Particle-hole [449]. particle-in-cell [1460]. particles [860, 398, 864, 629, 943, 481, 721]. partition [223, 466, 151, 1091]. partitions [618]. pass [1466]. passage [214]. passage-like [214]. passing [152]. passive [142]. past [491]. patch [1030, 12, 80]. path [132, 334]. paths [959, 1196]. Pathwise [1439]. pattern [162]. patterned [960]. patterns [405, 734]. Pauli [607, 1001, 429, 745, 721, 506]. PBW [1257]. PBW-type [1257]. PDE [154, 630, 1098, 455]. PDEs [1194]. peak [1348, 1524, 1384]. peakon [1247]. peakons [1382, 1246, 725]. peaks [1428]. Peeling [86]. pendulum [376]. pendulum-type [376]. Penrose [397]. percolation [1236, 1367, 926]. Perfect [514, 1357]. period [822]. Periodic [1270, 270, 1269, 308, 1365, 569, 761, 356, 311, 1231, 1382, 544, 565, 3, 385, 341, 388, 206, 949, 408, 726, 1123, 1490, 605, 711, 82, 929, 37, 903, 1249, 817, 427, 192, 803, 824, 1189, 1024, 1303, 1309]. periodically [458]. Periodicity [445]. permuted [968]. Permuting [312]. Persistence [870]. perspective [1272, 1281, 630, 83, 1203, 623]. perturbation [1010, 182, 1093, 1055, 663, 501, 1347, 936, 1234, 120]. Perturbations [561, 928, 673, 719, 497, 1287, 123, 258, 436, 347]. Perturbative [368, 1165, 647, 179, 242]. perturbed [110, 336, 1523, 1249, 1527, 735, 1361, 1058, 1521, 1310, 226]. perturbing [47]. Petersson [986]. Petviashvili [1170, 903, 735]. Pfaffian [1284]. Phase [916, 504, 213, 1052, 1096, 661, 339, 205, 798, 1137, 1427, 1040, 856, 277, 994, 300, 547, 176, 391, 337, 231, 770, 1469, 1324, 925, 899]. phase-field [1469]. phase-space [391]. phases [213, 1111, 696]. phenomena [520, 206, 770, 1345, 725]. phenomenon [1245, 1064]. Philippoff [1057]. phonon [475]. photon [944, 47]. photonic [73]. Phys [652, 1337, 653, 115, 385, 197, 195, 651, 447, 808, 970, 606, 448, 835, 1095, 54, 114, 196, 479, 1374, 1203, 1293, 865, 932, 1133, 384, 8, 607, 269, 1487, 80, 351, 6, 1336, 988]. Physical [438, 318, 148, 313]. physics [1391, 904, 33, 455]. pieces [647]. piecewise [60, 1353]. pipes [1309]. piston [22]. Pitaevskii [1071, 990, 454]. Planar [1271, 1043, 1235, 314, 942, 1457, 1489]. Planck [168, 1101, 1141, 486, 1379]. plane [223, 459, 78, 969, 685, 90, 310, 1168, 1503, 363, 180, 666, 1179, 618, 71]. plasma [1166, 104]. plasmas [89]. plasmon [183]. plasmon-polaritons [183]. plate [1507, 1431]. Plücker [461]. Pohozaev [311]. Poincaré [129, 1046]. Point [1517, 1060, 142, 827, 541, 110, 822, 115, 1291, 45, 296, 1266, 864, 282, 1236, 114, 1226, 1293, 885, 428, 753, 8, 1049, 5, 6, 1457]. point-potentials [296]. points [701, 1453, 1047, 590]. pointwise [616]. Poiseuille [1138]. Poisson [93, 1203, 1141, 1100, 638, 172, 1463, 123, 1171, 83, 488, 1275, 1245, 730, 301, 363, 203, 1146, 1100, 353, 260, 840]. Pol [1086]. polar [666]. polaritons [183]. polarization [1473, 377]. polaron [496]. pole [963]. Pollaczek [149]. pollution [436]. polyatomic [1330]. polygon [584]. polygonal [293]. polygons [603]. Polynomial [1056, 462, 431, 1407, 808, 10, 595, 522]. polynomials [516, 345, 986, 709, 440, 1035, 1022, 643, 145, 344, 526, 1073, 1065, 444, 468, 402, 338, 265, 77, 1196, 1523, 149, 1449, 697, 334, 424]. polytope [886, 781]. Polytropic [22]. Pontryagin [327, 366]. population [956]. porous [679, 93, 59, 52, 580, 306, 456, 1024]. porous-elastic [59]. portals [493]. posed [160]. posedness [993, 1362, 1424, 660, 577, 59, 55, 905, 1209, 1467, 994, 201, 1140, 975, 1343, 480, 534, 235, 871, 1025, 1426, 199, 1310, 778]. Position [424, 244, 247, 738, 1319]. position-dependent [244, 247, 738, 1319]. Position-momentum [424]. Positive [539, 468, 360, 191, 178, 877, 723, 951, 630, 99, 1520, 1285, 1235, 453, 550, 625, 1118, 779, 1464]. postselection [425]. Potential [153, 32, 175, 652, 756, 477, 1023, 675, 1145, 92, 752, 897, 312, 198, 651, 542, 25, 448, 1406, 246, 941, 918, 1171, 848, 693, 813, 1026, 1061, 863, 844, 1384, 796, 1488, 1249, 56, 280, 1353, 266, 573, 1392, 849, 203, 192, 803, 996, 1505, 1025, 418]. potential-Application [675]. Potentials [1218, 276, 190, 840, 810, 833, 877, 808, 970, 723, 882, 296, 472, 10, 605, 124, 786, 58, 383, 173, 274, 965, 79, 716, 935, 33, 131, 1514, 538, 464]. Potts [1337, 1279, 223]. POVMs [1223, 100]. Power [1417, 1300, 941, 19, 1347, 211]. power-law [1300, 19]. powers [663]. Prandtl [1364, 1339]. pre [507, 512]. pre-Leibniz [507]. pre-rackoid [512]. preferred [1130]. prefractal [290]. Preparation [252]. prepared [1376]. prescribed [401]. presence [1078, 470, 1480, 931, 1486, 571, 1423]. presentation [1332]. Preservation [1197]. preserving [233, 720, 1460, 795]. pressure [1402, 1438, 520, 1446, 1468, 1252, 1493, 936, 1248, 1302, 869]. pressureless [1302, 1426]. prevents [537]. Pridmore [1125]. Prime [1223, 429]. primitive [905, 61]. primordial [299]. principle [330, 297, 912, 893, 642, 1320, 545, 327, 506, 607, 285, 1514, 788]. principles [1481, 407, 649, 707, 829]. prize [904]. Probabilistic [490, 1053, 252]. probabilities [133, 547]. probability [322]. problem [1415, 836, 1011, 1030, 992, 355, 624, 752, 1124, 877, 723, 25, 448, 329, 261, 1330, 583, 387, 263, 83, 1203, 1002, 1209, 654, 1347, 302, 1141, 1138, 844, 22, 876, 453, 831, 658, 231, 1476, 12, 80, 1201, 1029, 531, 735, 869, 939, 871, 1067, 1423, 1173, 1464, 1380, 23, 278, 1181, 1027, 937, 988]. Problems [1306, 330, 304, 1481, 1023, 1271, 969, 1378, 1137, 1427, 866, 811, 85, 277, 1516, 1047, 1098, 294]. Proca [1476]. procedure [254]. process [144, 1126, 1013, 1018, 1323, 1335, 1128]. processes [946, 597, 44, 1508, 1054, 566, 1331]. processing [1005, 325]. product [766, 1122, 1182, 642, 1154, 1262, 1451, 621]. production [597, 57, 1143, 1366, 566]. products [513, 1436, 560]. profiles [1342, 158]. Progressive [1117]. Projectability [407]. projections [1106]. projective [1256]. prolate [1002, 753]. proliferation [1136]. proof [921, 587, 1453, 664]. Propagation [408, 1345, 1166, 1094, 163, 1269, 199]. propagator [477]. Proper [881, 1115]. Properties [1491, 478, 676, 350, 1474, 830, 68, 1034, 341, 498, 870, 1139, 434, 1409, 293, 1431, 1503, 1233]. property [1041, 209, 335]. proportional [687]. protocols [214, 250]. Pseudo [97, 1260, 898, 186, 323, 641, 1190, 703, 1091, 1025, 31]. pseudo-bosonic [323]. Pseudo-fermionic [97]. pseudo-forest [1091]. pseudo-fractional [1190, 703]. pseudo-generalization [186]. pseudo-hermitian [898]. pseudo-Hermiticity [1260, 641, 31]. pseudo-parabolic [1025]. pseudoclassical [135, 785]. pseudodifferential [1200]. pseudospectrum [1369]. PT [970, 1133, 1291, 786, 1076, 133, 547, 192, 803, 31]. PT-broken [547]. PT-symmetric [1133, 803, 1076, 133, 192]. PT-symmetrically [1291]. PT-symmetry [31]. Publisher [653, 448, 1293, 865, 932, 607, 269, 988]. Pullback [677, 1441, 471, 291]. pulselike [1273]. pumping [408, 74]. pure [741, 648, 1449, 748, 857]. purely [333, 1080].

q [342]. q-deformed [342]. Quadratic [695, 1369, 585, 221, 708, 333, 115, 916, 692, 726, 114, 1101, 1257, 8, 268, 5, 6, 1366]. quadratical [266]. quadratically [79, 716]. quadratures [42]. quadrupole [447, 7]. Qualitative [293, 1431, 1247]. quantitative [1375]. quantities [570, 708, 369]. Quantization [135, 763, 369, 1134, 525, 78, 851, 70, 64]. Quantized [1049, 1109, 785]. quantizer [322]. quantizer-dequantizer [322]. Quantum [1499, 1433, 467, 643, 29, 624, 364, 447, 7, 559, 1083, 1332, 326, 945, 348, 212, 327, 1107, 623, 176, 193, 185, 621, 351, 2, 138, 1181, 264, 134, 308, 1452, 251, 283, 317, 188, 423, 248, 65, 411, 850, 746, 722, 1039, 177, 101, 1161, 1110, 1222, 1000, 187, 1078, 1182, 548, 66, 136, 27, 1354, 500, 1075, 1180, 312, 494, 949, 792, 642, 626, 1312, 34, 246, 45, 255, 421, 743, 1095, 883, 425, 1225, 742, 258, 1409, 696, 1156, 83, 1203, 243, 620, 524, 1079, 509, 894, 1473, 588, 342, 442, 554, 1389, 349]. quantum [466, 1074, 426, 1158, 215, 1290, 1485, 1196, 1202, 300, 847, 1503, 1526, 194, 1449, 227, 133, 1508, 1112, 1486, 88, 666, 137, 391, 716, 1082, 1487, 739, 1098, 253, 747, 1221, 1179, 1295, 1080, 1394, 1395, 250, 32, 1146, 920, 1160, 744, 1237, 978, 89, 1184, 1454, 435, 622, 1092, 901, 1356, 247, 738, 390]. quantum-mechanical [666]. quantum-spin [45]. quantum-to-classical [1222]. quark [1372]. quark-lines [1372]. quarter [310]. quartic [477, 844, 567]. Quasi [1124, 1123, 1263, 1303, 766, 862, 298, 312, 667, 1129, 1142, 605, 82, 322, 125]. quasi-adiabatic [312]. quasi-duality [125]. quasi-Einstein [1129]. Quasi-factorization [1263]. quasi-geostropic [1142]. Quasi-homogeneous [1124]. quasi-linear [862]. quasi-locally [766]. Quasi-periodic [1123, 1303, 605, 82]. quasi-probability [322]. quasi-symmetry [298]. quasialgebras [1092]. quasigraded [968]. quasigroups [125]. quasilinear [912, 311, 1463, 723, 969, 657, 732, 1143, 817, 270, 843, 1464, 1430]. Quasinormal [874, 329]. Quasiperiodic [1406, 647]. Quasitriangular [1092, 1451]. quaternion [1230, 1407]. quaternion-valued [1230]. qubit [850, 67, 745]. qubits [1081]. qudits [429]. questions [1021]. Quiver [1294, 30]. Quivers [30].

R [508, 748, 353]. R12 [969]. R2 [538]. R3 [549, 909]. Rabinowitz [1137]. Racah [239, 1035, 32]. rackoid [512]. radial [959, 833, 501, 1, 479, 938, 1438]. Radially [161]. radiating [1489]. Radiation [678, 1260, 1315, 748, 176]. radiative [1459, 942, 460]. radius [601]. Ragnisco [1131]. Rajeev [945]. Raman [214]. Ramond [340]. Random [954, 93, 99, 956, 291, 823, 1087, 1418, 709, 1052, 1096, 800, 470, 53, 336, 959, 1182, 670, 1036, 949, 799, 682, 834, 1406, 1350, 1329, 282, 712, 310, 1409, 944, 620, 1168, 303, 374, 1495, 892, 1515, 707, 1238, 1059, 1112, 335, 1286, 1393, 815, 582, 632, 889, 1215, 901]. random-cluster [1329]. randomness [467, 177]. range [569, 1115, 954, 706, 741, 1236, 472, 688, 105, 848, 1417, 383, 907, 33]. Rank [1287, 928, 100]. Ranken [945]. rapid [635]. Rapidly [919]. rarefaction [352]. Rascle [1252, 1432]. Ratcliff [586]. Rate [737, 597, 1128, 1146]. rates [90, 1405, 1379, 1210, 1465, 89, 1489]. Rational [423, 522, 1035, 1522, 878, 157, 884]. rationality [1153]. ratios [1350]. Rayleigh [988, 937]. RBF [1242]. Reachability [1190, 703]. reacting [991, 1494]. reaction [90, 678, 1495, 873, 1331]. reaction-diffusion [90, 1495, 873]. real [330, 190, 110, 930, 826, 865, 804, 714, 1458, 1082, 1487, 705, 590]. real-matrix [590]. Realizability [635]. realization [135, 1255]. Realizations [241, 1406, 129, 1290, 228]. realm [1038]. Reciprocal [108]. Reconstruction [70, 157]. recovered [1325]. recovery [1264]. rectangles [895, 160, 831]. rectangular [970, 786, 997]. Recurrence [149, 502]. recurrences [1370]. Recurrent [358]. recursion [1333, 157]. recursions [1011]. Reduced [395, 1070, 902, 1, 479]. Reducibility [198, 258, 181]. Reducible [182]. Reduction [375, 295, 1439, 924, 324, 562]. Reductions [228, 1313]. Referee [15]. reference [1225]. reflection [642, 268, 625, 1118]. reflection-equation [268]. regardless [31]. regime [541, 773, 1071, 990, 645]. region [1052]. regions [101]. Regular [932, 888, 374, 1183, 922, 1056, 160, 1495, 1009, 1204, 689]. Regularity [837, 1385, 1207, 278, 993, 1114, 205, 1066, 52, 201, 975, 1379, 731, 1422, 1425, 779, 1240]. Regularization [280]. regularized [368, 1424]. Reiner [1057]. Reissner [691, 594]. Reissner-Nordström [594]. Reissner-Nordström-type [691]. related [154, 1106, 526, 205, 249, 560, 617, 985]. Relating [1355]. relation [502, 435]. Relations [682, 341, 1401, 9, 835, 1065, 444, 585, 215, 149, 1449, 1498]. relative [1263, 131, 704, 325, 566]. Relativistic [69, 629, 885, 405, 1121, 1326, 174, 890, 172, 734, 600, 121, 18, 288, 851, 1468, 977, 37, 209, 933, 1345, 360]. relaxation [858, 90]. relevant [1470]. Remarks [544, 301]. renewal [828]. renormalization [368, 1165, 719, 822, 634, 68, 897, 1475, 560, 209]. Renormalized [1427, 320]. Rényi [746, 1003, 215, 1005, 325, 1082, 1487]. repeaters [623]. Replica [1192, 1089, 224]. Representation [1230, 1415, 1333, 511, 966, 998, 666, 915]. representation-independent [666]. Representations [495, 126, 1151, 1398, 1106, 964, 1265, 128, 697, 79, 1081, 662]. repulsion [537, 816, 818, 732, 532]. repulsive [941, 611, 1245]. Reservoirs [681]. resistivity [1298]. resolvent [543]. Resonances [464, 774, 918]. Resonant [903]. resonator [774]. resource [177, 1219, 868]. respect [1253]. respecting [1033]. Response [652, 115, 196, 645, 1234]. rest [397]. rest-mass [397]. restricted [146, 1271, 1245, 1047, 1454]. Restriction [1084]. restrictions [766]. result [1142]. Results [294, 475, 517, 638, 1378, 610, 277, 1021, 202, 306, 1184, 566]. Resurgence [286]. retarded [1459]. Retraction [606]. reversal [566]. reverse [572]. review [1282]. revised [1448]. Revisit [206]. revisited [403, 1089, 1218, 388, 195, 236, 87, 196, 1159, 972, 363, 1397]. Revisiting [1415, 476, 847, 1005]. reweighting [1282]. Reynolds [736, 706]. Ricci [1211, 1250, 1360, 439, 1525]. Ricci-flat [439]. Riemann [269, 763, 754, 520, 355, 995, 866, 530, 1063, 1252, 1248, 189, 231, 94, 770, 39, 531, 775, 1448, 1423]. Riemann-type [531]. Riemannian [853, 1433, 1068, 1211, 70]. Riesenfeld [1291]. Riesz [1019]. Rigged [1285]. right [333, 431, 1011, 310]. right-hand [333, 431, 1011]. rigid [970, 241, 786]. rigidly [1357, 1402, 1438]. Rigorous [475, 1364, 822, 74, 1420, 1184]. Rindler [180]. ring [142]. ringdown [329]. rings [1276, 940, 601, 1138]. Rn [256, 867, 277, 689]. RNA [1387]. Robertson [272, 1129]. Robin [276, 304, 811, 47, 831, 690]. robust [1004]. robustness [1060]. Roch [39]. role [893, 219, 1272, 1281, 697]. roles [213]. root [1282, 360]. Rosen [700]. Rota [736, 1150, 999]. rotating [1048, 1357, 1402, 1438, 679, 1339, 74]. rotation [1480, 165]. rotation-two-component [165]. rotational [1067]. rotator [1288, 967]. rough [872, 274]. Rubinstein [1181]. rules [215]. Runge [1509].

Sachdev [1192, 242]. Saint [122]. same [331]. sample [889]. sandwich [395]. sandwiched [1003, 1005]. satisfy [932, 888]. satisfying [292, 880]. saturable [1176]. saturation [456, 841]. Scalar [1265, 1348, 1344, 355, 951, 150, 934, 1301, 372, 595, 1120, 102]. scale [1461, 758, 1155, 313]. scales [382, 1336]. Scaling [959, 71, 1364, 706]. Scattering [236, 691, 274, 475, 1472, 1346, 655, 611, 848, 546, 26, 33, 285, 1201, 735]. scenario [1473]. Schauder [864]. scheme [768]. Schlingemann [1221]. Schramm [263]. Schrieffer [846]. Schrödinger [808, 1482, 392, 135, 785, 1300, 276, 499, 569, 113, 1348, 311, 840, 1176, 762, 659, 497, 1145, 92, 916, 321, 810, 490, 1496, 660, 1463, 670, 1315, 208, 198, 111, 723, 1406, 741, 81, 236, 610, 995, 655, 282, 814, 501, 867, 727, 657, 535, 310, 1171, 10, 605, 124, 488, 487, 454, 693, 82, 654, 929, 428, 164, 987, 274, 721, 1384, 963, 1205, 130, 193, 776, 1304, 1249, 266, 79, 817, 935, 573, 441, 961, 181, 203, 690, 843, 192, 803, 360, 314, 996, 120, 155, 271, 1303, 1173, 1296, 1430]. Schrödinger-Poisson [840]. Schrödinger-quantized [785]. Schrödinger-type [321]. Schur [159, 125, 618, 705]. Schwartz [919, 1289]. Schwarz [1073, 340]. Schwarzschild [606, 1509, 515, 14, 874, 1119]. Schwarzschild-de [874]. Schwinger [213, 127, 393]. Scott [174]. sea [917, 1424]. sea-ice [1424]. search [507]. second [1071, 119, 228, 1441]. second-order [1441]. secretion [732]. sector [1315, 342]. sectors [423, 1062]. security [1161]. seed [445]. Segel [841]. segments [1013]. Seiberg [558]. Self [337, 1445, 1006, 1000, 489, 147, 25, 448, 148, 820, 426, 1014, 1466]. self-adjoint [147]. self-adjointness [426]. self-consistency [1000]. self-dual [1006, 489, 1466]. Self-intersection [1445]. Self-organization [337]. self-similar [1014]. Semantic [1161]. semclassical [281]. Semi [455, 1038, 722, 1065, 303, 183, 613, 815]. Semi-algebraic [455]. semi-classical [1038, 722, 1065]. semi-continuity [303, 815]. semi-infinite [183]. semi-martingales [613]. semicausal [1083]. Semiclassical [1001, 25, 448, 249, 85, 1164, 1096, 1114, 1354, 413, 733]. semiconductor [564]. Semicontinuity [1253]. semigroup [830]. semigroups [1083, 669]. semilinear [1421, 1026, 1208, 272, 976]. semimetal [1260]. semimetals [1501]. semisimple [1183]. semitoric [1479]. sensitivity [57, 532, 841]. Separability [1166, 980, 404]. separability/entanglement [980]. separable [917]. Separation [268, 646, 1149]. separations [757]. Separatrices [106]. Sequences [1475]. sequential [324]. series [1117, 317, 541, 1055, 47, 603]. sesquilinear [148]. set [137, 669]. sets [290, 326, 1007, 419, 1186, 1116, 505, 455]. settings [1264]. several [1273]. Shabat [1170]. shadow [1166]. shallow [720, 797]. shaped [492, 911].

Shapovalov [748]. Sharp [971, 90, 1236, 671, 297, 746]. shear [1099, 218]. sheaves [139]. sheets [183]. shells [727]. Sherrington [264]. shift [220]. shifts [676, 1449]. shock [1302]. shocks [1135, 1494, 60, 1248]. short [848, 649, 383, 829]. short-range [848]. short-time [649, 829]. SIC [1223, 100]. SIC-POVMs [1223]. side [177]. sides [333, 431, 1011]. Sierpinski [901]. sieved [145]. Sign [729, 118, 1023, 488, 730, 16, 314, 1173]. Sign-changing [729, 118, 1023, 488, 730, 16, 314, 1173]. signal [57, 1143, 532]. signal-dependent [1143, 532]. signature [722]. similar [1014]. Simons [996, 489, 574]. simple [1033, 289, 140, 715, 173]. simplex [584]. simplification [414, 768]. simulating [624]. sine [759, 518, 582]. sine-Gordon [518, 582]. single [740, 276, 1359, 67, 26, 1353]. single-entity [740]. single-qubit [67]. single-well [276]. singly [259]. Singular [928, 295, 457, 1022, 890, 1342, 1344, 1093, 1244, 498, 472, 124, 57, 813, 313, 1386, 658, 231, 131, 1527, 1025]. singularities [207, 645, 1386, 363, 571]. Singularity [1413, 117, 38, 752, 861, 161]. singularly [1058, 226]. SIR [908, 163, 1269]. site [1192]. Sitter [197, 251, 604, 11, 626, 18, 328, 874, 509, 186, 844, 1208, 698, 347, 982]. six [1043, 519]. size [242]. SK [1089, 957]. Skew [998, 560]. Skew-symmetric [998]. Skitovich [66]. Sklyanin [440, 1035]. Sklyanin-like [440]. slab [656]. slant [1034]. Slightly [679]. slip [614]. slow [41, 761, 219, 1046]. Smale [702, 853, 1121, 890, 451]. Small [168, 934, 1301, 1127, 499, 1231, 34, 874, 207, 1458, 1302, 1058]. small-amplitude [1231]. Small-time [168, 499]. smallest [226]. Smith [749, 1392]. Smoluchowski [161]. smooth [906, 1275, 60, 383]. Snyder [197, 11, 328]. Snyder-de [11]. Sobolev [619, 204, 762, 660, 386, 1142, 1213, 274]. soft [709]. soils [456]. solid [1389]. solids [1390]. solitary [421, 130, 271]. solitary-wave [130]. soliton [92, 995, 765, 859]. solitons [514, 1213, 553, 903, 771, 1525, 735]. Solution [1011, 808, 10, 230, 491, 862, 520, 311, 1145, 1191, 995, 812, 52, 861, 932, 888, 58, 1431, 994, 164, 1405, 907, 968, 1466, 1345, 1146, 155, 775, 536, 1489, 1381, 841, 1319]. Solutions [24, 588, 1430, 392, 91, 579, 653, 604, 1311, 719, 912, 420, 1348, 1363, 1357, 1402, 1438, 356, 166, 840, 539, 1023, 1230, 992, 1344, 489, 906, 1104, 910, 1136, 92, 584, 413, 927, 838, 1172, 692, 877, 1254, 808, 723, 726, 116, 969, 1427, 1406, 921, 386, 837, 777, 733, 328, 864, 867, 434, 842, 715, 1385, 357, 530, 451, 657, 162, 1171, 10, 117, 414, 1273, 488, 19, 487, 1125, 932, 888, 160, 20, 934, 1301, 1063, 1468, 612, 905, 1252, 275, 82, 57, 615, 1216, 1026, 1275, 1307, 1493, 37, 302, 728, 237, 60, 393, 730, 75, 1384, 656, 963, 1248, 609, 130]. solutions [732, 1488, 231, 167, 573, 416, 94, 469, 770, 1379, 1428, 294, 272, 158, 61, 16, 360, 314, 765, 578, 731, 1210, 868, 942, 1029, 1465, 538, 824, 1131, 460, 450, 353, 120, 315, 165, 483, 1204, 1024, 200, 1340, 452, 347, 418, 729, 1105, 1448, 1310, 1432, 1303, 1325, 614, 1173, 1464, 1296, 354, 23, 278, 161, 859]. solvability [667, 533, 939]. Solvable [431, 333, 284, 338, 1054, 1353, 1505]. solve [1482]. Solving [392]. Some [298, 1211, 891, 603, 301, 566, 960, 1413, 1271, 413, 465, 650, 985, 1526, 1233, 1464]. source [1344, 866, 387, 533, 868, 1462, 200]. sources [1266, 419]. Space [710, 614, 168, 1165, 41, 1106, 916, 667, 367, 397, 81, 1284, 655, 13, 54, 758, 150, 387, 1142, 801, 509, 1275, 977, 186, 240, 432, 1292, 1285, 280, 391, 337, 1167, 238, 1295, 1379, 925, 1489, 1309]. space-dependent [387]. space-periodic [1309]. Space-time [710, 614, 801, 509]. spaces [1187, 1242, 917, 220, 763, 204, 1034, 660, 1459, 575, 1358, 1517, 1084, 979, 820, 1113, 1495, 373, 1290, 274, 966, 285, 1009, 1324, 305, 689, 1240]. Spacetime [217, 359, 1509, 1166, 398, 626, 1267, 1510, 18, 1226, 1293, 934, 844, 931, 1208, 794, 272, 1119, 631, 86, 519]. spacetimes [516, 1227, 72, 514, 1129, 629, 952, 1164, 894, 1405, 748, 1464, 982]. spacing [883]. spanning [1116]. Spatial [58, 1510, 1168, 1120, 351, 2, 96]. spatially [1496, 864, 481, 934, 748, 907]. spatiotemporal [1325]. Special [989, 1528, 1006, 1199, 1340]. species [1136, 816, 533, 353]. spectra [190, 648]. Spectral [276, 411, 1474, 872, 895, 1506, 930, 346, 498, 834, 882, 1490, 1017, 1014, 44, 1256, 1501, 436, 192, 913, 1199, 836, 437, 1034, 670, 296, 1163, 714, 717, 978, 1278, 937, 988]. spectrum [175, 652, 1001, 1020, 990, 48, 1016, 635, 651, 741, 1390, 1226, 1293, 1189, 1423]. speed [621, 199]. sphere [1326, 142, 1088, 930, 261, 596, 342, 442, 796, 1179, 76]. spheres [887]. spherical [470, 1459, 727, 957, 1002, 796, 666, 857]. spherically [123, 1232, 453]. spheroidal [1002]. spheroids [753]. spiked [971]. Spin [785, 1510, 852, 943, 194, 264, 984, 946, 1162, 784, 1073, 45, 1178, 338, 511, 645, 957, 885, 348, 1250, 1077, 1389, 1074, 1007, 609, 791, 857, 347, 778, 390]. Spin-0 [1510]. spin-1 [645, 885, 1007]. spin-1/2 [645, 885]. spin-2 [1007]. spin-boson [1077]. spin-glass [264]. spindle [398]. spinor [509]. spins [1389]. spiral [675]. Split [715]. splitting [56, 366]. splittings [1451]. spontaneous [944]. SPT [819]. SqF4 [442]. square [1117, 244, 1282, 362, 360]. square-integrable [1117]. square-root [360]. squares [364, 1354, 745]. SSH [1217]. Stability [1099, 1224, 940, 459, 1246, 1349, 18, 1142, 1495, 1175, 1352, 202, 306, 564, 1046, 761, 376, 1298, 1382, 92, 751, 734, 1103, 1297, 121, 421, 1490, 104, 1213, 908, 1185, 63, 578, 1215]. Stabilization [456, 876]. stabilizer [429]. Stable [265, 419, 283, 219, 66, 408, 1013, 1366]. staggered [798]. standard [430, 463, 1498, 278, 1108]. standing [208]. Stark [1223, 918]. stars [123]. state [309, 1037, 423, 1071, 1387, 311, 1145, 1282, 496, 386, 236, 941, 856, 357, 657, 1171, 105, 1468, 591, 164, 1072, 770, 556, 505, 1302, 1434, 622, 483, 271, 1173]. states [1060, 175, 652, 1117, 244, 774, 1452, 766, 1472, 350, 569, 940, 1369, 178, 526, 548, 67, 321, 1180, 78, 667, 651, 1154, 668, 1396, 1126, 112, 1013, 663, 465, 1156, 290, 97, 894, 492, 693, 478, 1389, 323, 985, 648, 880, 664, 697, 1248, 334, 88, 671, 884, 819, 1080, 1394, 1395, 1455, 314, 564, 152, 435, 909, 738, 1319, 96]. static [1008, 1503]. Staticity [1511]. Stationary [1254, 1126, 611, 750, 1166, 991, 1357, 1402, 1438, 670, 809, 482, 501, 1385, 1274, 891, 656, 1489, 1478]. statistic [1527]. Statistical [337, 1194, 65, 1272, 1281, 828]. statistics [1371, 960, 1193, 1350, 883, 282, 319, 1017, 572, 334]. statistics-Interpolating [334]. steady [1030, 940, 1339, 1494, 261, 1305, 1138, 44, 1492, 564, 435]. Stealth [673]. steep [693, 203, 996, 418]. Steganography [250]. steps [309]. Stewartson [842]. Stieltje [344]. Stimulated [214]. Stinespring [253, 1152]. Stochastic [858, 1134, 675, 680, 535, 973, 310, 851, 324, 193, 450, 219, 273, 473, 474, 954, 1299, 751, 1147, 956, 434, 1439, 83, 1203, 981, 291, 374, 1495, 1270, 823, 1018, 1323, 1335, 1128, 1516, 307, 1098, 222, 141, 518, 750, 815, 1127, 1087, 1009, 582, 788, 632, 1046, 1324, 1445, 155, 689, 1429, 1478, 807]. Stokes [1214, 1206, 352, 237, 486, 1341, 1379, 1146, 1381, 161, 1362, 838, 1401, 575, 261, 812, 451, 303, 1044, 1274, 891, 60, 779, 1009, 939, 1204, 1024, 1340, 1426, 1310, 532, 354]. strain [779]. strained [321]. strange [1107]. stratified [279]. stretching [1057]. Strict [626, 1329, 369]. strictly [993]. string [515, 507, 555, 18]. strings [510, 256]. Strip [1255, 1298]. Striped [786, 970]. strips [1502]. Strong [992, 537, 302, 1280, 253, 1325, 23, 420, 752, 496, 451, 19, 932, 888, 615, 1493, 994, 184, 958, 942, 460, 1024, 1381, 354]. strongly [1010, 671, 590]. structural [158]. Structure [979, 1060, 1522, 763, 930, 27, 507, 706, 555, 1002]. structured [956]. Structures [361, 695, 1483, 409, 1411, 1444, 289, 804, 1177, 714, 70, 1451, 363, 886, 705]. Stuart [76]. stuck [1387]. Study [1402, 1438, 1101, 543, 1393]. Sturm [882, 1085, 77]. sub [1514]. sub-additive [1514]. subalgebra [1263]. subalgebra-relative [1263]. subalgebras [643]. Subcritical [1430, 579, 653]. subfactors [29]. sublinear [712]. subluminality [600]. submanifold [1013]. submanifolds [1084]. submultiplicative [1436]. subsolutions [1186]. subsonic [417]. subspaces [1182]. subsystems [281]. subtraction [621]. subwavelength [774]. SUC [159]. Sugawara [422]. sum [1415, 1387, 584, 189, 269]. summary [1294]. sums [409, 781]. Super [962, 644, 340, 342, 508, 650, 1200, 1500, 108, 997, 1032]. super- [508]. super-Galilean [1032]. super-Yangian [1500]. superalgebra [340, 28, 585, 1107, 1500]. superalgebras [593, 782, 997]. Supercanonical [1157]. superchannels [1083]. superconductivity [846]. superconductor [1338, 846]. superconductor-to-superinsulator [1338]. supercritical [208, 1378, 1488, 843, 120]. superfluid [377]. superhelices [832]. superinsulator [1338]. superintegrability [1456]. Superintegrable [685, 39, 432, 1292, 375]. superlinear [484]. supermechanics [1151]. superposed [1273]. Superposition [1168, 101]. superpositions [591]. superspace [365]. superstring [557]. Supersymmetric [211, 738, 134, 665, 187, 510, 343, 764, 361]. supersymmetry [1151, 494]. supported [872, 1287]. surface [763, 287, 608, 1142, 1057, 540]. surfaces [1006, 1104, 791, 1167, 1466, 457]. survey [1098]. susceptible [1325]. Susskind [668]. swarm [103]. swelling [306, 456, 876]. Swift [291, 815]. swirl [1422]. switching [184, 1280]. symbols [1372, 1388]. Symmetric [767, 1290, 1499, 190, 304, 1166, 643, 1089, 910, 446, 970, 1126, 866, 468, 123, 338, 362, 786, 1232, 77, 572, 1133, 1076, 133, 998, 453, 628, 794, 192, 1067, 161, 803]. symmetrically [1291]. Symmetries [251, 764, 721, 562, 115, 900, 1288, 720, 685, 114, 465, 972, 381, 1404, 316, 8, 1443, 5, 6, 399, 71, 449]. Symmetry [570, 13, 54, 599, 416, 593, 154, 298, 661, 1359, 513, 1169, 224, 1192, 1301, 938, 50, 433, 439, 1505, 925, 31]. symmetry-based [154]. Symplectic [281, 1410, 695, 295, 527, 159, 1177, 1256]. symplectic-Nijenhuis [1177]. symplectic/orthogonal [1256]. synchronized [1428]. Synchronous [503, 137, 920]. Synthesis [401]. synthetic [1316]. system [1482, 862, 248, 1099, 219, 172, 311, 840, 1299, 336, 1103, 1214, 1463, 692, 205, 577, 951, 1191, 503, 995, 1490, 1139, 866, 842, 1171, 59, 488, 934, 1301, 1209, 693, 57, 816, 1216, 1275, 1335, 699, 598, 1141, 157, 730, 1120, 209, 227, 133, 656, 818, 533, 122, 669, 739, 262, 80, 1295, 32, 1379, 314, 1210, 1127, 1425, 978, 1100, 353, 534, 939, 1247, 1046, 1426, 155, 729, 1310, 1441, 689, 532, 841, 1478, 161, 1240, 247, 937, 988]. Systems [879, 1187, 213, 135, 785, 308, 993, 438, 1410, 563, 619, 860, 411, 1522, 154, 946, 1479, 761, 333, 431, 1506, 549, 1497, 136, 1244, 798, 254, 367, 127, 289, 340, 421, 685, 1243, 758, 1385, 169, 1174, 207, 319, 696, 1156, 826, 865, 524, 293, 1270, 1026, 588, 17, 212, 1322, 553, 648, 1503, 432, 1292, 400, 1285, 506, 607, 1408, 776, 1051, 716, 1190, 747, 1186, 1234, 270, 703, 469, 12, 203, 1394, 1395, 1428, 399, 221, 39, 1197, 996, 1457, 435, 483, 1215, 1445, 271, 562, 923, 390]. Szego [1352]. Szego-type [1352]. Szekeres [699].

tactic [868]. Takiff [422]. tales [904]. Tangherlini [606, 14]. tau [462, 1370]. tau-functions [462]. Taub [722, 365]. technique [719]. Teleparallel [1359, 595]. temperature [420, 575, 45, 1520, 379, 957, 460, 899]. temperature-depending [420]. Temporal [19]. tension [608, 540]. tensor [971, 1039, 1106, 126, 1182, 757, 1401, 331, 241, 745, 880, 1458]. term [502, 477, 1344, 156, 109, 866, 357, 387, 202, 1140, 1283]. terms [456, 1393, 484, 200]. ternary [1251]. test [740]. Teukolsky [347]. th [1386, 1448]. their [1481, 364, 1287, 513, 574, 148, 128, 465, 1386, 1158, 595, 129, 998, 666, 1233]. theorem [1421, 772, 411, 344, 66, 482, 879, 864, 1156, 104, 1305, 896, 617, 892, 77, 1138, 656, 1508, 102, 1221, 958, 39, 1152, 1181]. theorems [709, 382, 336, 1075, 1274, 891, 38, 1336, 1375]. theoretic [623]. theoretical [281]. theories [567, 790, 510, 1219, 952, 85, 851, 38, 1416, 252, 1356]. theory [1096, 1165, 1111, 887, 1407, 1182, 1093, 295, 490, 635, 1045, 799, 626, 882, 1055, 592, 42, 691, 1243, 501, 30, 611, 767, 1450, 1390, 509, 894, 1400, 274, 50, 410, 194, 789, 998, 791, 153, 1098, 285, 1420, 1455, 1127, 1184, 636, 536, 974, 923]. There [980]. Thermal [771, 541, 313, 1051, 1220]. thermodiffusion [1103]. Thermodynamic [705, 853, 543, 858, 1013, 879, 1278, 1281]. thermodynamics [65, 1272, 1281, 289, 1328]. thermostated [259]. thermostatics [1368]. theta [1388]. thin [395, 473, 374, 1516, 978, 1478]. thin-sandwich [395]. third [706, 896]. third-order [706]. Thirring [201]. Thomas [56, 1029]. Thompson [1041]. Thorpe [1525]. three [502, 993, 557, 1227, 1136, 1154, 1266, 629, 55, 1021, 1072, 432, 1292, 1167, 765, 480, 1024, 614]. three-body [1021, 1072]. three-component [55, 765]. three-dimensional [1227, 1136, 629, 432, 1292, 1167, 480, 1024, 614]. three-term [502]. Threshold [1245, 309, 1099]. thresholds [1308]. Tight [832]. tiling [1418]. Tilings [674, 1122]. Time [1484, 477, 960, 84, 601, 136, 4, 323, 384, 1509, 382, 190, 862, 499, 168, 719, 1078, 356, 474, 762, 1114, 234, 92, 1496, 1172, 156, 1524, 949, 1251, 9, 835, 1126, 81, 40, 758, 387, 576, 207, 801, 97, 826, 865, 710, 509, 374, 1044, 977, 1245, 649, 1202, 829, 456, 732, 1054, 167, 849, 351, 1379, 1336, 2, 1283, 1029, 1465, 1024, 1445, 566, 913, 614, 354, 1240, 1309]. time-decay [1465]. time-delayed [40]. Time-dependent [477, 4, 323, 384, 92, 1172, 97, 826, 865, 1044, 1245, 1029]. time-domain [1524]. time-evolution [1078]. time-fractional [9, 835]. Time-frequency [1484]. time-inhomogeneous [566]. time-like [1509]. Time-optimal [136]. time-periodic [949, 1309]. time-space [758]. time-variable [1078]. time-varying [474, 374]. times [426]. Timoshenko [156, 1031]. Toda [1198, 1333, 749, 985, 553, 1149]. Toda- [1149]. Toeplitz [78]. Toledo [362]. tomogram [322]. tool [1093]. Topics [929]. Topological [1383, 791, 1455, 191, 971, 1037, 1111, 166, 73, 1338, 1411, 429, 1258, 1156, 171, 157, 1470, 337, 1217, 1420, 366, 1160, 1356]. Topologically [565, 228]. Topology [1316, 1267, 821, 1167]. tori [259, 182, 780, 458, 987, 102, 1492, 1234, 1457]. toric [169, 1434]. toroidal [495, 1312, 760, 1492]. torsion [629, 38, 595, 789]. tortoise [848]. torus [1097, 1049, 978, 1434]. total [1309]. Townsend [635]. toy [897]. Trace [1452, 113, 1041, 468, 1264, 102, 131, 138]. tracers [142]. traces [674]. tracial [1264]. tracks [1038]. traffic [1302, 1432]. trajectory [1253]. transfer [214]. transform [783, 1484, 296, 649]. Transformation [571, 1195, 111, 994]. transformations [570, 1410, 1291, 1320, 972, 714, 641, 266, 79, 716, 108, 889]. transforms [1459, 249, 759, 322, 829]. Transition [133, 547, 1338, 339, 524, 176]. transitions [798, 1040, 300, 926, 899]. transitivity [373]. Translating [969]. translation [763, 429, 666]. translations [780]. transmission [177, 1323]. transmutation [255]. Transonic [1494]. transport [415, 644, 688, 605, 1519, 521, 1074, 212, 1470, 74]. transportation [1202]. transpose [1317]. transposes [178]. transverse [264, 45]. trapezoids [583]. Trapped [543, 244, 990]. trapping [1488]. traps [602, 492]. Travel [426]. traveling [701, 18, 908, 911, 578, 534, 1325]. traversable [1437]. tree [300]. trees [544, 68, 1475]. Triangles [602]. triangular [863, 125]. triangulations [346]. trigonometric [469, 107]. triple [930, 1497, 963]. triples [1199, 1163, 714]. Trivial [743, 1095, 1404, 1133, 1076]. trivialization [971]. tropical [169]. Truncated [1396]. Tu [1131]. tubes [701]. Tunneling [1392]. turbulence [43, 104, 613]. turbulent [773, 275]. Twelve [904].

Twisted [736, 555, 495, 640, 48, 714, 1377, 672]. twistor [1517]. Two [1154, 727, 1228, 108, 1372, 541, 1365, 110, 676, 1052, 860, 84, 1078, 333, 431, 1299, 1288, 967, 1136, 1124, 205, 808, 408, 109, 1291, 45, 140, 1442, 429, 655, 777, 1236, 1192, 10, 124, 957, 293, 454, 816, 60, 393, 119, 170, 194, 133, 176, 533, 231, 819, 770, 1102, 869, 353, 152, 534, 315, 165, 1247, 23, 1356]. two-body [1124]. two-component [1299, 777, 152, 534, 1247]. two-cut [1052]. Two-dimensional [1228, 676, 860, 1078, 1288, 967, 1291, 429, 124, 170, 194, 819, 1102, 315, 23, 1356]. Two-electron [1154]. two-layer [1442]. two-level [408, 133, 176]. Two-parameter [727]. two-periodic [1365]. two-phase [205, 231, 770]. two-point [541, 110, 45, 1236]. two-site [1192]. two-species [1136, 816, 533, 353]. two-spin [957]. two-term [109]. Tye [1401]. type [396, 392, 1421, 382, 676, 297, 713, 643, 115, 311, 376, 1023, 1342, 1346, 446, 1244, 1411, 927, 482, 877, 111, 870, 834, 340, 296, 114, 357, 647, 795, 19, 737, 55, 372, 813, 896, 1216, 1274, 891, 1347, 17, 1257, 77, 1014, 26, 445, 748, 8, 963, 193, 1488, 122, 1049, 262, 1149, 484, 5, 6, 1336, 16, 578, 531, 1454, 483, 1239, 1375, 199, 418, 1173, 1086, 1027, 321, 691, 764, 1352]. types [128, 149].

ubiquitous [502]. Uglov [1022]. Uhlenbeck [830, 1018, 1323, 1335, 1128]. Ultra [1470]. Ultra-generalized [1470]. Ultradiscrete [400, 749]. ultrametricity [335]. ultraspherical [145]. unavoidable [111]. unbounded [369, 420, 473, 1351, 198, 1516, 453, 935, 222, 750, 460, 161]. Uncertainty [649, 829, 297, 1355, 545, 252]. underlying [1002]. unextended [113]. Unextendible [1262]. uniaxially [321]. unified [1513]. Uniform [267, 205, 1188, 1003, 747, 1361, 873, 11, 197, 663, 879, 224, 1160, 1237, 839]. uniformly [303]. unifying [1371]. Uniqueness [912, 1367, 798, 492, 1384]. unit [1326, 1399, 402, 64]. unitaries [949]. unitarity [590]. Unitary [501, 246, 767, 637, 1238, 253, 505, 1521, 529]. units [1223]. Universal [101, 68, 845, 317, 1264, 348, 1239]. Universality [847]. unknown [1019]. unlabeled [1419]. unravelings [83, 1203]. unstable [1013, 1439, 419]. unsteady [1102]. Upper [688, 815, 1071, 1236, 310, 303, 540, 170]. upper-right [310]. use [666]. using [719, 808, 1094, 296, 832, 10, 186, 342, 1190, 794, 703]. USp [345]. Uvarov [808, 10].

V [911, 1521]. V-shaped [911]. vacuum [1008, 992, 577, 629, 1207, 352, 1209, 302, 1248, 180, 453, 628, 1302, 1343, 480, 871, 23]. Vaisman [50]. valence [1389]. valence-bond-solid [1389]. Validity [1339, 383]. value [1011, 969, 85, 1141, 953]. valued [1230, 1388, 99, 907, 646]. Vandermonde [1116]. Vanishing [1276, 936, 122, 723, 1468, 935, 1423]. vapor [994]. Variable [204, 304, 1078, 296, 1109, 478, 119, 419, 975]. Variable-order [204]. variables [13, 54, 933, 1112, 268, 646, 1149, 1393]. variants [1388]. variation [1122]. Variational [143, 407, 788, 330, 1481, 912, 1320, 1079, 1138, 1514]. varieties [929]. Various [1079, 169]. varying [474, 374, 1324]. vector [396, 1509, 634, 567, 256, 861, 885, 1492]. vectors [1022, 439]. velocity [706, 1066]. Venant [122]. Verification [676]. version [1014, 1283]. vertex [495, 1130, 1153, 1069, 27, 140, 128]. vertical [905]. very [601]. VI [149]. via [134, 1242, 709, 78, 513, 995, 657, 402, 362, 521, 1317, 958]. view [885]. viewpoint [799]. Villain [790]. Violation [740]. Virasoro [1485]. virial [476]. viscoelastic [1342, 612, 1493, 1431, 200, 1031]. viscoelasticity [517]. Viscosity [1186, 1276, 420, 838, 1442, 352, 615, 933, 122, 975, 61, 942, 450, 464, 582]. Viscous [60, 600, 608, 19, 167, 1295, 460, 1309]. Vlasov [486, 1379, 773, 172, 1460, 1141, 209, 486, 1341, 1100, 353]. Vogel [715]. Voigt [237]. Volterra [1392, 1335]. volume [986, 818]. volume-filling [818]. volumes [760]. Vortex [1030, 12, 80, 315, 1276, 601, 1444, 510, 1517, 1138, 1049, 377, 538, 1457]. vortic [705]. vortices [142, 166, 1176, 805, 76]. vortices-antivortices [166]. Vorticity [1167, 601, 1049]. Vries [811]. vs [1218, 1288, 816, 1316, 847, 641].

W [361]. W-algebras [361]. Waals [339, 1507]. walk [251, 336, 892, 1237]. Walker [1129, 272]. walks [800, 624, 1409, 1196, 1286, 901]. wall [1162, 500, 312, 1486]. Wannier [385, 3, 171, 1470]. Wasserstein [892]. Wasserstein-1 [892]. water [720, 797, 994]. Wave [163, 1064, 491, 912, 1010, 356, 182, 1342, 1344, 1299, 233, 734, 206, 701, 600, 1246, 421, 947, 18, 866, 608, 414, 352, 274, 1202, 1405, 419, 227, 748, 1208, 1318, 130, 272, 141, 332, 631, 750, 824, 1269, 200, 1325, 1429]. wave-like [748]. wave-particle [947]. Wavefunctions [173, 1154, 666]. waveguide [48]. wavelet [297, 1484]. waves [1231, 780, 208, 90, 1490, 680, 1168, 908, 1480, 257, 202, 700, 616, 534, 271, 71]. way [1049]. WDVV [469]. Weak [1029, 1260, 773, 1136, 837, 777, 812, 1385, 932, 888, 20, 905, 1213, 1307, 641, 427, 61, 731, 1465, 165, 1204, 1024, 841, 278]. weakly [1020, 388, 1429]. Wei [249]. weight [568, 118, 1523, 226]. weighted [1187, 1365, 51, 762, 1150, 1333, 1213, 553, 976]. weights [814]. Weil [986]. Weil-Petersson [986]. Weizsäcker [56]. welding [263]. Well [1362, 59, 235, 244, 276, 1424, 1023, 660, 577, 55, 905, 1209, 693, 1467, 994, 1061, 1070, 153, 1353, 1140, 203, 975, 1343, 996, 480, 534, 871, 1025, 1426, 199, 418, 1310]. Well-posedness [1362, 59, 235, 660, 577, 55, 905, 1209, 1467, 994, 1140, 975, 1343, 480, 534, 871, 1025, 1426, 199, 1310]. Werner [1221]. Weyl [1260, 1153, 109, 478, 1290, 781, 125, 131]. Which [898, 1220, 932, 888]. white [1174]. Whitham [534]. whole [977]. Wick [1436]. Wick-ordered [1436]. wide [927]. Wiechert [404]. Wigner [1372, 919, 711, 245, 787]. Wilczek [504]. Willmore [974]. Wilson [1035, 95, 186, 342, 32]. Wilsonian [897]. Winding [1350]. wires [771]. wiretap [1161]. wise [984]. without [1298, 1136, 1297, 1137, 608, 905, 1492, 399, 540, 1308, 452]. Witt [1095, 743].

Witten [1111, 558]. WKB [448, 143, 25, 633]. Wong [1147, 632, 1215, 689]. Woods [830]. Work [435]. wormholes [1437]. Wronskian [79].

X [280]. X2 [317]. XX [45]. XXZ [784, 645]. XY [565, 300]. XY-interactions [300].

Yaglom [739]. Yamabe [514]. Yang [570, 1006, 1134, 1314, 1349, 1068, 715, 343, 1373, 958, 968, 915, 594, 1092]. Yangian [128, 1500, 618, 1277]. Yangians [997, 1294]. Yau [139]. Ye [1192, 242]. Yukawa [1399].

Zakai [1147, 632, 1215, 689]. Zakharov [1490, 160, 1170]. Zamolodchikov [292]. Zee [504]. Zehnder [1518]. Zeno [847]. Zero [397, 1477, 352, 437, 520, 3, 385, 286, 1226, 1293, 598, 963, 1248, 1234, 1422, 23, 1206]. zero-average [1234]. zero-gravity [1226, 1293]. zero-Hausdorff [437]. Zero-Hopf [1477, 598]. Zero-Mach [1206]. zero-pressure [520]. Zero-viscosity-capillarity [352]. zero/nonzero [963]. zeros [145, 223, 859]. zeta [863, 1132, 899]. zigzag [1471, 759]. zilch [403]. ZN [1434]. Zoll [198]. Zwanzig [807]. Zybin [581].

References

Johnson:1979:RRC

[1] Bruce R. Johnson and Joseph O. Hirschfelder. The radial reduced Coulomb Green’s function. Journal of Mathematical Physics, 20(12): 2484–2501, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

      resource/1/jmapaq/v20/i12/p2484_s1. See erratum [479].

VonZuben:2000:QTS

[2] Francis S. G. Von Zuben. Quantum time and spatial localization: an analysis of the Hegerfeldt paradox. Journal of Mathematical Physics, 41 (9):6093–6115, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [351].

DeNittis:2011:ELW

[3] G. De Nittis and M. Lein. Exponentially localized Wannier functions in periodic zero flux magnetic fields. Journal of Mathematical Physics, 52(11):112103, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

      resource/1/jmapaq/v52/i11/p112103_s1. See erratum [385].

Macedo:2012:TDC

[4] D. X. Macedo and I. Guedes. Time-dependent coupled harmonic oscillators. Journal of Mathematical Physics, 53(5):052101, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/
p052101_s1. See [384].

Tiwari:2013:CLP

[5] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0. Journal of Mathematical Physics, 54(5): 053506, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [6], comments [8, 114], and response [115].

Tiwari:2014:ECL

[6] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Erratum: “Classification of Lie point symmetries for quadratic Liénard type equation ẍ+f(x)ẋ^{2}+g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 55(5):059901, May 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5] and comment [8].

Fonseca:2015:QAM

[7] I. C. Fonseca and K. Bakke. Quantum aspects of a moving magnetic quadrupole moment interacting with an electric field. Journal of Mathematical Physics, 56(6):062107, June 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [447].

Paliathanasis:2016:CCL

[8] A. Paliathanasis and P. G. L. Leach. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)] and its erratum [j. math. phys. 55, 059901 (2014)]. Journal of Mathematical Physics, 57(2):024101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5].

Giusti:2018:DRT

[9] Andrea Giusti. Dispersion relations for the time-fractional Cattaneo– Maxwell heat equation. Journal of Mathematical Physics, 59(1):013506, January 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [835].

Karayer:2018:SSE

[10] H. Karayer, D. Demirhan, and F. Büyükkiliç. Solution of Schrödinger equation for two different potentials using extended Nikiforov–Uvarov method and polynomial solutions of biconfluent Heun equation. Journal of Mathematical Physics, 59(5):053501, May 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [808].

Falek:2019:BOU

[11] M. Falek, M. Merad, and M. Moumni. Bosonic oscillator under a uniform magnetic field with Snyder-de Sitter algebra. Journal of Mathematical Physics, 60(1):013505, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [197].

Sulaiman:2019:VPP

[12] Samira Alamin Sulaiman. Vortex patch problem for Euler–Boussinesq systems. Journal of Mathematical Physics, 60(1):013506, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [80].

Helland:2019:SSC

[13] Inge S. Helland. Symmetry in a space of conceptual variables. Journal of Mathematical Physics, 60(5):052101, May 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [54].

Fox:2019:MHS

[14] Matthew S. Fox. Multipole hair of Schwarzschild–Tangherlini black holes. Journal of Mathematical Physics, 60(10):102502, October 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See retraction [606].

Solovej:2020:RA

[15] Jan Philip Solovej. Referee acknowledgment for 2019. Journal of Mathematical Physics, 61(1):010201, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:LES

[16] Da-Bin Wang. Least energy sign-changing solutions of Kirchhoff-type equation with critical growth. Journal of Mathematical Physics, 61(1): 011501, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maia:2020:CHC

[17] B. B. V. Maia and O. H. Miyagaki. On a class of Hamiltonian Choquardtype elliptic systems. Journal of Mathematical Physics, 61(1):011502, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2020:STW

[18] Chun-Lei He, Shou-Jun Huang, and Changhua Wei. Stability of traveling wave for the relativistic string equation in de Sitter spacetime. Journal of Mathematical Physics, 61(1):011503, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2020:TDS

[19] Jae-Myoung Kim. Temporal decay of strong solutions to the magnetohydrodynamics with power-law type nonlinear viscous fluid. Journal of Mathematical Physics, 61(1):011504, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2020:MWS

[20] Jun Ik Lee, Jae-Myoung Kim, Yun-Ho Kim, and Jongrak Lee. Multiplicity of weak solutions to non-local elliptic equations involving the fractional p(x)-Laplacian. Journal of Mathematical Physics, 61(1):011505, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2020:GEA

[21] Qiuju Xu, Zhong Tan, and Huaqiao Wang. Global existence and asymptotic behavior for the 3D compressible magneto-micropolar fluids in a bounded domain. Journal of Mathematical Physics, 61(1):011506, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qu:2020:HMN

[22] Aifang Qu, Hairong Yuan, and Qin Zhao. High Mach number limit of one-dimensional piston problem for non-isentropic compressible Euler equations: Polytropic gas. Journal of Mathematical Physics, 61(1): 011507, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhong:2020:SSC

[23] Xin Zhong. Strong solutions to the Cauchy problem of two-dimensional compressible non-isothermal nematic liquid crystal flows with vacuum and zero heat conduction. Journal of Mathematical Physics, 61(1): 011508, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bao:2020:SLL

[24] Sijia Bao, Denis Constales, Hendrik De Bie, and Teppo Mertens. Solutions for the Lévy–Leblond or parabolic Dirac equation and its generalizations. Journal of Mathematical Physics, 61(1):011509, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujiie:2020:SWP

[25] Setsuro Fujiié and Spyridon Kamvissis. Semiclassical WKB problem for the non-self-adjoint Dirac operator with analytic potential. Journal of Mathematical Physics, 61(1):011510, January 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [448].

Morioka:2020:NSE

[26] Hisashi Morioka and Naotaka Shoji. Non-scattering energies for acoustictype equations on manifolds with a single flat end. Journal of Mathematical Physics, 61(1):011511, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeSole:2020:SQV

[27] Alberto De Sole, Matteo Gardini, and Victor G. Kac. On the structure of quantum vertex algebras. Journal of Mathematical Physics, 61(1): 011701, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isaac:2020:GGL

[28] Phillip S. Isaac, N. I. Stoilova, and Joris Van der Jeugt. The Z_{2} × Z_{2}- graded general linear Lie superalgebra. Journal of Mathematical Physics, 61(1):011702, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

De:2020:QDI

[29] Sandipan De. Quantum double inclusions associated to a family of Kac algebra subfactors. Journal of Mathematical Physics, 61(1):011703, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2020:QHB

[30] Zhi Hu and Pengfei Huang. Quivers from Higgs bundles over P^{1} and quiver gauge theory. Journal of Mathematical Physics, 61(1):011704, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:PSE

[31] Ruili Zhang, Hong Qin, and Jianyuan Xiao. PT-symmetry entails pseudo-Hermiticity regardless of diagonalizability. Journal of Mathematical Physics, 61(1):012101, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taiwo:2020:PFW

[32] T. J. Taiwo. Potential function of the Wilson–Racah quantum system. Journal of Mathematical Physics, 61(1):012102, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szmytkowski:2020:ACS

[33] Radosław Szmytkowski. Analytical calculations of scattering lengths for a class of long-range potentials of interest for atomic physics. Journal of Mathematical Physics, 61(1):012103, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ge:2020:HAF

[34] Yimin Ge and Vedran Dunjko. A hybrid algorithm framework for small quantum computers with application to finding Hamiltonian cycles. Journal of Mathematical Physics, 61(1):012201, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penna:2020:OM

[35] Robert Penna. SDiff(S^{2}) and the orbit method. Journal of Mathematical Physics, 61(1):012301, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigin:2020:X

[36] Boris Feigin and Sergei Gukov. VOA[M_{4}]. Journal of Mathematical Physics, 61(1):012302, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2020:PSR

[37] Jaume Llibre and Amar Makhlouf. On the periodic solutions of the relativistic driven harmonic oscillator. Journal of Mathematical Physics, 61(1):012501, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luz:2020:STI

[38] Paulo Luz and Filipe C. Mena. Singularity theorems and the inclusion of torsion in affine theories of gravity. Journal of Mathematical Physics, 61 (1):012502, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:2020:SSR

[39] A. V. Tsiganov. Superintegrable systems and Riemann–Roch theorem. Journal of Mathematical Physics, 61(1):012701, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2020:EDL

[40] Seung-Yeal Ha, Doheon Kim, Dohyun Kim, Hansol Park, and Woojoo Shim. Emergent dynamics of the Lohe matrix ensemble on a network under time-delayed interactions. Journal of Mathematical Physics, 61 (1):012702, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burby:2020:GCD

[41] J. W. Burby. Guiding center dynamics as motion on a formal slow manifold in loop space. Journal of Mathematical Physics, 61(1):012703, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grillo:2020:EHJ

[42] Sergio Grillo and Edith Padrón. Extended Hamilton–Jacobi theory, contact manifolds, and integrability by quadratures. Journal of Mathematical Physics, 61(1):012901, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flandoli:2020:ECM

[43] Franco Flandoli and Dejun Luo. Energy conditional measures and 2D turbulence. Journal of Mathematical Physics, 61(1):013101, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Murphy:2020:SAC

[44] N. B. Murphy, E. Cherkaev, J. Zhu, J. Xin, and K. M. Golden. Spectral analysis and computation for homogenization of advection diffusion processes in steady flows. Journal of Mathematical Physics, 61(1):013102, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gohmann:2020:HTA

[45] Frank Göhmann, Karol K. Kozlowski, and Junji Suzuki. Hightemperature analysis of the transverse dynamical two-point correlation function of the XX quantum-spin chain. Journal of Mathematical Physics, 61(1):013301, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cutimanco:2020:EEH

[46] Miguel Cutimanco and Vasilisa Shramchenko. Explicit examples of Hurwitz Frobenius manifolds in genus one. Journal of Mathematical Physics, 61(1):013501, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Machida:2020:BSP

[47] Manabu Machida and Gen Nakamura. Born series for the photon diffusion equation perturbing the Robin boundary condition. Journal of Mathematical Physics, 61(1):013502, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Briet:2020:ACS

[48] Ph. Briet, J. Dittrich, and D. Krejcirík. Absolute continuity of the spectrum in a twisted Dirichlet–Neumann waveguide. Journal of Mathematical Physics, 61(1):013503, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zuniga-Galindo:2020:NAC

[49] W. A. Zúñiga-Galindo and Sergii M. Torba. Non-Archimedean Coulomb gases. Journal of Mathematical Physics, 61(1):013504, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mori:2020:DAV

[50] Haruka Mori, Shin Sasaki, and Kenta Shiozawa. Doubled aspects of Vaisman algebroid and gauge symmetry in double field theory. Journal of Mathematical Physics, 61(1):013505, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertola:2020:GWH

[51] M. Bertola, J. Harnad, and B. Runov. Generating weighted Hurwitz numbers. Journal of Mathematical Physics, 61(1):013506, January 2020.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kieu:2020:ESG

[52] Thinh Kieu. Existence of a solution for generalized Forchheimer flow in porous media with minimal regularity conditions. Journal of Mathematical Physics, 61(1):013507, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2020:ELC

[53] Shuhua Chang, Tiefeng Jiang, and Yongcheng Qi. Eigenvalues of large chiral non-Hermitian random matrices. Journal of Mathematical Physics, 61(1):013508, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helland:2020:ESS

[54] Inge S. Helland. Erratum: “Symmetry in a space of conceptual variables” [J. Math. Phys. 60(5), 052101 (2019)]. Journal of Mathematical Physics, 61(1):019901, January 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13].

Li:2020:LWP

[55] Zhigang Li and Yuxi Hu. Local well-posedness and blow-up criteria for a three-component Camassa–Holm type equation. Journal of Mathematical Physics, 61(2):021501, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salazar:2020:MST

[56] Lorena Aguirre Salazar, Stan Alama, and Lia Bronsard. Mass splitting in the Thomas–Fermi–Dirac–von Weizsäcker model with background potential. Journal of Mathematical Physics, 61(2):021502, February 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2020:GSF

[57] Dongmei Liu. Global solutions in a fully parabolic chemotaxis system with singular sensitivity and nonlinear signal production. Journal of Mathematical Physics, 61(2):021503, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2020:SBS

[58] Yu-Chu Lin, Haitao Wang, and Kung-Chien Wu. Spatial behavior of the solution to the linearized Boltzmann equation with hard potentials.

Journal of Mathematical Physics, 61(2):021504, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khochemane:2020:WPG

[59] Houssem Eddine Khochemane, Abdelhak Djebabla, Salah Zitouni, and Lamine Bouzettouta. Well-posedness and general decay of a nonlinear damping porous-elastic system with infinite memory. Journal of Mathematical Physics, 61(2):021505, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2020:VLC

[60] Shixiang Ma. Viscous limits of the compressible Navier–Stokes equations to piecewise smooth solutions with two interacting out-going shocks. Journal of Mathematical Physics, 61(2):021506, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:GEW

[61] Fengchao Wang, Changsheng Dou, and Quansen Jiu. Global existence of weak solutions to 3D compressible primitive equations with degenerate viscosity. Journal of Mathematical Physics, 61(2):021507, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaygorodov:2020:DFA

[62] Ivan Kaygorodov and Yury Volkov. Degenerations of Filippov algebras. Journal of Mathematical Physics, 61(2):021701, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nam:2020:ISA

[63] Phan Thành Nam and Nicolas Rougerie. Improved stability for 2D attractive Bose gases. Journal of Mathematical Physics, 61(2):021901, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delOlmo:2020:CIQ

[64] M. A. del Olmo and J. P. Gazeau. Covariant integral quantization of the unit disk. Journal of Mathematical Physics, 61(2):022101, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bebiano:2020:TNH

[65] N. Bebiano, J. da Providência, and J. P. da Providência. Toward non-Hermitian quantum statistical thermodynamics. Journal of Mathematical Physics, 61(2):022102, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cuesta:2020:SQD

[66] Javier Cuesta. A stable quantum Darmois–Skitovich theorem. Journal of Mathematical Physics, 61(2):022201, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dahlberg:2020:CSQ

[67] Axel Dahlberg, Jonas Helsen, and Stephanie Wehner. Counting singlequbit Clifford equivalent graph states is #P-complete. Journal of Mathematical Physics, 61(2):022202, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clavier:2020:LRU

[68] Pierre Clavier, Li Guo, Sylvie Paycha, and Bin Zhang. Locality and renormalization: Universal properties and integrals on trees. Journal of Mathematical Physics, 61(2):022301, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dossa:2020:RDP

[69] Finagnon A. Dossa and Gabriel Y. H. Avossevou. Relativistic dynamics for a particle carrying a non-Abelian charge in a non-Abelian background electromagnetic field. Journal of Mathematical Physics, 61(2):022302, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:2020:RQR

[70] Shahn Majid. Reconstruction and quantization of Riemannian structures. Journal of Mathematical Physics, 61(2):022501, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:SCS

[71] P.-M. Zhang, M. Cariglia, M. Elbistan, and P. A. Horvathy. Scaling and conformal symmetries for plane gravitational waves. Journal of Mathematical Physics, 61(2):022502, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chernov:2020:KHC

[72] V. Chernov, G. Martin, and I. Petkova. Khovanov homology and causality in spacetimes. Journal of Mathematical Physics, 61(2):022503, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeNittis:2020:EEM

[73] Giuseppe De Nittis and Max Lein. Equivalence of electric, magnetic, and electromagnetic Chern numbers for topological photonic crystals. Journal of Mathematical Physics, 61(2):022901, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pachev:2020:RBH

[74] B. Pachev, J. P. Whitehead, G. Fantuzzi, and I. Grooms. Rigorous bounds on the heat transport of rotating convection with Ekman pumping. Journal of Mathematical Physics, 61(2):023101, February 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Michalak:2020:ESF

[75] M. Michalak and B. K. Shivamoggi. Exact solutions for the fluid impulse for incompressible and compressible flows. Journal of Mathematical Physics, 61(2):023102, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoon:2020:SVH

[76] Jongbin Yoon, Habin Yim, and Sun-Chul Kim. Stuart vortices on a hyperbolic sphere. Journal of Mathematical Physics, 61(2):023103, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masjed-Jamei:2020:ISO

[77] Mohammad Masjed-Jamei, Zahra Moalemi, and Nasser Saad. Incomplete symmetric orthogonal polynomials of finite type generated by a generalized Sturm–Liouville theorem. Journal of Mathematical Physics, 61 (2):023501, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durdevich:2020:CSM

[78] Micho Durdevich and Stephen Bruce Sontz. Coherent states for the Manin plane via Toeplitz quantization. Journal of Mathematical Physics, 61(2):023502, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2020:HOD

[79] Axel Schulze-Halberg. Higher-order Darboux transformations and Wronskian representations for Schrödinger equations with quadratically energy-dependent potentials. Journal of Mathematical Physics, 61(2): 023503, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sulaiman:2020:EVP

[80] Samira Alamin Sulaiman. Erratum: “Vortex patch problem for Euler– Boussinesq system” [J. Math. Phys. 60, 013506 (2019)]. Journal of Mathematical Physics, 61(2):029901, February 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [12].

Gorka:2020:ABT

[81] Przemysław Górka, Humberto Prado, and Daniel J. Pons. The asymptotic behavior of the time fractional Schrödinger equation on Hilbert space. Journal of Mathematical Physics, 61(3):031501, March 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2020:EAP

[82] Shujuan Liu. The existence of almost-periodic solutions for 1-dimensional nonlinear Schrödinger equation with quasi-periodic forcing. Journal of Mathematical Physics, 61(3):031502, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keys:2020:PSM

[83] Dustin Keys and Jan Wehr. Poisson stochastic master equation unravelings and the measurement problem: a quantum stochastic calculus perspective. Journal of Mathematical Physics, 61(3):032101, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1203].

Bruschi:2020:TET

[84] David Edward Bruschi. Time evolution of two harmonic oscillators with cross-Kerr interactions. Journal of Mathematical Physics, 61(3):032102, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juarez-Aubry:2020:STI

[85] Benito A. Juárez-Aubry, Tonatiuh Miramontes, and Daniel Sudarsky. Semiclassical theories as initial value problems. Journal of Mathematical Physics, 61(3):032301, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xuan:2020:PDF

[86] Pham Truong Xuan. Peeling of Dirac field on Kerr spacetime. Journal of Mathematical Physics, 61(3):032501, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ismail:2020:AIM

[87] Mourad E. H. Ismail and Nasser Saad. The asymptotic iteration method revisited. Journal of Mathematical Physics, 61(3):033501, March 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [195] and response [196].

Rougerie:2020:HQH

[88] Nicolas Rougerie and Jakob Yngvason. Holomorphic quantum Hall states in higher Landau levels. Journal of Mathematical Physics, 61(4):041101, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xi:2020:DRC

[89] Xiaoyu Xi, Xueke Pu, and Boling Guo. Decay rates of the compressible Hall-magnetohydrodynamic model for quantum plasmas. Journal of Mathematical Physics, 61(4):041501, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hadadifard:2020:SRR

[90] Fazel Hadadifard and Atanas G. Stefanov. Sharp relaxation rates for plane waves of general reaction-diffusion models. Journal of Mathematical Physics, 61(4):041502, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bian:2020:GSM

[91] Xingzhi Bian and Liping Luan. Global solutions to a model with Dirichlet boundary conditions for interface motion by interface diffusion. Journal of Mathematical Physics, 61(4):041503, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deng:2020:ASM

[92] Qingquan Deng and Xiaohua Yao. Asymptotic stability of multi-soliton solutions for nonlinear Schrödinger equations with time-dependent potential. Journal of Mathematical Physics, 61(4):041504, April 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeGregorio:2020:RFC

[93] Alessandro De Gregorio and Enzo Orsingher. Random flights connecting porous medium and Euler–Poisson–Darboux equations. Journal of Mathematical Physics, 61(4):041505, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2020:GLB

[94] Yating Song and Lihui Guo. General limiting behavior of Riemann solutions to the non-isentropic Euler equations for modified Chaplygin gas. Journal of Mathematical Physics, 61(4):041506, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frappat:2020:DPO

[95] Luc Frappat, Julien Gaboriaud, Eric Ragoucy, and Luc Vinet. The dual pair (U_{q}(∫⊓(∞, ∞)), ≀_{⨿^{∞/∈}}(∈\)), q-oscillators, and Askey–Wilson algebras. Journal of Mathematical Physics, 61(4):041701, April 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delasCuevas:2020:MSO

[96] Gemma de las Cuevas and Tim Netzer. Mixed states in one spatial dimension: Decompositions and correspondence with nonnegative matrices. Journal of Mathematical Physics, 61(4):041901, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koussa:2020:PFC

[97] W. Koussa, M. Attia, and M. Maamache. Pseudo-fermionic coherent states with time-dependent metric. Journal of Mathematical Physics, 61(4):042101, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chitambar:2020:EMB

[98] Eric Chitambar, Julio I. de Vicente, Mark W. Girard, and Gilad Gour. Entanglement manipulation beyond local operations and classical communication. Journal of Mathematical Physics, 61(4):042201, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heinosaari:2020:RPO

[99] Teiko Heinosaari, Maria Anastasia Jivulescu, and Ion Nechita. Random positive operator valued measures. Journal of Mathematical Physics, 61 (4):042202, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pandey:2020:EBR

[100] Satish K. Pandey, Vern I. Paulsen, Jitendra Prakash, and Mizanur Rahaman. Entanglement breaking rank and the existence of SIC POVMs. Journal of Mathematical Physics, 61(4):042203, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2020:USC

[101] Holger Boche, Gisbert Janßen, and Sajad Saeedinaeeni. Universal superposition codes: Capacity regions of compound quantum broadcast channel with confidential messages. Journal of Mathematical Physics, 61(4):042204, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ponge:2020:CTT

[102] Raphaël Ponge. Connes’s trace theorem for curved noncommutative tori: Application to scalar curvature. Journal of Mathematical Physics, 61(4): 042301, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2020:EBF

[103] Seung-Yeal Ha, Seungsu Hwang, Dohyun Kim, Sun-Chul Kim, and Chanho Min. Emergent behaviors of a first-order particle swarm model on the hyperboloid. Journal of Mathematical Physics, 61(4):042701, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kondo:2020:ETG

[104] Shintaro Kondo and Ryusuke Numata. Existence theorem and global asymptotical stability for low-dimensional dynamical models of plasma turbulence. Journal of Mathematical Physics, 61(4):042702, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerimov:2020:GSC

[105] Azer Kerimov. Ground state criteria in one-dimensional antiferromagnetic Ising model with long range interaction. Journal of Mathematical Physics, 61(4):043301, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez:2020:SHJ

[106] Gabriel Álvarez, Luis Martínez Alonso, Elena Medina, and Juan Luis Vázquez. Separatrices in the Hamilton–Jacobi formalism of inflaton models. Journal of Mathematical Physics, 61(4):043501, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zabrodin:2020:KHT

[107] A. Zabrodin. KP hierarchy and trigonometric Calogero–Moser hierarchy. Journal of Mathematical Physics, 61(4):043502, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tian:2020:TSC

[108] Kai Tian, Q. P. Liu, and Wen Jun Yue. Two super Camassa–Holm equations: Reciprocal transformations and applications. Journal of Mathematical Physics, 61(4):043503, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frank:2020:ETT

[109] Rupert L. Frank and Simon Larson. On the error in the two-term Weyl formula for the Dirichlet Laplacian. Journal of Mathematical Physics, 61(4):043504, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baker:2020:NRE

[110] C. Baker and B. Mityagin. Non-real eigenvalues of the harmonic oscillator perturbed by an odd, two-point interaction. Journal of Mathematical Physics, 61(4):043505, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fialkovsky:2020:MTS

[111] Ignat Fialkovsky and Maria Perel. Mode transformation for a Schrödinger type equation: Avoided and unavoidable level crossings. Journal of Mathematical Physics, 61(4):043506, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gontier:2020:ESO

[112] David Gontier. Edge states in ordinary differential equations for dislocations. Journal of Mathematical Physics, 61(4):043507, April 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Campoamor-Stursberg:2020:TFC

[113] R. Campoamor-Stursberg. Trace formulas for the Casimir operators of the unextended Schrödinger algebra S(N). Journal of Mathematical Physics, 61(4):043508, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iacono:2020:CCL

[114] Roberto Iacono. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 61(4):044101, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5, 115].

Chandrasekar:2020:RCC

[115] V. K. Chandrasekar, A. K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Response to “Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]” [j. math. phys. 61, 044101 (2020)]. Journal of Mathematical Physics, 61(4):044102, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5].

Gao:2020:LNS

[116] Fengshuang Gao and Yuxia Guo. Localized nodal solutions for p-Laplacian equations with critical exponents. Journal of Mathematical Physics, 61(5):051501, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kato:2020:SSL

[117] Keiichi Kato, Masaki Kawamoto, and Koichiro Nanbu. Singularity for solutions of linearized KdV equations. Journal of Mathematical Physics, 61(5):051502, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonaldo:2020:CEE

[118] L. M. M. Bonaldo, E. J. Hurtado, and O. H. Miyagaki. A class of elliptic equations involving nonlocal integrodifferential operators with sign-changing weight functions. Journal of Mathematical Physics, 61(5): 051503, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mansour:2020:SOD

[119] Toufik Mansour and Matthias Schork. On the second order differential equation involving two ordinary and one para-Grassmann variable. Journal of Mathematical Physics, 61(5):051504, May 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2020:NSF

[120] Tao Yang. Normalized solutions for the fractional Schrödinger equation with a focusing nonlocal L^{2}-critical or L^{2}-supercritical perturbation. Journal of Mathematical Physics, 61(5):051505, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia-Perciante:2020:ISR

[121] A. L. García-Perciante and O. Reula. On the illposedness and stability of the relativistic heat equation. Journal of Mathematical Physics, 61(5): 051506, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahoo:2020:VVL

[122] Manas R. Sahoo and Abhrojyoti Sen. Vanishing viscosity limit for a system of balance laws with general type initial data arising from 1D Saint–Venant model. Journal of Mathematical Physics, 61(5):051507, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jang:2020:LAB

[123] Juhi Jang and Tetu Makino. Linearized analysis of barotropic perturbations around spherically symmetric gaseous stars governed by the Euler– Poisson equations. Journal of Mathematical Physics, 61(5):051508, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karuhanga:2020:NET

[124] Martin Karuhanga and Eugene Shargorodsky. On negative eigenvalues of two-dimensional Schrödinger operators with singular potentials. Journal of Mathematical Physics, 61(5):051509, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2020:SWQ

[125] Guodong Shi and Shuanhong Wang. Schur–Weyl quasi-duality and (co)triangular Hopf quasigroups. Journal of Mathematical Physics, 61 (5):051701, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2020:RLB

[126] Liang Chang. Representations of the loop braid groups from braided tensor categories. Journal of Mathematical Physics, 61(5):051702, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foissy:2020:MDS

[127] Loïc Foissy. Multigraded Dyson–Schwinger systems. Journal of Mathematical Physics, 61(5):051703, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:2020:YDC

[128] Naihuan Jing, Fan Yang, and Ming Liu. Yangian doubles of classical types and their vertex representations. Journal of Mathematical Physics, 61(5):051704, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meljanac:2020:GHA

[129] Stjepan Meljanac, Tea Martinić-Bila’c, and Sasa Kresić-Jurić. Generalized Heisenberg algebra applied to realizations of the orthogonal, Lorentz, and Poincaré algebras and their dual extensions. Journal of Mathematical Physics, 61(5):051705, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rego-Monteiro:2020:GNS

[130] M. A. Rego-Monteiro. Generalized nonlinear Schrödinger equation: Conservation of energy and solitary-wave solutions. Journal of Mathematical Physics, 61(5):052101, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ullmann:2020:GWA

[131] J. Ullmann. Generalization of Weyl’s asymptotic formula for the relative trace of singular potentials. Journal of Mathematical Physics, 61(5): 052102, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morales-Ruiz:2020:DGA

[132] Juan J. Morales-Ruiz. A differential Galois approach to path integrals. Journal of Mathematical Physics, 61(5):052103, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohlsson:2020:TPT

[133] Tommy Ohlsson and Shun Zhou. Transition probabilities in the two-level quantum system with PT-symmetric non-Hermitian Hamiltonians. Journal of Mathematical Physics, 61(5):052104, May 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2020:GES

[134] N. Aizawa, K. Amakawa, and S. Doi. Z_{2}^{n}-graded extensions of supersymmetric quantum mechanics via Clifford algebras. Journal of Mathematical Physics, 61(5):052105, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allen:2020:QPS

[135] Theodore J. Allen, Donald Spector, and Christopher Wilson. Quantization of pseudoclassical systems in the Schrödinger realization. Journal of Mathematical Physics, 61(5):052106, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DAlessandro:2020:TOC

[136] Domenico D’Alessandro, Benjamin A. Sheller, and Zhifei Zhu. Timeoptimal control of quantum lambda systems in the KP configuration. Journal of Mathematical Physics, 61(5):052107, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Russell:2020:GSS

[137] Travis B. Russell. Geometry of the set of synchronous quantum correlations. Journal of Mathematical Physics, 61(5):052201, May 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamagata:2020:QMM

[138] K. Yamagata. Quantum monotone metrics induced from trace nonincreasing maps and additive noise. Journal of Mathematical Physics, 61(5):052202, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2020:CYG

[139] Qiuye Jia and Hai Lin. Calabi–Yau generalized complete intersections and aspects of cohomology of sheaves. Journal of Mathematical Physics, 61(5):052301, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grady:2020:CEV

[140] J. Connor Grady, Ching Hung Lam, James E. Tener, and Hiroshi Yamauchi. Classification of extremal vertex operator algebras with two simple modules. Journal of Mathematical Physics, 61(5):052302, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:ABS

[141] Renhai Wang and Yangrong Li. Asymptotic behavior of stochastic discrete wave equations with nonlinear noise and damping. Journal of Mathematical Physics, 61(5):052701, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andrade:2020:DBP

[142] Jaime Andrade, Stefanella Boatto, and Claudio Vidal. Dynamics and bifurcation of passive tracers advected by a ring of point vortices on a sphere. Journal of Mathematical Physics, 61(5):052702, May 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burby:2020:VNW

[143] J. W. Burby and D. E. Ruiz. Variational nonlinear WKB in the Eulerian frame. Journal of Mathematical Physics, 61(5):053101, May 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Costa:2020:DAF

[144] L. T. Costa and A. D. Ramos. Dynamic aspects of the flip-annihilation process. Journal of Mathematical Physics, 61(5):053301, May 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castillo:2020:EIZ

[145] K. Castillo, M. N. de Jesus, and J. Petronilho. An electrostatic interpretation of the zeros of sieved ultraspherical polynomials. Journal of Mathematical Physics, 61(5):053501, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnal:2020:ARF

[146] Didier Arnal, Mouna Chaabouni, and Mabrouka Hfaiedh. Analyticity of a restricted formality. Journal of Mathematical Physics, 61(5):053502, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chittaro:2020:CEI

[147] Francesca Carlotta Chittaro and Paolo Mason. On the conicity of eigenvalues intersections for parameter-dependent self-adjoint operators. Journal of Mathematical Physics, 61(5):053503, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2020:NSA

[148] Hiroshi Inoue. Non-self-adjoint Hamiltonians defined by sesquilinear forms and their physical applications. Journal of Mathematical Physics, 61(5):053504, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2020:RRM

[149] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials. VI. Meixner–Pollaczek and continuous Hahn types. Journal of Mathematical Physics, 61(5):053505, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huguet:2020:MSF

[150] E. Huguet, J. Queva, and J. Renaud. Massive scalar field on (A)dS space from a massless conformal field in R^{6}. Journal of Mathematical Physics, 61(5):053506, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motegi:2020:CPF

[151] Kohei Motegi. A class of partition functions associated with E_{τ,η}gl_{3} by Izergin–Korepin analysis. Journal of Mathematical Physics, 61(5): 053507, May 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2020:GST

[152] Liangshun Xu. Ground states of two-component Bose–Einstein condensates passing an obstacle. Journal of Mathematical Physics, 61(6): 061501, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2020:PWT

[153] Tarek Saanouni. Potential well theory for the focusing fractional Choquard equation. Journal of Mathematical Physics, 61(6):061502, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bluman:2020:NSB

[154] George W. Bluman, Rafael de la Rosa, María Santos Bruzón, and María Luz Gandarias. A new symmetry-based method for constructing nonlocally related PDE systems from admitted multi-parameter groups. Journal of Mathematical Physics, 61(6):061503, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:GSB

[155] Qi Zhang, Jinqiao Duan, and Yong Chen. Global solution and blow-up of the stochastic nonlinear Schrödinger system. Journal of Mathematical Physics, 61(6):061504, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DosSantos:2020:LTD

[156] M. J. Dos Santos, M. M. Freitas, A. J. A. Ramos, D. S. Almeida Júnior, and L. R. S. Rodrigues. Long-time dynamics of a nonlinear Timoshenko beam with discrete delay term and nonlinear damping. Journal of Mathematical Physics, 61(6):061505, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marchal:2020:IDR

[157] Olivier Marchal and Nicolas Orantin. Isomonodromic deformations of a rational differential system and reconstruction with the topological recursion: The ∫↕_{∈} case. Journal of Mathematical Physics, 61(6):061506, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:APS

[158] Yu-Zhu Wang and Yanshuo Li. Asymptotic profiles of solutions to the generalized double dispersion equation with structural damping. Journal of Mathematical Physics, 61(6):061507, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2020:GSS

[159] Fang Huang and Na Wang. Generalized symplectic Schur functions and SUC hierarchy. Journal of Mathematical Physics, 61(6):061508, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Larkin:2020:DRS

[160] N. A. Larkin. Decay of regular solutions for the critical 2D Zakharov– Kuznetsov equation posed on rectangles. Journal of Mathematical Physics, 61(6):061509, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2020:RSS

[161] Limei Zhu, Bingyuan Huang, and Jinrui Huang. Radially symmetric solutions for Navier–Stokes–Smoluchowski system: Global existence in unbounded annular domain and center singularity. Journal of Mathematical Physics, 61(6):061510, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalantarova:2020:GBS

[162] H. Kalantarova, V. Kalantarov, and O. Vantzos. Global behavior of solutions to chevron pattern equations. Journal of Mathematical Physics, 61(6):061511, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2020:WPN

[163] Weixin Wu, Long Zhang, and Zhidong Teng. Wave propagation in a nonlocal dispersal SIR epidemic model with nonlinear incidence and nonlocal distributed delays. Journal of Mathematical Physics, 61(6):061512, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mao:2020:ECG

[164] Anmin Mao and Qian Zhang. Existence and concentration of ground state solution to a nonlocal Schrödinger equation. Journal of Mathematical Physics, 61(6):061513, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2020:WSR

[165] Li Yang, Chunlai Mu, Shouming Zhou, and Xinyu Tu. On the weak solutions for the rotation-two-component Camassa–Holm equation. Journal of Mathematical Physics, 61(6):061514, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:ETS

[166] Shouxin Chen and Manman Yang. Existence of the topological solutions arising in vortices-antivortices equation. Journal of Mathematical Physics, 61(6):061515, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2020:LTB

[167] Weixuan Shi and Jiang Xu. The large-time behavior of solutions in the critical L^{p} framework for compressible viscous and heat-conductive gas flows. Journal of Mathematical Physics, 61(6):061516, June 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bilal:2020:STE

[168] Adel Bilal. Small-time expansion of the Fokker–Planck kernel for space and time dependent diffusion and drift coefficients. Journal of Mathematical Physics, 61(6):061517, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jerby:2020:LGS

[169] Yochay Jerby. On Landau–Ginzburg systems, co-tropical geometry, and D^{⌊}(X) of various toric Fano manifolds. Journal of Mathematical Physics, 61(6):061701, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mayer:2020:FET

[170] Simon Mayer and Robert Seiringer. The free energy of the twodimensional dilute Bose gas. II. Upper bound. Journal of Mathematical Physics, 61(6):061901, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ludewig:2020:GWB

[171] Matthias Ludewig and Guo Chuan Thiang. Good Wannier bases in Hilbert modules associated to topological insulators. Journal of Mathematical Physics, 61(6):061902, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:CMF

[172] Li Chen, Xin Li, Peter Pickl, and Qitao Yin. Combined mean field limit and non-relativistic limit of Vlasov–Maxwell particle system to Vlasov– Poisson system. Journal of Mathematical Physics, 61(6):061903, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mitnik:2020:WEA

[173] Darío M. Mitnik and Santiago A. H. Mitnik. Wavefunctions from energies: Applications in simple potentials. Journal of Mathematical Physics, 61(6):062101, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bley:2020:MSC

[174] Gonzalo A. Bley and Søren Fournais. The magnetic Scott correction for relativistic matter at criticality. Journal of Mathematical Physics, 61(6):062102, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2020:BSP

[175] A. D. Alhaidari and H. Bahlouli. Bound states and the potential parameter spectrum. Journal of Mathematical Physics, 61(6):062103, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [651] and response [652].

Quezada:2020:QPT

[176] L. F. Quezada, A. Martín-Ruiz, and A. Frank. Quantum phase transition of two-level atoms interacting with a finite radiation field. Journal of Mathematical Physics, 61(6):062104, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2020:MTC

[177] Holger Boche, Minglai Cai, and Ning Cai. Message transmission over classical quantum channels with a jammer with side information: Correlation as resource, common randomness generation. Journal of Mathematical Physics, 61(6):062201, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2020:EES

[178] Jinwon Choi, Young-Hoon Kiem, and Seung-Hyeok Kye. Entangled edge states of corank one with positive partial transposes. Journal of Mathematical Physics, 61(6):062202, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goswami:2020:NPE

[179] Abhishek Goswami. A new perturbative expansion for fermionic functional integrals. Journal of Mathematical Physics, 61(6):062301, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajeev:2020:ERV

[180] Karthik Rajeev and T. Padmanabhan. Exploring the Rindler vacuum and the Euclidean plane. Journal of Mathematical Physics, 61(6):062302, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2020:RSE

[181] Yingte Sun. Reducibility of Schrödinger equation at high frequencies. Journal of Mathematical Physics, 61(6):062701, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:RKT

[182] Yin Chen, Jiansheng Geng, and Shuaishuai Xue. Reducible KAM tori for higher dimensional wave equations under nonlocal and forced perturbation. Journal of Mathematical Physics, 61(6):062702, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:2020:EPP

[183] Dionisios Margetis. Edge plasmon-polaritons on isotropic semi-infinite conducting sheets. Journal of Mathematical Physics, 61(6):062901, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2020:CLE

[184] Nhu N. Nguyen and George Yin. A class of Langevin equations with Markov switching involving strong damping and fast switching. Journal of Mathematical Physics, 61(6):063301, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saidi:2020:QLO

[185] El Hassan Saidi. Quantum line operators from Lax pairs. Journal of Mathematical Physics, 61(6):063501, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lotfizadeh:2020:GDO

[186] M. Lotfizadeh. Gauged Dirac operator on the q-deformed fuzzy Euclidean anti-de Sitter space using the pseudo-generalization of q-deformed Ginsparg–Wilson algebra. Journal of Mathematical Physics, 61(6): 063502, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruce:2020:DGS

[187] Andrew James Bruce and Steven Duplij. Double-graded supersymmetric quantum mechanics. Journal of Mathematical Physics, 61(6):063503, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aziziheris:2020:QGQ

[188] K. Aziziheris, H. Fakhri, and S. Laheghi. The quantum group SL_{q}^{∗}(2) and quantum algebra U_{q}(sl_{2}^{∗}) based on a new associative multiplication on 2 × 2 matrices. Journal of Mathematical Physics, 61(6):063504, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scanlon:2020:BIL

[189] Thomas Scanlon and Roman Sverdlov. Berezin integral as a limit of Riemann sum. Journal of Mathematical Physics, 61(6):063505, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [269].

Bagchi:2020:NFN

[190] Bijan Bagchi and Jianke Yang. New families of non-parity-time-symmetric complex potentials with all-real spectra. Journal of Mathematical Physics, 61(6):063506, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alpay:2020:EPD

[191] Daniel Alpay and Ismael L. Paiva. On the extension of positive definite kernels to topological algebras. Journal of Mathematical Physics, 61(6): 063507, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Veliev:2020:SAS

[192] O. A. Veliev. Spectral analysis of the Schrödinger operator with a PTsymmetric periodic optical potential. Journal of Mathematical Physics, 61(6):063508, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ren:2020:QBN

[193] Suling Ren, Caishi Wang, and Yuling Tang. Quantum Bernoulli noises approach to stochastic Schrödinger equation of exclusion type. Journal of Mathematical Physics, 61(6):063509, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novak:2020:SDT

[194] Sebastian Novak and Ingo Runkel. Spin from defects in two-dimensional quantum field theory. Journal of Mathematical Physics, 61(6):063510, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2020:CAI

[195] Francisco M. Fernández. Comment on “The asymptotic iteration method revisited” [j. math. phys. 61, 033501 (2020)]. Journal of Mathematical Physics, 61(6):064101, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [87] and response [196].

Ismail:2020:RCA

[196] Mourad E. H. Ismail and Nasser Saad. Response to “Comment on ‘The asymptotic iteration method revisited”’ [j. math. phys. 61, 064101 (2020)]. Journal of Mathematical Physics, 61(6):064102, June 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [87, 195].

Falek:2020:EBO

[197] M. Falek, M. Merad, and M. Moumni. Erratum: “Bosonic oscillator under a uniform magnetic field with Snyder–de Sitter algebra” [j. math. phys. 60, 013505 (2019)]. Journal of Mathematical Physics, 61(6):069901, June 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11].

Feola:2020:RSE

[198] Roberto Feola, Benoît Grébert, and Trung Nguyen. Reducibility of Schrödinger equation on a Zoll manifold with unbounded potential. Journal of Mathematical Physics, 61(7):071501, July 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:GWP

[199] Zaiyun Zhang, Zhenhai Liu, Youjun Deng, Chuangxia Huang, Shiyou Lin, and Wen Zhu. Global well-posedness and infinite propagation speed for the N-abc family of Camassa–Holm type equation with both dissipation and dispersion. Journal of Mathematical Physics, 61(7):071502, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2020:GEN

[200] Jiali Yu, Yadong Shang, and Huafei Di. Global existence, nonexistence, and decay of solutions for a viscoelastic wave equation with nonlinear boundary damping and source terms. Journal of Mathematical Physics, 61(7):071503, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Selberg:2020:IPT

[201] Sigmund Selberg and Achenef Tesfahun. Ill-posedness of the Thirring model below the critical regularity. Journal of Mathematical Physics, 61(7):071504, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliveira:2020:SIR

[202] Rafael L. Oliveira and Higidio P. Oquendo. Stability and instability results for coupled waves with delay term. Journal of Mathematical Physics, 61(7):071505, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2020:SPS

[203] Juntao Sun and Tsung fang Wu. On Schrödinger–Poisson systems under the effect of steep potential well (2 < p < 4). Journal of Mathematical Physics, 61(7):071506, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2020:VOF

[204] Yi Cheng, Bin Ge, and Ravi P. Agarwal. Variable-order fractional Sobolev spaces and nonlinear elliptic equations with variable exponents. Journal of Mathematical Physics, 61(7):071507, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2020:URT

[205] Jishan Fan, Gen Nakamura, and Tong Tang. Uniform regularity for a two-phase model with magneto field and a related system. Journal of Mathematical Physics, 61(7):071508, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2020:RWB

[206] Xiaofang Dong. Revisit to wave breaking phenomena for the periodic Dullin–Gottwald–Holm equation. Journal of Mathematical Physics, 61 (7):071509, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2020:FFT

[207] Zhentao Jin and Yi Zhou. Formation of finite-time singularities for nonlinear hyperbolic systems with small initial disturbances. Journal of Mathematical Physics, 61(7):071510, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:2020:EIN

[208] Binhua Feng, Jiajia Ren, and Qingxuan Wang. Existence and instability of normalized standing waves for the fractional Schrödinger equations in the L^{2}-supercritical case. Journal of Mathematical Physics, 61(7): 071511, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2020:ECR

[209] Thanh-Nhan Nguyen and Minh-Phuong Tran. An endpoint case of the renormalization property for the relativistic Vlasov–Maxwell system. Journal of Mathematical Physics, 61(7):071512, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Girardot:2020:AFA

[210] Théotime Girardot. Average field approximation for almost bosonic anyons in a magnetic field. Journal of Mathematical Physics, 61(7): 071901, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ouellet:2020:SGP

[211] Mathieu Ouellet and Sébastien Tremblay. Supersymmetric generalized power functions. Journal of Mathematical Physics, 61(7):072101, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Merkli:2020:QET

[212] Marco Merkli, Gennady P. Berman, and Avadh Saxena. Quantum electron transport in degenerate donor-acceptor systems. Journal of Mathematical Physics, 61(7):072102, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akhilesh:2020:GPF

[213] K. S. Akhilesh, Arvind, S. Chaturvedi, K. S. Mallesh, and N. Mukunda. Geometric phases for finite-dimensional systems — the roles of Bargmann invariants, null phase curves, and the Schwinger–Majorana SU(2) framework. Journal of Mathematical Physics, 61(7):072103, July 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Groenland:2020:SRA

[214] Koen Groenland, Carla Groenland, and Reinier Kramer. Stimulated Raman adiabatic passage-like protocols for amplitude transfer generalize to many bipartite graphs. Journal of Mathematical Physics, 61(7):072201, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McKinlay:2020:DRQ

[215] Alexander McKinlay and Marco Tomamichel. Decomposition rules for quantum Rényi mutual information with an application to information exclusion relations. Journal of Mathematical Physics, 61(7):072202, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Battle:2020:AMH

[216] Guy Battle. Across the mass hyperboloid. Journal of Mathematical Physics, 61(7):072301, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2020:SG

[217] Ian Anderson and Charles Torre. Spacetime groups. Journal of Mathematical Physics, 61(7):072501, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herfray:2020:ASI

[218] Yannick Herfray. Asymptotic shear and the intrinsic conformal geometry of null-infinity. Journal of Mathematical Physics, 61(7):072502, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chao:2020:RSM

[219] Ying Chao, Pingyuan Wei, and Jinqiao Duan. The role of slow manifolds in parameter estimation for a multiscale stochastic system with α-stable Lévy noise. Journal of Mathematical Physics, 61(7):072701, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ban:2020:EDS

[220] Jung-Chao Ban, Chih-Hung Chang, and Nai-Zhu Huang. Entropy dimension of shift spaces on monoids. Journal of Mathematical Physics, 61(7):072702, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsamparlis:2020:QFI

[221] Michael Tsamparlis and Antonios Mitsopoulos. Quadratic first integrals of autonomous conservative dynamical systems. Journal of Mathematical Physics, 61(7):072703, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shu:2020:ABN

[222] Ji Shu, Xin Huang, and Jian Zhang. Asymptotic behavior for nonautonomous fractional stochastic Ginzburg–Landau equations on unbounded domains. Journal of Mathematical Physics, 61(7):072704, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2020:PZP

[223] Shu-Chiuan Chang, Roland K. W. Roeder, and Robert Shrock. q-plane zeros of the Potts partition function on diamond hierarchical graphs. Journal of Mathematical Physics, 61(7):073301, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itoi:2020:ARS

[224] C. Itoi and Y. Utsunomiya. Absence of replica symmetry breaking in disordered FKG–Ising models under uniform field. Journal of Mathematical Physics, 61(7):073302, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carrozza:2020:LEH

[225] Sylvain Carrozza, Frank Ferrari, Adrian Tanasa, and Guillaume Valette. On the large D expansion of Hermitian multi-matrix models. Journal of Mathematical Physics, 61(7):073501, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2020:SEL

[226] Mengkun Zhu, Yang Chen, and Chuanzhong Li. The smallest eigenvalue of large Hankel matrices generated by a singularly perturbed Laguerre weight. Journal of Mathematical Physics, 61(7):073502, July 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odzijewicz:2020:ICQ

[227] A. Odzijewicz and E. Wawreniuk. An integrable (classical and quantum) four-wave mixing Hamiltonian system. Journal of Mathematical Physics, 61(7):073503, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ucgun:2020:RTM

[228] Filiz Çagatay Uçgun, Ogul Esen, and Hasan Gümral. Reductions of topologically massive gravity II. First order realizations of second order Lagrangians. Journal of Mathematical Physics, 61(7):073504, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferapontov:2020:CIL

[229] E. V. Ferapontov, I. T. Habibullin, M. N. Kuznetsova, and V. S. Novikov. On a class of 2D integrable lattice equations. Journal of Mathematical Physics, 61(7):073505, July 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kieu:2020:SMF

[230] Thinh Kieu. Solution of the mixed formulation for generalized Forchheimer flows of isentropic gases. Journal of Mathematical Physics, 61 (8):081501, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2020:SLS

[231] Chun Shen. The singular limits of solutions to the Riemann problem for the liquid-gas two-phase isentropic flow model. Journal of Mathematical Physics, 61(8):081502, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liao:2020:CNP

[232] Menglan Liao. A class of nonlinear parabolic equations with anisotropic nonstandard growth conditions. Journal of Mathematical Physics, 61 (8):081503, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheviakov:2020:ICL

[233] A. F. Cheviakov, V. A. Dorodnitsyn, and E. I. Kaptsov. Invariant conservation law-preserving discretizations of linear and nonlinear wave equations. Journal of Mathematical Physics, 61(8):081504, August 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christov:2020:LTA

[234] Ivan C. Christov, Akif Ibraguimov, and Rahnuma Islam. Long-time asymptotics of non-degenerate non-linear diffusion equations. Journal of Mathematical Physics, 61(8):081505, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yan:2020:WPH

[235] Fangchi Yan. Well-posedness of a higher dispersion KdV equation on the half-line. Journal of Mathematical Physics, 61(8):081506, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2020:SBG

[236] Zihua Guo and Jia Shen. Scattering below the ground state for the 2D non-linear Schrödinger and Klein–Gordon equations revisited. Journal of Mathematical Physics, 61(8):081507, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyu:2020:DCS

[237] Wenbin Lyu, Liqing Lu, and Shaohua Wu. Decay characterization of the solutions to the Navier–Stokes–Voigt equations with damping. Journal of Mathematical Physics, 61(8):081508, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Staffilani:2020:HLM

[238] Gigliola Staffilani and Xueying Yu. On the high-low method for NLS on the hyperbolic space. Journal of Mathematical Physics, 61(8):081509, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:2020:HRH

[239] Geoffroy Bergeron, Nicolas Crampé, Satoshi Tsujimoto, Luc Vinet, and Alexei Zhedanov. The Heun–Racah and Heun–Bannai–Ito algebras. Journal of Mathematical Physics, 61(8):081701, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neergaard:2020:FSD

[240] K. Neergård. Fock space dualities. Journal of Mathematical Physics, 61 (8):081702, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hartglass:2020:RRC

[241] Michael Hartglass and Roberto Hernández Palomares. Realizations of rigid c^{∗}-tensor categories as bimodules over GJS c^{∗}-algebras. Journal of Mathematical Physics, 61(8):081703, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lucas:2020:NPD

[242] Andrew Lucas. Non-perturbative dynamics of the operator size distribution in the Sachdev–Ye–Kitaev model. Journal of Mathematical Physics, 61(8):081901, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kryukov:2020:MCQ

[243] Alexey A. Kryukov. Mathematics of the classical and the quantum. Journal of Mathematical Physics, 61(8):082101, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amir:2020:CSP

[244] Naila Amir and Shahid Iqbal. Coherent states of position-dependent mass trapped in an infinite square well. Journal of Mathematical Physics, 61 (8):082102, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schwonnek:2020:WDA

[245] René Schwonnek and Reinhard F. Werner. The Wigner distribution of n arbitrary observables. Journal of Mathematical Physics, 61(8):082103, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geniet:2020:QHU

[246] Paul Geniet. On a quantum Hamiltonian in a unitary magnetic field with axisymmetric potential. Journal of Mathematical Physics, 61(8):082104, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daCosta:2020:DQC

[247] Bruno G. da Costa, Ignacio S. Gomez, and Mariela Portesi. κ-deformed quantum and classical mechanics for a system with position-dependent effective mass. Journal of Mathematical Physics, 61(8):082105, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bebiano:2020:QSN

[248] N. Bebiano, J. da Providência, S. Nishiyama, and J. P. da Providência. A quantum system with a non-Hermitian Hamiltonian. Journal of Mathematical Physics, 61(8):082106, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerrero:2020:SIW

[249] Julio Guerrero and Manuel Berrondo. Semiclassical interpretation of Wei–Norman factorization for SU (1, 1) and its related integral transforms. Journal of Mathematical Physics, 61(8):082107, August 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tahmasbi:2020:SPQ

[250] Mehrdad Tahmasbi and Matthieu R. Bloch. Steganography protocols for quantum channels. Journal of Mathematical Physics, 61(8):082201, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Apadula:2020:SDQ

[251] Luca Apadula, Alessandro Bisio, Giacomo Mauro D’Ariano, and Paolo Perinotti. Symmetries of the Dirac quantum walk and emergence of the de Sitter group. Journal of Mathematical Physics, 61(8):082202, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takakura:2020:PUI

[252] Ryo Takakura and Takayuki Miyadera. Preparation uncertainty implies measurement uncertainty in a class of generalized probabilistic theories. Journal of Mathematical Physics, 61(8):082203, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirokov:2020:SCQ

[253] M. E. Shirokov. Strong convergence of quantum channels: Continuity of the Stinespring dilation and discontinuity of the unitary dilation. Journal of Mathematical Physics, 61(8):082204, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2020:GFP

[254] Felix Finster and Sebastian Kindermann. A gauge fixing procedure for causal fermion systems. Journal of Mathematical Physics, 61(8):082301, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorsky:2020:DTQ

[255] Alexander Gorsky, Peter Koroteev, Olesya Koroteeva, and Arkady Vainshtein. On dimensional transmutation in 1+1D quantum hydrodynamics. Journal of Mathematical Physics, 61(8):082302, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ida:2020:CIC

[256] Daisuke Ida. Complete integrability of cohomogeneity-one strings in rn,1 and canonical form of Killing vector algebra. Journal of Mathematical Physics, 61(8):082501, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morand:2020:EGC

[257] Kevin Morand. Embedding Galilean and carrollian geometries. i. gravitational waves. Journal of Mathematical Physics, 61(8):082502, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jian:2020:RQH

[258] Wenwen Jian. Reducibility of the quantum harmonic oscillator in ddimensions with finitely differentiable perturbations. Journal of Mathematical Physics, 61(8):082701, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Butler:2020:ITC

[259] Leo T. Butler. Invariant tori for a class of singly thermostated Hamiltonians. Journal of Mathematical Physics, 61(8):082702, August 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoshida:2020:DLP

[260] Zensho Yoshida and Philip J. Morrison. Deformation of Lie–Poisson algebras and chirality. Journal of Mathematical Physics, 61(8):082901, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galdi:2020:PSB

[261] Giovanni P. Galdi. On the problem of steady bifurcation of a falling sphere in a Navier–Stokes liquid. Journal of Mathematical Physics, 61 (8):083101, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Site:2020:LTE

[262] Luigi Delle Site and Rupert Klein. Liouville-type equations for the nparticle distribution functions of an open system. Journal of Mathematical Physics, 61(8):083102, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katori:2020:CWP

[263] Makoto Katori and Shinji Koshida. Conformal welding problem, flow line problem, and multiple Schramm–Loewner evolution. Journal of Mathematical Physics, 61(8):083301, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adhikari:2020:FEQ

[264] Arka Adhikari and Christian Brennecke. Free energy of the quantum Sherrington–Kirkpatrick spin-glass model with transverse field. Journal of Mathematical Physics, 61(8):083302, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurasov:2020:SPC

[265] P. Kurasov and P. Sarnak. Stable polynomials and crystalline measures. Journal of Mathematical Physics, 61(8):083501, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2020:GSE

[266] Axel Schulze-Halberg. Generalized Schrödinger equations with quadratical energy-dependence in the potential: Darboux transformations and application to the Heun class. Journal of Mathematical Physics, 61(8): 083502, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bakaleinikov:2020:UAE

[267] Leonid Bakaleinikov and Alexander Silbergleit. Uniform asymptotic expansion of Legendre functions. Journal of Mathematical Physics, 61(8): 083503, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2020:SVQ

[268] T. Skrypnyk. Separation of variables for quadratic algebras: Algebras of Maillet and reflection-equation algebras. Journal of Mathematical Physics, 61(8):083504, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scanlon:2020:PNB

[269] Thomas Scanlon and Roman Sverdlov. Publisher’s note: “Berezin integral as a limit of Riemann sum” [j. math. phys. 61, 063505 (2020)]. Journal of Mathematical Physics, 61(8):089901, August 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [189].

Silva:2020:PAP

[270] Edcarlos D. Silva, J. C. de Albuquerque, and Maxwell L. Silva. Periodic and asymptotically periodic quasilinear elliptic systems. Journal of Mathematical Physics, 61(9):091501, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:GSS

[271] Guoqing Zhang, Anjie Zuo, and Sanyang Liu. Ground state solitary waves for nonlocal nonlinear Schrödinger systems. Journal of Mathematical Physics, 61(9):091502, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsutaya:2020:BSS

[272] Kimitoshi Tsutaya and Yuta Wakasugi. Blow up of solutions of semilinear wave equations in Friedmann–Lemaître–Robertson–Walker spacetime. Journal of Mathematical Physics, 61(9):091503, September 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:ENS

[273] Yong Chen and Lixia Ran. The effect of a noise on the stochastic modified Camassa–Holm equation. Journal of Mathematical Physics, 61(9): 091504, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mizutani:2020:STH

[274] Haruya Mizutani. Scattering theory in homogeneous Sobolev spaces for Schrödinger and wave equations with rough potentials. Journal of Mathematical Physics, 61(9):091505, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2020:DTS

[275] Zhaoxia Liu and Pigong Han. Decay for turbulent solutions of the magneto-hydrodynamic equations in an exterior domain. Journal of Mathematical Physics, 61(9):091506, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashbaugh:2020:SGD

[276] Mark S. Ashbaugh and Derek Kielty. Spectral gaps of 1-D Robin Schrödinger operators with single-well potentials. Journal of Mathematical Physics, 61(9):091507, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2020:MRD

[277] Wulong Liu and Guowei Dai. Multiplicity results for double phase problems in RN. Journal of Mathematical Physics, 61(9):091508, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2020:RWS

[278] Jianfeng Zhou and Zhong Tan. Regularity of weak solutions to a class of nonlinear problem with non-standard growth conditions. Journal of Mathematical Physics, 61(9):091509, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Makino:2020:AOS

[279] Tetu Makino. On adiabatic oscillations of a stratified atmosphere on the flat earth. Journal of Mathematical Physics, 61(9):091510, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samar:2020:RXP

[280] M. I. Samar and V. M. Tkachuk. Regularization of 1/ x^{2} potential in general case of deformed space with minimal length. Journal of Mathematical Physics, 61(9):092101, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGosson:2020:SCG

[281] Maurice A. de Gosson. Symplectic coarse-grained classical and semclassical evolution of subsystems: New theoretical approach. Journal of Mathematical Physics, 61(9):092102, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hislop:2020:ESS

[282] Peter D. Hislop, Werner Kirsch, and M. Krishna. Eigenvalue statistics for Schrödinger operators with random point interactions on R^{d}, d = 1, 2, 3. Journal of Mathematical Physics, 61(9):092103, September 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Asch:2020:SQC

[283] Joachim Asch, Olivier Bourget, and Alain Joye. On stable quantum currents. Journal of Mathematical Physics, 61(9):092104, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galtbayar:2020:SMB

[284] A. Galtbayar, A. Jensen, and K. Yajima. A solvable model of the breakdown of the adiabatic approximation. Journal of Mathematical Physics, 61(9):092105, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taira:2020:LAP

[285] Kouichi Taira. Limiting absorption principle on L^{p}-spaces and scattering theory. Journal of Mathematical Physics, 61(9):092106, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fauvet:2020:RHM

[286] Frédéric Fauvet, Frédéric Menous, and Julien Quéva. Resurgence and holonomy of the φ^{2k} model in zero dimension. Journal of Mathematical Physics, 61(9):092301, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dedushenko:2020:CAL

[287] Mykola Dedushenko and Martin Fluder. Chiral algebra, localization, modularity, surface defects, and all that. Journal of Mathematical Physics, 61(9):092302, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiessling:2020:GRH

[288] Michael K.-H. Kiessling, A. Shadi Tahvildar-Zadeh, and Ebru Toprak. On general-relativistic hydrogen and hydrogenic ions. Journal of Mathematical Physics, 61(9):092303, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gay-Balmaz:2020:DSN

[289] François Gay-Balmaz and Hiroaki Yoshimura. Dirac structures in nonequilibrium thermodynamics for simple open systems. Journal of Mathematical Physics, 61(9):092701, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawaharada:2020:NNS

[290] Akane Kawaharada and Takao Namiki. Number of nonzero states in prefractal sets generated by cellular automata. Journal of Mathematical Physics, 61(9):092702, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2020:RPA

[291] Yongjun Li, Hongqing Wu, and Tinggang Zhao. Random pullback attractor of a non-autonomous local modified stochastic Swift–Hohenberg equation with multiplicative noise. Journal of Mathematical Physics, 61 (9):092703, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:2020:NCB

[292] Adam Doliwa and Rinat M. Kashaev. Non-commutative birational maps satisfying Zamolodchikov equation, and Desargues lattices. Journal of Mathematical Physics, 61(9):092704, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leyvraz:2020:QPS

[293] F. Leyvraz. Qualitative properties of systems of two complex homogeneous ODE’s: a connection to polygonal billiards. Journal of Mathematical Physics, 61(9):092705, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tibboel:2020:REM

[294] Pieter Tibboel. Results on equality of masses for choreographic solutions of n-body problems. Journal of Mathematical Physics, 61(9):092901, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diez:2020:SSC

[295] Tobias Diez and Gerd Rudolph. Singular symplectic cotangent bundle reduction of gauge field theory. Journal of Mathematical Physics, 61(9): 092902, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grinevich:2020:CAP

[296] P. G. Grinevich and R. G. Novikov. Creation and annihilation of pointpotentials using Moutard-type transform in spectral variable. Journal of Mathematical Physics, 61(9):093501, September 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banouh:2020:SCW

[297] Hicham Banouh and Anouar Ben Mabrouk. A sharp Clifford wavelet Heisenberg-type uncertainty principle. Journal of Mathematical Physics, 61(9):093502, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burby:2020:SMQ

[298] J. W. Burby, N. Kallinikos, and R. S. MacKay. Some mathematics for quasi-symmetry. Journal of Mathematical Physics, 61(9):093503, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Humi:2020:EPI

[299] Mayer Humi. On the evolution of a primordial interstellar gas cloud. Journal of Mathematical Physics, 61(9):093504, September 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mukhamedov:2020:QMC

[300] Farrukh Mukhamedov, Abdessatar Barhoumi, Abdessatar Souissi, and Soueidy El Gheteb. A quantum Markov chain approach to phase transitions for quantum Ising model with competing XY-interactions on a Cayley tree. Journal of Mathematical Physics, 61(9):093505, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ontani:2020:SRO

[301] Riccardo Ontani. Some remarks on the operator formalism for nonlocal Poisson brackets. Journal of Mathematical Physics, 61(9):093506, September 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2020:SSC

[302] Boqiang Lü, Xiang Wang, and Xin Zhong. Strong solutions to the 2D Cauchy problem of nonhomogeneous magnetohydrodynamic equations with vacuum. Journal of Mathematical Physics, 61(10):101501, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2020:LUU

[303] Fuzhi Li and Dongmei Xu. Local uniformly upper semi-continuity of random attractor for g-Navier–Stokes equation. Journal of Mathematical Physics, 61(10):101502, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bahrouni:2020:RFP

[304] Anouar Bahrouni, Vicentiu D. Radulescu, and Patrick Winkert. Robin fractional problems with symmetric variable growth. Journal of Mathematical Physics, 61(10):101503, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:2020:EEB

[305] Lingzhong Zeng. Estimates for the eigenvalues of the bi-drifting Laplacian on complete metric measure spaces. Journal of Mathematical Physics, 61(10):101504, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramos:2020:SRE

[306] A. J. A. Ramos, M. M. Freitas, D. S. Almeida, Jr., A. S. Noé, and M. J. Dos Santos. Stability results for elastic porous media swelling with nonlinear damping. Journal of Mathematical Physics, 61(10):101505, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qiu:2020:ABS

[307] Zhaoyang Qiu, Yanbin Tang, and Huaqiao Wang. Asymptotic behavior for the 1D stochastic Landau–Lifshitz–Bloch equation. Journal of Mathematical Physics, 61(10):101506, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ammari:2020:CPB

[308] Kaïs Ammari and Alessandro Duca. Controllability of periodic bilinear quantum systems on infinite graphs. Journal of Mathematical Physics, 61 (10):101507, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Assaad:2020:MST

[309] W. Assaad. Magnetic steps on the threshold of the normal state. Journal of Mathematical Physics, 61(10):101508, October 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaikina:2020:SNS

[310] Elena I. Kaikina and Norma Sotelo-Garcia. Stochastic nonlinear Schrödinger equation on an upper-right quarter plane with Dirichlet random boundary. Journal of Mathematical Physics, 61(10):101509, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:GSS

[311] Jianqing Chen and Qian Zhang. Ground state solution of Neharipohozaev type for periodic quasilinear Schrödinger system. Journal of Mathematical Physics, 61(10):101510, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duca:2020:PQE

[312] Alessandro Duca, Romain Joly, and Dmitry Turaev. Permuting quantum eigenmodes by a quasi-adiabatic motion of a potential wall. Journal of Mathematical Physics, 61(10):101511, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luo:2020:MSN

[313] Tao Luo and Yan-Lin Wang. Multi-scale nonlinear singular limit for thermal non-equilibrium gas flow with multiple non-equilibrium modes for analytic data in multi-dimensions with physical boundaries. Journal of Mathematical Physics, 61(10):101512, October 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:SCS

[314] Wenbo Wang, Quanqing Li, and Yongkun Li. The sign-changing solutions and ground states for planar Schrödinger–Newton system with an exponential critical growth. Journal of Mathematical Physics, 61(10): 101513, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2020:VST

[315] Jianfu Yang and Jinge Yang. Vortex solutions in two-dimensional Bose– Einstein condensates with attraction. Journal of Mathematical Physics, 61(10):101514, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Opanasenko:2020:GSC

[316] Stanislav Opanasenko and Roman O. Popovych. Generalized symmetries and conservation laws of (1 + 1)-dimensional Klein–Gordon equation. Journal of Mathematical Physics, 61(10):101515, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avetisyan:2020:AXK

[317] M. Y. Avetisyan and R. L. Mkrtchyan. On (ad)n(X2)k series of universal quantum dimensions. Journal of Mathematical Physics, 61(10):101701, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farnsworth:2020:GPO

[318] S. Farnsworth. The geometry of physical observables. Journal of Mathematical Physics, 61(10):101702, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapustin:2020:HCS

[319] Anton Kapustin and Nikita Sopenko. Hall conductance and the statistics of flux insertions in gapped interacting lattice systems. Journal of Mathematical Physics, 61(10):101901, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lampart:2020:RBF

[320] Jonas Lampart. The renormalized Bogoliubov-Fröhlich Hamiltonian. Journal of Mathematical Physics, 61(10):101902, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz-Bautista:2020:STC

[321] Erik Díaz-Bautista. Schrödinger-type 2D coherent states of magnetized uniaxially strained graphene. Journal of Mathematical Physics, 61(10): 102101, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manko:2020:ITB

[322] V. I. Man’ko and L. A. Markovich. Integral transforms between tomogram and quasi-probability functions based on quantizer-dequantizer operators formalism. Journal of Mathematical Physics, 61(10):102102, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mana:2020:TDP

[323] Naima Mana, Omar Zaidi, and Mustapha Maamache. Time-dependent pseudo-bosonic coherent states. Journal of Mathematical Physics, 61 (10):102103, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Menguturk:2020:SSR

[324] Levent Ali Mengütürk and Murat Cahit Mengütürk. Stochastic sequential reduction of commutative Hamiltonians. Journal of Mathematical Physics, 61(10):102104, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:ECD

[325] Haonan Zhang. Equality conditions of data processing inequality for α − z Rényi relative entropies. Journal of Mathematical Physics, 61 (10):102201, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kornell:2020:QS

[326] Andre Kornell. Quantum sets. Journal of Mathematical Physics, 61 (10):102202, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mulero-Martinez:2020:QPP

[327] J. I. Mulero-Martínez and J. Molina-Vilaplana. Quantum Pontryagin principle under continuous measurements. Journal of Mathematical Physics, 61(10):102203, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hemame:2020:ESD

[328] Zoubir Hemame, Mokhtar Falek, and Mustafa Moumni. Exact solutions of D-dimensional Klein–Gordon oscillator with Snyder–de Sitter algebra. Journal of Mathematical Physics, 61(10):102301, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gajic:2020:MPQ

[329] Dejan Gajic and Claude Warnick. A model problem for quasinormal ringdown of asymptotically flat or extremal black holes. Journal of Mathematical Physics, 61(10):102501, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almeida:2020:FVP

[330] Ricardo Almeida and Natália Martins. Fractional variational principle of Herglotz for a new class of problems with dependence on the boundaries and a real parameter. Journal of Mathematical Physics, 61(10):102701, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2020:CAL

[331] Seung-Yeal Ha and Hansol Park. Complete aggregation of the Lohe tensor model with the same free flow. Journal of Mathematical Physics, 61 (10):102702, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:COW

[332] Fei Wang, Jun-Min Wang, and Liang-Liang Li. Chaotic oscillations of wave equations due to nonlinear boundary condition. Journal of Mathematical Physics, 61(10):102703, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2020:NAS

[333] F. Calogero, R. Conte, and F. Leyvraz. New algebraically solvable systems of two autonomous first-order ordinary differential equations with purely quadratic right-hand sides. Journal of Mathematical Physics, 61 (10):102704, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramakrishna:2020:ACS

[334] Satish Ramakrishna. Algebra, coherent states, generalized Hermite polynomials, and path integrals for fractional statistics-interpolating from fermions to bosons. Journal of Mathematical Physics, 61(10):103301, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roldan:2020:UPR

[335] J. Roldan and R. Vila. On the ultrametricity property in random field Ising models. Journal of Mathematical Physics, 61(10):103302, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coletti:2020:LTR

[336] Cristian F. Coletti, Lucas R. de Lima, Renato J. Gava, and Denis A. Luiz. Limit theorems for a random walk with memory perturbed by a dynamical system. Journal of Mathematical Physics, 61(10):103303, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sato:2020:SMN

[337] Naoki Sato. Statistical mechanics with non-integrable topological constraints: Self-organization in knotted phase space. Journal of Mathematical Physics, 61(10):103304, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalvoda:2020:NFS

[338] Tomás Kalvoda and Frantisek Stampach. New family of symmetric orthogonal polynomials and a solvable model of a kinetic spin chain. Journal of Mathematical Physics, 61(10):103305, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Erignoux:2020:NFO

[339] Clément Erignoux and Alessandro Giuliani. Nematic first order phase transition for liquid crystals in the van der Waals–Kac limit. Journal of Mathematical Physics, 61(10):103306, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ge:2020:SCH

[340] Yanyan Ge. Super Camassa–Holm-type systems associated to the Kuper– Ramond–Schwarz superalgebra. Journal of Mathematical Physics, 61 (10):103501, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Do:2020:GPD

[341] Ngoc Do, Peter Kuchment, and Frank Sottile. Generic properties of dispersion relations for discrete periodic operators. Journal of Mathematical Physics, 61(10):103502, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lotfizadeh:2020:SQD

[342] M. Lotfizadeh. Super quantum Dirac operator on the q-deformed super fuzzy sphere in instanton sector using quantum super Ginsparg– Wilson algebra. Journal of Mathematical Physics, 61(10):103503, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levin:2020:OSK

[343] A. Levin, M. Olshanetsky, and A. Zotov. Odd supersymmetric Kronecker elliptic function and Yang–Baxter equations. Journal of Mathematical Physics, 61(10):103504, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castillo:2020:STC

[344] K. Castillo, F. R. Rafaeli, and A. Suzuki. Stieltjes’ theorem for classical discrete orthogonal polynomials. Journal of Mathematical Physics, 61 (10):103505, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez:2020:MLD

[345] E. Alvarez and N. C. Snaith. Moments of the logarithmic derivative of characteristic polynomials from SO ( N ) and USp (2 N ). Journal of Mathematical Physics, 61(10):103506, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chebbi:2020:SGD

[346] Yassin Chebbi. Spectral gap of the discrete Laplacian on triangulations. Journal of Mathematical Physics, 61(10):103507, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2020:AST

[347] Yao-Zhong Zhang. Analytic solutions of the Teukolsky equation for massless perturbations of any spin in de Sitter background. Journal of Mathematical Physics, 61(10):103508, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2020:QGM

[348] Chuanzhong Li and Bao Shou. Quantum Gaudin model, spin chains, and universal characters. Journal of Mathematical Physics, 61(10):103509, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:2020:DQG

[349] S. Majid and A. Pachoł. Digital quantum groups. Journal of Mathematical Physics, 61(10):103510, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banerjee:2020:BPS

[350] R. Banerjee and M. Niedermaier. Bonus properties of states of low energy. Journal of Mathematical Physics, 61(10):103511, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taillebois:2020:CQT

[351] E. R. F. Taillebois and A. T. Avelar. Comment on “Quantum time and spatial localization: An analysis of the Hegerfeldt paradox” [j. math. phys. 41, 6093 (2000)]. Journal of Mathematical Physics, 61(10):104101, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2].

Li:2020:ZVC

[352] Yeping Li and Peicheng Zhu. Zero-viscosity-capillarity limit toward rarefaction wave with vacuum for the Navier–Stokes–Korteweg equations of compressible fluids. Journal of Mathematical Physics, 61(11):111501, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiong:2020:CST

[353] Hang Xiong and Xianwen Zhang. On classical solutions of the two-species Vlasov–Poisson system in r^{3} with infinite charge. Journal of Mathematical Physics, 61(11):111502, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhong:2020:GEL

[354] Xin Zhong. Global existence and large time behavior of strong solutions for 3D nonhomogeneous heat conducting Navier–Stokes equations. Journal of Mathematical Physics, 61(11):111503, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2020:RPS

[355] Hongjun Cheng and Hanchun Yang. Riemann problem for the 2D scalar conservation law involving linear fluxes with discontinuous coefficients. Journal of Mathematical Physics, 61(11):111504, November 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chatzikaleas:2020:TPS

[356] Athanasios Chatzikaleas. On time periodic solutions to the conformal cubic wave equation on the Einstein cylinder. Journal of Mathematical Physics, 61(11):111505, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2020:GSS

[357] Huifang Jia. Ground state solutions for the nonlinear Kirchhoff type equations with lower term. Journal of Mathematical Physics, 61(11): 111506, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2020:RMF

[358] Ming Cheng. Recurrent motion in the fractional complex Ginzburg– Landau equation. Journal of Mathematical Physics, 61(11):111507, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Capoferri:2020:SDM

[359] Matteo Capoferri and Dmitri Vassiliev. Spacetime diffeomorphisms as matter fields. Journal of Mathematical Physics, 61(11):111508, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2020:PSR

[360] Youjun Wang and Yimin Zhang. Positive solutions for a relativistic nonlinear Schrödinger equation with square-root nonlinearity. Journal of Mathematical Physics, 61(11):111509, November 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suh:2020:SSC

[361] Uhi Rinn Suh. Structures of (supersymmetric) classical w-algebras. Journal of Mathematical Physics, 61(11):111701, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2020:TIL

[362] Inkang Kim and Genkai Zhang. Toledo invariant of lattices in SU(2,1) via symmetric square. Journal of Mathematical Physics, 61(11):111702, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qi:2020:PCP

[363] Zihao Qi and Guodong Zhou. Poisson cohomology of plane Poisson structures with isolated singularities revisited. Journal of Mathematical Physics, 61(11):111703, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelasCuevas:2020:QMS

[364] Gemma De las Cuevas, Tom Drescher, and Tim Netzer. Quantum magic squares: Dilations and their limitations. Journal of Mathematical Physics, 61(11):111704, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smilga:2020:EDT

[365] A. V. Smilga. An eight-dimensional Taub–NUT-like hyper-Kähler metric in harmonic superspace formalism. Journal of Mathematical Physics, 61(11):112301, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thuillier:2020:TMH

[366] F. Thuillier. 3D topological models and Heegaard splitting. II. Pontryagin duality and observables. Journal of Mathematical Physics, 61(11): 112302, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2020:LAC

[367] Felix Finster and Marco Oppio. Local algebras for causal fermion systems in Minkowski space. Journal of Mathematical Physics, 61(11):112303, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borji:2020:PRL

[368] Majdouline Borji and Christoph Kopper. Perturbative renormalization of the lattice regularized φ_{4}^{4} with flow equations. Journal of Mathematical Physics, 61(11):112304, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Browning:2020:CLU

[369] Thomas L. Browning, Benjamin H. Feintzeig, Robin Gates-Redburg, Jonah Librande, and Rory Soiffer. Classical limits of unbounded quantities by strict quantization. Journal of Mathematical Physics, 61(11): 112305, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2020:GBC

[370] Pak Tung Ho. The Gauss–Bonnet–Chern mass under geometric flows. Journal of Mathematical Physics, 61(11):112501, November 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciric:2020:AEC

[371] Marija Dimitrijević Ćirić, Grigorios Giotopoulos, Voja Radovanović, and Richard J. Szabo. L_{∞}-algebras of Einstein–Cartan–Palatini gravity. Journal of Mathematical Physics, 61(11):112502, November 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2020:MHM

[372] Zhuobin Liang and Xiao Zhang. Metrics of Horowitz–Myers type with the negative constant scalar curvature. Journal of Mathematical Physics, 61(11):112503, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2020:FDA

[373] Guo Liu, Tianxiu Lu, Xiaofang Yang, and Anwar Waseem. Further discussion about transitivity and mixing of continuous maps on compact metric spaces. Journal of Mathematical Physics, 61(11):112701, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2020:RRA

[374] Dingshi Li, Lin Shi, and Junyilang Zhao. Regular random attractors for non-autonomous stochastic evolution equations with time-varying delays on thin domains. Journal of Mathematical Physics, 61(11):112702, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:2020:RDC

[375] A. V. Tsiganov. Reduction of divisors for classical superintegrable GL (3) magnetic chain. Journal of Mathematical Physics, 61(11):112703, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2020:LSI

[376] Lu Chen and Jianhua Shen. Lagrange stability for impulsive pendulumtype equations. Journal of Mathematical Physics, 61(11):112704, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saluto:2020:CHF

[377] Lidia Saluto and David Jou. Coupling of heat flux and vortex polarization in superfluid helium. Journal of Mathematical Physics, 61(11):113101, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goto:2020:DEM

[378] Shin itiro Goto and Hideitsu Hino. Diffusion equations from master equations — a discrete geometric approach. Journal of Mathematical Physics, 61(11):113301, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koukiou:2020:FLT

[379] Flora Koukiou. Freezing and low temperature entropy: The case of meanfield Gaussian model. Journal of Mathematical Physics, 61(11):113302, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stachura:2020:NEB

[380] Piotr Stachura, Wiesław Pusz, and Jacek Wojtkiewicz. Non-existence of Bose–Einstein condensation in Bose–Hubbard model in dimensions 1 and 2. Journal of Mathematical Physics, 61(11):113303, November 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loring:2020:CFH

[381] Terry A. Loring and Fredy Vides. Computing Floquet Hamiltonians with symmetries. Journal of Mathematical Physics, 61(11):113501, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anerot:2020:NTT

[382] Baptiste Anerot, Jacky Cresson, Khaled Hariz Belgacem, and Frederic Pierret. Noether’s-type theorems on time scales. Journal of Mathematical Physics, 61(11):113502, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1336].

Mietzsch:2020:VLD

[383] Nicco Mietzsch. The validity of the local density approximation for smooth short range interaction potentials. Journal of Mathematical Physics, 61(11):113503, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moya-Cessa:2020:CTD

[384] H. M. Moya-Cessa and J. Récamier. Comment on “Time-dependent coupled harmonic oscillators” [j. math. phys. 53, 052101 (2012)]. Journal of Mathematical Physics, 61(11):114101, November 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4].

DeNittis:2020:EEL

[385] Giuseppe De Nittis and Max Lein. Erratum: “Exponentially localized Wannier functions in periodic zero flux magnetic fields” [j. math. phys. 52, 112103 (2011)]. Journal of Mathematical Physics, 61(11):119901, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [3].

Guo:2020:MBS

[386] Lun Guo and Qi Li. Multiple bound state solutions for fractional Choquard equation with Hardy–Littlewood–Sobolev critical exponent.

Journal of Mathematical Physics, 61(12):121501, December 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2020:ISD

[387] Suzhen Jiang and Yujiang Wu. An inverse space-dependent source problem for a multi-term time fractional diffusion equation. Journal of Mathematical Physics, 61(12):121502, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2020:BIW

[388] Xiaofang Dong. Blow-up issues for the weakly dissipative periodic bfamily of equations revisited. Journal of Mathematical Physics, 61(12): 121503, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mishra:2020:OHL

[389] Satyendra Kumar Mishra and Anita Naolekar. O-operators on hom-Lie algebras. Journal of Mathematical Physics, 61(12):121701, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeVen:2020:CLM

[390] Christiaan J. F. van de Ven. The classical limit of mean-field quantum spin systems. Journal of Mathematical Physics, 61(12):121901, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Santos:2020:HFN

[391] Jonas F. G. Santos. Heat flow and noncommutative quantum mechanics in phase-space. Journal of Mathematical Physics, 61(12):122101, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhaidari:2020:SSE

[392] A. D. Alhaidari. Solving Schrödinger equation by mapping it into a Heuntype equation with known solutions. Journal of Mathematical Physics, 61(12):122102, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maioli:2020:ESL

[393] Alan C. Maioli and Alexandre G. M. Schmidt. Exact solutions for the Lippmann–Schwinger equation in two dimensions and invisibility conditions. Journal of Mathematical Physics, 61(12):122103, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lucas:2020:OGB

[394] Andrew Lucas and Andrew Osborne. Operator growth bounds in a cartoon matrix model. Journal of Mathematical Physics, 61(12):122301, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avalos:2020:RTS

[395] R. Avalos and J. H. Lira. Reduced thin-sandwich equations on manifolds Euclidean at infinity and on closed manifolds: Existence and multiplicity. Journal of Mathematical Physics, 61(12):122501, December 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aadne:2020:NKV

[396] Matthew Terje Aadne. Nil-Killing vector fields and type III deformations. Journal of Mathematical Physics, 61(12):122502, December 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gasperin:2020:ZRM

[397] E. Gasperín and J. A. Valiente Kroon. Zero rest-mass fields and the Newman–Penrose constants on flat space. Journal of Mathematical Physics, 61(12):122503, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flathmann:2020:GEP

[398] Kai Flathmann and Noa Wassermann. Geodesic equations for particles and light in the black spindle spacetime. Journal of Mathematical Physics, 61(12):122504, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsamparlis:2020:FIH

[399] Michael Tsamparlis and Antonios Mitsopoulos. First integrals of holonomic systems without Noether symmetries. Journal of Mathematical Physics, 61(12):122701, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohmori:2020:UBO

[400] Shousuke Ohmori and Yoshihiro Yamazaki. Ultradiscrete bifurcations for one dimensional dynamical systems. Journal of Mathematical Physics, 61(12):122702, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2020:SLE

[401] Alexander Figotin. Synthesis of lossless electric circuits based on prescribed Jordan forms. Journal of Mathematical Physics, 61(12):122703, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Johnson:2020:EEU

[402] K. Johnson and B. Simanek. Electrostatic equilibria on the unit circle via Jacobi polynomials. Journal of Mathematical Physics, 61(12):122901, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aghapour:2020:ZEC

[403] Sajad Aghapour, Lars Andersson, and Kjell Rosquist. The zilch electromagnetic conservation law revisited. Journal of Mathematical Physics, 61(12):122902, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLenaghan:2020:CSH

[404] Raymond G. McLenaghan, Giovanni Rastelli, and Carlos Valero. Complete separability of the Hamilton–Jacobi equation for the charged particle orbits in a Liénard–Wiechert field. Journal of Mathematical Physics, 61(12):122903, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agliari:2020:RHM

[405] Elena Agliari, Alberto Fachechi, and Chiara Marullo. The relativistic Hopfield model with correlated patterns. Journal of Mathematical Physics, 61(12):123301, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nazarov:2020:FNN

[406] Fedor Nazarov and Mikhail Sodin. Fluctuations in the number of nodal domains. Journal of Mathematical Physics, 61(12):123302, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krupka:2020:VPP

[407] Demeter Krupka. Variational principles: Projectability onto Grassmann fibrations. Journal of Mathematical Physics, 61(12):123501, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Filipkovska:2020:PEF

[408] M. S. Filipkovska and V. P. Kotlyarov. Propagation of electric field generated by periodic pumping in a stable medium of two-level atoms of the Maxwell–Bloch model. Journal of Mathematical Physics, 61(12): 123502, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burrows:2020:AMF

[409] Antony Burrows, Shaun Cooper, Elke Pahl, and Peter Schwerdtfeger. Analytical methods for fast converging lattice sums for cubic and hexagonal close-packed structures. Journal of Mathematical Physics, 61(12): 123503, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mori:2020:MDA

[410] Haruka Mori and Shin Sasaki. More on doubled aspects of algebroids in double field theory. Journal of Mathematical Physics, 61(12):123504, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beckus:2020:SCA

[411] Siegfried Beckus, Jean Bellissard, and Giuseppe De Nittis. Spectral continuity for aperiodic quantum systems: Applications of a folklore theorem. Journal of Mathematical Physics, 61(12):123505, December 2020. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caudrelier:2020:HMD

[412] Vincent Caudrelier and Matteo Stoppato. Hamiltonian multiform description of an integrable hierarchy. Journal of Mathematical Physics, 61(12):123506, December 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2021:ESS

[413] Yanheng Ding, Qi Guo, and Yuanyang Yu. Existence of semiclassical solutions for some critical Dirac equation. Journal of Mathematical Physics, 62(1):011501, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kenig:2021:ASS

[414] Carlos E. Kenig. Asymptotic simplification for solutions of the energy critical nonlinear wave equation. Journal of Mathematical Physics, 62 (1):011502, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bal:2021:HHT

[415] Guillaume Bal, Andrew Lucas, and Mitchell Luskin. Homogenization of hydrodynamic transport in Dirac fluids. Journal of Mathematical Physics, 62(1):011503, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singla:2021:SCE

[416] Komal Singla and R. K. Gupta. Symmetry classification and exact solutions of (3 + 1)-dimensional fractional nonlinear incompressible nonhydrostatic coupled Boussinesq equations. Journal of Mathematical Physics, 62(1):011504, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:CSC

[417] Xiaohui Wang. Compressible subsonic cavity flow in a nozzle. Journal of Mathematical Physics, 62(1):011505, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:ECB

[418] Jian Zhang and Zhenluo Lou. Existence and concentration behavior of solutions to Kirchhoff type equation with steep potential well and critical growth. Journal of Mathematical Physics, 62(1):011506, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nhan:2021:SUS

[419] Le Cong Nhan and Le Xuan Truong. Stable and unstable sets for damped nonlinear wave equations with variable exponent sources. Journal of Mathematical Physics, 62(1):011507, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2021:GES

[420] Yuebo Cao, Yi Peng, and Ying Sun. Global existence of strong solutions to MHD with density-depending viscosity and temperature-depending heat-conductivity in unbounded domains. Journal of Mathematical Physics, 62(1):011508, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goudon:2021:NIS

[421] T. Goudon and L. Vivion. Numerical investigation of solitary wave stability for quantum dissipative systems. Journal of Mathematical Physics, 62(1):011509, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molev:2021:CES

[422] A. I. Molev. Casimir elements and Sugawara operators for Takiff algebras. Journal of Mathematical Physics, 62(1):011701, January 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bachmann:2021:RIQ

[423] Sven Bachmann, Alex Bols, Wojciech De Roeck, and Martin Fraas. Rational indices for quantum ground state sectors. Journal of Mathematical Physics, 62(1):011901, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stan:2021:PMD

[424] A. I. Stan, G. Popa, and R. Dutta. Position-momentum decomposition of linear operators defined on algebras of polynomials. Journal of Mathematical Physics, 62(1):012101, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2021:MQM

[425] Le Bin Ho and Yasushi Kondo. Multiparameter quantum metrology with postselection measurements. Journal of Mathematical Physics, 62 (1):012102, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-Perez:2021:DSA

[426] Armando Martínez-Pérez and Gabino Torres-Vega. Discrete selfadjointness and quantum dynamics. travel times. Journal of Mathematical Physics, 62(1):012103, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szczygielski:2021:MDU

[427] Krzysztof Szczygielski. Markovian dynamics under weak periodic coupling. Journal of Mathematical Physics, 62(1):012104, January 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lotoreichik:2021:FKI

[428] Vladimir Lotoreichik and Alessandro Michelangeli. Faber–Krahn inequalities for Schrödinger operators with point and with Coulomb interactions. Journal of Mathematical Physics, 62(1):012105, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haah:2021:CTI

[429] Jeongwan Haah. Classification of translation invariant topological Pauli stabilizer codes for prime dimensional qudits on two-dimensional lattices. Journal of Mathematical Physics, 62(1):012201, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besnard:2021:ESM

[430] Fabien Besnard. A U(1)_{B−L}-extension of the standard model from noncommutative geometry. Journal of Mathematical Physics, 62(1):012301, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2021:SST

[431] F. Calogero and F. Payandeh. Solvable systems of two coupled first-order ODEs with homogeneous cubic polynomial right-hand sides. Journal of Mathematical Physics, 62(1):012701, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nucci:2021:MSS

[432] M. C. Nucci and R. Campoamor-Stursberg. Maximally superintegrable systems in flat three-dimensional space are linearizable. Journal of Mathematical Physics, 62(1):012702, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pina:2021:FBC

[433] E. Piña. Five-body central configurations and symmetry. Journal of Mathematical Physics, 62(1):012901, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2021:APS

[434] Hai Huang and Xianlong Fu. Asymptotic properties of solutions for impulsive neutral stochastic functional integro-differential equations. Journal of Mathematical Physics, 62(1):013301, January 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamaga:2021:WRN

[435] Kazuki Yamaga. Work relation in non-equilibrium steady states of onedimensional quantum lattice systems. Journal of Mathematical Physics, 62(1):013302, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stepanenko:2021:SIP

[436] Alexei Stepanenko. Spectral inclusion and pollution for a class of dissipative perturbations. Journal of Mathematical Physics, 62(1):013501, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:2021:GZH

[437] Silas L. Carvalho and César R. de Oliveira. Generic zero-Hausdorff and one-packing spectral measures. Journal of Mathematical Physics, 62(1): 013502, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aremua:2021:DOF

[438] Isiaka Aremua, Mahouton Norbert Hounkonnou, and Ezinvi Baloïtcha. Density operator formulation for magnetic systems: Physical and mathematical aspects. Journal of Mathematical Physics, 62(1):013503, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sheftel:2021:NSC

[439] M. B. Sheftel. Nonlocal symmetry of CMA generates ASD Ricci-flat metric with no Killing vectors. Journal of Mathematical Physics, 62 (1):013504, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:2021:SLA

[440] Geoffroy Bergeron, Julien Gaboriaud, Luc Vinet, and Alexei Zhedanov. Sklyanin-like algebras for (q-)linear grids and (q-)para-Krawtchouk polynomials. Journal of Mathematical Physics, 62(1):013505, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skripka:2021:LEF

[441] A. Skripka. Lipschitz estimates for functions of Dirac and Schrödinger operators. Journal of Mathematical Physics, 62(1):013506, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lotfizadeh:2021:DOQ

[442] M. Lotfizadeh. Dirac operator on the quantum fuzzy four-sphere SqF4. Journal of Mathematical Physics, 62(1):013507, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnad:2021:BEL

[443] J. Harnad and A. Yu. Orlov. Bilinear expansions of lattices of KP τfunctions in BKP τ-functions: a fermionic approach. Journal of Mathematical Physics, 62(1):013508, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2021:DOR

[444] Choon-Lin Ho and Ryu Sasaki. Discrete orthogonality relations for multi-indexed Laguerre and Jacobi polynomials. Journal of Mathematical Physics, 62(1):013509, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nobe:2021:PLI

[445] Atsushi Nobe and Junta Matsukidaira. Periodicity, linearizability, and integrability in seed mutations of type AN(1). Journal of Mathematical Physics, 62(1):013510, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

De:2021:SET

[446] Debabrata De and Kunal Mukherjee. On symmetric embedding of type I factors. Journal of Mathematical Physics, 62(1):013511, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2021:CQA

[447] Francisco M. Fernández. Comment on: “Quantum aspects of a moving magnetic quadrupole moment interacting with an electric field” [j. math. phys. 56, 062107 (2015)]. Journal of Mathematical Physics, 62(1):014101, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [7].

Fujiie:2021:PNS

[448] Setsuro Fujiié and Spyridon Kamvissis. Publisher’s note: “Semiclassical WKB problem for the non-self-adjoint Dirac operator with analytic potential” [j. math. phys. 61, 011510 (2020)]. Journal of Mathematical Physics, 62(1):019901, January 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [25].

Zirnbauer:2021:PHS

[449] Martin R. Zirnbauer. Particle-hole symmetries in condensed matter. Journal of Mathematical Physics, 62(2):021101, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2021:SVS

[450] Jinbiao Wu. Stochastic viscosity solutions for stochastic integral-partial differential equations. Journal of Mathematical Physics, 62(2):021501, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2021:GES

[451] Chunyin Jin. Global existence of strong solutions to the kinetic Cucker– Smale model coupled with the Stokes equations. Journal of Mathematical Physics, 62(2):021502, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhai:2021:GSD

[452] Xiaoping Zhai. Global solutions to the n-dimensional incompressible Oldroyd-B model without damping mechanism. Journal of Mathematical Physics, 62(2):021503, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rickard:2021:VBP

[453] Calum Rickard. The vacuum boundary problem for the spherically symmetric compressible Euler equations with positive density and unbounded entropy. Journal of Mathematical Physics, 62(2):021504, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:DNS

[454] Yongsheng Li and Fangyan Yao. Derivation of the nonlinear Schrödinger equation with a general nonlinearity and Gross–Pitaevskii hierarchy in one and two dimensions. Journal of Mathematical Physics, 62(2):021505, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:SASa

[455] W.-M. Wang. Semi-algebraic sets method in PDE and mathematical physics. Journal of Mathematical Physics, 62(2):021506, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramos:2021:SSP

[456] A. J. A. Ramos, D. S. Almeida Júnior, M. M. Freitas, A. S. Noé, and M. J. Dos Santos. Stabilization of swelling porous elastic soils with fluid saturation and delay time terms. Journal of Mathematical Physics, 62 (2):021507, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valero-Valdes:2021:SGO

[457] Carlos Valero-Valdés. Singular geometrical optics for differential operators on surfaces. Journal of Mathematical Physics, 62(2):021508, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2021:EFD

[458] Shujuan Liu and Guanghua Shi. The existence of full-dimensional invariant tori for an almost-periodically forced nonlinear beam equation. Journal of Mathematical Physics, 62(2):021509, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2021:SBL

[459] Shijin Ding, Zhilin Lin, and Dongjuan Niu. Stability of the boundary layer expansion for the 3D plane parallel MHD flow. Journal of Mathematical Physics, 62(2):021510, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2021:GSS

[460] Yunshun Wu, Li Xiao, and Rongfeng Yu. Global strong solutions for viscous radiative gas with degenerate temperature dependent heat conductivity in one-dimensional unbounded domains. Journal of Mathematical Physics, 62(2):021511, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balogh:2021:IGP

[461] F. Balogh, J. Harnad, and J. Hurtubise. Isotropic Grassmannians, Plücker and Cartan maps. Journal of Mathematical Physics, 62(2): 021701, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kac:2021:PTF

[462] Victor G. Kac, Natasha Rozhkovskaya, and Johan van de Leur. Polynomial tau-functions of the KP, BKP, and the s-component KP hierarchies.

Journal of Mathematical Physics, 62(2):021702, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasnov:2021:CSM

[463] Kirill Krasnov. SO(9) characterization of the standard model gauge group. Journal of Mathematical Physics, 62(2):021703, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiong:2021:RVL

[464] Haoren Xiong. Resonances as viscosity limits for exponentially decaying potentials. Journal of Mathematical Physics, 62(2):022101, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

John:2021:CPF

[465] Tiju Cherian John and K. R. Parthasarathy. A common parametrization for finite mode Gaussian states, their symmetries, and associated contractions with some applications. Journal of Mathematical Physics, 62 (2):022102, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mann:2021:EAA

[466] Ryan L. Mann and Tyler Helmuth. Efficient algorithms for approximating quantum partition functions. Journal of Mathematical Physics, 62 (2):022201, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhojraj:2021:QAR

[467] Tejas Bhojraj. Quantum algorithmic randomness. Journal of Mathematical Physics, 62(2):022202, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huber:2021:PMT

[468] Felix Huber. Positive maps and trace polynomials from the symmetric group. Journal of Mathematical Physics, 62(2):022203, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stedman:2021:EST

[469] Richard Stedman and Ian A. B. Strachan. Extended ∨-systems and trigonometric solutions to the WDVV equations. Journal of Mathematical Physics, 62(2):022301, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cammarota:2021:NDR

[470] Valentina Cammarota, Domenico Marinucci, and Igor Wigman. Nodal deficiency of random spherical harmonics in presence of boundary. Journal of Mathematical Physics, 62(2):022701, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2021:PAL

[471] Kaixuan Zhu, Yongqin Xie, Feng Zhou, and Qiyuan Zhou. Pullback attractors for p-Laplacian equations with delays. Journal of Mathematical Physics, 62(2):022702, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jian:2021:ALL

[472] Wenwen Jian, Jia Shi, and Xiaoping Yuan. Anderson localization for long-range operators with singular potentials. Journal of Mathematical Physics, 62(2):022703, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2021:ABN

[473] Zhang Chen and Lingyu Li. Asymptotic behavior of non-autonomous stochastic complex Ginzburg–Landau equations on unbounded thin domains. Journal of Mathematical Physics, 62(2):022704, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2021:EAS

[474] Pengyu Chen and Xuping Zhang. Existence of attractors for stochastic diffusion equations with fractional damping and time-varying delay. Journal of Mathematical Physics, 62(2):022705, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amarel:2021:RRE

[475] J. Amarel, D. Belitz, and T. R. Kirkpatrick. Rigorous results for the electrical conductivity due to electron-phonon scattering. Journal of Mathematical Physics, 62(2):023301, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jansen:2021:RGA

[476] Sabine Jansen. Revisiting Groeneveld’s approach to the virial expansion. Journal of Mathematical Physics, 62(2):023302, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bohacik:2021:TDP

[477] J. Bohácik, P. Presnajder, and P. Augustín. Time-dependent propagator for an-harmonic oscillator with quartic term in potential. Journal of Mathematical Physics, 62(2):023501, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lugiewicz:2021:PSW

[478] Piotr Lugiewicz, Lech Jakóbczyk, and Andrzej Frydryszak. Properties of states on Weyl algebra with variable multiplication law. Journal of Mathematical Physics, 62(2):023502, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Johnson:2021:ERR

[479] Bruce R. Johnson. Erratum: “The radial reduced Coulomb Green’s function” [j. math. phys. 20, 2484 (1979)]. Journal of Mathematical Physics, 62(2):029901, February 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1].

Xu:2021:LWP

[480] Qingmei Xu and Xin Zhong. Local well-posedness to the threedimensional barotropic compressible magnetohydrodynamic equations with vacuum. Journal of Mathematical Physics, 62(3):031501, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2021:SHB

[481] Ho Lee. The spatially homogeneous Boltzmann equation for massless particles in an FLRW background. Journal of Mathematical Physics, 62 (3):031502, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2021:LTT

[482] Huiying Fan and Meng Wang. The Liouville type theorem for the stationary magnetohydrodynamic equations. Journal of Mathematical Physics, 62(3):031503, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:NGS

[483] Zuo Yang. Normalized ground state solutions for Kirchhoff type systems. Journal of Mathematical Physics, 62(3):031504, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2021:IKT

[484] Juntao Sun, Kuan-Hsiang Wang, and Tsung fang Wu. On indefinite Kirchhoff-type equations under the combined effect of linear and superlinear terms. Journal of Mathematical Physics, 62(3):031505, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kang:2021:ECN

[485] Lingping Kang, Xuemei Deng, and Qunyi Bie. Energy conservation for the nonhomogeneous incompressible ideal Hall–MHD equations. Journal of Mathematical Physics, 62(3):031506, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pinheiro:2021:NSV

[486] Cristyan Pinheiro and Gabriela Planas. On the α-Navier–Stokes–Vlasov and the α-Navier–Stokes–Vlasov–Fokker–Planck equations. Journal of Mathematical Physics, 62(3):031507, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2021:IMS

[487] Jae-Myoung Kim and Jung-Hyun Bae. Infinitely many solutions of fractional Schrödinger–Maxwell equations. Journal of Mathematical Physics, 62(3):031508, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khoutir:2021:LES

[488] Sofiane Khoutir. Least energy sign-changing solutions for a class of Schrödinger–Poisson system on bounded domains. Journal of Mathematical Physics, 62(3):031509, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chern:2021:CSS

[489] Jann-Long Chern, Zhi-You Chen, and Hung-Ying Shen. Classification of solutions for self-dual Chern–Simons CP (1) model. Journal of Mathematical Physics, 62(3):031510, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dinh:2021:PCT

[490] Van Duong Dinh. Probabilistic Cauchy theory for the mass-critical fourth-order nonlinear Schrödinger equation. Journal of Mathematical Physics, 62(3):031511, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Al-Mahdi:2021:GEA

[491] Adel M. Al-Mahdi, Mohammad M. Al-Gharabli, Mohammad Kafini, and Shadi Al-Omari. On the global existence and asymptotic behavior of the solution of a nonlinear wave equation with past history. Journal of Mathematical Physics, 62(3):031512, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:EUG

[492] Yan Li and Yong Luo. Existence and uniqueness of ground states for attractive Bose–Einstein condensates in box-shaped traps. Journal of Mathematical Physics, 62(3):031513, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jeltema:2021:LDP

[493] Tesla Jeltema, Stefano Profumo, and Jaryd F. Ulbricht. Lowestdimensional portals to SU(N) exotics. Journal of Mathematical Physics, 62(3):031701, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fakhri:2021:LCF

[494] H. Fakhri and S. Laheghi. Left-covariant first order differential calculus on quantum Hopf supersymmetry algebra. Journal of Mathematical Physics, 62(3):031702, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bakalov:2021:RTT

[495] Bojko Bakalov and Samantha Kirk. Representations of twisted toroidal Lie algebras from twisted modules over vertex algebras. Journal of Mathematical Physics, 62(3):031703, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghanta:2021:GSP

[496] Rohan Ghanta. Ground state of the polaron hydrogenic atom in a strong magnetic field. Journal of Mathematical Physics, 62(3):031901, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Correggi:2021:MPA

[497] Michele Correggi and Davide Fermi. Magnetic perturbations of anyonic and Aharonov–Bohm Schrödinger operators. Journal of Mathematical Physics, 62(3):032101, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Facchi:2021:SPS

[498] Paolo Facchi, Marilena Ligabò, and Davide Lonigro. Spectral properties of the singular Friedrichs–Lee Hamiltonian. Journal of Mathematical Physics, 62(3):032102, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beschastnyi:2021:OST

[499] Ivan Beschastnyi, Ugo Boscain, and Mario Sigalotti. An obstruction to small-time controllability of the bilinear Schrödinger equation. Journal of Mathematical Physics, 62(3):032103, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2021:IWF

[500] Jianping Dong and Ying Lu. Infinite wall in the fractional quantum mechanics. Journal of Mathematical Physics, 62(3):032104, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoffmann:2021:UCS

[501] Scott E. Hoffmann. Unitary, continuum, stationary perturbation theory for the radial Schrödinger equation. Journal of Mathematical Physics, 62(3):032105, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amore:2021:UTT

[502] Paolo Amore and Francisco M. Fernández. An ubiquitous three-term recurrence relation. Journal of Mathematical Physics, 62(3):032106, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goldberg:2021:SLC

[503] Adina Goldberg. Synchronous linear constraint system games. Journal of Mathematical Physics, 62(3):032201, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aguilar:2021:NEW

[504] Pedro Aguilar, Chryssomalis Chryssomalakos, and Edgar Guzmán-González. Noise effects on the Wilczek–Zee geometric phase. Journal of Mathematical Physics, 62(3):032202, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:LUC

[505] C. H. Wang, J. T. Yuan, Y. H. Yang, and G. F. Mu. Local unitary classification of generalized Bell state sets in C^{5} ⊗ C^{5}. Journal of Mathematical Physics, 62(3):032203, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reuvers:2021:GPC

[506] Robin Reuvers. Generalized Pauli constraints in large systems: The Pauli principle dominates. Journal of Mathematical Physics, 62(3):032204, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [607].

Dereli:2021:MCG

[507] Tekin Dereli and Keremcan Dogan. Metric-connection geometries on pre-Leibniz algebroids: a search for geometrical structure in string models. Journal of Mathematical Physics, 62(3):032301, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kreutzer:2021:CAE

[508] Lars T. Kreutzer. Canonical analysis of E6(6)(R) invariant five dimensional (super-)gravity. Journal of Mathematical Physics, 62(3):032302, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Letsios:2021:ESQ

[509] Vasileios A. Letsios. The eigenmodes for spinor quantum field theory in global de Sitter space-time. Journal of Mathematical Physics, 62(3): 032303, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gudnason:2021:BVS

[510] Sven Bjarke Gudnason, Minoru Eto, and Muneto Nitta. 1/2-BPS vortex strings in N = 2 supersymmetric U(1)^{N} gauge theories. Journal of Mathematical Physics, 62(3):032304, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khan:2021:CSR

[511] Abu Mohammad Khan. Continuous spin representation from contraction of the conformal algebra. Journal of Mathematical Physics, 62(3):032305, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ikeda:2021:GAD

[512] Noriaki Ikeda and Shin Sasaki. Global aspects of doubled geometry and pre-rackoid. Journal of Mathematical Physics, 62(3):032306, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Futaki:2021:HMS

[513] Masahiro Futaki and Hiroshige Kajiura. Homological mirror symmetry of CPn and their products via Morse homotopy. Journal of Mathematical Physics, 62(3):032307, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

De:2021:PFS

[514] U. C. De, S. K. Chaubey, and S. Shenawy. Perfect fluid spacetimes and Yamabe solitons. Journal of Mathematical Physics, 62(3):032501, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Collingbourne:2021:GLI

[515] Sam C. Collingbourne. The Gregory-laflamme instability of the Schwarzschild black string exterior. Journal of Mathematical Physics, 62(3):032502, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allen:2021:DLP

[516] Samantha Allen and Jacob H. Swenberg. Do link polynomials detect causality in globally hyperbolic spacetimes? Journal of Mathematical Physics, 62(3):032503, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aouadi:2021:DNR

[517] Moncef Aouadi, Imed Mahfoudhi, and Taoufik Moulahi. Decay and numerical results in nonsimple viscoelasticity. Journal of Mathematical Physics, 62(3):032701, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:EAS

[518] Pin Wang and Guanggan Chen. Effective approximation of stochastic sine-Gordon equation with a fast oscillation. Journal of Mathematical Physics, 62(3):032702, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yaremko:2021:EFS

[519] Yurij Yaremko. Electrodynamics in flat spacetime of six dimensions. Journal of Mathematical Physics, 62(3):032901, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaturvedi:2021:PCC

[520] Rahul Kumar Chaturvedi and L. P. Singh. The phenomena of concentration and cavitation in the Riemann solution for the isentropic zeropressure dusty gasdynamics. Journal of Mathematical Physics, 62(3): 033101, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:HMT

[521] Wuchen Li. Hessian metric via transport information geometry. Journal of Mathematical Physics, 62(3):033301, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valls:2021:RLC

[522] Clàudia Valls. Rational limit cycles on generalized bernouilli polynomial equations. Journal of Mathematical Physics, 62(3):033501, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muranova:2021:EIO

[523] Anna Muranova. Effective impedance over ordered fields. Journal of Mathematical Physics, 62(3):033502, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langley:2021:GLH

[524] R. S. Langley. A generalization of the Lagrange–Hamilton formalism with application to non-conservative systems and the quantum to classical transition. Journal of Mathematical Physics, 62(3):033503, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Domanski:2021:DQC

[525] Ziemowit Domański. Deformation quantization on the cotangent bundle of a Lie group. Journal of Mathematical Physics, 62(3):033504, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:2021:MDQ

[526] Won Sang Chung and Abdullah Algin. Multi-dimensional q-deformed bosonic Newton oscillators and the related q-calculus, q-coherent states, and Hermite q-polynomials. Journal of Mathematical Physics, 62(3): 033505, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Esen:2021:HJF

[527] Ogul Esen, Manuel de León, Cristina Sardón, and Marcin Zajsc. Hamilton–Jacobi formalism on locally conformally symplectic manifolds. Journal of Mathematical Physics, 62(3):033506, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avetisyan:2021:GHF

[528] Z. Avetisyan. Global hyperbolicity and factorization in cosmological models. Journal of Mathematical Physics, 62(3):033507, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2021:PIF

[529] Mengkun Zhu, Dan Wang, and Yang Chen. Painlevé IV, σ-form, and the deformed Hermite unitary ensembles. Journal of Mathematical Physics, 62(3):033508, March 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2021:LBR

[530] Weifeng Jiang, Tong Li, Zhen Wang, and Shutian Fang. The limiting behavior of the Riemann solutions of non-isentropic modified Chaplygin gas dynamics. Journal of Mathematical Physics, 62(4):041501, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2021:CPG

[531] Long Wei and Yang Wang. The Cauchy problem for a generalized Riemann-type hydrodynamical equation. Journal of Mathematical Physics, 62(4):041502, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zheng:2021:DAR

[532] Pan Zheng and Robert Willie. Dynamics in an attraction-repulsion Navier–Stokes system with signal-dependent motility and sensitivity. Journal of Mathematical Physics, 62(4):041503, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ren:2021:GST

[533] Guoqiang Ren. Global solvability in a two-species chemotaxis system with logistic source. Journal of Mathematical Physics, 62(4):041504, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2021:WPB

[534] Fei Xu, Yong Zhang, and Fengquan Li. The well-posedness, blow-up, and traveling waves for a two-component Fornberg–Whitham system. Journal of Mathematical Physics, 62(4):041505, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juarez-Campos:2021:SSE

[535] Beatriz Juarez-Campos. Stochastic Schrödinger equation with Dirichlet noise boundary conditions. Journal of Mathematical Physics, 62(4): 041506, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:MSD

[536] Xiangqin Zhang. Monotonicity of solution to the dark monopole equations in non-Abelian gauge field theory. Journal of Mathematical Physics, 62(4):041507, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2021:SDE

[537] Ke Lin and Tian Xiang. Strong damping effect of chemo-repulsion prevents blow-up. Journal of Mathematical Physics, 62(4):041508, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2021:IMM

[538] Juncheng Wei and Yuanze Wu. Infinitely many multi-vortex solutions of the magnetic Ginzburg–Landau equation with external potentials in R2. Journal of Mathematical Physics, 62(4):041509, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2021:PSK

[539] Peng Chen and Xiaochun Liu. Positive solutions for Kirchhoff equation in exterior domains. Journal of Mathematical Physics, 62(4):041510, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xie:2021:AED

[540] Binqiang Xie and Lan Zeng. Almost exponential decay of the Boussinesq equations without surface tension on the upper free boundary. Journal of Mathematical Physics, 62(4):041511, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babenko:2021:TFF

[541] Constantin Babenko, Frank Göhmann, Karol K. Kozlowski, and Junji Suzuki. A thermal form factor series for the longitudinal two-point function of the Heisenberg–Ising chain in the antiferromagnetic massive regime. Journal of Mathematical Physics, 62(4):041901, April 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Filoche:2021:EPM

[542] Marcel Filoche, Svitlana Mayboroda, and Terence Tao. The effective potential of an M-matrix. Journal of Mathematical Physics, 62(4):041902, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bahns:2021:TBT

[543] Dorothea Bahns and Detlev Buchholz. Trapped bosons, thermodynamic limit, and condensation: a study in the framework of resolvent algebras. Journal of Mathematical Physics, 62(4):041903, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christiansen:2021:RPJ

[544] Jacob S. Christiansen, Barry Simon, and Maxim Zinchenko. Remarks on periodic Jacobi matrices on trees. Journal of Mathematical Physics, 62(4):042101, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lian:2021:HUP

[545] Pan Lian. Heisenberg’s uncertainty principle associated with the Caputo fractional derivative. Journal of Mathematical Physics, 62(4):042102, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loran:2021:DFL

[546] Farhang Loran and Ali Mostafazadeh. Dynamical formulation of lowenergy scattering in one dimension. Journal of Mathematical Physics, 62(4):042103, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohlsson:2021:TPF

[547] Tommy Ohlsson and Shun Zhou. Transition probabilities for flavor eigenstates of non-Hermitian Hamiltonians in the PT-broken phase. Journal of Mathematical Physics, 62(4):042104, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cruz-Prado:2021:NDQ

[548] H. Cruz-Prado, G. Marmo, D. Schuch, and O. Castaños. Nonlinear description of quantum dynamics: Generalized coherent states. Journal of Mathematical Physics, 62(4):042105, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chong:2021:DLB

[549] Jacky Chong. Dynamics of large boson systems with attractive interaction and a derivation of the cubic focusing NLS equation in R3. Journal of Mathematical Physics, 62(4):042106, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tehrani:2021:CIM

[550] Alireza Tehrani and Rajesh Pereira. The coherent information on the manifold of positive definite density matrices. Journal of Mathematical Physics, 62(4):042201, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:2021:CMI

[551] Yangping Jing, Chi-Kwong Li, Edward Poon, and Chengyang Zhang. Coherence measures induced by norm functions. Journal of Mathematical Physics, 62(4):042202, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caro:2021:NCM

[552] Matthias C. Caro and Benedikt R. Graswald. Necessary criteria for Markovian divisibility of linear maps. Journal of Mathematical Physics, 62(4):042203, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mograby:2021:HST

[553] Gamal Mograby, Maxim Derevyagin, Gerald V. Dunne, and Alexander Teplyaev. Hamiltonian systems, Toda lattices, solitons, Lax pairs on weighted Z-graded graphs. Journal of Mathematical Physics, 62(4): 042204, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loulidi:2021:CDQ

[554] Faedi Loulidi and Ion Nechita. The compatibility dimension of quantum measurements. Journal of Mathematical Physics, 62(4):042205, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiorenza:2021:TCI

[555] Domenico Fiorenza, Hisham Sati, and Urs Schreiber. Twisted cohomotopy implies twisted string structure on M5-branes. Journal of Mathematical Physics, 62(4):042301, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takaesu:2021:FOE

[556] Toshimitsu Takaesu. The first order expansion of a ground state energy of the φ^{4} model with cutoffs. Journal of Mathematical Physics, 62(4): 042302, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bettadapura:2021:BCT

[557] Kowshik Bettadapura and Hai Lin. Boundary contributions to three loop superstring amplitudes. Journal of Mathematical Physics, 62(4): 042303, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamilton:2021:LSW

[558] Mark J. D. Hamilton. A lift of the Seiberg–Witten equations to Kaluza– Klein five-manifolds. Journal of Mathematical Physics, 62(4):042304, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giannakis:2021:QDC

[559] Dimitrios Giannakis. Quantum dynamics of the classical harmonic oscillator. Journal of Mathematical Physics, 62(4):042701, April 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Koch:2021:GMR

[560] Hans Koch. Golden mean renormalization for the almost Mathieu operator and related skew products. Journal of Mathematical Physics, 62 (4):042702, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2021:PCE

[561] Alexander Figotin. Perturbations of circuit evolution matrices with Jordan blocks. Journal of Mathematical Physics, 62(4):042703, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2021:SCM

[562] Manuel de León, Manuel Lainz, and Asier López-Gordón. Symmetries, constants of the motion, and reduction of mechanical systems with external forces. Journal of Mathematical Physics, 62(4):042901, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baez:2021:OSC

[563] John C. Baez, David Weisbart, and Adam M. Yassine. Open systems in classical mechanics. Journal of Mathematical Physics, 62(4):042902, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2021:SLN

[564] Yunshun Wu and Yong Wang. Stability on large non-constant steady states of semiconductor equations. Journal of Mathematical Physics, 62 (4):043101, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cosco:2021:TIM

[565] Clément Cosco and Assaf Shapira. Topologically induced metastability in a periodic XY chain. Journal of Mathematical Physics, 62(4):043301, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:SNR

[566] Wei Zhang. Some new results on relative entropy production, time reversal, and optimal control of time-inhomogeneous diffusion processes. Journal of Mathematical Physics, 62(4):043302, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Erbin:2021:CEV

[567] Harold Erbin, Vincent Lahoche, and Mohamed Tamaazousti. Constructive expansion for vector field theories I. Quartic models in low dimensions. Journal of Mathematical Physics, 62(4):043501, April 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amakawa:2021:CLW

[568] K. Amakawa and N. Aizawa. A classification of lowest weight irreducible modules over Z_{2}^{2}-graded extension of osp(1|2). Journal of Mathematical Physics, 62(4):043502, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boumaza:2021:IDS

[569] H. Boumaza and O. Lafitte. Integrated density of states: From the finite range to the periodic Airy–Schrödinger operator. Journal of Mathematical Physics, 62(4):043503, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Accornero:2021:STE

[570] Luca Accornero and Marcella Palese. Symmetry transformations of extremals and higher conserved quantities: Invariant Yang–Mills connections. Journal of Mathematical Physics, 62(4):043504, April 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Villanueva:2021:FPI

[571] Lloyd L. Villanueva and Eric A. Galapon. Finite-part integration in the presence of competing singularities: Transformation equations for the hypergeometric functions arising from finite-part integration. Journal of Mathematical Physics, 62(4):043505, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maurya:2021:FLE

[572] Shambhu Nath Maurya and Koushik Saha. Fluctuations of linear eigenvalue statistics of reverse circulant and symmetric circulant matrices with independent entries. Journal of Mathematical Physics, 62(4):043506, April 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shuai:2021:EMS

[573] Wei Shuai. Existence and multiplicity of solutions for logarithmic Schrödinger equations with potential. Journal of Mathematical Physics, 62(5):051501, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huh:2021:BCS

[574] Hyungjin Huh and Jihyun Yim. L^{∞} bounds for Chern–Simons gauged equations in R^{1+1} and their applications. Journal of Mathematical Physics, 62(5):051502, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2021:BCF

[575] Jie Fan, Quansen Jiu, Yanqing Wang, and Yuelong Xiao. Blow up criterion for the 2D full compressible Navier–Stokes equations involving temperature in critical spaces. Journal of Mathematical Physics, 62(5): 051503, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2021:GEL

[576] Zaihong Jiang, Qingning Zhang, and Mingxuan Zhu. Global existence and long time behavior of the generalized Camassa–Holm equation with k + 1 degree nonlinearities. Journal of Mathematical Physics, 62(5): 051504, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2021:LWP

[577] Jishan Fan and Yong Zhou. Local well-posedness for the isentropic compressible MHD system with vacuum. Journal of Mathematical Physics, 62(5):051505, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:GSN

[578] Yang Wang, Xinyue Cao, Zhaohai Ma, and Xiong Li. Global stability of noncritical traveling front solutions of Fisher-type equations with degenerate nonlinearity. Journal of Mathematical Physics, 62(5):051506, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boer:2021:EMS

[579] Eduardo de S. Böer and Olímpio H. Miyagaki. Existence and multiplicity of solutions for the fractional p-Laplacian Choquard logarithmic equation involving a nonlinearity with exponential critical and subcritical growth. Journal of Mathematical Physics, 62(5):051507, May 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [653].

Paranamana:2021:GMF

[580] Pushpi Paranamana, Eugenio Aulisa, and Magdalena Toda. Geometric model of the fracture as a manifold immersed in porous media. Journal of Mathematical Physics, 62(5):051508, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarrico:2021:MPA

[581] C. O. R. Sarrico. The movement of a particle according to the Gurevich– Zybin dark matter model. Journal of Mathematical Physics, 62(5): 051509, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:DIM

[582] Shuang Yang and Yangrong Li. Dynamics and invariant measures of multi-stochastic sine-Gordon lattices with random viscosity and nonlinear noise. Journal of Mathematical Physics, 62(5):051510, May 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hezari:2021:DIP

[583] Hamid Hezari, Z. Lu, and J. Rowlett. The Dirichlet isospectral problem for trapezoids. Journal of Mathematical Physics, 62(5):051511, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:2021:GPS

[584] Aristophanes Dimakis and Igor G. Korepanov. Grassmannianparameterized solutions to direct-sum polygon and simplex equations. Journal of Mathematical Physics, 62(5):051701, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kojima:2021:QRD

[585] Takeo Kojima. Quadratic relations of the deformed W -superalgebra W_{⨿,⊔}(∫↕(∈|∞)). Journal of Mathematical Physics, 62(5):051702, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dali:2021:KLB

[586] Béchir Dali and Rafik Khalfi. On Kirillov’s lemma and the Bensonratcliff invariant. Journal of Mathematical Physics, 62(5):051703, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hainzl:2021:APB

[587] Christian Hainzl. Another proof of BEC in the GP-limit. Journal of Mathematical Physics, 62(5):051901, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lopez:2021:SLD

[588] P. C. López, R. Santos-Silva, and A. García. Solutions to linear dissipative quantum systems. Journal of Mathematical Physics, 62(5):052101, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sadurni:2021:HDA

[589] Emerson Sadurní, Francois Leyvraz, Thomas Stegmann, Thomas H. Seligman, and Douglas J. Klein. Hidden duality and accidental degeneracy in cycloacene and Möbius cycloacene. Journal of Mathematical Physics, 62(5):052102, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Znojil:2021:EPD

[590] Miloslav Znojil. Exceptional points and domains of unitarity for a class of strongly non-Hermitian real-matrix Hamiltonians. Journal of Mathematical Physics, 62(5):052103, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2021:CSS

[591] Hai Lin and Yuwei Zhu. Coherent state superpositions, entanglement, and gauge/gravity correspondence. Journal of Mathematical Physics, 62(5):052301, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grewcoe:2021:DFT

[592] Clay James Grewcoe and Larisa Jonke. Double field theory algebroid and curved L_{∞}-algebras. Journal of Mathematical Physics, 62(5):052302, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Acik:2021:GSS

[593] Özgür Açik and Ümit Ertem. Generalized symmetry superalgebras. Journal of Mathematical Physics, 62(5):052303, May 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:YMM

[594] Xiangqin Zhang and Yisong Yang. Yang–Mills monopoles in extremal Reissner-Nordström black hole metric. Journal of Mathematical Physics, 62(5):052304, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McNutt:2021:TGC

[595] D. D. McNutt, A. A. Coley, and R. J. van den Hoogen. Teleparallel geometries not characterized by their scalar polynomial torsion invariants. Journal of Mathematical Physics, 62(5):052501, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2021:EDL

[596] Seung-Yeal Ha, Myeongju Kang, and Hansol Park. Emergent dynamics of the Lohe Hermitian sphere model with frustration. Journal of Mathematical Physics, 62(5):052701, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Budhiraja:2021:LDE

[597] Amarjit Budhiraja, Yong Chen, and Lihu Xu. Large deviations of the entropy production rate for a class of Gaussian processes. Journal of Mathematical Physics, 62(5):052702, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2021:ZHB

[598] Jaume Llibre and Yuzhou Tian. The zero-Hopf bifurcations of a fourdimensional hyperchaotic system. Journal of Mathematical Physics, 62 (5):052703, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Massa:2021:SCL

[599] Enrico Massa and Enrico Pagani. Symmetry and conservation laws in non-holonomic mechanics. Journal of Mathematical Physics, 62(5): 052901, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freistuhler:2021:DSM

[600] Heinrich Freistühler, Moritz Reintjes, and Blake Temple. Decay and subluminality of modes of all wave numbers in the relativistic dynamics of viscous and heat conductive fluids. Journal of Mathematical Physics, 62(5):053101, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cavallaro:2021:TEV

[601] Guido Cavallaro and Carlo Marchioro. Time evolution of vortex rings with large radius and very concentrated vorticity. Journal of Mathematical Physics, 62(5):053102, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giorgadze:2021:TEI

[602] G. Giorgadze and G. Khimshiashvili. Triangles and electrostatic ion traps. Journal of Mathematical Physics, 62(5):053501, May 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molinari:2021:SNB

[603] Luca Guido Molinari. Some Neumann–Bessel series and the Laplacian on polygons. Journal of Mathematical Physics, 62(5):053502, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boldyreva:2021:ESK

[604] Maria N. Boldyreva and Alexey A. Magazev. Exact solutions of Klein– Gordon equations in external electromagnetic fields on 3D de Sitter background. Journal of Mathematical Physics, 62(5):053503, May 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karpeshina:2021:BTS

[605] Yulia Karpeshina, Leonid Parnovski, and Roman Shterenberg. Ballistic transport for Schrödinger operators with quasi-periodic potentials. Journal of Mathematical Physics, 62(5):053504, May 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fox:2021:RMH

[606] Matthew S. Fox. Retraction: “Multipole hair of Schwarzschild– Tangherlini black holes” [J. Math. Phys. 60, 102502 (2019)]. Journal of Mathematical Physics, 62(5):059901, May 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [14].

Reuvers:2021:PNG

[607] Robin Reuvers. Publisher’s note: “Generalized Pauli constraints in large systems: The Pauli principle dominates” [J. Math. Phys. 62, 032204 (2021)]. Journal of Mathematical Physics, 62(5):059902, May 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [506].

Huang:2021:CVH

[608] Yongting Huang and Tao Luo. Compressible viscous heat-conducting surface wave without surface tension. Journal of Mathematical Physics, 62(6):061501, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prokofev:2021:ESM

[609] V. Prokofev and A. Zabrodin. Elliptic solutions to matrix KP hierarchy and spin generalization of elliptic Calogero–Moser model. Journal of Mathematical Physics, 62(6):061502, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2021:EMR

[610] Ya-Hong Guo, Hong-Rui Sun, and Na Cui. Existence and multiplicity results for the fractional magnetic Schrödinger equations with critical growth. Journal of Mathematical Physics, 62(6):061503, June 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itakura:2021:SST

[611] Kyohei Itakura. Stationary scattering theory for repulsive Hamiltonians. Journal of Mathematical Physics, 62(6):061504, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:EDB

[612] Donghao Li, Hongwei Zhang, and Qingying Hu. Energy decay and blow-up of solutions for a viscoelastic equation with nonlocal nonlinear boundary dissipation. Journal of Mathematical Physics, 62(6):061505, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nzissila:2021:BSM

[613] Florent Nzissila, Octave Moutsinga, and Fulgence Eyi Obiang. Backward semi-martingales into Burgers turbulence. Journal of Mathematical Physics, 62(6):061506, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2021:STD

[614] Xiaopeng Zhao. Space-time decay of solutions to three-dimensional MHD equations with Hall and ion-slip effects. Journal of Mathematical Physics, 62(6):061507, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2021:GEE

[615] Lingxi Liu and Xin Zhong. Global existence and exponential decay of strong solutions for 2D nonhomogeneous micropolar fluids with densitydependent viscosity. Journal of Mathematical Physics, 62(6):061508, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schlag:2021:PDW

[616] W. Schlag. On pointwise decay of waves. Journal of Mathematical Physics, 62(6):061509, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Logunov:2021:EAI

[617] A. Logunov and H. Papazov. An elliptic adaptation of ideas of Carleman and Domar from complex analysis related to Levinson’s loglog theorem. Journal of Mathematical Physics, 62(6):061510, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:AYS

[618] Na Wang and Linjie Shi. Affine Yangian and Schur functions on plane partitions of 4. Journal of Mathematical Physics, 62(6):061701, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bardet:2021:MLS

[619] Ivan Bardet, Ángela Capel, Angelo Lucia, David Pérez-García, and Cambyse Rouzé. On the modified logarithmic Sobolev inequality for the heat-bath dynamics for 1D systems. Journal of Mathematical Physics, 62(6):061901, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kukulski:2021:GRQ

[620] Ryszard Kukulski, Ion Nechita, Łukasz Pawela, Zbigniew Puchała, and Karol Życzkowski. Generating random quantum channels. Journal of Mathematical Physics, 62(6):062201, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:2021:QAC

[621] Masuo Suzuki. Quantum analysis on the convergence speed of exponential product formulas — differential-subtraction and exchangeintegration method on concise norm bounds. Journal of Mathematical Physics, 62(6):062202, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamagata:2021:MLD

[622] K. Yamagata. Maximum logarithmic derivative bound on quantum state estimation as a dual of the Holevo bound. Journal of Mathematical Physics, 62(6):062203, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pereg:2021:QBC

[623] Uzi Pereg, Christian Deppe, and Holger Boche. Quantum broadcast channels with cooperating decoders: an information-theoretic perspective on quantum repeaters. Journal of Mathematical Physics, 62(6): 062204, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Debbasch:2021:QWS

[624] F. Debbasch. Quantum walks simulating non-commutative geometry in the Landau problem. Journal of Mathematical Physics, 62(6):062205, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tlas:2021:CEI

[625] T. Tlas. A construction of Euclidean invariant, reflection positive measures on a compactification of distributions. Journal of Mathematical Physics, 62(6):062301, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frob:2021:SDQ

[626] M. B. Fröb and A. Much. Strict deformations of quantum field theory in de Sitter spacetime. Journal of Mathematical Physics, 62(6):062302, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaninsky:2021:DLC

[627] J. Káninský. Discrete linear canonical evolution. Journal of Mathematical Physics, 62(6):062303, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ripley:2021:SHF

[628] Justin L. Ripley. A symmetric hyperbolic formulation of the vacuum Einstein equations in affine-null coordinates. Journal of Mathematical Physics, 62(6):062501, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:2021:RPT

[629] J. Herrera, M. de la Rosa, and R. M. Rubio. Relativistic particles with torsion in three-dimensional non-vacuum spacetimes. Journal of Mathematical Physics, 62(6):062502, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2021:BLP

[630] Xiaokai He, Xiaoning Wu, and Naqing Xie. From bending of light to positive mass: a non-PDE perspective. Journal of Mathematical Physics, 62(6):062503, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:NWE

[631] Jinhua Wang. Nonlinear wave equation in a cosmological Kaluza Klein spacetime. Journal of Mathematical Physics, 62(6):062504, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:WZA

[632] Yuan Yang, Ji Shu, and Xiaohu Wang. Wong–Zakai approximations and random attractors of non-autonomous stochastic discrete complex Ginzburg–Landau equations. Journal of Mathematical Physics, 62(6): 062701, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nwaigwe:2021:CWA

[633] Dwight Nwaigwe. On the convergence of WKB approximations of the damped Mathieu equation. Journal of Mathematical Physics, 62(6): 062702, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiba:2021:NFC

[634] Hayato Chiba. Normal forms of C^{∞} vector fields based on the renormalization group. Journal of Mathematical Physics, 62(6):062703, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fabregas:2021:RRD

[635] A. Ruiz de Zarate Fabregas, N. L. Dias, and D. G. Alfaro Vigo. Realizability of the rapid distortion theory spectrum: the mechanism behind the Kelvin–Townsend equations. Journal of Mathematical Physics, 62 (6):063101, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:ATD

[636] Kui Zhang and Gustavo Didier. Asymptotic theory for the detection of mixing in anomalous diffusion. Journal of Mathematical Physics, 62(6): 063301, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyu:2021:ALE

[637] Shulin Lyu, Chao Min, and Yang Chen. Asymptotics of the largest eigenvalue distribution of the Laguerre unitary ensemble. Journal of Mathematical Physics, 62(6):063302, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caceres:2021:ERP

[638] Manuel O. Cáceres. Exact results on Poisson noise, Poisson flights, and Poisson fluctuations. Journal of Mathematical Physics, 62(6):063303, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tan:2021:CAE

[639] Chengliang Tan and Jilong Zhang. Calculation of algebraic entropies of d-P_{IV} and d-P_{V}. Journal of Mathematical Physics, 62(6):063501, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berkolaiko:2021:CMT

[640] Gregory Berkolaiko and Laura Booton. Convergence of moments of twisted COE matrices. Journal of Mathematical Physics, 62(6):063502, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saidani:2021:CVC

[641] Samira Saidani and Sid-Ahmed Yahiaoui. Complementarity vs coordinate transformations: Mapping between pseudo-Hermiticity and weak pseudo-Hermiticity. Journal of Mathematical Physics, 62(6):063503, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedmann:2021:ERF

[642] Tamar Friedmann and Quincy Webb. Euler’s reflection formula, infinite product formulas, and the correspondence principle of quantum mechanics. Journal of Mathematical Physics, 62(6):063504, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bourgine:2021:QWS

[643] Jean-Emile Bourgine. Quantum W_{1+∞} subalgebras of BCD type and symmetric polynomials. Journal of Mathematical Physics, 62(6):063505, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eder:2021:SFB

[644] Konstantin Eder. Super fiber bundles, connection forms, and parallel transport. Journal of Mathematical Physics, 62(6):063506, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlowski:2021:SDR

[645] Karol K. Kozlowski. On singularities of dynamic response functions in the massless regime of the XXZ spin-1/2 chain. Journal of Mathematical Physics, 62(6):063507, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2021:CVC

[646] T. Skrypnyk. On a class of gl(n) ⊗ gl(n)-valued classical r-matrices and separation of variables. Journal of Mathematical Physics, 62(6):063508, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kachkovskiy:2021:PDM

[647] Ilya Kachkovskiy, Stanislav Krymski, Leonid Parnovski, and Roman Shterenberg. Perturbative diagonalization for Maryland-type quasiperiodic operators with flat pieces. Journal of Mathematical Physics, 62(6): 063509, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mouayn:2021:CSS

[648] Zouhaïr Mouayn and Hashim A. Yamani. Coherent states of systems with pure continuous energy spectra. Journal of Mathematical Physics, 62(6):063510, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mejjaoli:2021:UPA

[649] Hatem Mejjaoli and Firdous A. Shah. Uncertainty principles associated with the directional short-time Fourier transform. Journal of Mathematical Physics, 62(6):063511, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuznetsova:2021:CMZ

[650] Zhanna Kuznetsova and Francesco Toppan. Classification of minimal Z_{2} × Z_{2}-graded Lie (super)algebras and some applications. Journal of Mathematical Physics, 62(6):063512, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2021:CBS

[651] Francisco M. Fernández. Comment on “Bound states and the potential parameter spectrum” [J. Math. Phys. 61, 062103 (2020)]. Journal of Mathematical Physics, 62(6):064101, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [175] and response [652].

Alhaidari:2021:RCB

[652] A. D. Alhaidari and H. Bahlouli. Response to “Comment on ‘Bound states and the potential parameter spectrum”’ [J. Math. Phys. 62, 064101 (2021)]. Journal of Mathematical Physics, 62(6):064102, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [175, 651].

Boer:2021:PNE

[653] Eduardo de S. Böer and Olímpio H. Miyagaki. Publisher’s note: “Existence and multiplicity of solutions for the fractional p-Laplacian Choquard logarithmic equation involving a nonlinearity with exponential critical and subcritical growth” [J. Math. Phys. 62, 051507 (2021)]. Journal of Mathematical Physics, 62(6):069901, June 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [579].

Liu:2021:CPF

[654] Xuan Liu and Ting Zhang. The Cauchy problem for the fourth-order Schrödinger equation in H^{s}. Journal of Mathematical Physics, 62(7): 071501, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayashi:2021:MSH

[655] Nakao Hayashi and Pavel I. Naumkin. Modified scattering for the higherorder anisotropic nonlinear Schrödinger equation in two space dimensions. Journal of Mathematical Physics, 62(7):071502, July 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2021:LTS

[656] Xinghong Pan. Liouville theorem of D-solutions to the stationary magnetohydrodynamics system in a slab. Journal of Mathematical Physics, 62(7):071503, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:2021:NGS

[657] Yongtao Jing and Haidong Liu. Nonexistence of ground state solutions for generalized quasilinear Schrödinger equations via dual approach. Journal of Mathematical Physics, 62(7):071504, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saoudi:2021:SEP

[658] Kamel Saoudi, Akasmika Panda, and Debajyoti Choudhuri. A singular elliptic problem involving fractional p-Laplacian and a discontinuous critical nonlinearity. Journal of Mathematical Physics, 62(7):071505, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2021:DMN

[659] Mi-Ran Choi, Younghoon Kang, and Young-Ran Lee. On dispersion managed nonlinear Schrödinger equations with lumped amplification. Journal of Mathematical Physics, 62(7):071506, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dodson:2021:GWP

[660] Benjamin Dodson, Avraham Soffer, and Thomas Spencer. Global wellposedness for the cubic nonlinear Schrödinger equation with initial data lying in L^{p}-based Sobolev spaces. Journal of Mathematical Physics, 62 (7):071507, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choreno:2021:AAB

[661] E. Choreño, R. Valencia, and D. Ojeda-Guillén. Algebraic approach and Berry phase of a Hamiltonian with a general SU ( 1, 1 ) symmetry. Journal of Mathematical Physics, 62(7):071701, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tao:2021:RCE

[662] W.-Q. Tao. On representations of the centrally extended Heisenberg double of SL_{2}. Journal of Mathematical Physics, 62(7):071702, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haldane:2021:GIP

[663] F. D. M. Haldane. Gauge-invariant perturbation expansion in powers of electric charge for the density-of-states of a network model for chargedparticle motion in a uniform background magnetic flux density. Journal of Mathematical Physics, 62(7):071901, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oldenburg:2021:GSB

[664] Jakob Oldenburg. On ground states of the Bogoliubov energy functional: a direct proof. Journal of Mathematical Physics, 62(7):071902, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagchi:2021:DHS

[665] Bijan Bagchi and Rahul Ghosh. Dirac Hamiltonian in a supersymmetric framework. Journal of Mathematical Physics, 62(7):072101, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rushka:2021:CTO

[666] Michael Rushka, M. A. Esrick, W. N. Mathews, Jr., and J. K. Freericks. Converting translation operators into plane polar and spherical coordinates and their use in determining quantum-mechanical wavefunctions in a representation-independent fashion. Journal of Mathematical Physics, 62(7):072102, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Escobar-Ruiz:2021:FBH

[667] Adrian M. Escobar-Ruiz, Alexander V. Turbiner, and Willard Miller, Jr. Four-body (an)harmonic oscillator in d-dimensional space: S-states, (quasi)-exact-solvability, hidden algebra sl(7). Journal of Mathematical Physics, 62(7):072103, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gazeau:2021:GSG

[668] J.-P. Gazeau, V. Hussin, J. Moran, and K. Zelaya. Generalized Susskind– Glogower coherent states. Journal of Mathematical Physics, 62(7): 072104, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shahbeigi:2021:LCS

[669] Fereshte Shahbeigi, David Amaro-Alcalá, Zbigniew Puchała, and Karol Zyczkowski. Log-convex set of Lindblad semigroups acting on N-level system. Journal of Mathematical Physics, 62(7):072105, July 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duerinckx:2021:NSA

[670] Mitia Duerinckx and Christopher Shirley. A new spectral analysis of stationary random Schrödinger operators. Journal of Mathematical Physics, 62(7):072106, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sanchez-Mendoza:2021:SBI

[671] Daniel Sánchez-Mendoza. Sharp bounds for the integrated density of states of a strongly disordered 1D Anderson–Bernoulli model. Journal of Mathematical Physics, 62(7):072107, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weinzierl:2021:CIN

[672] Stefan Weinzierl. On the computation of intersection numbers for twisted cocycles. Journal of Mathematical Physics, 62(7):072301, July 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernardo:2021:SBH

[673] Reginald Christian Bernardo and Ian Vega. Stealth black hole perturbations in kinetic gravity braiding. Journal of Mathematical Physics, 62(7):072501, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Trevino:2021:TT

[674] Rodrigo Treviño. Tilings and traces. Journal of Mathematical Physics, 62(7):072701, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cresson:2021:SMN

[675] Jacky Cresson, Laurent Nottale, and Thierry Lehner. Stochastic modification of Newtonian dynamics and induced potential-application to spiral galaxies and the dark potential. Journal of Mathematical Physics, 62(7):072702, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ban:2021:VMP

[676] Jung-Chao Ban, Wen-Guei Hu, Song-Sun Lin, and Yin-Heng Lin. Verification of mixing properties in two-dimensional shifts of finite type. Journal of Mathematical Physics, 62(7):072703, July 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2021:PAM

[677] Xiaoya Song and Yangmin Xiong. Pullback attractors for 2D MHD equations with delays. Journal of Mathematical Physics, 62(7):072704, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoang:2021:RRH

[678] Vu Hoang, Maria Radosz, Angel Harb, Aaron DeLeon, and Alan Baza. Radiation reaction in higher-order electrodynamics. Journal of Mathematical Physics, 62(7):072901, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2021:SCF

[679] Emine Celik, Luan Hoang, and Thinh Kieu. Slightly compressible Forchheimer flows in rotating porous media. Journal of Mathematical Physics, 62(7):073101, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holm:2021:SEW

[680] Darryl D. Holm and Ruiao Hu. Stochastic effects of waves on currents in the ocean mixed layer. Journal of Mathematical Physics, 62(7):073102, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeMasi:2021:RFL

[681] A. De Masi, I. Merola, and E. Presutti. Reservoirs, Fick law, and the Darken effect. Journal of Mathematical Physics, 62(7):073301, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2021:RBM

[682] Peter J. Forrester and Anas A. Rahman. Relations between moments for the Jacobi and Cauchy random matrix ensembles. Journal of Mathematical Physics, 62(7):073302, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ceniceros:2021:ECI

[683] Jose Ceniceros, Mohamed Elhamdadi, and Alireza Mashaghi. Enhancement of the coloring invariant for folded molecular chains. Journal of Mathematical Physics, 62(7):073501, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ortenzi:2021:BHB

[684] Giovanni Ortenzi and Marco Pedroni. Boussinesq hierarchy and bi-Hamiltonian geometry. Journal of Mathematical Physics, 62(7):073502, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gubbiotti:2021:SSN

[685] G. Gubbiotti and M. C. Nucci. Superintegrable systems in non-Euclidean plane: Hidden symmetries leading to linearity. Journal of Mathematical Physics, 62(7):073503, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2021:LCD

[686] Salih Celik. Left-covariant differential calculi on GL_{q}(2). Journal of Mathematical Physics, 62(7):073504, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2021:CEP

[687] P. J. Forrester and G. Mazzuca. The classical β ensembles with β proportional to 1/N: From loop equations to Dyson’s disordered chain. Journal of Mathematical Physics, 62(7):073505, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Jitomirskaya:2021:UBT

[688] Svetlana Jitomirskaya and Wencai Liu. Upper bounds on transport exponents for long-range operators. Journal of Mathematical Physics, 62 (7):073506, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2021:WZA

[689] Wenqiang Zhao. Wong–Zakai approximations of the non-autonomous stochastic FitzHugh–Nagumo system on RN in higher regular spaces. Journal of Mathematical Physics, 62(8):081501, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tin:2021:DNS

[690] Phan Van Tin. On the derivative nonlinear Schrödinger equation on the half line with Robin boundary condition. Journal of Mathematical Physics, 62(8):081502, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hafner:2021:STD

[691] Dietrich Häfner, Mokdad Mokdad, and Jean-Philippe Nicolas. Scattering theory for Dirac fields inside a Reissner-Nordström-type black hole. Journal of Mathematical Physics, 62(8):081503, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Esfahani:2021:DSF

[692] Amin Esfahani. Dynamics of solutions of a fractional NLS system with quadratic interaction. Journal of Mathematical Physics, 62(8):081504, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2021:GSL

[693] Ying-Chieh Lin, Kuan-Hsiang Wang, and Tsung fang Wu. Ground states for a linearly coupled indefinite Schrödinger system with steep potential well. Journal of Mathematical Physics, 62(8):081505, August 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radobolja:2021:GA

[694] G. Radobolja. Galilean W_{3} algebra. Journal of Mathematical Physics, 62(8):081701, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagheri:2021:QSS

[695] Zahra Bagheri and Esmaeil Peyghan. Quadratic and symplectic structures on 3-(Hom)-ρ-Lie algebras. Journal of Mathematical Physics, 62 (8):081702, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapustin:2021:CIP

[696] Anton Kapustin, Nikita Sopenko, and Bowen Yang. A classification of invertible phases of bosonic quantum lattice systems in one dimension. Journal of Mathematical Physics, 62(8):081901, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pierobon:2021:RHL

[697] Gianfranco Pierobon, Gianfranco Cariolaro, and Giuseppe Dattoli. On the role of Hermite-like polynomials in the Fock representations of Gaussian states. Journal of Mathematical Physics, 62(8):082101, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schlue:2021:OFS

[698] Volker Schlue. Optical functions in de Sitter. Journal of Mathematical Physics, 62(8):082501, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2021:DSS

[699] Jaume Llibre and Claudia Valls. Dynamics of the Szekeres system. Journal of Mathematical Physics, 62(8):082502, August 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penna:2021:ERW

[700] Robert F. Penna. Einstein–Rosen waves and the Geroch group. Journal of Mathematical Physics, 62(8):082503, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2021:EPD

[701] Alexander Figotin. Exceptional points of degeneracy in traveling wave tubes. Journal of Mathematical Physics, 62(8):082701, August 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahn:2021:EBC

[702] Hyunjin Ahn, Seung-Yeal Ha, Hansol Park, and Woojoo Shim. Emergent behaviors of Cucker–Smale flocks on the hyperboloid. Journal of Mathematical Physics, 62(8):082702, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sousa:2021:RFD

[703] J. Vanterler da C. Sousa, M. Vellappandi, V. Govindaraj, and Gastão S. F. Frederico. Reachability of fractional dynamical systems using ψ-Hilfer pseudo-fractional derivative. Journal of Mathematical Physics, 62 (8):082703, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

You:2021:RCN

[704] Zhongli You, Michal Feckan, and JinRong Wang. On the relative controllability of neutral delay differential equations. Journal of Mathematical Physics, 62(8):082704, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2021:TVS

[705] Jian-Zhou Zhu. Thermodynamic and vortic structures of real Schur flows. Journal of Mathematical Physics, 62(8):083101, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Djenidi:2021:MTO

[706] L. Djenidi and R. A. Antonia. Modeling the third-order velocity structure function in the scaling range at finite Reynolds numbers. Journal of Mathematical Physics, 62(8):083102, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nguyen:2021:LDP

[707] Nhu N. Nguyen and George Yin. Large deviation principles for Langevin equations in random environment and applications. Journal of Mathematical Physics, 62(8):083301, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonnemain:2021:LCC

[708] Thibault Bonnemain, Thierry Gobron, and Denis Ullmo. Lax connection and conserved quantities of quadratic mean field games. Journal of Mathematical Physics, 62(8):083302, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andraus:2021:LTS

[709] Sergio Andraus, Kilian Hermann, and Michael Voit. Limit theorems and soft edge of freezing random matrix models via dual orthogonal polynomials. Journal of Mathematical Physics, 62(8):083303, August 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenzi:2021:STF

[710] E. K. Lenzi and L. R. Evangelista. Space-time fractional diffusion equations in d-dimensions. Journal of Mathematical Physics, 62(8):083304, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lauritsen:2021:FWC

[711] Asbjørn Bækgaard Lauritsen. Floating Wigner crystal and periodic jellium configurations. Journal of Mathematical Physics, 62(8):083305, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jain:2021:CLR

[712] Vishesh Jain, Indrajit Jana, Kyle Luh, and Sean O’Rourke. Circular law for random block band matrices with genuinely sublinear bandwidth. Journal of Mathematical Physics, 62(8):083306, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernard:2021:HOL

[713] Pierre-Antoine Bernard, Nicolas Crampé, Dounia Shaaban Kabakibo, and Luc Vinet. Heun operator of Lie type and the modified algebraic Bethe ansatz. Journal of Mathematical Physics, 62(8):083501, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magee:2021:GTS

[714] Adam M. Magee and Ludwik Dabrowski. Gauge transformations of spectral triples with twisted real structures. Journal of Mathematical Physics, 62(8):083502, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isaev:2021:SCO

[715] A. P. Isaev and S. O. Krivonos. Split Casimir operator for simple Lie algebras, solutions of Yang–Baxter equations, and Vogel parameters. Journal of Mathematical Physics, 62(8):083503, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2021:CDT

[716] Axel Schulze-Halberg. Characterization of Darboux transformations for quantum systems with quadratically energy-dependent potentials. Journal of Mathematical Physics, 62(8):083504, August 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stepanov:2021:GFS

[717] Yuriy Stepanov, Hendrik Herrmann, and Thomas Guhr. Generic features in the spectral decomposition of correlation matrices. Journal of Mathematical Physics, 62(8):083505, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:CHF

[718] Yi Yang, Lumin Geng, and Jipeng Cheng. CKP hierarchy and free bosons. Journal of Mathematical Physics, 62(8):083506, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braga:2021:ANI

[719] Gastão A. Braga, Jussara M. Moreira, and Camila F. Souza. Asymptotics for nonlinear integral equations with a generalized heat kernel using renormalization group technique II: Marginal perturbations and logarithmic corrections to the time decay of solutions. Journal of Mathematical Physics, 62(8):083507, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorodnitsyn:2021:DSW

[720] V. A. Dorodnitsyn and E. I. Kaptsov. Discrete shallow water equations preserving symmetries and conservation laws. Journal of Mathematical Physics, 62(8):083508, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikitin:2021:SSP

[721] A. G. Nikitin. Symmetries of the Schrödinger–Pauli equation for neutral particles. Journal of Mathematical Physics, 62(8):083509, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berkowitz:2021:AES

[722] Daniel Berkowitz. Applying the Euclidean-signature semi-classical method to the quantum Taub models with a cosmological constant and aligned electromagnetic field. Journal of Mathematical Physics, 62(8): 083510, August 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueiredo:2021:ELE

[723] Giovany Figueiredo, Sandra Moreira Neto, and Ricardo Ruviaro. Existence of least energy positive and nodal solutions for a quasilinear Schrödinger problem with potentials vanishing at infinity. Journal of Mathematical Physics, 62(9):091501, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Watson:2021:EFM

[724] Alexander B. Watson and Mitchell Luskin. Existence of the first magic angle for the chiral model of bilayer graphene. Journal of Mathematical Physics, 62(9):091502, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:BPP

[725] Shaojie Yang, Xuanxuan Han, and Tingting Wang. Blow-up phenomena and peakons for the gFQXL/gCH–mCH equation. Journal of Mathematical Physics, 62(9):091503, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujiwara:2021:GES

[726] Kazumasa Fujiwara and Vladimir Georgiev. On global existence of L^{2} solutions for 1D periodic NLS with quadratic nonlinearity. Journal of Mathematical Physics, 62(9):091504, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2021:TPL

[727] Chen Jia, Zhimin Zhang, and Lewei Zhao. Two-parameter localization for eigenfunctions of a Schrödinger operator in balls and spherical shells. Journal of Mathematical Physics, 62(9):091505, September 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luo:2021:GCS

[728] Zhaonan Luo, Zhijun Qiao, and Zhaoyang Yin. Globally conservative solutions for the modified Camassa–Holm (MOCH) equation. Journal of Mathematical Physics, 62(9):091506, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:SCS

[729] Qi Zhang. Sign-changing solutions for a kind of Klein–Gordon–Maxwell system. Journal of Mathematical Physics, 62(9):091507, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meng:2021:LES

[730] Yuxi Meng, Xingrui Zhang, and Xiaoming He. Least energy signchanging solutions for a class of fractional Kirchhoff–Poisson system. Journal of Mathematical Physics, 62(9):091508, September 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Wang:2021:RWS

[731] Yanqing Wang, Baoquan Yuan, Jiefeng Zhao, and Daoguo Zhou. On the regularity of weak solutions of the MHD equations in BMO^{−1} and Ḃ_{∞,∞}^{−1}. Journal of Mathematical Physics, 62(9):091509, September 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ren:2021:LTB

[732] Guoqiang Ren and Bin Liu. Large time behavior of solutions to a quasilinear attraction-repulsion chemotaxis model with nonlinear secretion. Journal of Mathematical Physics, 62(9):091510, September 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2021:LNS

[733] Rui He and Xiangqing Liu. Localized nodal solutions for semiclassical Choquard equations. Journal of Mathematical Physics, 62(9):091511, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2021:NRL

[734] Min Ding and Lang He. Non-relativistic limits and stability of composite wave patterns to the relativistic Euler equations. Journal of Mathematical Physics, 62(9):091512, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2021:DSP

[735] Derchyi Wu. The direct scattering problem for perturbed Kadomtsev– Petviashvili multi line solitons. Journal of Mathematical Physics, 62(9): 091513, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:2021:TRB

[736] Apurba Das. Twisted Rota–Baxter operators and Reynolds operators on Lie algebras and NS–Lie algebras. Journal of Mathematical Physics, 62 (9):091701, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2021:RCT

[737] Jinyeop Lee. Rate of convergence toward Hartree type equations for mixture condensates with factorized initial data. Journal of Mathematical Physics, 62(9):091901, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daCosta:2021:SQM

[738] Bruno G. da Costa, Genilson A. C. da Silva, and Ignacio S. Gomez. Supersymmetric quantum mechanics and coherent states for a deformed oscillator with position-dependent effective mass. Journal of Mathematical Physics, 62(9):092101, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shea:2021:GGY

[739] Meredith Shea. Generalized Gelfand-yaglom formula for a discretized quantum mechanic system. Journal of Mathematical Physics, 62(9): 092102, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aerts:2021:VBC

[740] Diederik Aerts and Massimiliano Sassoli de Bianchi. Violation of the Bell–CHSH inequality and marginal laws in a single-entity Bell-test experiment. Journal of Mathematical Physics, 62(9):092103, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golenia:2021:BPA

[741] Sylvain Golénia and Marc-Adrien Mandich. Bands of pure absolutely continuous spectrum for lattice Schrödinger operators with a more general long range condition. Journal of Mathematical Physics, 62(9): 092104, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holevo:2021:AIG

[742] A. S. Holevo. Accessible information of a general quantum Gaussian ensemble. Journal of Mathematical Physics, 62(9):092201, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haah:2021:CQC

[743] Jeongwan Haah. Clifford quantum cellular automata: Trivial group in 2D and Witt group in 3D. Journal of Mathematical Physics, 62(9):092202, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1095].

Walter:2021:HMC

[744] Michael Walter and Freek Witteveen. Hypergraph min-cuts from quantum entropies. Journal of Mathematical Physics, 62(9):092203, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Hermes:2021:DPD

[745] Alexander Müller-Hermes. Decomposable Pauli diagonal maps and tensor squares of qubit maps. Journal of Mathematical Physics, 62(9): 092204, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergh:2021:LSQ

[746] Bjarne Bergh, Robert Salzmann, and Nilanjana Datta. The α → 1 limit of the sharp quantum Rényi divergence. Journal of Mathematical Physics, 62(9):092205, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirokov:2021:UCB

[747] M. E. Shirokov. Uniform continuity bounds for characteristics of multipartite quantum systems. Journal of Mathematical Physics, 62(9): 092206, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osetrin:2021:EMP

[748] Konstantin Osetrin, Evgeny Osetrin, and Altair Filippov. Exact models of pure radiation in r^{2} gravity for spatially homogeneous wave-like Shapovalov spacetimes type II. Journal of Mathematical Physics, 62(9): 092501, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kobayashi:2021:UTL

[749] Katsuki Kobayashi and Satoshi Tsujimoto. The ultradiscrete Toda lattice and the Smith normal form of bidiagonal matrices. Journal of Mathematical Physics, 62(9):092701, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:SASb

[750] Xiaohu Wang, Kening Lu, and Bixiang Wang. Stationary approximations of stochastic wave equations on unbounded domains with critical exponents. Journal of Mathematical Physics, 62(9):092702, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dhanalakshmi:2021:ESI

[751] K. Dhanalakshmi and P. Balasubramaniam. Exponential stability of impulsive fractional neutral stochastic differential equations. Journal of Mathematical Physics, 62(9):092703, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deng:2021:GES

[752] Yanxia Deng and Slim Ibrahim. Global existence and singularity of Hill’s lunar problem with strong potential. Journal of Mathematical Physics, 62(9):092901, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majic:2021:IPC

[753] Matt Majic. Images of point charges in conducting ellipses and prolate spheroids. Journal of Mathematical Physics, 62(9):092902, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boritchev:2021:IRN

[754] Alexandre Boritchev, Daniel Eceizabarrena, and Victor Vilaça Da Rocha. Intermittency of Riemann’s non-differentiable function through the fourth-order flatness. Journal of Mathematical Physics, 62(9):093101, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yassin:2021:NMB

[755] Houssam Yassin. Normal modes with boundary dynamics in geophysical fluids. Journal of Mathematical Physics, 62(9):093102, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Assi:2021:DML

[756] I. A. Assi, A. D. Alhaidari, and H. Bahlouli. Deformed Morse-like potential. Journal of Mathematical Physics, 62(9):093501, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelasCuevas:2021:ATD

[757] Gemma De las Cuevas, Andreas Klingler, and Tim Netzer. Approximate tensor decompositions: Disappearance of many separations. Journal of Mathematical Physics, 62(9):093502, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hovhannisyan:2021:LSN

[758] Gro Hovhannisyan. Lax systems and nonlinear equations on a timespace scale. Journal of Mathematical Physics, 62(9):093503, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hrivnak:2021:DCS

[759] Jirí Hrivnák and Lenka Motlochová. Discrete cosine and sine transforms generalized to honeycomb lattice II. Zigzag boundaries. Journal of Mathematical Physics, 62(9):093504, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pfefferle:2021:GFM

[760] David Pfefferlé, Lyle Noakes, and David Perrella. Gauge freedom in magnetostatics and the effect on helicity in toroidal volumes. Journal of Mathematical Physics, 62(9):093505, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burby:2021:NSS

[761] J. W. Burby and E. Hirvijoki. Normal stability of slow manifolds in nearly periodic Hamiltonian systems. Journal of Mathematical Physics, 62(9): 093506, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2021:LTA

[762] Qiaoyuan Cheng, Yiling Yang, and Engui Fan. Long-time asymptotic behavior of a mixed Schrödinger equation with weighted Sobolev initial data. Journal of Mathematical Physics, 62(9):093507, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biswas:2021:DQM

[763] Indranil Biswas. Deformation quantization of moduli spaces of Higgs bundles on a Riemann surface with translation structure. Journal of Mathematical Physics, 62(9):093508, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:SMC

[764] Chuanzhong Li and Qingling Shi. Symmetries of the multi-component supersymmetric (ABC)-type KP hierarchies. Journal of Mathematical Physics, 62(9):093509, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:DMS

[765] Zi-Yi Wang, Shou-Fu Tian, and Jia Cheng. The ¯∂-dressing method and soliton solutions for the three-component coupled Hirota equations. Journal of Mathematical Physics, 62(9):093510, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aragones-Soria:2021:CRG

[766] Y. Aragonés-Soria, J. Åberg, C-Y. Park, and M. J. Kastoryano. Classical restrictions of generic matrix product states are quasi-locally Gibbsian. Journal of Mathematical Physics, 62(9):093511, September 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jonnadula:2021:SFT

[767] Bhargavi Jonnadula, Jonathan P. Keating, and Francesco Mezzadri. Symmetric function theory and unitary invariant ensembles. Journal of Mathematical Physics, 62(9):093512, September 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Lemoula:2021:KSS

[768] Romuald K. K. Lemoula, Brice A. Kamdem, Victor K. Kuetche, Raïssa S. Noule, Jean J. Defo, and Saliou Youssoufa. Kruskal’s simplification scheme in ferrite dynamics. Journal of Mathematical Physics, 62(9): 093513, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:ECI

[769] Zhipeng Zhang. Energy conservation for the incompressible inhomogeneous Euler–Korteweg equations in a bounded domain. Journal of Mathematical Physics, 62(10):101501, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2021:IPC

[770] Meina Sun. The intrinsic phenomena of cavitation and concentration in Riemann solutions for the isentropic two-phase model with the logarithmic equation of state. Journal of Mathematical Physics, 62(10):101502, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sciacca:2021:TSA

[771] M. Sciacca, F. X. Alvarez, D. Jou, and J. Bafaluy. Thermal solitons along wires with flux-limited lateral exchange. Journal of Mathematical Physics, 62(10):101503, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beck:2021:FHI

[772] T. Beck and D. Jerison. The Friedland–Hayman inequality and Caffarelli’s contraction theorem. Journal of Mathematical Physics, 62(10): 101504, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bardos:2021:DLV

[773] Claude Bardos and Nicolas Besse. Diffusion limit of the Vlasov equation in the weak turbulent regime. Journal of Mathematical Physics, 62(10): 101505, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ammari:2021:BSC

[774] Habib Ammari, Bryn Davies, Erik Orvehed Hiltunen, Hyundae Lee, and Sanghyeon Yu. Bound states in the continuum and Fano resonances in subwavelength resonator arrays. Journal of Mathematical Physics, 62 (10):101506, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2021:IDR

[775] Tingting Zhang. The invariant domain of Riemann solution for 1D nonisentropic gas dynamics equations. Journal of Mathematical Physics, 62 (10):101507, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saanouni:2021:ICN

[776] Tarek Saanouni and Radhia Ghanmi. Inhomogeneous coupled non-linear Schrödinger systems. Journal of Mathematical Physics, 62(10):101508, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2021:GCW

[777] Cheng He and Changzheng Qu. Global conservative weak solutions for the two-component Novikov equation. Journal of Mathematical Physics, 62(10):101509, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhong:2021:EIP

[778] Penghong Zhong, Ye Chen, and Ganshan Yang. The estimates of the illposedness index of the (deformed-) continuous Heisenberg spin equation.

Journal of Mathematical Physics, 62(10):101510, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2021:FNS

[779] Fan Wu. Fractional Navier–Stokes regularity criterion involving the positive part of the intermediate eigenvalue of the strain matrix. Journal of Mathematical Physics, 62(10):101511, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dereli:2021:BWN

[780] Tekin Dereli and Todor Popov. Bloch waves and non-commutative tori of magnetic translations. Journal of Mathematical Physics, 62(10):101701, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rasmussen:2021:DFW

[781] Jørgen Rasmussen and Mark A. Walton. Demazure formula for A_{n} Weyl polytope sums. Journal of Mathematical Physics, 62(10):101702, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rao:2021:GCO

[782] S. Eswara Rao. Generalized Casimir operators for Lie superalgebras. Journal of Mathematical Physics, 62(10):101703, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atakishiyeva:2021:AII

[783] Mesuma Atakishiyeva, Natig Atakishiyev, and Alexei Zhedanov. An algebraic interpretation of the intertwining operators associated with the discrete Fourier transform. Journal of Mathematical Physics, 62(10): 101704, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fischbacher:2021:EEB

[784] Christoph Fischbacher and Oluwadara Ogunkoya. Entanglement entropy bounds in the higher spin XXZ chain. Journal of Mathematical Physics, 62(10):101901, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allen:2021:SSQ

[785] Theodore J. Allen. Spin in Schrödinger-quantized pseudoclassical systems. Journal of Mathematical Physics, 62(10):102101, October 2021.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulkarni:2021:SRR

[786] Shailesh Kulkarni and Rajeev K. Pathak. “Striped” rectangular rigid box with Hermitian and non-Hermitian PT symmetric potentials. Journal of Mathematical Physics, 62(10):102102, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [970].

Truong:2021:MEW

[787] T. T. Truong. Moyal equation-Wigner distribution functions for anharmonic oscillators. Journal of Mathematical Physics, 62(10):102103, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2021:VPS

[788] Jianhao M. Yang. Variational principle for stochastic mechanics based on information measures. Journal of Mathematical Physics, 62(10):102104, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okay:2021:CTT

[789] Cihan Okay. Commutative d-torsion K-theory and its applications. Journal of Mathematical Physics, 62(10):102201, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorantla:2021:MVF

[790] Pranay Gorantla, Ho Tat Lam, Nathan Seiberg, and Shu-Heng Shao. A modified Villain formulation of fractons and other exotic theories. Journal of Mathematical Physics, 62(10):102301, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Runkel:2021:TFT

[791] Ingo Runkel and Lóránt Szegedy. Topological field theory on r-spin surfaces and the Arf-invariant. Journal of Mathematical Physics, 62(10): 102302, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freidel:2021:QGD

[792] Laurent Freidel, Christophe Goeller, and Etera R. Livine. The quantum gravity disk: Discrete current algebra. Journal of Mathematical Physics, 62(10):102303, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2021:CEF

[793] Peng Zhao, David Hilditch, and Juan A. Valiente Kroon. The conformal Einstein field equations and the local extension of future null infinity. Journal of Mathematical Physics, 62(10):102501, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sheykin:2021:GEB

[794] A. A. Sheykin, M. V. Markov, and S. A. Paston. Global embedding of BTZ spacetime using generalized method of symmetric embeddings construction. Journal of Mathematical Physics, 62(10):102502, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamiya:2021:CPD

[795] R. Kamiya, M. Kanki, T. Mase, and T. Tokihiro. Coprimeness-preserving discrete KdV type equation on an arbitrary dimensional lattice. Journal of Mathematical Physics, 62(10):102701, October 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ogle:2021:NCS

[796] C. Ogle, O. Costin, and M. Bevis. Non-convergence of the spherical harmonic expansion of gravitational potential below the Brillouin sphere: The continuous case. Journal of Mathematical Physics, 62(10):102901, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Le:2021:EWD

[797] Thi Thai Le. Effect of water depth on Kelvin–Helmholtz instability in a shallow water flow. Journal of Mathematical Physics, 62(10):103101, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2021:LPT

[798] Roberto Fernández, Manuel González-Navarrete, Eugene Pechersky, and Anatoly Yambartsev. Lack of phase transitions in staggered magnetic systems. A comparison of uniqueness criteria. Journal of Mathematical Physics, 62(10):103301, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2021:DDL

[799] Peter J. Forrester. Dyson’s disordered linear chain from a random matrix theory viewpoint. Journal of Mathematical Physics, 62(10):103302, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Bercu:2021:AAR

[800] Bernard Bercu and Fabien Montégut. Asymptotic analysis of random walks on ice and graphite. Journal of Mathematical Physics, 62(10): 103303, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostrobij:2021:GDE

[801] P. Kostrobij, M. Tokarchuk, B. Markovych, and I. Ryzha. Generalized diffusion equation with nonlocality of space-time: Analytical and numerical analysis. Journal of Mathematical Physics, 62(10):103304, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takatsu:2021:CCC

[802] Asuka Takatsu. Change the coefficients of conditional entropies in extensivity. Journal of Mathematical Physics, 62(10):103305, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Veliev:2021:SOP

[803] O. A. Veliev. On the Schrödinger operator with a periodic PT-symmetric matrix potential. Journal of Mathematical Physics, 62(10):103501, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levin:2021:GPH

[804] A. Levin, M. Olshanetsky, and A. Zotov. Generalizations of parabolic Higgs bundles, real structures, and integrability. Journal of Mathematical Physics, 62(10):103502, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2021:CVA

[805] Xiaosen Han, Genggeng Huang, and Yisong Yang. Coexisting vortices and antivortices generated by dually gauged harmonic maps. Journal of Mathematical Physics, 62(10):103503, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weisbart:2021:IGF

[806] David Weisbart. On infinitesimal generators and Feynman–Kac integrals of adelic diffusion. Journal of Mathematical Physics, 62(10):103504, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2021:HMZ

[807] Yuanran Zhu and Daniele Venturi. Hypoellipticity and the Mori–Zwanzig formulation of stochastic differential equations. Journal of Mathematical Physics, 62(10):103505, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2021:CSS

[808] Francisco M. Fernández. Comment on “Solution of Schrödinger equation for two different potentials using extended Nikiforov–Uvarov method and polynomial solutions of biconfluent Heun equation” [J. Math. Phys. 59, 053501 (2018)]. Journal of Mathematical Physics, 62(10):104101, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [10].

Fan:2021:ABS

[809] Huiying Fan and Meng Wang. Asymptotic behavior of the stationary magnetohydrodynamic equations in an exterior domain. Journal of Mathematical Physics, 62(11):111501, November 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dietze:2021:DEN

[810] Charlotte Dietze. Dispersive estimates for nonlinear Schrödinger equations with external potentials. Journal of Mathematical Physics, 62(11): 111502, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2021:NRP

[811] A. Alexandrou Himonas, Carlos Madrid, and Fangchi Yan. The Neumann and Robin problems for the Korteweg-de Vries equation on the half-line. Journal of Mathematical Physics, 62(11):111503, November 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2021:LWS

[812] Xiangdi Huang and Wei Yan. Local weak solution of the isentropic compressible Navier–Stokes equations. Journal of Mathematical Physics, 62(11):111504, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2021:FKT

[813] Senli Liu and Haibo Chen. Fractional Kirchhoff-type equation with singular potential and critical exponent. Journal of Mathematical Physics, 62(11):111505, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2021:SPL

[814] Ky Ho, Yun-Ho Kim, and Jongrak Lee. Schrödinger p·-Laplace equations in R^{N} involving indefinite weights and critical growth. Journal of Mathematical Physics, 62(11):111506, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2021:USC

[815] Jintao Wang, Chunqiu Li, Lu Yang, and Mo Jia. Upper semi-continuity of random attractors and existence of invariant measures for nonlocal stochastic Swift–Hohenberg equation with multiplicative noise. Journal of Mathematical Physics, 62(11):111507, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2021:BVB

[816] Aichao Liu and Binxiang Dai. Blow-up vs boundedness in a two-species attraction-repulsion chemotaxis system with two chemicals. Journal of Mathematical Physics, 62(11):111508, November 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Severo:2021:APQ

[817] Uberlandio B. Severo and Diogo de S. Germano. Asymptotically periodic quasilinear Schrödinger equations with critical exponential growth. Journal of Mathematical Physics, 62(11):111509, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peng:2021:ARC

[818] Hongyun Peng. On the attraction-repulsion chemotaxis system with volume-filling effect. Journal of Mathematical Physics, 62(11):111510, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sopenko:2021:ITD

[819] Nikita Sopenko. An index for two-dimensional SPT states. Journal of Mathematical Physics, 62(11):111901, November 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2021:AAN

[820] Hiroshi Inoue. An algebraic approach of non-self-adjoint Hamiltonians in Krein spaces. Journal of Mathematical Physics, 62(11):112101, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morchio:2021:MTO

[821] G. Morchio and F. Strocchi. Manifold topology, observables, and gauge group. Journal of Mathematical Physics, 62(11):112102, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burbanks:2021:RCA

[822] Andrew D. Burbanks, Andrew H. Osbaldestin, and Judi A. Thurlby. Rigorous computer-assisted bounds on the period doubling renormalization fixed point and eigenfunctions in maps with critical point of degree 4. Journal of Mathematical Physics, 62(11):112701, November 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2021:RIM

[823] Xianming Liu. Random invariant manifolds of stochastic evolution equations driven by Gaussian and non-Gaussian noises. Journal of Mathematical Physics, 62(11):112702, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2021:MPS

[824] Hui Wei, Mu Ma, and Shuguan Ji. Multiple periodic solutions for an asymptotically linear wave equation with x-dependent coefficients. Journal of Mathematical Physics, 62(11):112703, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramond:2021:NMI

[825] Paul Ramond and Jérôme Perez. New methods of isochrone mechanics. Journal of Mathematical Physics, 62(11):112704, November 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Kumar:2021:DIT

[826] Narender Kumar, S. B. Bhardwaj, Vinod Kumar, Ram Mehar Singh, and Fakir Chand. Dynamical invariants for time-dependent real and complex Hamiltonian systems. Journal of Mathematical Physics, 62(11):112705, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [865].

Arustamyan:2021:NEB

[827] Nickolas Arustamyan, Christopher Cox, Erik Lundberg, Sean Perry, and Zvi Rosen. On the number of equilibria balancing Newtonian point masses with a central force. Journal of Mathematical Physics, 62(11): 112901, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zamparo:2021:CFR

[828] Marco Zamparo. Critical fluctuations in renewal models of statistical mechanics. Journal of Mathematical Physics, 62(11):113301, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poria:2021:UPF

[829] Anirudha Poria. Uncertainty principles for the Fourier and the short-time Fourier transforms. Journal of Mathematical Physics, 62(11):113501, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bikram:2021:CPO

[830] Panchugopal Bikram and Rajeeb R. Mohanta. Contractivity properties of Ornstein–Uhlenbeck semigroup for mixed q-Araki–Woods von Neumann algebras. Journal of Mathematical Physics, 62(11):113502, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudnick:2021:RPR

[831] Zeév Rudnick and Igor Wigman. The Robin problem on rectangles. Journal of Mathematical Physics, 62(11):113503, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Huh:2021:TCB

[832] Youngsik Huh, Hyoungjun Kim, and Seungsang Oh. Tight conformation of 2-bridge knots using superhelices. Journal of Mathematical Physics, 62(11):113504, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fabbri:2021:ARI

[833] Luca Fabbri and Andre G. Campos. Angular-radial integrability of Coulomb-like potentials in Dirac equations. Journal of Mathematical Physics, 62(11):113505, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fu:2021:STC

[834] Zhengqi Fu and Xiong Li. Spectral type of a class of random Jacobi operators. Journal of Mathematical Physics, 62(11):113506, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giusti:2021:EDR

[835] Andrea Giusti. Erratum: “Dispersion relations for the time-fractional Cattaneo–Maxwell heat equation” [J. Math. Phys. 59, 013506 (2018)]. Journal of Mathematical Physics, 62(11):119901, November 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [9].

Buoso:2021:SAB

[836] Davide Buoso, Paolo Luzzini, Luigi Provenzano, and Joachim Stubbe. On the spectral asymptotics for the buckling problem. Journal of Mathematical Physics, 62(12):121501, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2021:RCA

[837] Zhengguang Guo, Yu Wang, and Yeping Li. Regularity criteria of axisymmetric weak solutions to the 3D MHD equations. Journal of Mathematical Physics, 62(12):121502, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2021:ASC

[838] Jianwei Dong and Litao Zhang. Analytical solutions to the 1D compressible isothermal Navier–Stokes equations with density-dependent viscosity. Journal of Mathematical Physics, 62(12):121503, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2021:NUD

[839] Yanghai Yu and Xiaolei Yang. Non-uniform dependence on initial data for the 2D MHD–Boussinesq equations. Journal of Mathematical Physics, 62(12):121504, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2021:MSC

[840] Yongpeng Chen and Miaomiao Niu. The multiplicity of solutions for the critical Schrödinger-Poisson system with competing potentials. Journal of Mathematical Physics, 62(12):121505, December 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zheng:2021:GWS

[841] Pan Zheng. Global weak solution in a p-Laplacian Keller–Segel system with nonlinear sensitivity and saturation effect. Journal of Mathematical Physics, 62(12):121506, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2021:NCL

[842] Juan Huang, Yulin Li, and Yunya Yang. A new criterion and the limit of blowup solutions for a generalized Davey–Stewartson system. Journal of Mathematical Physics, 62(12):121507, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tu:2021:CQS

[843] Kaifei Tu and Yongkuan Cheng. On a class of quasilinear Schrödinger equations with the supercritical growth. Journal of Mathematical Physics, 62(12):121508, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakamura:2021:CPK

[844] Makoto Nakamura. The Cauchy problem for the Klein–Gordon equation under the quartic potential in the de Sitter spacetime. Journal of Mathematical Physics, 62(12):121509, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alhussein:2021:HCU

[845] H. Alhussein and P. Kolesnikov. On the Hochschild cohomology of universal enveloping associative conformal algebras. Journal of Mathematical Physics, 62(12):121701, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rybakov:2021:ASN

[846] Filipp N. Rybakov and Egor Babaev. The absence of superconductivity in the next-to-leading order Ginzburg–Landau functional for Bardeen– Cooper–Schrieffer superconductor. Journal of Mathematical Physics, 62(12):121901, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Na:2021:RQZ

[847] Kyungsun Na. Revisiting quantum Zeno effect and anti-Zeno effect: Universality vs non-universality. Journal of Mathematical Physics, 62(12): 122101, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2021:LRP

[848] Wen-Du Li and Wu-Sheng Dai. Long-range potential scattering: Converting long-range potential to short-range potential by tortoise coordinate. Journal of Mathematical Physics, 62(12):122102, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soto-Eguibar:2021:BPT

[849] Francisco Soto-Eguibar, Felipe A. Asenjo, Sergio A. Hojman, and Héctor M. Moya-Cessa. Bohm potential for the time dependent harmonic oscillator. Journal of Mathematical Physics, 62(12):122103, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigi:2021:IQH

[850] Salman Beigi. Improved quantum hypercontractivity inequality for the qubit depolarizing channel. Journal of Mathematical Physics, 62(12): 122201, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuipers:2021:SQR

[851] Folkert Kuipers. Stochastic quantization of relativistic theories. Journal of Mathematical Physics, 62(12):122301, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kennon:2021:SMM

[852] Aaron Kennon. Spin(7)-manifolds and multisymplectic geometry. Journal of Mathematical Physics, 62(12):122302, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahn:2021:EBD

[853] Hyunjin Ahn, Seung-Yeal Ha, and Woojoo Shim. Emergent behaviors of the discrete thermodynamic Cucker–Smale model on complete Riemannian manifolds. Journal of Mathematical Physics, 62(12):122701, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burby:2021:INF

[854] J. W. Burby, N. Duignan, and J. D. Meiss. Integrability, normal forms, and magnetic axis coordinates. Journal of Mathematical Physics, 62(12): 122901, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2021:CLD

[855] Manuel de León, Manuel Laínz, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Constrained Lagrangian dissipative contact dynamics. Journal of Mathematical Physics, 62(12):122902, December 2021. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hittmeir:2021:DSI

[856] Sabine Hittmeir, Rupert Klein, Annette Müller, and Peter Névir. The dynamic state index with moisture and phase changes. Journal of Mathematical Physics, 62(12):123101, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Subag:2021:CCS

[857] Eliran Subag and Ofer Zeitouni. Concentration of the complexity of spherical pure p-spin models at arbitrary energies. Journal of Mathematical Physics, 62(12):123301, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Breiten:2021:SGD

[858] Tobias Breiten, Carsten Hartmann, Lara Neureither, and Upanshu Sharma. Stochastic gradient descent and fast relaxation to thermodynamic equilibrium: a stochastic control approach. Journal of Mathematical Physics, 62(12):123302, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2021:NFG

[859] Jin-Yan Zhu and Yong Chen. A new form of general soliton solutions and multiple zeros solutions for a higher-order Kaup–Newell equation. Journal of Mathematical Physics, 62(12):123501, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barth:2021:AEE

[860] Simon Barth, Andreas Bitter, and Semjon Vugalter. The absence of the Efimov effect in systems of one- and two-dimensional particles. Journal of Mathematical Physics, 62(12):123502, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kolupaiev:2021:ASS

[861] Oleksii Kolupaiev. Anomalous singularity of the solution of the vector Dyson equation in the critical case. Journal of Mathematical Physics, 62(12):123503, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bao:2021:LBB

[862] Aiguo Bao and Xianfa Song. Lower bound for the blowup time of the solution to a quasi-linear parabolic system. Journal of Mathematical Physics, 62(12):123504, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naber:2021:ZFT

[863] M. G. Naber. The zeta function for the triangular potential. Journal of Mathematical Physics, 62(12):123505, December 2021. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hennig:2021:EES

[864] Dirk Hennig and Nikos I. Karachalios. Existence of exponentially spatially localized breather solutions for lattices of nonlinearly coupled particles: Schauder’s fixed point theorem approach. Journal of Mathematical Physics, 62(12):123506, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kumar:2021:PND

[865] Narender Kumar, S. B. Bhardwaj, Vinod Kumar, Ram Mehar Singh, and Fakir Chand. Publisher’s note: “Dynamical invariants for timedependent real and complex Hamiltonian systems” [j. math. phys. 62, 112705 (2021)]. Journal of Mathematical Physics, 62(12):129901, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [826].

He:2022:RPW

[866] Fen He, Tingting Chen, Qingling Zhang, and Zhen Wang. Riemann problems and wave interactions for a non-symmetric Keyfitz–Kranzer system with a source term. Journal of Mathematical Physics, 63(1): 011501, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2022:CSN

[867] Congcong Hu and Weiming Liu. Construction of solutions for the nonlinear magnetic Schrödinger equation in RN. Journal of Mathematical Physics, 63(1):011502, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2022:GSD

[868] Jianping Wang. Global solutions of a doubly tactic resource consumption model with logistic source. Journal of Mathematical Physics, 63 (1):011503, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2022:TDG

[869] Wei Xiao and WeiLi Li. The two dimensional gas expansion problem of the Euler equations for a class of pressure laws. Journal of Mathematical Physics, 63(1):011504, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freire:2022:PPC

[870] Igor Leite Freire. Persistence properties of a Camassa–Holm type equation with (n + 1)-order non-linearities. Journal of Mathematical Physics, 63(1):011505, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2022:GWP

[871] Xinying Yang and Xin Zhong. Global well-posedness and decay estimates to the 3D Cauchy problem of nonhomogeneous magneto-micropolar fluid equations with vacuum. Journal of Mathematical Physics, 63(1):011506, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benhellal:2022:SAD

[872] Badreddine Benhellal. Spectral analysis of Dirac operators with delta interactions supported on the boundaries of rough domains. Journal of Mathematical Physics, 63(1):011507, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2022:UAR

[873] Xiangming Zhu and Chengkui Zhong. Uniform attractors for reactiondiffusion equations with a larger class of external forces. Journal of Mathematical Physics, 63(1):011508, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hintz:2022:QMS

[874] Peter Hintz and YuQing Xie. Quasinormal modes of small Schwarzschildde Sitter black holes. Journal of Mathematical Physics, 63(1):011509, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ndogmo:2022:ICF

[875] J. C. Ndogmo. Integrable classes of a family of evolution equations. Journal of Mathematical Physics, 63(1):011510, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramos:2022:EBE

[876] A. J. A. Ramos, T. A. Apalara, M. M. Freitas, and M. L. Araújo. Equivalence between exponential stabilization and boundary observability for swelling problem. Journal of Mathematical Physics, 63(1):011511, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2022:MCP

[877] Haining Fan and Xiaochun Liu. On the multiplicity and concentration of positive solutions to a Kirchhoff-type problem with competing potentials. Journal of Mathematical Physics, 63(1):011512, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Butorac:2022:HAA

[878] Marijana Butorac and Slaven Kozić. On the Heisenberg algebra associated with the rational R-matrix. Journal of Mathematical Physics, 63 (1):011701, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Henheik:2022:ATT

[879] Joscha Henheik and Stefan Teufel. Adiabatic theorem in the thermodynamic limit: Systems with a uniform gap. Journal of Mathematical Physics, 63(1):011901, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ogata:2022:DBC

[880] Yoshiko Ogata. A derivation of braided C^{∗}-tensor categories from gapped ground states satisfying the approximate Haag duality. Journal of Mathematical Physics, 63(1):011902, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buchholz:2022:PC

[881] Detlev Buchholz. Proper condensates. Journal of Mathematical Physics, 63(1):011903, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fulsche:2022:STS

[882] Robert Fulsche and Medet Nursultanov. Spectral theory for Sturm– Liouville operators with measure potentials through Otelbaev’s function. Journal of Mathematical Physics, 63(1):012101, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haynes:2022:LSS

[883] Alan Haynes and Roland Roeder. Level spacing statistics for the multidimensional quantum harmonic oscillator: Algebraic case. Journal of Mathematical Physics, 63(1):012102, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Sanchez-Mendoza:2022:IDS

[884] Daniel Sánchez-Mendoza. The integrated density of states of the 1D discrete Anderson–Bernoulli model at rational energies. Journal of Mathematical Physics, 63(1):012103, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2022:RMP

[885] Heon Lee. Relativistic massive particle with spin-1/2: a vector bundle point of view. Journal of Mathematical Physics, 63(1):012201, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajchel-Mieldzioc:2022:AGS

[886] Grzegorz Rajchel-Mieldzioć, Kamil Korzekwa, Zbigniew Puchała, and Karol Zyczkowski. Algebraic and geometric structures inside the Birkhoff polytope. Journal of Mathematical Physics, 63(1):012202, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cercle:2022:LCF

[887] Baptiste Cerclé. Liouville conformal field theory on even-dimensional spheres. Journal of Mathematical Physics, 63(1):012301, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kyriakopoulos:2022:RIS

[888] E. Kyriakopoulos. Regular interior solutions to the solution of Kerr which satisfy the weak and the strong energy conditions. Journal of Mathematical Physics, 63(1):012501, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [932].

Yang:2022:DIA

[889] Kexiang Yang, Ercai Chen, and Xiaoyao Zhou. Dynamical intricacy and average sample complexity for random bundle transformations. Journal of Mathematical Physics, 63(1):012701, January 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Byeon:2022:AFD

[890] Junhyeok Byeon, Seung-Yeal Ha, and Jeongho Kim. Asymptotic flocking dynamics of a relativistic Cucker–Smale flock under singular communications. Journal of Mathematical Physics, 63(1):012702, January 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2022:SLT

[891] Pan Liu and Genqian Liu. Some Liouville-type theorems for the stationary density-dependent Navier–Stokes equations. Journal of Mathematical Physics, 63(1):013101, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2022:WDC

[892] Xiaohui Ma, Mohamed El Machkouri, and Xiequan Fan. On Wasserstein-1 distance in the central limit theorem for elephant random walk. Journal of Mathematical Physics, 63(1):013301, January 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbonaro:2022:RPI

[893] Bruno Carbonaro. The role of the principle of inertia in KTAP models. Journal of Mathematical Physics, 63(1):013302, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lewandowski:2022:HSB

[894] Max Lewandowski. Hadamard states for bosonic quantum field theory on globally hyperbolic spacetimes. Journal of Mathematical Physics, 63 (1):013501, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Briet:2022:SOD

[895] Philippe Briet and David Krejcirík. Spectral optimization of Dirac rectangles. Journal of Mathematical Physics, 63(1):013502, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2022:ATT

[896] Yixuan Liu, Guoliang Shi, and Jun Yan. Ambarzumyan-type theorem for third order linear measure differential equations. Journal of Mathematical Physics, 63(1):013503, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Derezinski:2022:MAR

[897] Jan Dereziński and Oskar Grocholski. Momentum approach to the 1/r^{2} potential as a toy model of the Wilsonian renormalization. Journal of Mathematical Physics, 63(1):013504, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feinberg:2022:WMC

[898] Joshua Feinberg and Miloslav Znojil. Which metrics are consistent with a given pseudo-hermitian matrix? Journal of Mathematical Physics, 63 (1):013505, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zuniga-Galindo:2022:GLZ

[899] W. A. Zúñiga-Galindo, B. A. Zambrano-Luna, and E. León-Cardenal. Graphs, local zeta functions, log-Coulomb gases, and phase transitions at finite temperature. Journal of Mathematical Physics, 63(1):013506, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Charbonneau:2022:CND

[900] Benoit Charbonneau, Anuk Dayaprema, C. J. Lang, Ákos Nagy, and Haoyang Yu. Construction of Nahm data and BPS monopoles with continuous symmetries. Journal of Mathematical Physics, 63(1):013507, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:2022:LQR

[901] Kai Zhao and Wei-Shih Yang. The localization of quantum random walks on Sierpinski gaskets. Journal of Mathematical Physics, 63(1):013508, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jecko:2022:AER

[902] Thierry Jecko. On the analyticity of electronic reduced densities for molecules. Journal of Mathematical Physics, 63(1):013509, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rao:2022:RCB

[903] Jiguang Rao, Jingsong He, and Boris A. Malomed. Resonant collisions between lumps and periodic solitons in the Kadomtsev–Petviashvili I equation. Journal of Mathematical Physics, 63(1):013510, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:2022:TTM

[904] Barry Simon. Twelve tales in mathematical physics: an expanded Heineman prize lecture. Journal of Mathematical Physics, 63(2):021101, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2022:GWPa

[905] Jinkai Li and Guozhi Yuan. Global well-posedness of z-weak solutions to the primitive equations without vertical diffusivity. Journal of Mathematical Physics, 63(2):021501, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiang:2022:GEL

[906] Yuan-Jen Chiang and Changhua Wei. Global existence and lifespan of smooth solutions of biwave maps. Journal of Mathematical Physics, 63 (2):021502, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qi:2022:MVS

[907] Kunlun Qi. Measure valued solution to the spatially homogeneous Boltzmann equation with inelastic long-range interactions. Journal of Mathematical Physics, 63(2):021503, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2022:GST

[908] Kai-Kai Liu and Yun-Rui Yang. Global stability of traveling waves for a SIR model with nonlocal dispersal and delay. Journal of Mathematical Physics, 63(2):021504, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2022:NGS

[909] Penghui Zhang and Zhiqing Han. Normalized ground states for Kirchhoff equations in R3 with a critical nonlinearity. Journal of Mathematical Physics, 63(2):021505, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DAprile:2022:NSB

[910] Teresa D’Aprile. Non-symmetric blowing-up solutions for a class of Liouville equations in the ball. Journal of Mathematical Physics, 63(2): 021506, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2022:VST

[911] Luyi Ma, Hong-Tao Niu, and Zhi-Cheng Wang. V-shaped traveling fronts of fractional Allen–Cahn equations. Journal of Mathematical Physics, 63 (2):021507, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cai:2022:UCS

[912] Hong Cai, Geng Chen, Yi Du, and Yannan Shen. Uniqueness of conservative solutions to a one-dimensional general quasilinear wave equation through variational principle. Journal of Mathematical Physics, 63(2): 021508, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2022:SAL

[913] Hong-Yi Zhang and Yu-Feng Zhang. Spectral analysis and long-time asymptotics of complex mKdV equation. Journal of Mathematical Physics, 63(2):021509, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazorchuk:2022:SHC

[914] Volodymyr Mazorchuk and Rafael Mrden. ∫↕_{∈}-Harish-Chandra modules for ∫↕_{∈}[ltimes]L(△). Journal of Mathematical Physics, 63(2):021701, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2022:YBA

[915] Na Wang and Ke Wu. Yang–Baxter algebra and MacMahon representation. Journal of Mathematical Physics, 63(2):021702, February 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dessano:2022:PSS

[916] H. Dessano, G. X. A. Petronilo, R. G. G. Amorim, and A. E. Santana. Phase space Schrödinger equation and quadratic friction. Journal of Mathematical Physics, 63(2):022101, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bachelot:2022:DSN

[917] Alain Bachelot. The Dirac sea for the non-separable Hilbert spaces. Journal of Mathematical Physics, 63(2):022102, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kameoka:2022:CAP

[918] Kentaro Kameoka. Complex absorbing potential method for Stark resonances. Journal of Mathematical Physics, 63(2):022103, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez:2022:RDW

[919] Felipe Hernández and C. Jess Riedel. Rapidly decaying Wigner functions are Schwartz functions. Journal of Mathematical Physics, 63(2):022104, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vidick:2022:ASQ

[920] Thomas Vidick. Almost synchronous quantum correlations. Journal of Mathematical Physics, 63(2):022201, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gicquaud:2022:ESL

[921] Romain Gicquaud. Existence of solutions to the Lichnerowicz equation: a new proof. Journal of Mathematical Physics, 63(2):022501, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2022:RAE

[922] Ming Wang, Zhiming Liu, and Jianhua Huang. Regular attractor of the β-evolution equation with fractional damping on R^{n}. Journal of Mathematical Physics, 63(2):022701, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2022:GHJ

[923] Manuel de León, Manuel Lainz, and Asier López-Gordón. Geometric Hamilton–Jacobi theory for systems with external forces. Journal of Mathematical Physics, 63(2):022901, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maggi:2022:DRE

[924] Rocco Maggi, Elisa Ercolessi, Paolo Facchi, Giuseppe Marmo, Saverio Pascazio, and Francesco V. Pepe. Dimensional reduction of electromagnetism. Journal of Mathematical Physics, 63(2):022902, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoshida:2022:KOF

[925] Zensho Yoshida and Philip J. Morrison. The kinetic origin of the fluid helicity-A symmetry in the kinetic phase space. Journal of Mathematical Physics, 63(2):023101, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woolls:2022:HCP

[926] Michael Woolls and Leonid P. Pryadko. Homology-changing percolation transitions on finite graphs. Journal of Mathematical Physics, 63(2): 023301, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Domrin:2022:MSW

[927] A. V. Domrin, M. A. Shumkin, and B. I. Suleimanov. Meromorphy of solutions for a wide class of ordinary differential equations of Painlevé type. Journal of Mathematical Physics, 63(2):023501, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Astaburuaga:2022:SRO

[928] M. A. Astaburuaga, V. H. Cortés, C. Fernández, and R. Del Río. Singular rank one perturbations. Journal of Mathematical Physics, 63(2):023502, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2022:TFV

[929] Wencai Liu. Topics on Fermi varieties of discrete periodic Schrödinger operators. Journal of Mathematical Physics, 63(2):023503, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakraborty:2022:STR

[930] Anwesha Chakraborty, Partha Nandi, and Biswajit Chakraborty. Spectral triple with real structure on fuzzy sphere. Journal of Mathematical Physics, 63(2):023504, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Obukhov:2022:AIM

[931] Valeriy V. Obukhov. Algebras of integrals of motion for the Hamilton– Jacobi and Klein–Gordon–Fock equations in spacetime with fourparameter groups of motions in the presence of an external electromagnetic field. Journal of Mathematical Physics, 63(2):023505, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kyriakopoulos:2022:PNR

[932] E. Kyriakopoulos. Publisher’s note: “Regular interior solutions to the solution of Kerr which satisfy the weak and the strong energy conditions” [J. Math. Phys. 63, 012501 (2022)]. Journal of Mathematical Physics, 63 (2):029901, February 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [888].

Olbermann:2022:GVC

[933] H. Olbermann. Godunov variables and convex entropy for relativistic fluid dynamics with bulk viscosity. Journal of Mathematical Physics, 63 (3):031501, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2022:SSE

[934] Ho Lee and Jiho Lee. Small solutions of the Einstein–Boltzmann-scalar field system in a spatially flat FLRW spacetime. Journal of Mathematical Physics, 63(3):031502, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2022:PLK

[935] Zhiheng Shi and Sihua Liang. On the p-Laplacian Kirchhoff–Schrödinger equation with potentials vanishing or unbounded at infinity in R^{3}. Journal of Mathematical Physics, 63(3):031503, March 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nie:2022:VFP

[936] Yongsheng Nie, Jin Hao, and Hanchun Yang. Vanishing flux perturbation, pressure, and magnetic field limit in a Chaplygin magnetogasdynamics. Journal of Mathematical Physics, 63(3):031504, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deHoop:2022:IPR

[937] Maarten V. de Hoop and Alexei Iantchenko. Inverse problem for the Rayleigh system with spectral data. Journal of Mathematical Physics, 63(3):031505, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [988].

Lu:2022:GEI

[938] Peng Lu and Yi Zhou. Global existence of ideal invicid compressible and heat-conductive fluids with radial symmetry. Journal of Mathematical Physics, 63(3):031506, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2022:GSC

[939] Fuyi Xu, Ai Huang, and Peng Fu. The global solvability of the Cauchy problem for a multi-dimensional chemotaxis-Navier–Stokes system modeling coral fertilization. Journal of Mathematical Physics, 63(3):031507, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbou:2022:SSS

[940] G. Carbou, M. Moussaoui, and R. Rachi. Stability of steady states in ferromagnetic rings. Journal of Mathematical Physics, 63(3):031508, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamano:2022:ECI

[941] Masaru Hamano and Masahiro Ikeda. Equivalence of conditions on initial data below the ground state to NLS with a repulsive inverse power potential. Journal of Mathematical Physics, 63(3):031509, March 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2022:GSS

[942] Teng Wang, Xiang Wang, and Rong Zhang. Global strong solutions to planar radiative magnetohydrodynamic equations with densitydependent viscosity and degenerate heat-conductivity. Journal of Mathematical Physics, 63(3):031510, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasnov:2022:SPO

[943] Kirill Krasnov. Spin(11, 3), particles, and octonions. Journal of Mathematical Physics, 63(3):031701, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kraisler:2022:CSE

[944] Joseph Kraisler and John C. Schotland. Collective spontaneous emission and kinetic equations for one-photon light in random media. Journal of Mathematical Physics, 63(3):031901, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krishnaswami:2022:QRR

[945] Govind S. Krishnaswami and T. R. Vishnu. Quantum Rajeev–Ranken model as an anharmonic oscillator. Journal of Mathematical Physics, 63 (3):032101, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borasi:2022:FDS

[946] Luigi Borasi. Finite dimensional systems of free fermions and diffusion processes on spin groups. Journal of Mathematical Physics, 63(3):032102, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamza:2022:OWP

[947] Danish Ali Hamza, Tasawar Abbas, Muhammad Asjad, Muhammad Imran, and Rameez ul Islam. Observing wave-particle behavior of entangled atoms in cavity assisted delayed-choice experiment. Journal of Mathematical Physics, 63(3):032103, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shekarriz:2022:NCO

[948] Mohammad H. Shekarriz and Karl Svozil. Noncontextual coloring of orthogonality hypergraphs. Journal of Mathematical Physics, 63(3):

032104, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farshi:2022:MLR

[949] Tom Farshi, Daniele Toniolo, Carlos E. González-Guillén, Álvaro M. Alhambra, and Lluis Masanes. Mixing and localization in random timeperiodic quantum circuits of Clifford unitaries. Journal of Mathematical Physics, 63(3):032201, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cameron:2022:AFH

[950] Peter Cameron and Piotr T. Chruściel. Asymptotic flatness in higher dimensions. Journal of Mathematical Physics, 63(3):032501, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fournodavlos:2022:FDF

[951] Grigorios Fournodavlos. Future dynamics of FLRW for the masslessscalar field system with positive cosmological constant. Journal of Mathematical Physics, 63(3):032502, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohmann:2022:MFF

[952] Manuel Hohmann, Christian Pfeifer, and Nicoleta Voicu. Mathematical foundations for field theories on Finsler spacetimes. Journal of Mathematical Physics, 63(3):032503, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pires:2022:IVF

[953] L. Pires. The initial-value formulation of the λ–R model. Journal of Mathematical Physics, 63(3):032504, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2022:RAS

[954] Yiju Chen and Xiaohu Wang. Random attractors for stochastic discrete complex Ginzburg–Landau equations with long-range interactions. Journal of Mathematical Physics, 63(3):032701, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Omiste:2022:ICM

[955] Juan J. Omiste, Rosario González-Férez, and Rafael Ortega. An introduction to classical monodromy: Applications to molecules in external fields. Journal of Mathematical Physics, 63(3):032702, March 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2022:RAS

[956] Wenjie Hu and Quanxin Zhu. Random attractors for a stochastic agestructured population model. Journal of Mathematical Physics, 63(3): 032703, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landon:2022:FEF

[957] Benjamin Landon. Free energy fluctuations of the two-spin spherical SK model at critical temperature. Journal of Mathematical Physics, 63 (3):033301, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:2022:SGL

[958] Barry Simon. The strong Gauss–Lucas theorem and analyticity of correlation functions via the Lee–Yang theorem. Journal of Mathematical Physics, 63(3):033302, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Coletti:2022:SLF

[959] Cristian F. Coletti and Leon A. Valencia. Scaling limit for a family of coalescing radial random paths absorbed at the origin. Journal of Mathematical Physics, 63(3):033303, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bose:2022:TDF

[960] Arup Bose, Shambhu Nath Maurya, and Koushik Saha. Time dependent fluctuations of linear eigenvalue statistics of some patterned matrices. Journal of Mathematical Physics, 63(3):033304, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spohn:2022:HEA

[961] Herbert Spohn. Hydrodynamic equations for the Ablowitz–Ladik discretization of the nonlinear Schrödinger equation. Journal of Mathematical Physics, 63(3):033305, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Chen:2022:SMK

[962] Huizhan Chen, Jipeng Cheng, and Zhiwei Wu. Super modified KP hierarchy in Kac–van de Leur construction. Journal of Mathematical Physics, 63(3):033501, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peng:2022:DTP

[963] Wei-Qi Peng and Yong Chen. Double and triple pole solutions for the Gerdjikov–Ivanov type of derivative nonlinear Schrödinger equation with zero/nonzero boundary conditions. Journal of Mathematical Physics, 63 (3):033502, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durand:2022:FOM

[964] Loyal Durand. Fractional operators and multi-integral representations for associated Legendre functions. Journal of Mathematical Physics, 63 (3):033503, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Said:2022:FPA

[965] Salem Ben Saïd and Selma Negzaoui. Flett potentials associated with differential-difference Laplace operators. Journal of Mathematical Physics, 63(3):033504, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Najafizade:2022:RDO

[966] Amene Najafizade, Hossein Panahi, Won Sang Chung, and Hassan Hassanabadi. A representation of the Dunkl oscillator model on curved spaces: Factorization approach. Journal of Mathematical Physics, 63 (3):033505, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DAchille:2022:DTD

[967] Matteo D’Achille, Aernout C. D. van Enter, and Arnaud Le Ny. Decimations for two-dimensional Ising and rotator models. Journal of Mathematical Physics, 63(3):033506, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2022:GSP

[968] T. Skrypnyk. On the general solution of the permuted classical Yang– Baxter equation and quasigraded Lie algebras. Journal of Mathematical Physics, 63(3):033507, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2022:TSC

[969] Ya Gao, Jing-Hua Li, and Jing Mao. Translating solutions for a class of quasilinear parabolic initial boundary value problems in Lorentz– Minkowski plane R12. Journal of Mathematical Physics, 63(3):033508, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2022:C

[970] Francisco M. Fernández. Comment on ‘Striped’ rectangular rigid box with Hermitian and non-Hermitian PT symmetric potentials“ [J. Math. Phys. 62, 102102 (2021)]. Journal of Mathematical Physics, 63(3): 034101, March 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [786].

Auffinger:2022:SCA

[971] Antonio Auffinger, Gerard Ben Arous, and Zhehua Li. Sharp complexity asymptotics and topological trivialization for the (p, k) spiked tensor model. Journal of Mathematical Physics, 63(4):043303, April 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL https://pubs.aip.org/aip/jmp/article/63/4/043303/
2843034/Sharp-complexity-asymptotics-and-topological. Special

collection in honor of Freeman Dyson.

Li:2022:NSD

[972] Nianhua Li and Kai Tian. Nonlocal symmetries and Darboux transformations of the Camassa–Holm equation and modified Camassa–Holm equation revisited. Journal of Mathematical Physics, 63(4):041501, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juarez-Campos:2022:SGL

[973] B. Juarez-Campos, E. I. Kaikina, and A. V. Vázquez-Esquivel. Stochastic Ginzburg–Landau equation on a half-line driven by the multiplicative noise. Journal of Mathematical Physics, 63(4):041502, April 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2022:ATA

[974] Jingxuan Zhang. Adiabatic theory for the area-constrained Willmore flow. Journal of Mathematical Physics, 63(4):041503, April 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2022:LRG

[975] Weixian Sun and Zhuan Ye. Low regularity global well-posedness for 2D Boussinesq equations with variable viscosity. Journal of Mathematical Physics, 63(4):041504, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2022:SEE

[976] Cong Wang and Jiabao Su. The semilinear elliptic equations with double weighted critical exponents. Journal of Mathematical Physics, 63(4): 041505, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2022:GET

[977] Zhengrong Liu. Global existence and time decay of the relativistic BGK equation in the whole space. Journal of Mathematical Physics, 63(4): 041506, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Warze:2022:SGF

[978] Simone Warze and Amanda Young. The spectral gap of a fractional quantum Hall system on a thin torus. Journal of Mathematical Physics, 63(4):041901, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huo:2022:SOH

[979] Qinghai Huo and Guangbin Ren. Structure of octonionic Hilbert spaces with applications in the Parseval equality and Cayley–Dickson algebras. Journal of Mathematical Physics, 63(4):042101, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2022:TEI

[980] Kil-Chan Ha, Kyung Hoon Han, and Seung-Hyeok Kye. There exist infinitely many kinds of partial separability/entanglement. Journal of Mathematical Physics, 63(4):042201, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuipers:2022:ACS

[981] Folkert Kuipers. Analytic continuation of stochastic mechanics. Journal of Mathematical Physics, 63(4):042301, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:2022:BCI

[982] Christyan C. de Oliveira, Ricardo A. Mosna, and João Paulo M. Pitelli. Boundary conditions for isolated asymptotically anti-de Sitter spacetimes. Journal of Mathematical Physics, 63(4):042501, April 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lieb:2022:EFI

[983] Elliott H. Lieb and Siddhartha Sahi. On the extension of the FKG inequality to n functions. Journal of Mathematical Physics, 63(4):043301, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Albanese:2022:GSG

[984] Linda Albanese and Andrea Alessandrelli. On Gaussian spin glass with P -wise interactions. Journal of Mathematical Physics, 63(4):043302, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazzuca:2022:MDS

[985] G. Mazzuca. On the mean density of states of some matrices related to the beta ensembles and an application to the Toda lattice. Journal of Mathematical Physics, 63(4):043501, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anantharaman:2022:HGA

[986] Nalini Anantharaman and Laura Monk. A high-genus asymptotic expansion of Weil-petersson volume polynomials. Journal of Mathematical Physics, 63(4):043502, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McDonald:2022:CEN

[987] Edward McDonald and Raphaël Ponge. Cwikel estimates and negative eigenvalues of Schrödinger operators on noncommutative tori. Journal of Mathematical Physics, 63(4):043503, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deHoop:2022:PNI

[988] Maarten V. de Hoop and Alexei Iantchenko. Publisher’s note: “Inverse problem for the Rayleigh system with spectral data” [J. Math. Phys. 63, 031505 (2022)]. Journal of Mathematical Physics, 63(4):049901, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [937].

Dyatlov:2022:ISI

[989] Semyon Dyatlov, Svetlana Jitomirskaya, and Zeev Rudnick. Introduction to the special issue: In memory of Jean Bourgain. Journal of Mathematical Physics, 63(5):050401, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/5/050401/2845918/Introduction-to-the-Special-
Issue-In-memory-of.

Brennecke:2022:LES

[990] Christian Brennecke. The low energy spectrum of trapped bosons in the Gross–Pitaevskii regime. Journal of Mathematical Physics, 63(5): 051101, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
5/051101/2845929/The-low-energy-spectrum-of-trapped-bosons-
in-the.

Bulicek:2022:EAS

[991] Miroslav Bulíček, Ansgar Jüngel, Milan Pokorný, and Nicola Zamponi. Existence analysis of a stationary compressible fluid model for heat-conducting and chemically reacting mixtures. Journal of Mathematical Physics, 63(5):051501, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/051501/2845981/Existence-analysis-
of-a-stationary-compressible.

Chen:2022:SSC

[992] Hong Chen, Ziqi Wan, and Xin Zhong. Strong solutions to the 2D Cauchy problem of compressible non-isothermal nematic liquid crystal flows with vacuum at infinity. Journal of Mathematical Physics, 63(5):051502, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
051502/2845976/Strong-solutions-to-the-2D-Cauchy-problem-of.

An:2022:LRI

[993] Xinliang An, Haoyang Chen, and Silu Yin. Low regularity ill-posedness for non-strictly hyperbolic systems in three dimensions. Journal of Mathematical Physics, 63(5):051503, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/051503/2845944/Low-regularity-ill-
posedness-for-non-strictly.

Ma:2022:LWP

[994] Jieqiong Ma, Ruxu Lian, and Qingcun Zeng. Local well-posedness of strong solution to a climate dynamic model with phase transformation of water vapor. Journal of Mathematical Physics, 63(5):051504, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
051504/2845964/Local-well-posedness-of-strong-solution-to-a.

Guo:2022:SSH

[995] Yuli Guo, Weiguo Zhang, and Leilei Liu. The N-soliton solution to the higher-order nonlinear Schrödinger–Maxwell–Bloch system via the Riemann–Hilbert approach. Journal of Mathematical Physics, 63(5): 051505, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
5/051505/2845973/The-N-soliton-solution-to-the-higher-order.

Xiao:2022:CEC

[996] Yingying Xiao, Chuanxi Zhu, and Jianhua Chen. Combined effects of concave and convex nonlinearities for the generalized Chern– Simons–Schrödinger systems with steep potential well and 1 < p < 2 < q < 6. Journal of Mathematical Physics, 63(5):051506, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
051506/2845937/Combined-effects-of-concave-and-convex.

Ueda:2022:ASY

[997] Mamoru Ueda. Affine super Yangians and rectangular W -superalgebras. Journal of Mathematical Physics, 63(5):051701, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/5/051701/2845986/Affine-super-
Yangians-and-rectangular-W.

Peyghan:2022:SSH

[998] Esmaeil Peyghan and Leila Nourmohammadifar. Skew-symmetric homalgebroids and their representation theory. Journal of Mathematical Physics, 63(5):051702, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/051702/2845979/Skew-symmetric-hom-
algebroids-and-their.

Das:2022:DAR

[999] Apurba Das and Satyendra Kumar Mishra. The L_{∞}-deformations of associative Rota–Baxter algebras and homotopy Rota–Baxter operators. Journal of Mathematical Physics, 63(5):051703, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
051703/2845914/The-L-deformations-of-associative-Rota-Baxter.

Bru:2022:CDS

[1000] J.-B. Bru and W. de Siqueira Pedra. Classical dynamics from selfconsistency equations in quantum mechanics. Journal of Mathematical Physics, 63(5):052101, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/052101/2845920/Classical-dynamics-
from-self-consistency-equations.

Bon:2022:SSD

[1001] Enguerrand Lavigne Bon. Semiclassical spectrum of the Dirichlet–Pauli operator on an annulus. Journal of Mathematical Physics, 63(5):052102, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
052102/2845933/Semiclassical-spectrum-of-the-Dirichlet-Pauli.

Le:2022:ASU

[1002] Dai-Nam Le and Van-Hoang Le. Algebraic structure underlying spherical, parabolic, and prolate spheroidal bases of the ninedimensional MICZ–Kepler problem. Journal of Mathematical Physics, 63(5):052103, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/5/052103/2845954/Algebraic-structure-underlying-
spherical-parabolic.

Marwah:2022:UCB

[1003] Ashutosh Marwah and Frédéric Dupuis. Uniform continuity bound for sandwiched Rényi conditional entropy. Journal of Mathematical Physics, 63(5):052201, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/052201/2845950/Uniform-continuity-
bound-for-sandwiched-Renyi.

Koh:2022:FBI

[1004] Dax Enshan Koh, Guoming Wang, Peter D. Johnson, and Yudong Cao. Foundations for Bayesian inference with engineered likelihood functions for robust amplitude estimation. Journal of Mathematical Physics, 63(5): 052202, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
5/052202/2845960/Foundations-for-Bayesian-inference-with-engineered.

Wang:2022:REC

[1005] Jinzhao Wang and Henrik Wilming. Revisiting the equality conditions of the data-processing inequality for the sandwiched Rényi divergence. Journal of Mathematical Physics, 63(5):052203, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic).  URL https://pubs.aip.org/aip/jmp/article/63/
5/052203/2845970/Revisiting-the-equality-conditions-of-the-
data.

Araneda:2022:SDY

[1006] Bernardo Araneda. On self-dual Yang–Mills fields on special complex surfaces. Journal of Mathematical Physics, 63(5):052501, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic).  URL https://pubs.aip.org/aip/jmp/article/63/
5/052501/2845995/On-self-dual-Yang-Mills-fields-on-special-
complex.

Mohamed:2022:ACS

[1007] Mariem Magdy Ali Mohamed and Juan A. Valiente Kroon. Asymptotic charges for spin-1 and spin-2 fields at the critical sets of null infinity. Journal of Mathematical Physics, 63(5):052502, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
052502/2846010/Asymptotic-charges-for-spin-1-and-spin-2-fields-
at.

An:2022:NAF

[1008] Zhongshan An and Lan-Hsuan Huang. New asymptotically flat static vacuum metrics with near Euclidean boundary data. Journal of Mathematical Physics, 63(5):052503, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/5/052503/2846008/New-asymptotically-
flat-static-vacuum-metrics-with.

Xu:2022:AAD

[1009] Dongmei Xu and Fuzhi Li. Asymptotically autonomous dynamics for non-autonomous stochastic 2D g-Navier–Stokes equation in regular spaces. Journal of Mathematical Physics, 63(5):052701, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
052701/2846001/Asymptotically-autonomous-dynamics-for-non.

Chang:2022:CDB

[1010] Qingquan Chang and Dandan Li. Continuity of dynamical behaviors for strongly damped wave equations with perturbation. Journal of Mathematical Physics, 63(5):052702, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/5/052702/2846005/Continuity-of-dynamical-
behaviors-for-strongly.

Calogero:2022:SIV

[1011] F. Calogero and F. Payandeh. Solution of the initial-value problem of first-order nonlinear recursions with homogeneous right-hand sides. Journal of Mathematical Physics, 63(5):052703, May 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/5/052703/2845998/Solution-
of-the-initial-value-problem-of-first.

Turova:2022:EDC

[1012] Tatyana S. Turova. Exponential decay of correlations in the onedimensional Coulomb gas ensembles. Journal of Mathematical Physics, 63(5):053301, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/5/053301/2846004/Exponential-decay-of-correlations-
in-the-one.

Goto:2022:NTP

[1013] Shin itiro Goto. Nonequilibrium thermodynamic process with hysteresis and metastable states — a contact Hamiltonian with unstable and stable segments of a Legendre submanifold. Journal of Mathematical Physics, 63(5):053302, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/5/053302/2846031/Nonequilibrium-
thermodynamic-process-with.

Mograby:2022:SDS

[1014] Gamal Mograby, Radhakrishnan Balu, Kasso A. Okoudjou, and Alexander Teplyaev. Spectral decimation of a self-similar version of almost Mathieu-type operators. Journal of Mathematical Physics, 63(5):053501, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/5/053501/2846014/Spectral-decimation-of-
a-self-similar-version-of.

Guan:2022:MDH

[1015] Wenchuang Guan, Shen Wang, Weici Guo, and Jipeng Cheng. Modified DKP hierarchy as modified BKP hierarchy. Journal of Mathematical Physics, 63(5):053502, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/5/053502/2845999/Modified-DKP-hierarchy-
as-modified-BKP-hierarchy.

Cipolloni:2022:FSN

[1016] Giorgio Cipolloni. Fluctuations in the spectrum of non-Hermitian i.i.d. matrices. Journal of Mathematical Physics, 63(5):053503, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
053503/2845994/Fluctuations-in-the-spectrum-of-non-Hermitian-
i-i.

Khalkhali:2022:SSD

[1017] Masoud Khalkhali and Nathan Pagliaroli. Spectral statistics of Dirac ensembles. Journal of Mathematical Physics, 63(5):053504, May 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/5/
053504/2845996/Spectral-statistics-of-Dirac-ensembles.

Liu:2022:ASL

[1018] Qun Liu and Daqing Jiang. Analysis of a stochastic logistic model with diffusion and Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 63(5):053505, May 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/5/053505/2846024/Analysis-
of-a-stochastic-logistic-model-with.

Lewin:2022:CRG

[1019] Mathieu Lewin. Coulomb and Riesz gases: The known and the unknown. Journal of Mathematical Physics, 63(6):061101, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/6/061101/2845978/Coulomb-
and-Riesz-gases-The-known-and-the-unknown. Special collection in

honor of Freeman Dyson.

Bossmann:2022:LES

[1020] Lea Boßmann. Low-energy spectrum and dynamics of the weakly interacting Bose gas. Journal of Mathematical Physics, 63(6):061102, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061102/2845967/Low-energy-spectrum-and-dynamics-of-the-weakly.

Nam:2022:DBG

[1021] Phan Thành Nam, Julien Ricaud, and Arnaud Triay. Dilute Bose gas with three-body interaction: Recent results and open questions. Journal of Mathematical Physics, 63(6):061103, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061103/2845931/Dilute-Bose-gas-with-three-body-interaction-
Recent.

Bershtein:2022:NSV

[1022] Mikhail Bershtein and Angelina Vargulevich. NSR singular vectors from Uglov polynomials. Journal of Mathematical Physics, 63(6):061706, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061706/2846029/NSR-singular-vectors-from-Uglov-polynomials.

Chen:2022:LES

[1023] Xiao-Ping Chen and Chun-Lei Tang. Least energy sign-changing solutions for Kirchhoff-type problems with potential well. Journal of Mathematical Physics, 63(6):061501, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/6/061501/2845958/Least-energy-sign-changing-
solutions-for-Kirchhoff.

Ye:2022:BWS

[1024] Hailong Ye and Chunhua Jin. Bounded weak and strong time periodic solutions to a three-dimensional chemotaxis-Stokes model with porous medium diffusion. Journal of Mathematical Physics, 63(6):061502, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061502/2845941/Bounded-weak-and-strong-time-periodic-solutions-
to.

Yuan:2022:GWP

[1025] Wen-Shuo Yuan and Bin Ge. Global well-posedness for pseudoparabolic p-Laplacian equation with singular potential and logarithmic nonlinearity. Journal of Mathematical Physics, 63(6):061503, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061503/2845949/Global-well-posedness-for-pseudo-parabolic-p.

Liu:2022:ENS

[1026] Chuangye Liu and Xiaolong Yang. Existence of normalized solutions for semilinear elliptic systems with potential. Journal of Mathematical Physics, 63(6):061504, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/6/061504/2845962/Existence-
of-normalized-solutions-for-semilinear.

Zuo:2022:KTD

[1027] Jiabin Zuo and Juliana Honda Lopes. The Kirchhoff-type diffusion problem driven by a magnetic fractional Laplace operator. Journal of Mathematical Physics, 63(6):061505, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/6/061505/2845917/The-Kirchhoff-type-
diffusion-problem-driven-by-a.

Wang:2022:ILN

[1028] Jiawei Wang. Incompressible limit of nonisentropic Hookean elastodynamics. Journal of Mathematical Physics, 63(6):061506, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061506/2845926/Incompressible-limit-of-nonisentropic-Hookean.

Wang:2022:WSC

[1029] Shu Wang and Yabo Ren. Weak solutions to the Cauchy problem of the time-dependent Thomas–Fermi equations. Journal of Mathematical Physics, 63(6):061507, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/6/061507/2845969/Weak-solutions-
to-the-Cauchy-problem-of-the-time.

Cao:2022:VPP

[1030] Daomin Cao, Guolin Qin, and Changjun Zou. Vortex patch problem for steady lake equation. Journal of Mathematical Physics, 63(6):061508, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061508/2845915/Vortex-patch-problem-for-steady-lake-equation.

Zheng:2022:VTB

[1031] Xiangcheng Zheng, Yiqun Li, and Hong Wang. A viscoelastic Timoshenko beam: Model development, analysis, and investigation. Journal of Mathematical Physics, 63(6):061509, June 2022. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/6/061509/2845922/A-viscoelastic-
Timoshenko-beam-Model-development.

Xie:2022:FMS

[1032] Qiang Xie, Jiancai Sun, and Hengyun Yang. U(h)-free modules over the super-Galilean conformal algebras. Journal of Mathematical Physics, 63(6):061701, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/6/061701/2845930/U-h-free-
modules-over-the-super-Galilean-conformal.

Bisbo:2022:BID

[1033] Asmus K. Bisbo and Joris Van der Jeugt. Bases for infinite dimensional simple ≀∫√(∞|∈\)-modules respecting the branching

≀∫√(∞|∈\) ⊃ }↕(\). Journal of Mathematical Physics, 63(6):061702,

June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061702/2845936/Bases-for-infinite-dimensional-simple-osp-1-
2n.

Datt:2022:ECS

[1034] Gopal Datt and Bhawna Bansal Gupta. Essential commutativity and spectral properties of slant Hankel operators over Lebesgue spaces. Journal of Mathematical Physics, 63(6):061703, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/6/061703/2845946/Essential-
commutativity-and-spectral-properties-of.

Bergeron:2022:RSA

[1035] Geoffroy Bergeron, Julien Gaboriaud, Luc Vinet, and Alexei Zhedanov. The rational Sklyanin algebra and the Wilson and para-Racah polynomials. Journal of Mathematical Physics, 63(6):061704, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
061704/2845953/The-rational-Sklyanin-algebra-and-the-Wilson-
and.

Edelman:2022:CDC

[1036] Alan Edelman and Sungwoo Jeong. On the Cartan decomposition for classical random matrix ensembles. Journal of Mathematical Physics, 63 (6):061705, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
6/061705/2845972/On-the-Cartan-decomposition-for-classical-

random. Special collection in honor of Freeman Dyson.

Aza:2022:TIS

[1037] N. J. B. Aza, L. C. P. A. M. Müssnich, and A. F. Reyes-Lega. A Z_{2}-topological index as a Z_{2}-state index. Journal of Mathematical Physics, 63(6):061901, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/6/061901/2846036/A-Z2-topological-index-as-a-Z2-
state-index.

Benoist:2022:APT

[1038] Tristan Benoist, Martin Fraas, and Jürg Fröhlich. The appearance of particle tracks in detectors. II. The semi-classical realm. Journal of Mathematical Physics, 63(6):062101, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/6/062101/2846049/The-appearance-
of-particle-tracks-in-detectors-II.

Bluhm:2022:TNA

[1039] Andreas Bluhm and Ion Nechita. A tensor norm approach to quantum compatibility. Journal of Mathematical Physics, 63(6):062201, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/63/6/062201/2846040/A-
tensor-norm-approach-to-quantum-compatibility.

Girotti:2022:LDC

[1040] Federico Girotti, Merlijn van Horssen, Raffaella Carbone, and Mădălin Guţă. Large deviations, central limit, and dynamical phase transitions in the atom maser. Journal of Mathematical Physics, 63 (6):062202, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/6/062202/2846065/Large-deviations-central-limit-
and-dynamical-phase.

Carlen:2022:TIA

[1041] Eric A. Carlen and Elliott H. Lieb. A trace inequality of Ando, Hiai, and Okubo and a monotonicity property of the Golden–Thompson inequality. Journal of Mathematical Physics, 63(6):062203, June 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/63/6/062203/2846030/A-
trace-inequality-of-Ando-Hiai-and-Okubo-and-a. Special collec-

tion in honor of Freeman Dyson.

Chaubey:2022:CLM

[1042] S. K. Chaubey and Y. J. Suh. Characterizations of Lorentzian manifolds. Journal of Mathematical Physics, 63(6):062501, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
062501/2846013/Characterizations-of-Lorentzian-manifolds.

Fernandes:2022:PCC

[1043] Antonio Carlos Fernandes, Luis Fernando Mello, Lucas Ruiz dos Santos, and Claudio Vidal. Planar central configurations of six bodies. Journal of Mathematical Physics, 63(6):062701, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/6/062701/2846044/Planar-central-
configurations-of-six-bodies.

Liu:2022:DIN

[1044] Yang Liu, Shan Ma, and Chunyou Sun. Dynamics for 2D incompressible Navier–Stokes flow coupled with time-dependent Darcy flow. Journal of Mathematical Physics, 63(6):062702, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/6/062702/2846019/Dynamics-
for-2D-incompressible-Navier-Stokes-flow.

Figotin:2022:ATM

[1045] Alexander Figotin. Analytic theory of multicavity klystrons. Journal of Mathematical Physics, 63(6):062703, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/6/062703/2846071/Analytic-
theory-of-multicavity-klystrons.

Yang:2022:SPM

[1046] Min Yang and Guanggan Chen. Stability of the Poincaré maps for a stochastic fast–slow system. Journal of Mathematical Physics, 63(6):062704, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/6/062704/2846054/Stability-of-the-Poincare-maps-
for-a-stochastic.

Sanchez-Cerritos:2022:EPR

[1047] Juan Manuel Sánchez-Cerritos, Liang Ding, and Jinlong Wei. Equilibrium points in restricted problems on S^{2} and H^{2}. Journal of Mathematical Physics, 63(6):062705, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/6/062705/2846076/Equilibrium-points-
in-restricted-problems-on-S2.

Burdzy:2022:FAR

[1048] Krzysztof Burdzy, Mauricio Duarte, Carl-Erik Gauthier, C. Robin Graham, and Jaime San Martin. Fermi acceleration in rotating drums. Journal of Mathematical Physics, 63(6):062706, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
062706/2846022/Fermi-acceleration-in-rotating-drums.

Sakajo:2022:QPV

[1049] Takashi Sakajo and Vikas S. Krishnamurthy. Quantized point vortex equilibria in a one-way interaction model with a Liouville-type background vorticity on a curved torus. Journal of Mathematical Physics, 63 (6):063101, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/6/063101/2846058/Quantized-point-vortex-equilibria-in-a-
one-way.

Koukiou:2022:EFG

[1050] Flora Koukiou. Entropy and freezing in Gaussian models. Journal of Mathematical Physics, 63(6):063301, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/6/063301/2846034/Entropy-
and-freezing-in-Gaussian-models.

Saluto:2022:EFD

[1051] Lidia Saluto, Liliana Restuccia, and David Jou. Electric field dependence of thermal conductivity in bulk systems and nanosystems with charged mobile defects. Journal of Mathematical Physics, 63(6):063302, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
063302/2846097/Electric-field-dependence-of-thermal-conductivity.

Barhoumi:2022:ITC

[1052] Ahmad Barhoumi, Pavel Bleher, Alfredo Deaño, and Maxim Yattselev. Investigation of the two-cut phase region in the complex cubic ensemble of random matrices. Journal of Mathematical Physics, 63(6): 063303, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
6/063303/2846121/Investigation-of-the-two-cut-phase-region-

in-the. Special collection in honor of Freeman Dyson.

Ramos:2022:LIA

[1053] A. D. Ramos and C. D. B. da Silva. Locality of the interaction affects dynamics in probabilistic cellular automata. Journal of Mathematical Physics, 63(6):063304, June 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/6/063304/2846108/Locality-
of-the-interaction-affects-dynamics-in.

Sasaki:2022:ESD

[1054] Ryu Sasaki. Exactly solvable discrete time birth and death processes. Journal of Mathematical Physics, 63(6):063305, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/6/063305/2846104/Exactly-
solvable-discrete-time-birth-and-death.

Giscard:2022:EIS

[1055] P.-L. Giscard and A. Tamar. Elementary integral series for Heun functions: Application to black-hole perturbation theory. Journal of Mathematical Physics, 63(6):063501, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/6/063501/2846111/Elementary-integral-series-
for-Heun-functions.

Cartin:2022:KRG

[1056] Daniel Cartin. Knotted 4-regular graphs: Polynomial invariants and the Pachner moves. Journal of Mathematical Physics, 63(6):063502, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/6/
063502/2846119/Knotted-4-regular-graphs-Polynomial-invariants-
and.

Khan:2022:EOF

[1057] Sohail A. Khan, T. Hayat, A. Alsaedi, and B. Ahmad. Entropy optimized flow of hydromagnetic Reiner–Philippoff fluid over a stretching surface. Journal of Mathematical Physics, 63(6):063503, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/6/063503/2846117/Entropy-
optimized-flow-of-hydromagnetic-Reiner.

Yu:2022:ASP

[1058] Jianduo Yu, Chuanzhong Li, Mengkun Zhu, and Yang Chen. Asymptotics for a singularly perturbed GUE, Painlevé III, double-confluent Heun equations, and small eigenvalues. Journal of Mathematical Physics, 63(6):063504, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/6/063504/2846113/Asymptotics-for-a-singularly-perturbed-
GUE.

Pastur:2022:EDL

[1059] L. Pastur. Eigenvalue distribution of large random matrices arising in deep neural networks: Orthogonal case. Journal of Mathematical Physics, 63(6):063505, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/6/063505/2846100/Eigenvalue-distribution-of-

large-random-matrices. Special collection in honor of Freeman Dyson.

Adami:2022:ESR

[1060] Riccardo Adami, Filippo Boni, Raffaele Carlone, and Lorenzo Tentarelli. Existence, structure, and robustness of ground states of a NLSE in 3D with a point defect. Journal of Mathematical Physics, 63(7):071501, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071501/2843527/Existence-structure-and-robustness-of-ground.

Moreno:2022:MAP

[1061] M. Moreno, P. Rioseco, and H. Van Den Bosch. Mixing in anharmonic potential well. Journal of Mathematical Physics, 63(7):071502, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071502/2843589/Mixing-in-anharmonic-potential-well.

Cassano:2022:DCO

[1062] Biagio Cassano, Matteo Gallone, and Fabio Pizzichillo. Dirac–Coulomb operators with infinite mass boundary conditions in sectors. Journal of Mathematical Physics, 63(7):071503, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/7/071503/2843559/Dirac-Coulomb-operators-
with-infinite-mass.

Lei:2022:LBR

[1063] Zhoutong Lei and Zhiqiang Shao. The limit behavior of Riemann solutions to the Euler equations of compressible fluid flow for the modified Chaplygin gas. Journal of Mathematical Physics, 63(7):071504, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071504/2843598/The-limit-behavior-of-Riemann-solutions-to-the.

Zhu:2022:WBP

[1064] Mingxuan Zhu and Zhenda Li. Wave breaking phenomenon for the generalized FORQ equation. Journal of Mathematical Physics, 63(7):071505, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071505/2843532/Wave-breaking-phenomenon-for-the-generalized-
FORQ.

Han:2022:ARS

[1065] Pengju Han and Yang Chen. Asymptotic relations for semiclassical Laguerre orthogonal polynomials and the associated Hankel determinants. Journal of Mathematical Physics, 63(7):071506, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071506/2843567/Asymptotic-relations-for-semi-classical-Laguerre.

Guo:2022:RCL

[1066] Helin Guo and Lingling Zhao. On the regularity criteria for liquid crystal flows involving the gradient of one velocity component. Journal of Mathematical Physics, 63(7):071507, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/7/071507/2843582/On-the-regularity-criteria-
for-liquid-crystal.

Zhang:2022:CFB

[1067] Fan Zhang. Convexity of the free boundary for an axially symmetric rotational cavity flow problem. Journal of Mathematical Physics, 63(7):071508, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/7/071508/2843592/Convexity-of-the-free-boundary-
for-an-axially.

Huang:2022:AIY

[1068] Teng Huang. The analysis of inhomogeneous Yang–Mills connections on closed Riemannian manifold. Journal of Mathematical Physics, 63(7): 071509, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
7/071509/2843552/The-analysis-of-inhomogeneous-Yang-Mills.

Boyallian:2022:FVL

[1069] C. Boyallian and J. Guzman. Formal vertex laws associated to Lie conformal algebras. Journal of Mathematical Physics, 63(7):071701, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071701/2843544/Formal-vertex-laws-associated-to-Lie-conformal.

Olgiati:2022:RFB

[1070] Alessandro Olgiati. Reduced fluctuations for bosons in a double well. Journal of Mathematical Physics, 63(7):071901, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/7/071901/2843522/Reduced-
fluctuations-for-bosons-in-a-double-well.

Basti:2022:SOU

[1071] Giulia Basti. A second order upper bound on the ground state energy of a Bose gas beyond the Gross–Pitaevskii regime. Journal of Mathematical Physics, 63(7):071902, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/7/071902/2843530/A-second-order-upper-bound-

on-the-ground-state. Special collection in honor of Freeman Dyson.

Nam:2022:GSE

[1072] Phan Thành Nam, Julien Ricaud, and Arnaud Triay. Ground state energy of the low density Bose gas with three-body interactions. Journal of Mathematical Physics, 63(7):071903, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071903/2843539/Ground-state-energy-of-the-low-density-Bose-

gas. Special collection in honor of Freeman Dyson.

Garg:2022:DCM

[1073] Anupam Garg. The discrete Chebyshev–Meckler–Mermin–Schwarz polynomials and spin algebra. Journal of Mathematical Physics, 63(7):072101, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/7/072101/2843548/The-discrete-Chebyshev-Meckler-
Mermin-Schwarz.

Marcelli:2022:CSA

[1074] Giovanna Marcelli and Domenico Monaco. From charge to spin: Analogies and differences in quantum transport coefficients. Journal of Mathematical Physics, 63(7):072102, July 2022. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/7/072102/2843556/From-charge-to-
spin-Analogies-and-differences-in.

Doring:2022:CFT

[1075] Andreas Döring and Markus Frembs. Contextuality and the fundamental theorems of quantum mechanics. Journal of Mathematical Physics, 63 (7):072103, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
7/072103/2843510/Contextuality-and-the-fundamental-theorems-
of.

Mead:2022:NTP

[1076] Lawrence R. Mead, Sungwook Lee, and David Garfinkle. A non-trivial PT-symmetric continuum Hamiltonian and its eigenstates and eigenvalues. Journal of Mathematical Physics, 63(7):072104, July 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/63/7/072104/2843566/A-
non-trivial-PT-symmetric-continuum-Hamiltonian. See erratum

[1133].

Lonigro:2022:GSB

[1077] Davide Lonigro. Generalized spin-boson models with non-normalizable form factors. Journal of Mathematical Physics, 63(7):072105, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
072105/2843570/Generalized-spin-boson-models-with-non.

Buyukasik:2022:ETE

[1078] Şirin A. Büyükaşık and Zehra Çayiç. Exact time-evolution of a generalized two-dimensional quantum parametric oscillator in the presence of time-variable magnetic and electric fields. Journal of Mathematical Physics, 63(7):072106, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/7/072106/2843584/Exact-time-evolution-of-a-generalized-
two.

Lasser:2022:VVA

[1079] Caroline Lasser and Chunmei Su. Various variational approximations of quantum dynamics. Journal of Mathematical Physics, 63(7):072107, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
072107/2843668/Various-variational-approximations-of-quantum.

Svetlichnyy:2022:DQC

[1080] Pavel Svetlichnyy and T. A. B. Kennedy. Decay of quantum conditional mutual information for purely generated finitely correlated states. Journal of Mathematical Physics, 63(7):072201, July 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/7/072201/2843699/Decay-of-
quantum-conditional-mutual-information.

Svampa:2022:AAQ

[1081] Ilaria Svampa, Stefano Mancini, and Andreas Winter. An approach to p-adic qubits from irreducible representations of SO(3)_{p}. Journal of Mathematical Physics, 63(7):072202, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/7/072202/2843682/An-approach-to-
p-adic-qubits-from-irreducible.

Seo:2022:QRD

[1082] Yuki Seo. The quantum ♮_{α}-Rényi divergence of real order. Journal of Mathematical Physics, 63(7):072203, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/7/072203/2843692/The-quantum-Renyi-
divergence-of-real-order. See erratum [1487].

Hasenohrl:2022:QCD

[1083] Markus Hasenöhrl and Matthias C. Caro. Quantum and classical dynamical semigroups of superchannels and semicausal channels. Journal of Mathematical Physics, 63(7):072204, July 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/7/072204/2843710/Quantum-
and-classical-dynamical-semigroups-of.

Huguet:2022:RLO

[1084] E. Huguet, J. Queva, and J. Renaud. Restriction of Laplace operator on one-forms: From R^{n+2} and R^{n+1} ambient spaces to embedded (A)dS_{n} submanifolds. Journal of Mathematical Physics, 63(7): 072301, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/7/072301/2843658/Restriction-of-Laplace-operator-on-one-
forms-From.

Guo:2022:OMG

[1085] Shuyuan Guo, Gang Meng, Ping Yan, and Meirong Zhang. Optimal maximal gaps of Dirichlet eigenvalues of Sturm–Liouville operators. Journal of Mathematical Physics, 63(7):072701, July 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/7/072701/2843661/Optimal-
maximal-gaps-of-Dirichlet-eigenvalues-of.

Zhu:2022:CLH

[1086] Pengxian Zhu and Qigui Yang. Chaos of the 2D linear hyperbolic equation with general van der Pol type boundary condition. Journal of Mathematical Physics, 63(7):072702, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/7/072702/2843714/Chaos-of-the-2D-linear-
hyperbolic-equation-with.

Wen:2022:RAS

[1087] Lan Wen and Lu Yang. Random attractor for stochastic non-autonomous Berger equation with critical exponent. Journal of Mathematical Physics, 63(7):072703, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/7/072703/2843675/Random-attractor-for-stochastic-
non-autonomous.

Bodineau:2022:CED

[1088] Thierry Bodineau, Isabelle Gallagher, Laure Saint-Raymond, and Sergio Simonella. Cluster expansion for a dilute hard sphere gas dynamics. Journal of Mathematical Physics, 63(7):073301, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/7/073301/2843705/Cluster-
expansion-for-a-dilute-hard-sphere-gas.

Brennecke:2022:RSF

[1089] Christian Brennecke and Horng-Tzer Yau. The replica symmetric formula for the SK model revisited. Journal of Mathematical Physics, 63(7): 073302, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
7/073302/2843656/The-replica-symmetric-formula-for-the-SK-model.

Katkov:2022:MMH

[1090] Mikhail Katkov, Michelangelo Naim, Antonios Georgiou, and Misha Tsodyks. Mathematical models of human memory. Journal of Mathematical Physics, 63(7):073303, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/7/073303/2843660/Mathematical-models-of-

human-memory. Special collection in honor of Freeman Dyson. See erratum [1374].

Vanchurin:2022:DEP

[1091] Vitaly Vanchurin. Differential equation for partition functions and a duality pseudo-forest. Journal of Mathematical Physics, 63(7):073501, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
073501/2843666/Differential-equation-for-partition-functions-
and.

Zhang:2022:QHG

[1092] Senlin Zhang and Shuanhong Wang. Quasitriangular Hopf groupquasialgebras and generalized quantum Yang–Baxter equations. Journal of Mathematical Physics, 63(7):073502, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/7/073502/2843671/Quasitriangular-
Hopf-group-quasialgebras-and.

DelVecchio:2022:LIB

[1093] Simone Del Vecchio, Jürg Fröhlich, Alessandro Pizzo, and Stefano Rossi. Local iterative block-diagonalization of gapped Hamiltonians: a new tool in singular perturbation theory. Journal of Mathematical Physics, 63(7): 073503, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
7/073503/2843679/Local-iterative-block-diagonalization-of-gapped.

Special collection in honor of Freeman Dyson.

Garcia-Pelaez:2022:LPT

[1094] D. García-Peláez, C. S. López-Monsalvo, and A. Rubio Ponce. Light propagation through optical media using metric contact geometry. Journal of Mathematical Physics, 63(7):073504, July 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/7/073504/2843741/Light-propagation-
through-optical-media-using.

Haah:2022:ECQ

[1095] Jeongwan Haah. Erratum: “Clifford quantum cellular automata: Trivial group in 2D and Witt group in 3D” [J. Math. Phys. 62, 092202 (2021)]. Journal of Mathematical Physics, 63(7):079901, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/7/079901/2843740/Erratum-
Clifford-quantum-cellular-automata-Trivial. See [743].

Benedikter:2022:EDI

[1096] Niels Benedikter. Effective dynamics of interacting fermions from semiclassical theory to the random phase approximation. Journal of Mathematical Physics, 63(8):081101, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/8/081101/2845977/Effective-dynamics-
of-interacting-fermions-from.

Desiraju:2022:PCC

[1097] Harini Desiraju. Painlevé/ CFT correspondence on a torus. Journal of Mathematical Physics, 63(8):081102, August 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/8/081102/2845980/Painleve-
CFT-correspondence-on-a-torus.

Shen:2022:SPA

[1098] Hao Shen. A stochastic PDE approach to large N problems in quantum field theory: a survey. Journal of Mathematical Physics, 63(8):081103, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081103/2845992/A-stochastic-PDE-approach-to-large-N-problems-
in.

Bian:2022:STS

[1099] Dongfen Bian and Xueke Pu. Stability threshold for 2D shear flows of the Boussinesq system near Couette. Journal of Mathematical Physics, 63(8): 081501, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/081501/2845947/Stability-threshold-for-2D-shear-flows-of-
the.

Wu:2022:IEP

[1100] Weijun Wu, Fujun Zhou, and Yongsheng Li. Incompressible Euler– Poisson limit of the Vlasov–Poisson–Boltzmann system. Journal of Mathematical Physics, 63(8):081502, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/8/081502/2845990/Incompressible-
Euler-Poisson-limit-of-the-Vlasov.

Karaki:2022:SKF

[1101] Zeinab Karaki. Study of the Kramers–Fokker–Planck quadratic operator with a constant magnetic field. Journal of Mathematical Physics, 63(8):

081503, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/081503/2845982/Study-of-the-Kramers-Fokker-Planck-quadratic.

Wang:2022:BFB

[1102] Ya-Guang Wang and Shi-Yong Zhu. On back flow of boundary layers in two-dimensional unsteady incompressible heat conducting flow. Journal of Mathematical Physics, 63(8):081504, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/8/081504/2845942/On-back-
flow-of-boundary-layers-in-two-dimensional.

Djellali:2022:STB

[1103] Fayssal Djellali and Soraya Labidi. On the stability of a thermodiffusion Bresse system. Journal of Mathematical Physics, 63(8):081505, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081505/2845971/On-the-stability-of-a-thermodiffusion-Bresse.

Choi:2022:ESG

[1104] Nari Choi. The existence of solutions for the gravitational Maxwell gauged O(3) model on compact surfaces. Journal of Mathematical Physics, 63(8):081506, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/8/081506/2845985/The-existence-of-solutions-for-
the-gravitational.

Zhang:2022:MCS

[1105] Weiqiang Zhang, Yanyun Wen, and Peihao Zhao. Multiplicity and concentration of solutions for fractional Kirchhoff–Choquard equation with critical growth. Journal of Mathematical Physics, 63(8):081507, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081507/2845963/Multiplicity-and-concentration-of-solutions-
for.

Bytsko:2022:OPR

[1106] Andrei Bytsko. On orthogonal projections related to representations of the Hecke algebra on a tensor space. Journal of Mathematical Physics, 63(8):081701, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/8/081701/2845939/On-orthogonal-projections-related-
to.

Nazarov:2022:QBS

[1107] Maxim Nazarov. Quantum Berezinian for a strange Lie superalgebra. Journal of Mathematical Physics, 63(8):081702, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/8/081702/2845945/Quantum-
Berezinian-for-a-strange-Lie-superalgebra.

Manogue:2022:ODS

[1108] Corinne A. Manogue, Tevian Dray, and Robert A. Wilson. Octions: an E_{8} description of the Standard Model. Journal of Mathematical Physics, 63(8):081703, August 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/8/081703/2845955/Octions-
An-E8-description-of-the-Standard-Model.

Jakobsen:2022:AVC

[1109] Hans Plesner Jakobsen. Algebras of variable coefficient quantized differential operators. Journal of Mathematical Physics, 63(8):081704, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/081704/2845968/Algebras-of-variable-coefficient-quantized.

Boyallian:2022:AQC

[1110] Carina Boyallian and Vanesa Meinardi. An approach to quantum conformal algebra. Journal of Mathematical Physics, 63(8):081705, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081705/2845974/An-approach-to-quantum-conformal-algebra.

Bullivant:2022:CCD

[1111] Alex Bullivant and Clement Delcamp. Crossing with the circle in Dijkgraaf–Witten theory and applications to topological phases of matter. Journal of Mathematical Physics, 63(8):081901, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081901/2845993/Crossing-with-the-circle-in-Dijkgraaf-Witten.

Rademacher:2022:DRV

[1112] Simone Rademacher. Dependent random variables in quantum dynamics. Journal of Mathematical Physics, 63(8):081902, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081902/2845935/Dependent-random-variables-in-quantum-dynamics.

Lemm:2022:BEC

[1113] Marius Lemm and Oliver Siebert. Bose–Einstein condensation on hyperbolic spaces. Journal of Mathematical Physics, 63(8):081903, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
081903/2845987/Bose-Einstein-condensation-on-hyperbolic-spaces.

Chong:2022:GTS

[1114] J. J. Chong, L. Lafleche, and C. Saffirio. Global-in-time semiclassical regularity for the Hartree–Fock equation. Journal of Mathematical Physics, 63(8):081904, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/8/081904/2846026/Global-in-time-
semiclassical-regularity-for-the.

Buchholz:2022:PCL

[1115] Detlev Buchholz. Proper condensates and long range order. Journal of Mathematical Physics, 63(8):081905, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/8/081905/2846032/Proper-condensates-
and-long-range-order.

Sunko:2022:ESS

[1116] D. K. Sunko. Evaluation and spanning sets of confluent Vandermonde forms. Journal of Mathematical Physics, 63(8):082101, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
082101/2846027/Evaluation-and-spanning-sets-of-confluent.

Alhaidari:2022:PAB

[1117] A. D. Alhaidari. Progressive approximation of bound states by finite series of square-integrable functions. Journal of Mathematical Physics, 63(8):082102, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/8/082102/2846051/Progressive-approximation-of-bound-
states-by.

Tlas:2022:CEI

[1118] T. Tlas. On the construction of Euclidean invariant and reflection positive measures on the cylindrical compactification of distributions. Journal of Mathematical Physics, 63(8):082301, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
082301/2846018/On-the-construction-of-Euclidean-invariant-and.

Viennot:2022:FSS

[1119] David Viennot. Fuzzy Schwarzschild (2 + 1)-spacetime. Journal of Mathematical Physics, 63(8):082302, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/8/082302/2846064/Fuzzy-Schwarzschild-2-1-
spacetime.

Minucci:2022:MSF

[1120] Marica Minucci, Rodrigo Panosso Macedo, and Juan A. Valiente Kroon. The Maxwell-scalar field system near spatial infinity. Journal of Mathematical Physics, 63(8):082501, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/8/082501/2846039/The-Maxwell-scalar-
field-system-near-spatial.

Ahn:2022:NLR

[1121] Hyunjin Ahn, Seung-Yeal Ha, and Jeongho Kim. Nonrelativistic limits of the relativistic Cucker–Smale model and its kinetic counterpart. Journal of Mathematical Physics, 63(8):082701, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
082701/2846033/Nonrelativistic-limits-of-the-relativistic-Cucker.

Baake:2022:FDP

[1122] Michael Baake, Franz Gähler, and Jan Mazáč. Fibonacci direct product variation tilings. Journal of Mathematical Physics, 63(8):082702, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/082702/2846055/Fibonacci-direct-product-variation-tilings.

Ge:2022:QPB

[1123] Chuanfang Ge, Jiansheng Geng, and Yingfei Yi. Quasi-periodic breathers in Newton’s cradle. Journal of Mathematical Physics, 63(8):082703, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
082703/2846047/Quasi-periodic-breathers-in-Newton-s-cradle.

Deng:2022:QHT

[1124] Yanxia Deng, Slim Ibrahim, and Lingjun Qian. Quasi-homogeneous two-body problem. Journal of Mathematical Physics, 63(8):082901, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/082901/2846059/Quasi-homogeneous-two-body-problem.

Krpensky:2022:NCA

[1125] A. Krpensky, V. Hruska, and M. Bednarik. A new class of approximate analytical solutions of the Pridmore–Brown equation. Journal of Mathematical Physics, 63(8):083101, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/8/083101/2846021/A-new-class-of-
approximate-analytical-solutions-of.

Goldstein:2022:SSO

[1126] S. Goldstein, J. L. Lebowitz, and E. R. Speer. Stationary states of the one-dimensional discrete-time facilitated symmetric exclusion process. Journal of Mathematical Physics, 63(8):083301, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
083301/2846043/Stationary-states-of-the-one-dimensional-discrete.

Special collection in honor of Freeman Dyson.

Wang:2022:SML

[1127] Wei Wang, Guangying Lv, and Jinglong Wei. Small mass limit in mean field theory for stochastic N particle system. Journal of Mathematical Physics, 63(8):083302, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/8/083302/2846095/Small-mass-limit-in-mean-field-
theory-for.

Pirjol:2022:GRS

[1128] Dan Pirjol. Growth rate of a stochastic growth process driven by an exponential Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 63(8):083303, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/8/083303/2846101/Growth-rate-of-
a-stochastic-growth-process-driven.

Guler:2022:GQE

[1129] Sinem Güler and Uday Chand De. Generalized quasi-Einstein metrics and applications on generalized Robertson–Walker spacetimes. Journal of Mathematical Physics, 63(8):083501, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
083501/2846096/Generalized-quasi-Einstein-metrics-and.

Baradaran:2022:KNV

[1130] Marzieh Baradaran and Pavel Exner. Kagome network with vertex coupling of a preferred orientation. Journal of Mathematical Physics, 63(8): 083502, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/8/083502/2846107/Kagome-network-with-vertex-coupling-of-
a-preferred.

Wei:2022:FGS

[1131] Jiao Wei, Xianguo Geng, Xin Wang, and Yunyun Zhai. Finite genus solutions of the generalized Merola–Ragnisco–Tu lattice hierarchy. Journal of Mathematical Physics, 63(8):083503, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
8/083503/2846103/Finite-genus-solutions-of-the-generalized-
Merola.

Naber:2022:BZF

[1132] M. G. Naber, B. M. Bruck, and S. E. Costello. The Bessel zeta function. Journal of Mathematical Physics, 63(8):083504, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/8/
083504/2846099/The-Bessel-zeta-function.

Mead:2022:ENT

[1133] Lawrence R. Mead, Sungwook Lee, and David Garfinkle. Erratum: A non-trivial PT-symmetric continuum Hamiltonian and its eigenstates and eigenvalues [J. Math. Phys. 63, 072104 (2022)]. Journal of Mathematical Physics, 63(8):089901, August 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/8/089901/2846094/Erratum-
A-non-trivial-PT-symmetric-continuum. See [1076].

Chevyrev:2022:SQY

[1134] Ilya Chevyrev. Stochastic quantization of Yang–Mills. Journal of Mathematical Physics, 63(9):091101, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/9/091101/2844235/Stochastic-quantization-
of-Yang-Mills.

Christodoulou:2022:DSF

[1135] Demetrios Christodoulou. On the development of shocks in fluids. Journal of Mathematical Physics, 63(9):091102, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091102/2844195/On-the-development-of-shocks-in-fluids.

Dai:2022:GWS

[1136] Feng Dai and Bin Liu. Global weak solutions in a three-dimensional two-species cancer invasion haptotaxis model without cell proliferation. Journal of Mathematical Physics, 63(9):091501, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091501/2844210/Global-weak-solutions-in-a-three-dimensional-
two.

Ge:2022:DPP

[1137] Bin Ge, Jin-Wei Zhao, and Wen-Shuo Yuan. On double-phase problems without any growth and Ambrosetti–Rabinowitz conditions. Journal of Mathematical Physics, 63(9):091502, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/091502/2844217/On-double-
phase-problems-without-any-growth-and.

Mezadek:2022:CVP

[1138] A. Kainane Mezadek and D. Rebah. A constrained variational problem for an existence theorem of steady vortex rings in Poiseuille flow. Journal of Mathematical Physics, 63(9):091503, September 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/63/9/091503/2844207/A-
constrained-variational-problem-for-an-existence.

Han:2022:DPD

[1139] Pigong Han, Keke Lei, Chenggang Liu, and Xuewen Wang. Decay properties for a density-dependent liquid crystal system. Journal of Mathematical Physics, 63(9):091504, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/9/091504/2844240/Decay-properties-
for-a-density-dependent-liquid.

Shimizu:2022:LWP

[1140] Ikkei Shimizu. Local well-posedness of the Landau–Lifshitz equation with helicity term. Journal of Mathematical Physics, 63(9):091505, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091505/2844231/Local-well-posedness-of-the-Landau-Lifshitz.

Ma:2022:IBV

[1141] Xuan Ma and Fuli He. The initial boundary value problem for the Vlasov–Poisson–Fokker–Planck system. Journal of Mathematical Physics, 63(9):091506, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/9/091506/2844205/The-initial-boundary-value-
problem-for-the-Vlasov.

Jin:2022:SRS

[1142] Hong Sung Jin, Minkyu Kwak, and Bataa Lkhagvasuren. Stability result for the surface quasi-geostropic equations with horizontal dissipation in anisotropic Sobolev space. Journal of Mathematical Physics, 63(9):091507, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/091507/2844201/Stability-
result-for-the-surface-quasi-geostropic.

Ren:2022:GEQ

[1143] Guoqiang Ren. Global existence of a quasilinear chemotaxis model with signal-dependent motility and indirect signal production mechanism. Journal of Mathematical Physics, 63(9):091508, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091508/2844228/Global-existence-of-a-quasilinear-chemotaxis-
model.

Li:2022:LEB

[1144] Huicong Li and Jingyu Li. Lifespan of effective boundary conditions for the heat equation. Journal of Mathematical Physics, 63(9):091509, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/091509/2844190/Lifespan-of-effective-boundary-conditions-
for-the.

Cui:2022:GSS

[1145] Na Cui and Hong-Rui Sun. Ground state solution for a nonlinear fractional magnetic Schrödinger equation with indefinite potential. Journal of Mathematical Physics, 63(9):091510, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091510/2844197/Ground-state-solution-for-a-nonlinear-fractional.

Tong:2022:GEA

[1146] Leilei Tong and Yi Xia. Global existence and the algebraic decay rate of the solution for the quantum Navier–Stokes–Poisson equations in R^{3}. Journal of Mathematical Physics, 63(9):091511, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091511/2844203/Global-existence-and-the-algebraic-decay-rate-
of.

Gokhale:2022:WZA

[1147] Soham Gokhale and Utpal Manna. Wong–Zakai approximations for the stochastic Landau–Lifshitz–Bloch equations. Journal of Mathematical Physics, 63(9):091512, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/9/091512/2844206/Wong-Zakai-approximations-for-
the-stochastic.

Bagnoli:2022:CFI

[1148] Lucia Bagnoli and Fabrizio Caselli. Classification of finite irreducible conformal modules for K′_{4}. Journal of Mathematical Physics, 63(9):091701, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/9/091701/2844208/Classification-
of-finite-irreducible-conformal.

Skrypnyk:2022:AGC

[1149] T. Skrypnyk. Anisotropic Z_{n}-graded classical r-matrix, deformed A_{n} Toda- and Gaudin-type models, and separation of variables. Journal of Mathematical Physics, 63(9):091702, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/091702/2844214/Anisotropic-
Zn-graded-classical-r-matrix-deformed.

Das:2022:CDW

[1150] Apurba Das. Cohomology and deformations of weighted Rota–Baxter operators. Journal of Mathematical Physics, 63(9):091703, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091703/2844224/Cohomology-and-deformations-of-weighted-Rota.

Aizawa:2022:IRG

[1151] N. Aizawa and S. Doi. Irreducible representations of Z_{2}^{2}-graded N = ∈ supersymmetry algebra and Z_{2}^{2}-graded supermechanics. Journal of Mathematical Physics, 63(9):091704, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/091704/2844243/Irreducible-
representations-of-Z22-graded-N-2.

Verdon:2022:CST

[1152] Dominic Verdon. A covariant Stinespring theorem. Journal of Mathematical Physics, 63(9):091705, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/63/9/091705/2844177/A-
covariant-Stinespring-theorem.

Barron:2022:RGV

[1153] Katrina Barron, Karina Batistelli, Florencia Orosz Hunziker, Veronika Pedić Tomić, and Gaywalee Yamskulna. On rationality of C-graded vertex algebras and applications to Weyl vertex algebras under conformal flow. Journal of Mathematical Physics, 63(9):091706, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091706/2844308/On-rationality-of-C-graded-vertex-algebras-and.

Friesecke:2022:TEW

[1154] Gero Friesecke and Benedikt R. Graswald. Two-electron wavefunctions are matrix product states with bond dimension three. Journal of Mathematical Physics, 63(9):091901, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/9/091901/2844337/Two-electron-wavefunctions-
are-matrix-product.

Ludewig:2022:LSG

[1155] Matthias Ludewig and Guo Chuan Thiang. Large-scale geometry obstructs localization. Journal of Mathematical Physics, 63(9):091902, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
091902/2844302/Large-scale-geometry-obstructs-localization.

Kapustin:2022:LNT

[1156] Anton Kapustin and Nikita Sopenko. Local Noether theorem for quantum lattice systems and topological invariants of gapped states. Journal of Mathematical Physics, 63(9):091903, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/091903/2844344/Local-Noether-theorem-for-quantum-lattice-
systems.

Anzaldo-Meneses:2022:SCO

[1157] A. Anzaldo-Meneses. Supercanonical coordinates for orthosymplectic evolution operators. Journal of Mathematical Physics, 63(9): 092101, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/9/092101/2844330/Supercanonical-coordinates-for-
orthosymplectic.

Matsuda:2022:CQG

[1158] Junichiro Matsuda. Classification of quantum graphs on M_{2} and their quantum automorphism groups. Journal of Mathematical Physics, 63(9):092201, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/9/092201/2844353/Classification-
of-quantum-graphs-on-M2-and-their.

Kye:2022:CMR

[1159] Seung-Hyeok Kye. Choi matrices revisited. Journal of Mathematical Physics, 63(9):092202, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/9/092202/2844297/Choi-matrices-revisited.

Vieira:2022:GUC

[1160] Vandenberg Lopes Vieira and Orlando Stanley Juriaans. On geometrically uniform codes and topological quantum MDS codes. Journal of Mathematical Physics, 63(9):092203, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/092203/2844316/On-geometrically-
uniform-codes-and-topological.

Boche:2022:SSQ

[1161] Holger Boche, Minglai Cai, Christian Deppe, Roberto Ferrara, and Moritz Wiese. Semantic security for quantum wiretap channels. Journal of Mathematical Physics, 63(9):092204, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/9/092204/2844303/Semantic-
security-for-quantum-wiretap-channels.

Dedushenko:2022:CWS

[1162] Mykola Dedushenko and Davide Gaiotto. Correlators on the wall and ∫↕_{\} spin chain. Journal of Mathematical Physics, 63(9):092301, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/092301/2844342/Correlators-on-the-wall-and-sln-spin-chain.

Iseppi:2022:GFB

[1163] Roberta A. Iseppi. The ghost fields and the BV extension for finite spectral triples. Journal of Mathematical Physics, 63(9):092302, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/092302/2844294/The-ghost-fields-and-the-BV-extension-for-
finite.

Juarez-Aubry:2022:SGC

[1164] Benito A. Juárez-Aubry and Sujoy K. Modak. Semiclassical gravity with a conformally covariant field in globally hyperbolic spacetimes. Journal of Mathematical Physics, 63(9):092303, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
092303/2844300/Semiclassical-gravity-with-a-conformally-covariant.

Borji:2022:PRT

[1165] Majdouline Borji and Christoph Kopper. Perturbative renormalization of φ_{4}^{4} theory on the half space R^{+} × R^{3} with flow equations. Journal of Mathematical Physics, 63(9):092304, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/092304/2844349/Perturbative-renormalization-of-44-theory-
on-the.

Bezdekova:2022:LPP

[1166] Barbora Bezděková, Volker Perlick, and Jiří Bičák. Light propagation in a plasma on an axially symmetric and stationary spacetime: Separability of the Hamilton–Jacobi equation and shadow. Journal of Mathematical Physics, 63(9):092501, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/9/092501/2844306/Light-propagation-in-a-
plasma-on-an-axially.

Sato:2022:VES

[1167] Naoki Sato and Michio Yamada. Vorticity equation on surfaces with arbitrary topology embedded in three-dimensional Euclidean space. Journal of Mathematical Physics, 63(9):093101, September 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/9/093101/2844313/Vorticity-
equation-on-surfaces-with-arbitrary.

Lacroix-A-Chez-Toine:2022:SRP

[1168] Bertrand Lacroix-A-Chez-Toine, Yan V. Fyodorov, and Sirio Belga Fedeli. Superposition of random plane waves in high spatial dimensions: Random matrix approach to landscape complexity. Journal of Mathematical Physics, 63(9):093301, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/9/093301/2844320/Superposition-of-random-

plane-waves-in-high. Special collection in honor of Freeman Dyson.

Garban:2022:CSB

[1169] Christophe Garban and Thomas Spencer. Continuous symmetry breaking along the Nishimori line. Journal of Mathematical Physics, 63 (9):093302, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/9/093302/2844382/Continuous-symmetry-breaking-along-

the-Nishimori. Special collection in honor of Freeman Dyson.

Magnot:2022:ZSE

[1170] Jean-Pierre Magnot, Enrique G. Reyes, and Vladimir Rubtsov. On (t_{2}, t_{3})-Zakharov–Shabat equations of generalized Kadomtsev– Petviashvili hierarchies. Journal of Mathematical Physics, 63(9):093501, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/9/
093501/2844379/On-t2-t3-Zakharov-Shabat-equations-of-generalized.

Kang:2022:GSS

[1171] Jin-Cai Kang, Xiao-Ping Chen, and Chun-Lei Tang. Ground state solutions for Schrödinger–Poisson system with critical growth and nonperiodic potential. Journal of Mathematical Physics, 63(10):101501, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
101501/2843127/Ground-state-solutions-for-Schrodinger-Poisson.

Dong:2022:ASC

[1172] Jianwei Dong and Jingjing Li. Analytical solutions to the compressible Euler equations with time-dependent damping and free boundaries. Journal of Mathematical Physics, 63(10):101502, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
10/101502/2843109/Analytical-solutions-to-the-compressible-
Euler.

Zhao:2022:GSS

[1173] Yu-Xin Zhao, Xing-Ping Wu, and Chun-Lei Tang. Ground state signchanging solutions for Schrödinger–Kirchhoff-type problem with critical growth. Journal of Mathematical Physics, 63(10):101503, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
101503/2843116/Ground-state-sign-changing-solutions-for.

Ji:2022:ESC

[1174] Un Cig Ji. Evolution systems and continuity equations for white noise functionals. Journal of Mathematical Physics, 63(10):101504, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
101504/2843158/Evolution-systems-and-continuity-equations-for.

Ma:2022:SAB

[1175] Liangliang Ma, Lin Li, and Dongbing Liu. Stability of the 2D anisotropic Boussinesq equations with mixed partial dissipation. Journal of Mathematical Physics, 63(10):101505, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
101505/2843132/Stability-of-the-2D-anisotropic-Boussinesq.

Chen:2022:EVS

[1176] Shouxin Chen and Guange Su. Existence of vortices for Schrödinger equations with logarithmic and saturable nonlinearity. Journal of Mathematical Physics, 63(10):101506, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/10/101506/2843101/Existence-of-vortices-
for-Schrodinger-equations.

Liu:2022:CMN

[1177] Jiefeng Liu, Sihan Zhou, and Lamei Yuan. Conformal r-matrix-Nijenhuis structures, symplectic-Nijenhuis structures, and ON -structures. Journal of Mathematical Physics, 63(10):101701, October 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/10/101701/2843141/Conformal-
r-matrix-Nijenhuis-structures-symplectic.

Jones:2022:ISC

[1178] Nick G. Jones and Noah Linden. Integrable spin chains and the Clifford group. Journal of Mathematical Physics, 63(10):101901, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
101901/2843173/Integrable-spin-chains-and-the-Clifford-group.

Srivastava:2022:ADE

[1179] Atulit Srivastava and S. K. Soni. Algebraic derivation of the energy eigenvalues for the quantum oscillator defined on the sphere and the hyperbolic plane. Journal of Mathematical Physics, 63(10):102101, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102101/2843163/Algebraic-derivation-of-the-energy-eigenvalues-
for.

Duboscq:2022:LQG

[1180] Romain Duboscq and Olivier Pinaud. On local quantum Gibbs states. Journal of Mathematical Physics, 63(10):102102, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102102/2843152/On-local-quantum-Gibbs-states.

Zhou:2022:QEM

[1181] Li Zhou, Nengkun Yu, Shenggang Ying, and Mingsheng Ying. Quantum earth mover’s distance, a no-go quantum Kantorovich–Rubinstein theorem, and quantum marginal problem. Journal of Mathematical Physics, 63(10):102201, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/10/102201/2843097/Quantum-earth-mover-s-
distance-a-no-go-quantum.

Collins:2022:CER

[1182] Benoît Collins and Félix Parraud. Concentration estimates for random subspaces of a tensor product and application to quantum information theory. Journal of Mathematical Physics, 63(10):102202, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102202/2843106/Concentration-estimates-for-random-subspaces-
of-a.

Lorenzoni:2022:RNS

[1183] Paolo Lorenzoni and Sara Perletti. Regular non-semisimple Dubrovin– Frobenius manifolds. Journal of Mathematical Physics, 63(10):102301, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102301/2843112/Regular-non-semisimple-Dubrovin-Frobenius.

Xu:2022:RRA

[1184] Feng Xu. Rigorous results about entropies in quantum field theory. Journal of Mathematical Physics, 63(10):102302, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102302/2843119/Rigorous-results-about-entropies-in-quantum-
field.

Ma:2022:LSA

[1185] Zhichao Ma and Junxiang Xu. Lagrange stability for asymptotic linear Duffing equations. Journal of Mathematical Physics, 63(10):102701, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
10/102701/2843170/Lagrange-stability-for-asymptotic-linear-
Duffing.

Shu:2022:VSH

[1186] Xiang Shu, Jun Yan, and Kai Zhao. Viscosity subsolutions of Hamilton–Jacobi equations and invariant sets of contact Hamilton systems. Journal of Mathematical Physics, 63(10):102702, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102702/2843137/Viscosity-subsolutions-of-Hamilton-Jacobi.

Abdallah:2022:DNA

[1187] Ahmed Y. Abdallah, Heba N. Abu-Shaab, Taqwa M. Al-Khader, and Rania T. Wannan. Dynamics of non-autonomous first order lattice systems in weighted spaces. Journal of Mathematical Physics, 63(10):102703, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/10/102703/2843146/Dynamics-
of-non-autonomous-first-order-lattice.

He:2022:UAN

[1188] Jinfang He, Shan Ma, Chunyou Sun, and Lu Yang. Uniform attractors for nonautonomous 2D MHD equations with partial dissipation. Journal of Mathematical Physics, 63(10):102704, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102704/2843155/Uniform-attractors-for-nonautonomous-2D-MHD.

Wood:2022:SPA

[1189] William Wood. On the spectrum of the periodic Anderson–Bernoulli model. Journal of Mathematical Physics, 63(10):102705, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
10/102705/2843171/On-the-spectrum-of-the-periodic-Anderson-
Bernoulli.

Selvam:2022:RFD

[1190] A. Panneer Selvam and V. Govindaraj. Reachability of fractional dynamical systems with multiple delays in control using ψ-Hilfer pseudofractional derivative. Journal of Mathematical Physics, 63(10):102706, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
102706/2843166/Reachability-of-fractional-dynamical-systems-
with.

Frassek:2022:ESI

[1191] Rouven Frassek and Cristian Giardinà. Exact solution of an integrable non-equilibrium particle system. Journal of Mathematical Physics, 63(10):103301, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/10/103301/2843267/Exact-solution-of-
an-integrable-non-equilibrium.

Jia:2022:RSB

[1192] Yiyang Jia, Dario Rosa, and Jacobus J. M. Verbaarschot. Replica symmetry breaking for the integrable two-site Sachdev–Ye–Kitaev model. Journal of Mathematical Physics, 63(10):103302, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/10/103302/2843256/Replica-
symmetry-breaking-for-the-integrable-two. Special collection in

honor of Freeman Dyson.

Cipolloni:2022:DES

[1193] Giorgio Cipolloni, László Erdős, Dominik Schröder, and Yuanyuan Xu. Directional extremal statistics for Ginibre eigenvalues. Journal of Mathematical Physics, 63(10):103303, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/10/103303/2843234/Directional-extremal-statistics-
for-Ginibre.

Agliari:2022:NPA

[1194] Elena Agliari, Alberto Fachechi, and Chiara Marullo. Nonlinear PDEs approach to statistical mechanics of dense associative memories. Journal of Mathematical Physics, 63(10):103304, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103304/2843264/Nonlinear-PDEs-approach-to-statistical-mechanics.

Aktosun:2022:GMD

[1195] Tuncay Aktosun and Mehmet Unlu. A generalized method for the Darboux transformation. Journal of Mathematical Physics, 63(10):103501, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103501/2843241/A-generalized-method-for-the-Darboux.

Miki:2022:CQW

[1196] Hiroshi Miki, Satoshi Tsujimoto, and Luc Vinet. Classical and quantum walks on paths associated with exceptional Krawtchouk polynomials. Journal of Mathematical Physics, 63(10):103502, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103502/2843247/Classical-and-quantum-walks-on-paths-associated.

Xia:2022:PAI

[1197] Li-Li Xia, Meng-Meng Wu, and Long Bai. Preservation of adiabatic invariants and geometric numerical algorithm for disturbed nonholonomic systems. Journal of Mathematical Physics, 63(10):103503, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
10/103503/2843259/Preservation-of-adiabatic-invariants-and-
geometric.

Adler:2022:NAT

[1198] V. E. Adler and M. P. Kolesnikov. Non-Abelian Toda lattice and analogs of Painlevé III equation. Journal of Mathematical Physics, 63(10):

103504, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
10/103504/2843225/Non-Abelian-Toda-lattice-and-analogs-of-Painleve.

Besnard:2022:PMS

[1199] F. Besnard and S. Farnsworth. Particle models from special Jordan backgrounds and spectral triples. Journal of Mathematical Physics, 63(10):103505, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/10/103505/2843253/Particle-
models-from-special-Jordan-backgrounds.

Lee:2022:CMP

[1200] Gihyun Lee and Max Lein. A calculus for magnetic pseudodifferential super operators. Journal of Mathematical Physics, 63(10):103506, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103506/2843271/A-calculus-for-magnetic-pseudodifferential-super.

Tataris:2022:DGL

[1201] Andreas Tataris and Tristan van Leeuwen. A distributional Gelfand– Levitan–Marchenko equation for the Helmholtz scattering problem on the line. Journal of Mathematical Physics, 63(10):103507, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103507/2843217/A-distributional-Gelfand-Levitan-Marchenko.

Morato:2022:FTE

[1202] Laura M. Morato. Free time evolution of the quantum wave function and optimal transportation. Journal of Mathematical Physics, 63(10):103508, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/10/
103508/2843230/Free-time-evolution-of-the-quantum-wave-function.

Keys:2022:EPS

[1203] Dustin Keys and Jan Wehr. Erratum: “Poisson stochastic master equation unravelings and the measurement problem: a quantum stochastic calculus perspective” [J. Math. Phys. 61, 032101 (2020)]. Journal of Mathematical Physics, 63(10):109901, October 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/10/109901/2843251/Erratum-
Poisson-stochastic-master-equation. See [83].

Yang:2022:PRW

[1204] Jiaqi Yang. Partially regular weak solutions to the fractional Navier– Stokes equations with the critical dissipation. Journal of Mathematical Physics, 63(11):111501, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/11/111501/2846011/Partially-regular-
weak-solutions-to-the-fractional.

Read:2022:LSE

[1205] Larry Read, Bogusław Zegarliński, and Mengchun Zhang. Logarithmic Schrödinger equations in infinite dimensions. Journal of Mathematical Physics, 63(11):111502, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/111502/2846028/Logarithmic-Schrodinger-equations-
in-infinite.

Ju:2022:ZML

[1206] Qiangchang Ju and Jianjun Xu. Zero-Mach limit of the compressible Navier–Stokes–Korteweg equations. Journal of Mathematical Physics, 63(11):111503, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/11/111503/2846068/Zero-Mach-limit-of-the-compressible-
Navier-Stokes.

Kim:2022:RII

[1207] Jae-Myoung Kim. Regularity for 3D inhomogeneous incompressible MHD equations with vacuum. Journal of Mathematical Physics, 63(11):111504, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/11/111504/2846023/Regularity-for-3D-
inhomogeneous-incompressible-MHD.

Palmieri:2022:SWE

[1208] Alessandro Palmieri and Hiroyuki Takamura. On a semilinear wave equation in anti-de Sitter spacetime: The critical case. Journal of Mathematical Physics, 63(11):111505, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/11/111505/2846062/On-a-semilinear-wave-equation-
in-anti-de-Sitter.

Li:2022:GWPb

[1209] Huanyuan Li. Global well-posedness to the Cauchy problem of 2D nonhomogeneous Bénard system with large initial data and vacuum. Journal of Mathematical Physics, 63(11):111506, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/11/111506/2846007/Global-
well-posedness-to-the-Cauchy-problem-of-2D.

Wang:2022:EDR

[1210] Shuai Wang, Fei Chen, Yongye Zhao, and Chuanbao Wang. Estimation of decay rates to large-solutions of 3D compressible magnetohydrodynamic system. Journal of Mathematical Physics, 63 (11):111507, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/111507/2846017/Estimation-of-decay-rates-
to-large-solutions-of-3D.

Huang:2022:SGI

[1211] Guangyue Huang and Mingfang Zhu. Some geometric inequalities on Riemannian manifolds associated with the generalized modified Ricci curvature. Journal of Mathematical Physics, 63(11):111508, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
111508/2846053/Some-geometric-inequalities-on-Riemannian.

Rottensteiner:2022:HAO

[1212] David Rottensteiner and Michael Ruzhansky. Harmonic and anharmonic oscillators on the Heisenberg group. Journal of Mathematical Physics, 63(11):111509, November 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/11/111509/2846056/Harmonic-
and-anharmonic-oscillators-on-the.

Liu:2022:ASS

[1213] Anran Liu and Engui Fan. The asymptotic stability of solitons for the focusing mKdV equation with weak weighted Sobolev initial data. Journal of Mathematical Physics, 63(11):111510, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
111510/2846048/The-asymptotic-stability-of-solitons-for-the.

Dlotko:2022:NSC

[1214] Tomasz Dlotko. Navier–Stokes–Cahn–Hilliard system of equations. Journal of Mathematical Physics, 63(11):111511, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/11/111511/2846003/Navier-
Stokes-Cahn-Hilliard-system-of-equations.

Yang:2022:DSR

[1215] Shuang Yang, Yangrong Li, and Tomás Caraballo. Dynamical stability of random delayed FitzHugh–Nagumo lattice systems driven by nonlinear Wong–Zakai noise. Journal of Mathematical Physics, 63(11):111512, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/11/111512/2846009/Dynamical-stability-of-random-delayed-
FitzHugh.

Liu:2022:EMB

[1216] Weiming Liu. Existence of multi-bump solutions for a nonlinear Kirchhoff-type system. Journal of Mathematical Physics, 63(11):111513, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/11/111513/2846015/Existence-of-multi-bump-solutions-for-
a-nonlinear.

Shapiro:2022:CSM

[1217] Jacob Shapiro and Michael I. Weinstein. Is the continuum SSH model topological? Journal of Mathematical Physics, 63(11):111901, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
111901/2846020/Is-the-continuum-SSH-model-topological.

Curtright:2022:PVG

[1218] T. Curtright and S. Subedi. Potentials vs geometry, revisited. Journal of Mathematical Physics, 63(11):112101, November 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/11/112101/2846025/Potentials-
vs-geometry-revisited.

Heimendahl:2022:AOA

[1219] Arne Heimendahl, Markus Heinrich, and David Gross. The axiomatic and the operational approaches to resource theories of magic do not coincide. Journal of Mathematical Physics, 63 (11):112201, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/11/112201/2846042/The-axiomatic-and-the-
operational-approaches-to.

vomEnde:2022:WBH

[1220] Frederik vom Ende. Which bath Hamiltonians matter for thermal operations? Journal of Mathematical Physics, 63(11):112202, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112202/2846069/Which-bath-Hamiltonians-matter-for-thermal.

Shirokov:2022:OFK

[1221] M. E. Shirokov. Optimal form of the Kretschmann–Schlingemann– Werner theorem for energy-constrained quantum channels and operations. Journal of Mathematical Physics, 63(11):112203, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112203/2846038/Optimal-form-of-the-Kretschmann-Schlingemann.

Brannan:2022:QCG

[1222] Michael Brannan, Priyanga Ganesan, and Samuel J. Harris. The quantum-to-classical graph homomorphism game. Journal of Mathematical Physics, 63(11):112204, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/11/112204/2846145/The-quantum-to-classical-
graph-homomorphism-game.

Appleby:2022:SPS

[1223] Marcus Appleby, Ingemar Bengtsson, Markus Grassl, Michael Harrison, and Gary McConnell. SIC-POVMs from Stark units: Prime dimensions n^{2} + 3. Journal of Mathematical Physics, 63(11):112205, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112205/2846122/SIC-POVMs-from-Stark-units-Prime-dimensions-
n2-3.

Burgarth:2022:SCD

[1224] Daniel Burgarth, Paolo Facchi, and Robin Hillier. Stability and convergence of dynamical decoupling with finite amplitude controls. Journal of Mathematical Physics, 63(11):112206, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112206/2846127/Stability-and-convergence-of-dynamical-decoupling.

Hohn:2022:IQR

[1225] Philipp A. Höhn, Marius Krumm, and Markus P. Müller. Internal quantum reference frames for finite Abelian groups. Journal of Mathematical Physics, 63(11):112207, November 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/11/112207/2846139/Internal-
quantum-reference-frames-for-finite.

Kiessling:2022:DDS

[1226] Michael K.-H. Kiessling, Eric Ling, and A. Shadi Tahvildar-Zadeh. On the discrete Dirac spectrum of a point electron in the zerogravity Kerr–Newman spacetime. Journal of Mathematical Physics, 63(11):112301, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/11/112301/2846150/On-the-discrete-Dirac-
spectrum-of-a-point-electron. See publisher’s note [1293].

Boucetta:2022:KTD

[1227] Mohamed Boucetta, Aissa Meliani, and Abdelghani Zeghib. Kundt three-dimensional left invariant spacetimes. Journal of Mathematical Physics, 63(11):112501, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/112501/2846105/Kundt-three-dimensional-left-
invariant-spacetimes.

Panah:2022:TDL

[1228] B. Eslam Panah. Two-dimensional Lifshitz-like AdS black holes in F (r) gravity. Journal of Mathematical Physics, 63(11):112502, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112502/2846116/Two-dimensional-Lifshitz-like-AdS-black-holes-
in-F.

Chen:2022:GCM

[1229] Yu-Jie Chen, Shi-Lin Li, Yu-Zhu Chen, Wen-Du Li, and Wu-Sheng Dai. Gravitational constant model and correction. Journal of Mathematical Physics, 63(11):112503, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/11/112503/2846098/Gravitational-
constant-model-and-correction.

Chen:2022:LQV

[1230] Dan Chen, Michal Fečkan, and JinRong Wang. Linear quaternionvalued difference equations: Representation of solutions, controllability, and observability. Journal of Mathematical Physics, 63 (11):112701, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/11/112701/2846147/Linear-quaternion-
valued-difference-equations.

Chen:2022:ISA

[1231] Shuang Chen and Jinqiao Duan. Instability of small-amplitude periodic waves from fold-Hopf bifurcation. Journal of Mathematical Physics, 63(11):112702, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/11/112702/2846112/Instability-of-small-amplitude-
periodic-waves-from.

Littlejohn:2022:CPM

[1232] Robert Littlejohn, Philip Morrison, and Jeffrey Heninger. Charged particle motion in spherically symmetric distributions of magnetic monopoles. Journal of Mathematical Physics, 63(11):112703, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112703/2846136/Charged-particle-motion-in-spherically-symmetric.

Saba:2022:GJG

[1233] N. Saba and A. Boussayoud. Gaussian (p, q)-Jacobsthal and Gaussian (p, q)-Jacobsthal Lucas numbers and their some interesting properties. Journal of Mathematical Physics, 63(11):112704, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112704/2846102/Gaussian-p-q-Jacobsthal-and-Gaussian-p-q.

Si:2022:DRT

[1234] Wen Si and Xinyu Guan. Degenerate response tori in Hamiltonian systems with higher zero-average perturbation. Journal of Mathematical Physics, 63(11):112705, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/112705/2846109/Degenerate-response-tori-in-
Hamiltonian-systems.

Pina:2022:PCC

[1235] E. Piña. Planar central configurations with five different positive masses. Journal of Mathematical Physics, 63(11):112901, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
112901/2846115/Planar-central-configurations-with-five-different.

Hutchcroft:2022:SHU

[1236] Tom Hutchcroft. Sharp hierarchical upper bounds on the critical twopoint function for long-range percolation on Z^{d}. Journal of Mathematical Physics, 63(11):113301, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/113301/2846118/Sharp-hierarchical-upper-bounds-
on-the-critical.

Wang:2022:UMQ

[1237] Ce Wang. The uniform measure for quantum walk on hypercube: a quantum Bernoulli noises approach. Journal of Mathematical Physics, 63(11):113501, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/11/113501/2846125/The-uniform-measure-for-quantum-
walk-on-hypercube.

Palheta:2022:CRM

[1238] Pedro H. S. Palheta, Marcelo R. Barbosa, and Marcel Novaes. Commutators of random matrices from the unitary and orthogonal groups. Journal of Mathematical Physics, 63(11):113502, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/11/
113502/2846130/Commutators-of-random-matrices-from-the-unitary.

Yang:2022:MUC

[1239] Qianqian Yang and Chuanzhong Li. Multiparameter universal characters of B-type and integrable hierarchy. Journal of Mathematical Physics, 63(11):113503, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/11/113503/2846182/Multiparameter-universal-characters-
of-B-type-and.

Zhu:2022:LTA

[1240] Jin-Yan Zhu and Yong Chen. Long-time asymptotic behavior of the coupled dispersive AB system in low regularity spaces. Journal of Mathematical Physics, 63(11):113504, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/11/113504/2846187/Long-time-asymptotic-behavior-
of-the-coupled.

Nakata:2022:IDD

[1241] Kenta Nakata. Integrable delay-difference and delay-differential analogs of the KdV, Boussinesq, and KP equations. Journal of Mathematical Physics, 63(11):113505, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/11/113505/2846180/Integrable-delay-difference-
and-delay-differential.

Alpay:2022:AGR

[1242] Daniel Alpay, Fabrizio Colombo, Kamal Diki, and Irene Sabadini. An approach to the Gaussian RBF kernels via Fock spaces. Journal of Mathematical Physics, 63(11):113506, November 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/11/113506/2846177/An-approach-to-
the-Gaussian-RBF-kernels-via-Fock.

Henheik:2022:ATE

[1243] Joscha Henheik and Tom Wessel. On adiabatic theory for extended fermionic lattice systems. Journal of Mathematical Physics, 63(12):121101, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/121101/2846084/On-adiabatic-theory-for-extended-
fermionic-lattice.

Deng:2022:CSH

[1244] Shengbing Deng and Junwei Yu. On a class of singular Hamiltonian Choquard-type elliptic systems with critical exponential growth. Journal of Mathematical Physics, 63(12):121501, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/121501/2846061/On-a-class-
of-singular-Hamiltonian-Choquard-type.

Liu:2022:TPD

[1245] Jianli Liu and Kenan Zhang. Threshold phenomenon in n-dimensional repulsive restricted Euler–Poisson equations with time-dependent damping. Journal of Mathematical Physics, 63(12):121502, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/12/121502/2846041/Threshold-phenomenon-in-n-dimensional-
repulsive.

Gao:2022:SPC

[1246] Yu Gao, Hao Liu, and Tak Kwong Wong. Stability of peakons of the Camassa–Holm equation beyond wave breaking. Journal of Mathematical Physics, 63(12):121503, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/12/121503/2846080/Stability-of-peakons-of-the-
Camassa-Holm-equation.

Yang:2022:QAT

[1247] Shaojie Yang and Zhijun Qiao. Qualitative analysis for a two-component peakon system with cubic nonlinearity. Journal of Mathematical Physics, 63(12):121504, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/121504/2846089/Qualitative-analysis-for-a-two-
component-peakon.

Priyanka:2022:DSV

[1248] Priyanka and M. Zafar. Delta shocks and vacuum states in the Riemann solutions of Chaplygin Euler equations as pressure and magnetic field drop to zero. Journal of Mathematical Physics, 63(12):121505, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
63/12/121505/2846091/Delta-shocks-and-vacuum-states-in-the-
Riemann.

Safronov:2022:EPS

[1249] Oleg Safronov. Eigenvalues of a periodic Schrödinger operator perturbed by a fast decaying potential. Journal of Mathematical Physics, 63(12):121506, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/12/121506/2846050/Eigenvalues-of-a-periodic-
Schrodinger-operator.

Li:2022:EDO

[1250] Yuqiao Li. Eigenvalues of the Dirac operator on compact spin manifolds under Ricci flow. Journal of Mathematical Physics, 63(12):121507, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
121507/2846074/Eigenvalues-of-the-Dirac-operator-on-compact-
spin.

Freitas:2022:LTD

[1251] M. M. Freitas, R. Q. Caljaro, A. J. A. Ramos, and H. C. M. Rodrigues. Long-time dynamics of ternary mixtures with localized dissipation. Journal of Mathematical Physics, 63(12):121508, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
121508/2846066/Long-time-dynamics-of-ternary-mixtures-with.

Li:2022:RSA

[1252] Shiwei Li. Riemann solutions of the anti-Chaplygin pressure Aw– Rascle model with friction. Journal of Mathematical Physics, 63 (12):121509, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/121509/2846093/Riemann-solutions-of-the-anti-
Chaplygin-pressure.

Samprogna:2022:STA

[1253] Rodrigo A. Samprogna and Leonardo Pires. Semicontinuity of trajectory attractors with respect to exponents for p-Laplacian equation. Journal of Mathematical Physics, 63(12):121510, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
121510/2846037/Semicontinuity-of-trajectory-attractors-with.

Fang:2022:SSC

[1254] Li Fang, Yugang Zhao, and Zhenhua Guo. Stationary solutions to a compressible non-Newtonian fluid with general boundary conditions. Journal of Mathematical Physics, 63(12):121511, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
121511/2846045/Stationary-solutions-to-a-compressible-non.

Kim:2022:SBR

[1255] Jeong-Ah Kim and Dong-Uy Shin. Strip bundle realization of the crystals over U_{q}(D_{4}^{(3)}). Journal of Mathematical Physics, 63(12):121701, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/63/
12/121701/2846057/Strip-bundle-realization-of-the-crystals-
over-Uq.

Roy:2022:SDP

[1256] Sumit Roy. Spectral data for parabolic projective symplectic/ orthogonal Higgs bundles. Journal of Mathematical Physics, 63 (12):121702, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/12/121702/2846063/Spectral-data-for-
parabolic-projective-symplectic.

Marquette:2022:GQC

[1257] Ian Marquette, Luke Yates, and Peter D. Jarvis. Generalized quadratic commutator algebras of PBW-type. Journal of Mathematical Physics, 63(12):121703, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/12/121703/2846070/Generalized-quadratic-
commutator-algebras-of-PBW.

Jezequel:2022:EBE

[1258] Lucien Jezequel, Clément Tauber, and Pierre Delplace. Estimating bulk and edge topological indices in finite open chiral chains. Journal of Mathematical Physics, 63(12):121901, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
121901/2846078/Estimating-bulk-and-edge-topological-indices-
in.

Urban:2022:DFA

[1259] Roman Urban. On a diffusion on finite adeles and the Feynman– Kac integral. Journal of Mathematical Physics, 63(12):122101, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
122101/2846082/On-a-diffusion-on-finite-adeles-and-the-Feynman.

Bagchi:2022:AHR

[1260] Bijan Bagchi and Sauvik Sen. Artificial Hawking radiation, weak pseudo-Hermiticity, and Weyl semimetal blackhole analogy. Journal of Mathematical Physics, 63(12):122102, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/12/122102/2846087/Artificial-Hawking-
radiation-weak-pseudo.

Anshu:2022:CCN

[1261] Anurag Anshu and Nikolas P. Breuckmann. A construction of combinatorial NLTS. Journal of Mathematical Physics, 63(12):122201, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
122201/2846035/A-construction-of-combinatorial-NLTS.

Hu:2022:UPO

[1262] Mengyao Hu, Lin Chen, Fei Shi, Xiande Zhang, and Jordi Tura. Unextendible product operator basis. Journal of Mathematical Physics, 63(12):122202, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/122202/2846086/Unextendible-product-operator-
basis.

LaRacuente:2022:QFM

[1263] Nicholas LaRacuente. Quasi-factorization and multiplicative comparison of subalgebra-relative entropy. Journal of Mathematical Physics, 63(12):122203, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/122203/2846170/Quasi-factorization-and-multiplicative-
comparison.

Junge:2022:MTI

[1264] Marius Junge and Nicholas LaRacuente. Multivariate trace inequalities, p-fidelity, and universal recovery beyond tracial settings. Journal of Mathematical Physics, 63(12):122204, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
122204/2846146/Multivariate-trace-inequalities-p-fidelity-and.

Higuchi:2022:SFA

[1265] Atsushi Higuchi, Lasse Schmieding, and David Serrano Blanco. Scalar field in AdS_{2} and representations of S̃L(2, R). Journal of Mathematical Physics, 63(12):122301, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/12/122301/2846160/Scalar-
field-in-AdS2-and-representations-of-SL-2-R.

Henheik:2022:IBC

[1266] Joscha Henheik and Roderich Tumulka. Interior-boundary conditions for the Dirac equation at point sources in three dimensions. Journal of Mathematical Physics, 63(12):122302, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/122302/2846135/Interior-
boundary-conditions-for-the-Dirac.

Galloway:2022:CNC

[1267] Gregory J. Galloway and Eric Ling. The codimension 2 null cut locus with applications to spacetime topology. Journal of Mathematical Physics, 63(12):122501, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/12/122501/2846129/The-codimension-2-
null-cut-locus-with-applications.

Chai:2022:MAH

[1268] Xiaoxiang Chai and Xueyuan Wan. The mass of an asymptotically hyperbolic end and distance estimates. Journal of Mathematical Physics, 63(12):122502, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/122502/2846153/The-mass-of-an-asymptotically-
hyperbolic-end-and.

Wu:2022:PWP

[1269] Weixin Wu and Zhidong Teng. Periodic wave propagation in a diffusive SIR epidemic model with nonlinear incidence and periodic environment. Journal of Mathematical Physics, 63(12):122701, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
122701/2846167/Periodic-wave-propagation-in-a-diffusive-SIR.

Lin:2022:PMI

[1270] Yusen Lin, Yayu Li, and Dingshi Li. Periodic measures of impulsive stochastic neural networks lattice systems with delays. Journal of Mathematical Physics, 63(12):122702, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/122702/2846141/Periodic-
measures-of-impulsive-stochastic-neural.

Corbera:2022:PCC

[1271] M. Corbera, J. Llibre, and C. Valls. Planar central configurations of some restricted (4 + 1)-body problems. Journal of Mathematical Physics, 63(12):122901, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/122901/2846165/Planar-central-configurations-
of-some-restricted-4.

Esen:2022:RGSa

[1272] Oğul Esen, Miroslav Grmela, and Michal Pavelka. On the role of geometry in statistical mechanics and thermodynamics. I. Geometric perspective. Journal of Mathematical Physics, 63(12):122902, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
122902/2846172/On-the-role-of-geometry-in-statistical-mechanics.

Khare:2022:NSK

[1273] Avinash Khare and Avadh Saxena. Novel superposed kinklike and pulselike solutions for several nonlocal nonlinear equations. Journal of Mathematical Physics, 63(12):122903, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/12/122903/2846126/Novel-superposed-
kinklike-and-pulselike-solutions.

Liu:2022:LTT

[1274] Pan Liu. Liouville-type theorems for the stationary compressible barotropic and incompressible inhomogeneous Navier–Stokes equations. Journal of Mathematical Physics, 63(12):123101, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123101/2846131/Liouville-type-theorems-for-the-stationary.

Liu:2022:GES

[1275] Cunming Liu. Global existence of smooth solutions to a full Euler–Poisson system in one space dimension. Journal of Mathematical Physics, 63(12):123102, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/12/123102/2846140/Global-
existence-of-smooth-solutions-to-a-full.

Butta:2022:VVL

[1276] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Vanishing viscosity limit for concentrated vortex rings. Journal of Mathematical Physics, 63(12):123103, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/123103/2846144/Vanishing-viscosity-limit-for-
concentrated-vortex.

Wang:2022:DHK

[1277] Na Wang. (β-deformed) Hurwitz–Kontsevich model and affine Yangian of }↕(∞). Journal of Mathematical Physics, 63(12):123301, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123301/2846149/deformed-Hurwitz-Kontsevich-model-and-affine.

Weiss:2022:TLI

[1278] Daniel Alexander Weiss. The thermodynamic limit of an ideal Bose gas by asymptotic expansions and spectral ζ-functions. Journal of Mathematical Physics, 63(12):123302, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/12/123302/2846156/The-thermodynamic-
limit-of-an-ideal-Bose-gas-by.

Bet:2022:MDP

[1279] Gianmarco Bet, Anna Gallo, and Francesca R. Nardi. Metastability for the degenerate Potts model with negative external magnetic field under Glauber dynamics. Journal of Mathematical Physics, 63(12):123303, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/63/12/123303/2846162/Metastability-for-the-
degenerate-Potts-model-with. See erratum [1337].

Qian:2022:MDL

[1280] Hongjiang Qian and George Yin. Moderate deviations for the Langevin equations: Strong damping and fast Markovian switching. Journal of Mathematical Physics, 63(12):123304, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/123304/2846222/Moderate-
deviations-for-the-Langevin-equations.

Esen:2022:RGSb

[1281] Oğul Esen, Miroslav Grmela, and Michal Pavelka. On the role of geometry in statistical mechanics and thermodynamics. II. Thermodynamic perspective. Journal of Mathematical Physics, 63(12):123305, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123305/2846208/On-the-role-of-geometry-in-statistical-mechanics.

Donati:2022:RGR

[1282] Luca Donati, Marcus Weber, and Bettina G. Keller. A review of Girsanov reweighting and of square root approximation for building molecular Markov state models. Journal of Mathematical Physics, 63(12):123306, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/123306/2846228/A-review-
of-Girsanov-reweighting-and-of-square.

Wang:2022:LTA

[1283] Deng-Shan Wang and Xiaodong Zhu. Long-time asymptotics of the good Boussinesq equation with q_{xx}-term and its modified version. Journal of Mathematical Physics, 63(12):123501, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123501/2846230/Long-time-asymptotics-of-the-good-Boussinesq.

Hanisch:2022:FIL

[1284] Florian Hanisch and Matthias Ludewig. The fermionic integral on loop space and the Pfaffian line bundle. Journal of Mathematical Physics, 63(12):123502, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/63/12/123502/2846225/The-fermionic-integral-on-
loop-space-and-the.

Ohmori:2022:RHS

[1285] S. Ohmori and J. Takahashi. Rigged Hilbert space approach for non-Hermitian systems with positive definite metric. Journal of Mathematical Physics, 63(12):123503, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/12/123503/2846218/Rigged-
Hilbert-space-approach-for-non-Hermitian.

Shiozawa:2022:LBG

[1286] Yuichi Shiozawa. Limiting behaviors of generalized elephant random walks. Journal of Mathematical Physics, 63(12):123504, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123504/2846202/Limiting-behaviors-of-generalized-elephant-random.

Erman:2022:ROP

[1287] Fatih Erman, Sema Seymen, and O. Teoman Turgut. Rank one perturbations supported by hybrid geometries and their deformations. Journal of Mathematical Physics, 63(12):123505, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123505/2846231/Rank-one-perturbations-supported-by-hybrid.

DAchille:2022:DOT

[1288] Matteo D’Achille, Aernout C. D. van Enter, and Arnaud Le Ny. Decimations for one- and two-dimensional Ising and rotator models. II. Continuous vs discrete symmetries. Journal of Mathematical Physics, 63(12):123506, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/123506/2846226/Decimations-for-one-and-two-dimensional-
Ising-and.

Novikova:2022:CSD

[1289] E. M. Novikova. Coherent Schwartz distributions of the Heisenberg algebra and inverted oscillator. Journal of Mathematical Physics, 63(12):123507, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/63/12/123507/2846224/Coherent-Schwartz-
distributions-of-the-Heisenberg.

Meljanac:2022:SOW

[1290] Stjepan Meljanac, Zoran Škoda, and Saša Krešić-Jurić. Symmetric ordering and Weyl realizations for quantum Minkowski spaces. Journal of Mathematical Physics, 63(12):123508, December 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/63/12/123508/2846195/Symmetric-
ordering-and-Weyl-realizations-for.

Fring:2022:LRI

[1291] Andreas Fring and Rebecca Tenney. Lewis–Riesenfeld invariants for PTsymmetrically coupled oscillators from two-dimensional point transformations and Lie algebraic expansions. Journal of Mathematical Physics, 63(12):123509, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/63/12/123509/2846206/Lewis-Riesenfeld-invariants-for-
PT-symmetrically.

Nucci:2022:MSS

[1292] M. C. Nucci and R. Campoamor-Stursberg. Minimally superintegrable systems in flat three-dimensional space are also linearizable. Journal of Mathematical Physics, 63(12):123510, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
123510/2846216/Minimally-superintegrable-systems-in-flat-three.

Kiessling:2022:PND

[1293] Michael K.-H. Kiessling, Eric Ling, and A. Shadi Tahvildar-Zadeh. Publisher’s note: “On the discrete Dirac spectrum of a point electron in the zero-gravity Kerr–Newman spacetime” [J. Math. Phys. 63, 112301 (2022)]. Journal of Mathematical Physics, 63(12):129901, December 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/63/12/
129901/2846221/Publisher-s-Note-On-the-discrete-Dirac-spectrum-

of. See [1226].

Yamazaki:2023:QYC

[1294] Masahito Yamazaki. Quiver Yangians and crystal meltings: a concise summary. Journal of Mathematical Physics, 64(1):011101, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/1/011101/2870501/Quiver-Yangians-and-crystal-meltings-A-
concise.

Sun:2023:FVQ

[1295] Wenlong Sun, Yeping Li, and Xiaoying Han. The full viscous quantum hydrodynamic system in one dimensional space. Journal of Mathematical Physics, 64(1):011501, January 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/1/011501/2870486/The-full-
viscous-quantum-hydrodynamic-system-in.

Zheng:2023:BSC

[1296] Bowen Zheng, Tohru Ozawa, and Jian Zhai. Blow-up solutions for a class of divergence Schrödinger equations with intercritical inhomogeneous nonlinearity. Journal of Mathematical Physics, 64(1):011502, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011502/2870450/Blow-up-solutions-for-a-class-of-divergence.

Dong:2023:ASM

[1297] Lihua Dong and Xiaoxia Ren. Asymptotic stability of the 2D MHD equations without magnetic diffusion. Journal of Mathematical Physics, 64(1):011503, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/1/011503/2870493/Asymptotic-stability-of-the-2D-
MHD-equations.

Chen:2023:ASM

[1298] Dongxiang Chen, Xiaoli Li, and Xiaoli Chen. Asymptotic stability of the 2D MHD equations without resistivity on a flat strip domain. Journal of Mathematical Physics, 64(1):011504, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011504/2870489/Asymptotic-stability-of-the-2D-MHD-equations.

Chen:2023:GEW

[1299] Yajie Chen, Yingting Miao, and Shijie Shi. Global existence and wave breaking for a stochastic two-component Camassa–Holm system. Journal of Mathematical Physics, 64(1):011505, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011505/2870479/Global-existence-and-wave-breaking-for-a.

An:2023:CPL

[1300] Xiaoming An and Xian Yang. Convergence from power-law to logarithmlaw in nonlinear fractional Schrödinger equations. Journal of Mathematical Physics, 64(1):011506, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/1/011506/2870446/Convergence-from-power-
law-to-logarithm-law-in.

Lee:2023:SSE

[1301] Ho Lee, Jiho Lee, and Ernesto Nungesser. Small solutions of the Einstein– Boltzmann-scalar field system with Bianchi symmetry. Journal of Mathematical Physics, 64(1):011507, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/1/011507/2870473/Small-solutions-of-the-
Einstein-Boltzmann-scalar.

Wang:2023:FDS

[1302] Yixuan Wang and Meina Sun. Formation of delta shock and vacuum state for the pressureless hydrodynamic model under the small disturbance of traffic pressure. Journal of Mathematical Physics, 64(1): 011508, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
1/011508/2870497/Formation-of-delta-shock-and-vacuum-state-
for-the.

Zhang:2023:QPS

[1303] Min Zhang, Yi Wang, and Jie Rui. Quasi-periodic solutions for one dimensional Schrödinger equation with quasi-periodic forcing and Dirichlet boundary condition. Journal of Mathematical Physics, 64(1):011509, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011509/2870461/Quasi-periodic-solutions-for-one-dimensional.

Saanouni:2023:CIC

[1304] Tarek Saanouni and Radhia Ghanmi. Critical inhomogeneous coupled Schrödinger equations. Journal of Mathematical Physics, 64(1):011510, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
1/011510/2870444/Critical-inhomogeneous-coupled-Schrodinger.

Li:2023:NKT

[1305] Siran Li and Marshall Slemrod. From the Nash–Kuiper theorem of isometric embeddings to the Euler equations for steady fluid motions: Analogues, examples, and extensions. Journal of Mathematical Physics, 64(1):011511, January 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/1/011511/2870447/From-the-
Nash-Kuiper-theorem-of-isometric.

Bahrouni:2023:PIF

[1306] Sabri Bahrouni, Hichem Ounaies, and Olfa Elfalah. Problems involving the fractional g-Laplacian with lack of compactness. Journal of Mathematical Physics, 64(1):011512, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/1/011512/2870456/Problems-involving-the-
fractional-g-Laplacian-with.

Liu:2023:AEW

[1307] Bingchen Liu, Ke Li, and Fengjie Li. Asymptotic estimate of weak solutions in a fourth-order parabolic equation with logarithm. Journal of Mathematical Physics, 64(1):011513, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011513/2870504/Asymptotic-estimate-of-weak-solutions-in-a-fourth.

Yuan:2023:ETG

[1308] Yiyin Yuan, Shuai Tian, Jun Qing, and Shihui Zhu. Exact thresholds for global existence to the nonlinear beam equations with and without a damping. Journal of Mathematical Physics, 64(1):011514, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011514/2870466/Exact-thresholds-for-global-existence-to-the.

daVeiga:2023:FDD

[1309] Hugo Beirão da Veiga and Jiaqi Yang. Fully developed, doubly periodic, viscous flows in infinite space-periodic pipes under general time-periodic total fluxes. Journal of Mathematical Physics, 64(1):011515, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/1/011515/2870440/Fully-developed-doubly-
periodic-viscous-flows-in.

Zhang:2023:GWP

[1310] Jingjing Zhang and Ting Zhang. Global well-posedness of perturbed Navier–Stokes system around Landau solutions. Journal of Mathematical Physics, 64(1):011516, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/1/011516/2870454/Global-well-posedness-of-perturbed-
Navier-Stokes.

Braden:2023:CES

[1311] H. W. Braden, Sergey A. Cherkis, and Jason M. Quinones. Construction of exact solutions to Nahm’s equations for the multimonopole. Journal of Mathematical Physics, 64(1):011701, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011701/2870582/Construction-of-exact-solutions-to-Nahm-s.

Garbali:2023:CMQ

[1312] Alexandr Garbali and Andrei Neguţ. Computing the r-matrix of the quantum toroidal algebra. Journal of Mathematical Physics, 64(1):011702, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/1/011702/2870589/Computing-the-R-matrix-
of-the-quantum-toroidal.

Lee:2023:DRC

[1313] Gahng Sahn Lee, Arim Song, and Uhi Rinn Suh. Dirac reductions and classical W -algebras. Journal of Mathematical Physics, 64(1):011703, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
1/011703/2870613/Dirac-reductions-and-classical-W-algebras.

Gubarev:2023:CYB

[1314] Vsevolod Gubarev and Roman Kozlov. Conformal Yang–Baxter equation on Cur(sl_{2}(C)). Journal of Mathematical Physics, 64(1):011704, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011704/2870604/Conformal-Yang-Baxter-equation-on-Cur-sl2-C.

Falconi:2023:DMS

[1315] Marco Falconi and Nikolai Leopold. Derivation of the Maxwell– Schrödinger equations: a note on the infrared sector of the radiation field. Journal of Mathematical Physics, 64(1):011901, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011901/2870572/Derivation-of-the-Maxwell-Schrodinger-equations-
A.

Monaco:2023:TVL

[1316] Domenico Monaco and Thaddeus Roussigné. Topology vs localization in synthetic dimensions. Journal of Mathematical Physics, 64(1):011902, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
011902/2870597/Topology-vs-localization-in-synthetic-dimensions.

Mayer:2023:NLO

[1317] L. P. Mayer. Non-local order parameters for fermion chains via the partial transpose. Journal of Mathematical Physics, 64(1):011903, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/1/011903/2870617/Non-local-order-parameters-for-fermion-
chains-via.

Qureshi:2023:EWP

[1318] Izma Qureshi, Tasawar Abbas, Muhammad Imran, and Rameez ul Islam. Exploring wave–particle behaviors of entangled Bragg diffracted neutral atoms. Journal of Mathematical Physics, 64(1):012101, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
012101/2870592/Exploring-wave-particle-behaviors-of-entangled.

daCosta:2023:ESC

[1319] Bruno G. da Costa, Ignacio S. Gomez, and Biswanath Rath. Exact solution and coherent states of an asymmetric oscillator with positiondependent mass. Journal of Mathematical Physics, 64(1):012102, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
012102/2870607/Exact-solution-and-coherent-states-of-an.

Golovnev:2023:VPC

[1320] Alexey Golovnev. The variational principle, conformal and disformal transformations, and the degrees of freedom. Journal of Mathematical Physics, 64(1):012501, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/1/012501/2870561/The-variational-principle-conformal-
and-disformal.

Sakovich:2023:NDE

[1321] A. Sakovich and C. Sormani. The null distance encodes causality. Journal of Mathematical Physics, 64(1):012502, January 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/1/012502/2870555/The-null-
distance-encodes-causality.

Mitsopoulos:2023:CFI

[1322] Antonios Mitsopoulos and Michael Tsamparlis. Cubic first integrals of autonomous dynamical systems in E^{2} by an algorithmic approach. Journal of Mathematical Physics, 64(1):012701, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
012701/2870562/Cubic-first-integrals-of-autonomous-dynamical.

Liu:2023:ASH

[1323] Qun Liu. Analysis of a stochastic HIV model with cell-to-cell transmission and Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 64(1):012702, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/1/012702/2870575/Analysis-
of-a-stochastic-HIV-model-with-cell-to.

Yang:2023:LBC

[1324] Juan Yang, Jiaxin Gong, Longyu Wu, and Ji Shu. Limiting behavior of center manifolds for stochastic evolutionary equations with delay in varying phase spaces. Journal of Mathematical Physics, 64(1): 012703, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
1/012703/2870586/Limiting-behavior-of-center-manifolds-for.

Zhang:2023:STW

[1325] Ran Zhang and Hongyong Zhao. Strong traveling wave solutions for a nonlocal diffusive susceptible–infectious–recovered model with spatiotemporal delay. Journal of Mathematical Physics, 64(1):012704, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
012704/2870657/Strong-traveling-wave-solutions-for-a-nonlocal.

Ahn:2023:RFU

[1326] Hyunjin Ahn, Junhyeok Byeon, Seung-Yeal Ha, and Jaeyoung Yoon. On the relativistic flocks over the unit sphere and the hyperboloid in a bonding force field. Journal of Mathematical Physics, 64 (1):012705, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/1/012705/2870664/On-the-relativistic-flocks-
over-the-unit-sphere.

Zessin:2023:CDI

[1327] Hans Zessin. On cluster dynamics of infinite Newtonian dynamics. Journal of Mathematical Physics, 64(1):012706, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic).  URL https://pubs.aip.org/aip/jmp/article/64/
1/012706/2870649/On-cluster-dynamics-of-infinite-Newtonian-
dynamics.

Goto:2023:AGD

[1328] Shin itiro Goto. Affine geometric description of thermodynamics. Journal of Mathematical Physics, 64(1):013301, January 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/1/013301/2870677/Affine-geometric-
description-of-thermodynamics.

Hansen:2023:SMC

[1329] Ulrik Thinggaard Hansen and Frederik Ravn Klausen. Strict monotonicity, continuity, and bounds on the Kertész line for the randomcluster model on Z^{d}. Journal of Mathematical Physics, 64(1):013302, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
1/013302/2870674/Strict-monotonicity-continuity-and-bounds-
on-the.

Gamba:2023:CPB

[1330] Irene M. Gamba and Milana Pavić-Čolić. On the Cauchy problem for Boltzmann equation modeling a polyatomic gas. Journal of Mathematical Physics, 64(1):013303, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/1/013303/2870688/On-the-Cauchy-problem-for-Boltzmann-
equation.

delRazo:2023:CDM

[1331] Mauricio J. del Razo, Stefanie Winkelmann, Rupert Klein, and Felix Höfling. Chemical diffusion master equation: Formulations of reaction– diffusion processes on the molecular level. Journal of Mathematical Physics, 64(1):013304, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/1/013304/2870691/Chemical-diffusion-
master-equation-Formulations-of.

Huang:2023:QFP

[1332] Fang Huang and Chuanzhong Li. Quantum field presentation for generalized Hall–Littlewood functions. Journal of Mathematical Physics, 64(1):013501, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/1/013501/2870660/Quantum-field-presentation-for-
generalized-Hall.

Ding:2023:TFW

[1333] Xiang-Mao Ding and Xiang Li. 2D Toda τ functions, weighted Hurwitz numbers and the Cayley graph: Determinant representation and recursion formula. Journal of Mathematical Physics, 64 (1):013502, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/1/013502/2870667/2D-Toda-functions-weighted-
Hurwitz-numbers-and-the.

Figueroa-OFarrill:2023:LAC

[1334] José Figueroa-O’Farrill. Lie algebraic Carroll/ Galilei duality. Journal of Mathematical Physics, 64(1):013503, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/1/013503/2870683/Lie-algebraic-Carroll-
Galilei-duality.

Liu:2023:ASL

[1335] Qun Liu. Analysis of a stochastic Lotka–Volterra competitive system with infinite delays and Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 64(1):013504, January 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/1/013504/2870647/Analysis-
of-a-stochastic-Lotka-Volterra.

Torres:2023:CNT

[1336] Delfim F. M. Torres. Comment on “Noether’s-type theorems on time scales” [J. Math. Phys. 61, 113502 (2020)]. Journal of Mathematical Physics, 64(1):014101, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/1/014101/2870654/Comment-on-Noether-
s-type-theorems-on-time-scales. See [382].

Bet:2023:EMD

[1337] Gianmarco Bet, Anna Gallo, and Francesca R. Nardi. Erratum: “Metastability for the degenerate Potts model with negative external magnetic field under Glauber dynamics” [J. Math. Phys. 63, 123303 (2022)]. Journal of Mathematical Physics, 64(1):019901, January 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/1/
019901/2870659/Erratum-Metastability-for-the-degenerate-Potts.

See [1279].

Diamantini:2023:BTI

[1338] M. C. Diamantini and C. A. Trugenberger. Bosonic topological insulators at the superconductor-to-superinsulator transition. Journal of Mathematical Physics, 64(2):021101, February 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/2/021101/2873881/Bosonic-
topological-insulators-at-the.

Ding:2023:VPL

[1339] Shijin Ding and Cuiyu Wang. Validity of Prandtl layer expansions for steady magnetohydrodynamics over a rotating disk. Journal of Mathematical Physics, 64(2):021501, February 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/2/021501/2873964/Validity-
of-Prandtl-layer-expansions-for-steady.

Yu:2023:GSC

[1340] Yanghai Yu, Hui Wang, Jinlu Li, and Xiaolei Yang. Global solutions to the 3D compressible Navier–Stokes equations with a class of special initial data. Journal of Mathematical Physics, 64 (2):021502, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/021502/2873937/Global-solutions-to-the-3D-
compressible-Navier.

Shi:2023:AAC

[1341] Xinran Shi, Yunfei Su, and Lei Yao. Asymptotic analysis for 1D compressible Navier–Stokes–Vlasov equations with local alignment force. Journal of Mathematical Physics, 64(2):021503, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
021503/2873944/Asymptotic-analysis-for-1D-compressible-Navier.

Chen:2023:APS

[1342] Wenhui Chen and Yan Liu. Asymptotic profiles and singular limits for the viscoelastic damped wave equation with memory of type I. Journal of Mathematical Physics, 64(2):021504, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
021504/2873951/Asymptotic-profiles-and-singular-limits-for-
the.

Wang:2023:GWP

[1343] Fuguo Wang and Xin Zhong. Global well-posedness to the compressible nematic liquid crystal flows with large oscillations and vacuum in 3D exterior domains. Journal of Mathematical Physics, 64(2):021505, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
2/021505/2873957/Global-well-posedness-to-the-compressible-
nematic.

Chen:2023:DWS

[1344] Honghua Chen, Geng Lai, and Wancheng Sheng. Detonation wave solutions to the scalar combustion models with a singular source term. Journal of Mathematical Physics, 64(2):021506, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
021506/2873959/Detonation-wave-solutions-to-the-scalar-combustion.

Sun:2023:PPS

[1345] Baoyan Sun. Propagation phenomena of the solution for the relativistic BGK model. Journal of Mathematical Physics, 64(2):021507, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
2/021507/2873915/Propagation-phenomena-of-the-solution-for-
the.

Cho:2023:CCA

[1346] Yonggeun Cho, Seokchang Hong, and Tohru Ozawa. Charge conjugation approach to scattering for the Hartree type Dirac equations with chirality. Journal of Mathematical Physics, 64(2):021508, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
2/021508/2873890/Charge-conjugation-approach-to-scattering-
for-the.

Liu:2023:NPT

[1347] Zhongyuan Liu, Ziying Liu, and Wenhuan Xu. Non-power type perturbation for the critical Hénon problem. Journal of Mathematical Physics, 64(2):021509, February 2023. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/2/021509/2873956/Non-power-type-
perturbation-for-the-critical-Henon.

Cao:2023:MPS

[1348] Daomin Cao, Shanfa Lai, and Weilin Yu. Multi-peak solutions to the Schrödinger equations coupled with a neutral scalar field. Journal of Mathematical Physics, 64(2):021510, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/2/021510/2873877/Multi-peak-solutions-
to-the-Schrodinger-equations.

Han:2023:SEI

[1349] Xiaoli Han, Xishen Jin, and Yang Wen. Stability and energy identity for Yang–Mills–Higgs pairs. Journal of Mathematical Physics, 64(2):021511, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/2/021511/2873954/Stability-and-energy-identity-for-Yang-
Mills-Higgs.

Hahn:2023:WNS

[1350] Nico Hahn, Mario Kieburg, Omri Gat, and Thomas Guhr. Winding number statistics for chiral random matrices: Averaging ratios of determinants with parametric dependence. Journal of Mathematical Physics, 64(2):021901, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/2/021901/2873897/Winding-number-
statistics-for-chiral-random.

Cornean:2023:BEC

[1351] H. D. Cornean, M. Moscolari, and K. S. Sørensen. Bulk–edge correspondence for unbounded Dirac–Landau operators. Journal of Mathematical Physics, 64(2):021902, February 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/2/021902/2873934/Bulk-edge-
correspondence-for-unbounded-Dirac.

Muller:2023:SST

[1352] Peter Müller and Ruth Schulte. Stability of a Szegő-type asymptotics. Journal of Mathematical Physics, 64(2):022101, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
022101/2873875/Stability-of-a-Szego-type-asymptotics.

Sasaki:2023:ESP

[1353] Ryu Sasaki. Exactly solvable piecewise analytic double well potential V_{D}(x) = min[(x + d)^{2}, (x − d)^{2}] and its dual single well potential V_{S}(x) = max[(x + d)^{2}, (x − d)^{2}]. Journal of Mathematical Physics, 64(2):022102, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/022102/2873941/Exactly-solvable-piecewise-
analytic-double-well.

DelasCuevas:2023:MSL

[1354] Gemma De las Cuevas, Tim Netzer, and Inga Valentiner-Branth. Magic squares: Latin, semiclassical, and quantum. Journal of Mathematical Physics, 64(2):022201, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/022201/2873948/Magic-squares-Latin-semiclassical-
and-quantum.

DeBievre:2023:RIN

[1355] S. De Bièvre. Relating incompatibility, noncommutativity, uncertainty, and Kirkwood–Dirac nonclassicality. Journal of Mathematical Physics, 64(2):022202, February 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/2/022202/2873952/Relating-
incompatibility-noncommutativity.

Zhu:2023:HAT

[1356] Honglin Zhu. The Hermitian axiom on two-dimensional topological quantum field theories. Journal of Mathematical Physics, 64(2):022301, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/2/022301/2874008/The-Hermitian-axiom-on-two-dimensional-
topological.

Celerier:2023:FII

[1357] M.-N. Célérier. Fully integrated interior solutions of GR for stationary rigidly rotating cylindrical perfect fluids. Journal of Mathematical Physics, 64(2):022501, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/022501/2874048/Fully-integrated-interior-solutions-
of-GR-for.

Heefer:2023:MKS

[1358] Sjors Heefer, Christian Pfeifer, Jorn van Voorthuizen, and Andrea Fuster. On the metrizability of m-Kropina spaces with closed null one-form. Journal of Mathematical Physics, 64(2):022502, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
022502/2874027/On-the-metrizability-of-m-Kropina-spaces-with.

Coley:2023:TGS

[1359] A. A. Coley and R. J. van den Hoogen. Teleparallel geometry with a single affine symmetry. Journal of Mathematical Physics, 64(2):022503, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
2/022503/2874051/Teleparallel-geometry-with-a-single-affine.

Moore:2023:BER

[1360] Kenneth Moore and Eric Woolgar. Bakry-Émery Ricci curvature, Xminimal hypersurfaces, and near horizon geometries. Journal of Mathematical Physics, 64(2):022504, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/2/022504/2874038/Bakry-Emery-Ricci-curvature-
X-minimal.

Xie:2023:UAN

[1361] Yongqin Xie, Di Liu, Jiangwei Zhang, and Ximeng Liu. Uniform attractors for nonclassical diffusion equations with perturbed parameter and memory. Journal of Mathematical Physics, 64 (2):022701, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/2/022701/2874054/Uniform-attractors-for-
nonclassical-diffusion.

Chang:2023:WPD

[1362] Qingquan Chang and Dandan Li. Well-posedness and dynamics of 2D Navier–Stokes equations with moving boundary. Journal of Mathematical Physics, 64(2):022702, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/2/022702/2874013/Well-posedness-and-
dynamics-of-2D-Navier-Stokes.

Cao:2023:SCN

[1363] Nan Cao and Xianlong Fu. On solutions of a class of neutral evolution equations with discrete nonlocal conditions. Journal of Mathematical Physics, 64(2):022703, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/022703/2874058/On-solutions-of-a-class-of-
neutral-evolution.

Arslan:2023:RSL

[1364] Ali Arslan, Giovanni Fantuzzi, John Craske, and Andrew Wynn. Rigorous scaling laws for internally heated convection at infinite Prandtl number. Journal of Mathematical Physics, 64(2):023101, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
023101/2874060/Rigorous-scaling-laws-for-internally-heated.

Bain:2023:LCF

[1365] Emily Bain. Local correlation functions of the two-periodic weighted Aztec diamond in mesoscopic limit. Journal of Mathematical Physics, 64(2):023301, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/2/023301/2874072/Local-correlation-functions-of-
the-two-periodic.

Zamparo:2023:LDQ

[1366] Marco Zamparo and Massimiliano Semeraro. Large deviations for quadratic functionals of stable Gauss–Markov chains and entropy production. Journal of Mathematical Physics, 64(2):023302, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
023302/2874002/Large-deviations-for-quadratic-functionals-of.

Poghosyan:2023:UAP

[1367] Suren Poghosyan and Hans Zessin. Uniqueness and absence of percolation of classical gases. Journal of Mathematical Physics, 64(2):023303, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/2/023303/2874010/Uniqueness-and-absence-
of-percolation-of-classical.

Baez:2023:CT

[1368] John C. Baez, Owen Lynch, and Joe Moeller. Compositional thermostatics. Journal of Mathematical Physics, 64(2):023304, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
023304/2874017/Compositional-thermostatics.

Cerjan:2023:QPI

[1369] Alexander Cerjan, Terry A. Loring, and Fredy Vides. Quadratic pseudospectrum for identifying localized states. Journal of Mathematical Physics, 64(2):023501, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/2/023501/2874032/Quadratic-
pseudospectrum-for-identifying-localized.

Arthamonov:2023:TFI

[1370] S. Arthamonov, J. Harnad, and J. Hurtubise. Tau functions, infinite Grassmannians, and lattice recurrences. Journal of Mathematical Physics, 64(2):023502, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/2/023502/2874044/Tau-functions-infinite-
Grassmannians-and-lattice.

Akemann:2023:EGE

[1371] G. Akemann, M. Duits, and L. D. Molag. The elliptic Ginibre ensemble: a unifying approach to local and global statistics for higher dimensions. Journal of Mathematical Physics, 64(2):023503, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
023503/2874065/The-elliptic-Ginibre-ensemble-A-unifying-approach.

Special collection in honor of Freeman Dyson.

Alcock-Zeilinger:2023:WSS

[1372] Judith Alcock-Zeilinger, Stefan Keppeler, Simon Plätzer, and Malin Sjodahl. Wigner 6j symbols for SU(N): Symbols with at least two quark-lines. Journal of Mathematical Physics, 64(2): 023504, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/2/023504/2874159/Wigner-6j-symbols-for-SU-N-
Symbols-with-at-least.

Meljanac:2023:NYM

[1373] S. Meljanac and S. Mignemi. Noncommutative Yang model and its generalizations. Journal of Mathematical Physics, 64(2):023505, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/2/
023505/2874152/Noncommutative-Yang-model-and-its-generalizations.

Katkov:2023:EMM

[1374] Mikhail Katkov, Misha Tsodyks, Michelangelo Naim, and Antonis Georgiou. Erratum: “Mathematical models of human memory” [J. Math. Phys. 63, 073303 (2022)]. Journal of Mathematical Physics, 64(2):029901, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/2/029901/2874149/Erratum-Mathematical-models-of-
human-memory-J-Math. See [1090].

Yoldas:2023:QHE

[1375] Havva Yoldaş. On quantitative hypocoercivity estimates based on Harris-type theorems. Journal of Mathematical Physics, 64 (3):031101, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/3/031101/2881551/On-quantitative-hypocoercivity-
estimates-based-on.

Gu:2023:ILI

[1376] Xiaoyu Gu, Yaobin Ou, and Lu Yang. Incompressible limit of isentropic magnetohydrodynamic equations with ill-prepared data in bounded domains. Journal of Mathematical Physics, 64(3):031501, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/031501/2881547/Incompressible-limit-of-isentropic.

Watson:2023:BMD

[1377] Alexander B. Watson, Tianyu Kong, Allan H. MacDonald, and Mitchell Luskin. Bistritzer–MacDonald dynamics in twisted bilayer graphene. Journal of Mathematical Physics, 64(3):031502, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031502/2881557/Bistritzer-MacDonald-dynamics-in-twisted-bilayer.

Gao:2023:ERF

[1378] Liu Gao and Zhong Tan. Existence results for fractional Kirchhoff problems with magnetic field and supercritical growth. Journal of Mathematical Physics, 64(3):031503, March 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/3/031503/2881539/Existence-
results-for-fractional-Kirchhoff.

Tan:2023:GSE

[1379] Lihua Tan and Yingzhe Fan. Global solutions and exponential time decay rates to the Navier–Stokes–Vlasov–Fokker–Planck system in low regularity space. Journal of Mathematical Physics, 64(3):031504, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031504/2881534/Global-solutions-and-exponential-time-decay-
rates.

Zhao:2023:GEI

[1380] Zhiwen Zhao. Gradient estimates for the insulated conductivity problem: The case of m-convex inclusions. Journal of Mathematical Physics, 64(3):031505, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/3/031505/2881527/Gradient-estimates-for-the-
insulated-conductivity.

Zhao:2023:SSN

[1381] Xiaopeng Zhao. On the strong solution of 3D non-isothermal Navier–Stokes–Cahn–Hilliard equations. Journal of Mathematical Physics, 64(3):031506, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/3/031506/2881545/On-the-strong-solution-
of-3D-non-isothermal-Navier.

Chong:2023:OSP

[1382] Gezi Chong and Ying Fu. Orbital stability of periodic peakons for a new higher-order μ-Camassa–Holm equation. Journal of Mathematical Physics, 64(3):031507, March 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/3/031507/2881520/Orbital-
stability-of-periodic-peakons-for-a-new.

Bal:2023:TCC

[1383] Guillaume Bal. Topological charge conservation for continuous insulators. Journal of Mathematical Physics, 64(3):031508, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031508/2881542/Topological-charge-conservation-for-continuous.

Niu:2023:LUM

[1384] Yahui Niu, Shuying Tian, and Pingping Yang. Local uniqueness of multi-peak solutions to a class of Schrödinger equations with competing potential. Journal of Mathematical Physics, 64(3):031509, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031509/2881526/Local-uniqueness-of-multi-peak-solutions-to-
a.

Jarrin:2023:RWS

[1385] Oscar Jarrín. Regularity of weak solutions for the stationary Ericksen– Leslie and MHD systems. Journal of Mathematical Physics, 64(3): 031510, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/031510/2881543/Regularity-of-weak-solutions-for-the-stationary.

Ma:2023:SLM

[1386] Guorui Ma, Stephen S.-T. Yau, and Huaiqing Zuo. k-th singular locus moduli algebras of singularities and their derivation Lie algebras. Journal of Mathematical Physics, 64(3):031701, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031701/2881524/k-th-singular-locus-moduli-algebras-of.

Ceniceros:2023:RFO

[1387] Jose Ceniceros, Mohamed Elhamdadi, Brendan Magill, and Gabriana Rosario. RNA foldings, oriented stuck knots, and state sum invariants. Journal of Mathematical Physics, 64(3):031702, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031702/2881559/RNA-foldings-oriented-stuck-knots-and-state-
sum.

Coquereaux:2023:AIV

[1388] Robert Coquereaux. About integer-valued variants of the theta and 6j symbols. Journal of Mathematical Physics, 64(3):031703, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031703/2881531/About-integer-valued-variants-of-the-theta-and-
6j.

Maekawa:2023:AVB

[1389] Daisuke Maekawa and Hal Tasaki. The asymmetric valence-bond-solid states in quantum spin chains: The difference between odd and even spins. Journal of Mathematical Physics, 64(3):031901, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031901/2881535/The-asymmetric-valence-bond-solid-states-in.

Kellendonk:2023:BST

[1390] Johannes Kellendonk. Bragg spectrum, K-theory, and gap labeling of aperiodic solids. Journal of Mathematical Physics, 64(3):031902, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031902/2881615/Bragg-spectrum-K-theory-and-gap-labeling-of.

Frohlich:2023:GIA

[1391] Jürg Fröhlich. Gauge invariance and anomalies in condensed matter physics. Journal of Mathematical Physics, 64(3):031903, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
031903/2881632/Gauge-invariance-and-anomalies-in-condensed-
matter.

Singh:2023:TGS

[1392] Vibhav Narayan Singh, Mohammad Umar, Mohammad Hasan, and Bhabani Prasad Mandal. Tunneling from general Smith–Volterra– Cantor potential. Journal of Mathematical Physics, 64(3):032101, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032101/2881636/Tunneling-from-general-Smith-Volterra-Cantor.

Stan:2023:SRV

[1393] A. I. Stan, G. Popa, and R. Dutta. A study of random variables in terms of the number operator. Journal of Mathematical Physics, 64(3):032102, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032102/2881614/A-study-of-random-variables-in-terms-of-the-
number.

Szabados:2023:MCS

[1394] László B. Szabados. The “most classical” states of Euclidean invariant elementary quantum mechanical systems. Journal of Mathematical Physics, 64(3):032103, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/3/032103/2881596/The-most-classical-
states-of-Euclidean-invariant.

Szabados:2023:OFM

[1395] László B. Szabados. An odd feature of the “most classical” states of SU(2) invariant quantum mechanical systems. Journal of Mathematical Physics, 64(3):032104, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/3/032104/2881627/An-odd-feature-of-the-most-
classical-states-of-SU.

Giraldi:2023:TGC

[1396] Filippo Giraldi and Francesco Mainardi. Truncated generalized coherent states. Journal of Mathematical Physics, 64(3):032105, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032105/2881601/Truncated-generalized-coherent-states.

Ruskai:2023:LAR

[1397] Mary Beth Ruskai and Jon Yard. Local additivity revisited. Journal of Mathematical Physics, 64(3):032201, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/3/032201/2881610/Local-additivity-

revisited. Special collection in honor of Freeman Dyson.

Blanco:2023:DFA

[1398] David Serrano Blanco. Dirac field in AdS_{2} and representations of S̃L(2, R). Journal of Mathematical Physics, 64(3):032301, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032301/2881631/Dirac-field-in-AdS2-and-representations-of-SL-
2-R.

Goswami:2023:MGH

[1399] Abhishek Goswami. Mass gap in U(1) Higgs–Yukawa model on a unit lattice. Journal of Mathematical Physics, 64(3):032302, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032302/2881624/Mass-gap-in-U-1-Higgs-Yukawa-model-on-a-unit.

Mastropietro:2023:ACL

[1400] Vieri Mastropietro. Anomaly cancellation in the lattice effective electroweak theory. Journal of Mathematical Physics, 64(3):032303, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032303/2881594/Anomaly-cancellation-in-the-lattice-effective.

Escudero:2023:HDC

[1401] Valentina Guarín Escudero, Cristhiam Lopez-Arcos, and Alexander Quintero Vélez. Homotopy double copy and the Kawai–Lewellen–Tye relations for the non-abelian and tensor Navier–Stokes equations. Journal of Mathematical Physics, 64(3):032304, March 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/3/032304/2881598/Homotopy-
double-copy-and-the-Kawai-Lewellen-Tye.

Celerier:2023:SSRa

[1402] M.-N. Célérier. Study of stationary rigidly rotating anisotropic cylindrical fluids with new exact interior solutions of GR. II. More about axial pressure. Journal of Mathematical Physics, 64(3):032501, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032501/2881604/Study-of-stationary-rigidly-rotating-anisotropic.

Keeton:2023:MGL

[1403] Charles R. Keeton, Erik Lundberg, and Sean Perry. Multiplane gravitational lenses with an abundance of images. Journal of Mathematical Physics, 64(3):032502, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/3/032502/2881612/Multiplane-gravitational-lenses-
with-an-abundance.

McNutt:2023:FBA

[1404] D. D. McNutt, A. A. Coley, and R. J. van den Hoogen. A frame based approach to computing symmetries with non-trivial isotropy groups. Journal of Mathematical Physics, 64(3):032503, March 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/3/032503/2881713/A-frame-
based-approach-to-computing-symmetries.

Natario:2023:EFD

[1405] José Natário and Flavio Rossetti. Explicit formulas and decay rates for the solution of the wave equation in cosmological spacetimes. Journal of Mathematical Physics, 64(3):032504, March 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/3/032504/2881735/Explicit-
formulas-and-decay-rates-for-the-solution.

Geng:2023:QSN

[1406] Jiansheng Geng, Yingnan Sun, and W.-M. Wang. Quasiperiodic solutions to nonlinear random Schrödinger equations at fixed potential realizations. Journal of Mathematical Physics, 64(3):032701, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
032701/2881732/Quasiperiodic-solutions-to-nonlinear-random.

Chen:2023:GTH

[1407] Desu Chen, Chao Wang, and Zhien Li. General theory of the higherorder linear quaternion q-difference equations through the quaternion determinant algorithm and the characteristic polynomial. Journal of Mathematical Physics, 64(3):032702, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/3/032702/2881728/General-theory-
of-the-higher-order-linear.

Ryan:2023:WAL

[1408] Joseph Ryan. When action is not least for systems with actiondependent Lagrangians. Journal of Mathematical Physics, 64(3):032901, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/032901/2881724/When-action-is-not-least-for-systems-with-
action.

Kang:2023:MPP

[1409] Yuan Bao Kang. Markov properties of partially open quantum random walks. Journal of Mathematical Physics, 64(3):033301, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
033301/2881684/Markov-properties-of-partially-open-quantum-
random.

Azuaje:2023:CCT

[1410] R. Azuaje and A. M. Escobar-Ruiz. Canonical and canonoid transformations for Hamiltonian systems on (co)symplectic and (co)contact manifolds. Journal of Mathematical Physics, 64(3):033501, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
033501/2881689/Canonical-and-canonoid-transformations-for.

Dinar:2023:LDB

[1411] Yassir Dinar. Low-dimensional bihamiltonian structures of topological type. Journal of Mathematical Physics, 64(3):033502, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
033502/2881704/Low-dimensional-bihamiltonian-structures-of.

DeNittis:2023:NLF

[1412] Giuseppe De Nittis and Kiyonori Gomi. A new light on the FKMM invariant and its consequences. Journal of Mathematical Physics, 64(3): 033503, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/033503/2881719/A-new-light-on-the-FKMM-invariant-and-its.

Carstea:2023:SAB

[1413] A. S. Carstea. Singularity analysis and bilinear approach to some Bogoyavlensky equations. Journal of Mathematical Physics, 64(3):033504, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/3/
033504/2881718/Singularity-analysis-and-bilinear-approach-to-
some.

DellAtti:2023:CDN

[1414] Marta Dell’Atti and Pierandrea Vergallo. Classification of degenerate non-homogeneous Hamiltonian operators. Journal of Mathematical Physics, 64(3):033505, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/3/033505/2881679/Classification-
of-degenerate-non-homogeneous.

Alhaidari:2023:RCP

[1415] A. D. Alhaidari. Revisiting the Coulomb problem: a novel representation of the confluent hypergeometric function as an infinite sum of discrete Bessel functions. Journal of Mathematical Physics, 64(3): 033506, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/033506/2881686/Revisiting-the-Coulomb-problem-A-novel.

Rivas:2023:NCF

[1416] Xavier Rivas. Nonautonomous k-contact field theories. Journal of Mathematical Physics, 64(3):033507, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/3/033507/2881697/Nonautonomous-k-contact-
field-theories.

Liu:2023:PLL

[1417] Wencai Liu. Power law logarithmic bounds of moments for long range operators in arbitrary dimension. Journal of Mathematical Physics, 64(3): 033508, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/033508/2881706/Power-law-logarithmic-bounds-of-moments-for-
long.

Adler:2023:DIR

[1418] Mark Adler and Pierre van Moerbeke. Double interlacing in random tiling models. Journal of Mathematical Physics, 64(3):033509, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
3/033509/2881737/Double-interlacing-in-random-tiling-models.

Special collection in honor of Freeman Dyson.

Osada:2023:EUD

[1419] Hirofumi Osada and Shota Osada. Ergodicity of unlabeled dynamics of Dyson’s model in infinite dimensions. Journal of Mathematical Physics, 64(3):043505, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/4/043505/2877773/Ergodicity-of-unlabeled-dynamics-
of-Dyson-s-model?searchresult=1. Special collection in honor of

Freeman Dyson.

Tasaki:2023:RIT

[1420] Hal Tasaki. Rigorous index theory for one-dimensional interacting topological insulators. Journal of Mathematical Physics, 64(4):041903, April 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
4/041903/2883460/Rigorous-index-theory-for-one-dimensional.

An:2023:LTT

[1421] Xiaowei An, Huixia He, and Xianfa Song. Liouville type theorem and Morse indices of a semilinear elliptic equation with nonlocal nonlinearities. Journal of Mathematical Physics, 64(5):051501, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
051501/2887571/Liouville-type-theorem-and-Morse-indices-of-
a.

Wang:2023:RCA

[1422] Peng Wang and Zhengguang Guo. A regularity criterion for the 3D axisymmetric Boussinesq equations with non-zero swirl. Journal of Mathematical Physics, 64(5):051502, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/5/051502/2889010/A-regularity-criterion-
for-the-3D-axisymmetric.

Zhang:2023:RHP

[1423] Xiao-Fan Zhang and Shou-Fu Tian. Riemann–Hilbert problem for the Fokas–Lenells equation in the presence of high-order discrete spectrum with non-vanishing boundary conditions. Journal of Mathematical Physics, 64(5):051503, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/5/051503/2889011/Riemann-
Hilbert-problem-for-the-Fokas-Lenells.

Chatta:2023:LWP

[1424] Soufiane Chatta, Boualem Khouider, and M’hamed Kesri. Linear well posedness of regularized equations of sea-ice dynamics. Journal of Mathematical Physics, 64(5):051504, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/5/051504/2889492/Linear-well-posedness-of-
regularized-equations-of.

Wang:2023:RCG

[1425] Jinhuan Wang, Wei Tan, and Yongsheng Nie. A regularity criterion for the 3D generalized MHD system involving partial components. Journal of Mathematical Physics, 64(5):051505, May 1, 2023. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/64/5/051505/2890217/A-
regularity-criterion-for-the-3D-generalized-MHD.

Zhai:2023:OWP

[1426] Xiaoping Zhai, Yiren Chen, Yongsheng Li, and Yongye Zhao. Optimal well-posedness for the pressureless Euler–Navier–Stokes system. Journal of Mathematical Physics, 64(5):051506, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/5/051506/2890427/Optimal-
well-posedness-for-the-pressureless-Euler.

Ge:2023:RNN

[1427] Bin Ge, Qinghai Cao, and Yu Zhang. Renormalized non-negative solutions for the double phase Dirichlet problems with L^{1} data. Journal of Mathematical Physics, 64(5):051507, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/5/051507/2890776/Renormalized-non-
negative-solutions-for-the-double.

Tang:2023:MMI

[1428] Zhongwei Tang, Lushun Wang, and Huafei Xie. Multiple mixed interior and boundary peaks synchronized solutions for nonlinear coupled elliptic systems. Journal of Mathematical Physics, 64(5):051508, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
5/051508/2891376/Multiple-mixed-interior-and-boundary-peaks.

Zhao:2023:EDC

[1429] Jin-Wei Zhao, Bin Ge, and Lu Liu. Effective dynamics for a class of stochastic weakly damped wave equation with a fast oscillation. Journal of Mathematical Physics, 64(5):051509, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
051509/2891412/Effective-dynamics-for-a-class-of-stochastic.

dosSantos:2023:SQS

[1430] Gelson C. G. dos Santos, Laila C. Fontinele, Rubia G. Nascimentoa, and Suellen Cristina Q. Arrudab. Solutions for a quasilinear Schrödinger equation: Subcritical and critical cases. Journal of Mathematical Physics, 64(5):051510, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/5/051510/2892276/Solutions-
for-a-quasilinear-Schrodinger-equation.

Liu:2023:QPS

[1431] Yang Liu, Byungsoo Moon, Vicenţiu D. Rădulescu, Runzhang Xu, and Chao Yang. Qualitative properties of solution to a viscoelastic Kirchhoff-like plate equation. Journal of Mathematical Physics, 64(5): 051511, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
5/051511/2892545/Qualitative-properties-of-solution-to-a.

Zhang:2023:LSA

[1432] Yu Zhang and Shuai Fan. Limits of solutions to the Aw–Rascle traffic flow model with generalized Chaplygin gas by flux approximation. Journal of Mathematical Physics, 64(5):051512, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/5/051512/2892935/Limits-of-
solutions-to-the-Aw-Rascle-traffic-flow.

Argota-Quiroz:2023:QRG

[1433] J. N. Argota-Quiroz and S. Majid. Quantum Riemannian geometry of the discrete interval and q-deformation. Journal of Mathematical Physics, 64(5):051701, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/5/051701/2890221/Quantum-Riemannian-geometry-of-
the-discrete.

Watanabe:2023:GSD

[1434] Haruki Watanabe, Meng Cheng, and Yohei Fuji. Ground state degeneracy on torus in a family of ZN ZN ZN toric code. Journal of Mathematical Physics, 64(5):051901, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/5/051901/2891377/Ground-state-degeneracy-on-
torus-in-a-family-of-ZN.

Bruzda:2023:BCC

[1435] W. Bruzda. Block-circulant complex Hadamard matrices. Journal of Mathematical Physics, 64(5):052201, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/5/052201/2890042/Block-circulant-
complex-Hadamard-matrices.

Geisler:2023:ASN

[1436] Jakob Geisler. Absence of submultiplicative norms for Wick-ordered operator products. Journal of Mathematical Physics, 64(5):052301, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
052301/2892934/Absence-of-submultiplicative-norms-for-Wick.

Bhattacharya:2023:CFP

[1437] Subhra Bhattacharya and Subhasis Nalui. Complexity factor parameterization for traversable wormholes. Journal of Mathematical Physics, 64(5):052501, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/5/052501/2887577/Complexity-factor-parameterization-
for-traversable.

Celerier:2023:SSRb

[1438] M.-N. Célérier. Study of stationary rigidly rotating anisotropic cylindrical fluids with new exact interior solutions of GR. IV. Radial pressure. Journal of Mathematical Physics, 64(5):052502, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/5/052502/2890038/Study-of-
stationary-rigidly-rotating-anisotropic.

Ju:2023:PUI

[1439] Xuewei Ju. Pathwise unstable invariant manifolds reduction for stochastic evolution equations driven by nonlinear noise. Journal of Mathematical Physics, 64(5):052701, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/5/052701/2887914/Pathwise-
unstable-invariant-manifolds-reduction.

Chandre:2023:CNB

[1440] Cristel Chandre and Atsushi Horikoshi. Classical Nambu brackets in higher dimensions. Journal of Mathematical Physics, 64(5):052702, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
052702/2889494/Classical-Nambu-brackets-in-higher-dimensions.

Zhang:2023:PEA

[1441] Jiangwei Zhang, Zhiming Liu, and Jianhua Huang. Pullback exponential attractors for second-order lattice system with nonstandard growth condition. Journal of Mathematical Physics, 64(5):052703, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
052703/2890040/Pullback-exponential-attractors-for-second-order.

Guan:2023:BAE

[1442] Yi Guan. Boundedness of atmospheric Ekman flows with two-layer eddy viscosity. Journal of Mathematical Physics, 64(5):052704, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
052704/2891413/Boundedness-of-atmospheric-Ekman-flows-with-
two.

Perrella:2023:EGS

[1443] David Perrella, Nathan Duignan, and David Pfefferlé. Existence of global symmetries of divergence-free fields with first integrals. Journal of Mathematical Physics, 64(5):052705, May 1, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/5/052705/2892544/Existence-
of-global-symmetries-of-divergence-free.

Flandoli:2023:NBV

[1444] Franco Flandoli and Ruojun Huang. Noise based on vortex structures in 2D and 3D. Journal of Mathematical Physics, 64(5):053101, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/5/053101/2890029/Noise-based-on-vortex-structures-in-2D-
and-3D.

Yu:2023:SIL

[1445] Qian Yu. Self-intersection local time derivative for systems of nonlinear stochastic heat equations. Journal of Mathematical Physics, 64(5): 053301, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
5/053301/2889501/Self-intersection-local-time-derivative-for.

Fialho:2023:APN

[1446] Paula M. S. Fialho, Bernardo N. B. de Lima, Aldo Procacci, and Benedetto Scoppola. On the analyticity of the pressure for a nonideal gas with high density boundary conditions. Journal of Mathematical Physics, 64(5):053302, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/5/053302/2890222/On-the-analyticity-
of-the-pressure-for-a-non-ideal.

Ma:2023:DFM

[1447] Shilin Ma and Dafeng Zuo. Dubrovin–Frobenius manifolds associated with B_{n} and the constrained KP hierarchy. Journal of Mathematical Physics, 64(5):053501, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/5/053501/2887605/Dubrovin-Frobenius-
manifolds-associated-with-Bn.

Zhang:2023:ETO

[1448] Yongshuai Zhang, Deqin Qiu, and Jingsong He. Explicit N-th order solutions of Fokas–Lenells equation based on revised Riemann– Hilbert approach. Journal of Mathematical Physics, 64(5):053502, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
053502/2890041/Explicit-Nth-order-solutions-of-Fokas-Lenells.

Odake:2023:DOR

[1449] Satoru Odake. Discrete orthogonality relations for the multi-indexed orthogonal polynomials in discrete quantum mechanics with pure imaginary shifts. Journal of Mathematical Physics, 64(5):053503, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/5/
053503/2890117/Discrete-orthogonality-relations-for-the-multi.

Joseph:2023:GCC

[1450] Collin Mark Joseph and Ralf Meyer. Geometric construction of classes in van Daele’s K-theory. Journal of Mathematical Physics, 64(5):053504, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/5/053504/2891378/Geometric-construction-
of-classes-in-van-Daele-s-K.

Montani:2023:LOS

[1451] H. Montani. Lagrangian and orthogonal splittings, quasitriangular Lie bialgebras, and almost complex product structures. Journal of Mathematical Physics, 64(5):053505, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/5/053505/2891856/Lagrangian-and-
orthogonal-splittings.

Aniello:2023:TCO

[1452] Paolo Aniello, Stefano Mancini, and Vincenzo Parisi. Trace class operators and states in p-adic quantum mechanics. Journal of Mathematical Physics, 64(5):053506, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/5/053506/2892280/Trace-class-operators-and-states-
in-p-adic-quantum.

Malkoun:2023:NPA

[1453] Joseph Malkoun. A new proof of Atiyah’s conjecture on configurations of four points. Journal of Mathematical Physics, 64(5):053507, May 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
5/053507/2892933/A-new-proof-of-Atiyah-s-conjecture-on.

Xu:2023:NTR

[1454] Yongjun Xu and Jialei Chen. A new type restricted quantum group. Journal of Mathematical Physics, 64(6):061701, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/6/
061701/2893747/A-new-type-restricted-quantum-group.

Thiang:2023:TES

[1455] Guo Chuan Thiang. Topological edge states of 1D chains and index theory. Journal of Mathematical Physics, 64(6):061901, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/6/
061901/2893748/Topological-edge-states-of-1D-chains-and-index.

Kubu:2023:CFO

[1456] Ondřej Kubů and Libor Šnobl. Cylindrical first-order superintegrability with complex magnetic fields. Journal of Mathematical Physics, 64(6): 062101, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
6/062101/2893746/Cylindrical-first-order-superintegrability-
with.

Xiong:2023:ENC

[1457] Xuanqing Xiong and Qihuai Liu. Explicit nondegeneracy conditions of KAM tori for the planar N-point vortex systems. Journal of Mathematical Physics, 64(6):062701, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/6/062701/2893744/Explicit-nondegeneracy-
conditions-of-KAM-tori-for.

Sasakura:2023:EAE

[1458] Naoki Sasakura. Exact analytic expressions of real tensor eigenvalue distributions of Gaussian tensor model for small N. Journal of Mathematical Physics, 64(6):063501, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/6/063501/2893745/Exact-analytic-expressions-
of-real-tensor.

Durand:2023:MFT

[1459] Loyal Durand. Mehler–Fock transforms and retarded radiative Green functions in hyperbolic and spherical spaces. Journal of Mathematical Physics, 64(6):063502, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/6/063502/2893751/Mehler-Fock-transforms-and-
retarded-radiative.

Ji:2023:APE

[1460] Lijie Ji, Zhiguo Yang, Zhuoning Li, Dong Wu, Shi Jin, and Zhenli Xu. An asymptotic-preserving and energy-conserving particle-in-cell method for Vlasov–Maxwell equations. Journal of Mathematical Physics, 64(6): 063503, June 1, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
6/063503/2893958/An-asymptotic-preserving-and-energy-conserving.

E:2023:MLA

[1461] Weinan E, Huan Lei, Pinchen Xie, and Linfeng Zhang. Machine learning-assisted multi-scale modeling. Journal of Mathematical Physics, 64(7):071101, July 28, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/7/071101/2900787/Machine-
learning-assisted-multi-scale-modeling.

Xu:2023:GBF

[1462] Lu Xu, Li Yang, and Qiao Xin. Global boundedness for a food chain model with general logistic source. Journal of Mathematical Physics, 64(7):071501, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/7/071501/2900785/Global-boundedness-for-
a-food-chain-model-with.

Du:2023:QSP

[1463] Yao Du, Jiabao Su, and Cong Wang. The quasilinear Schrödinger– Poisson system. Journal of Mathematical Physics, 64(7):071502, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
071502/2901627/The-quasilinear-Schrodinger-Poisson-system.

Zhao:2023:GBP

[1464] Zhongzi Zhao and Ruyun Ma. Global bifurcation of positive solutions of a quasilinear indefinite Neumann problem in some FLRW spacetimes. Journal of Mathematical Physics, 64(7):071503, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
071503/2901816/Global-bifurcation-of-positive-solutions-of-
a.

Wang:2023:GET

[1465] Yinxia Wang, Zehua Luo, and Dan Li. Global existence and timedecay rates of solutions to the generalized Boussinesq equation with weak damping. Journal of Mathematical Physics, 64(7):071504, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
071504/2902202/Global-existence-and-time-decay-rates-of-solutions.

Sohn:2023:MPS

[1466] Juhee Sohn. Mountain pass solution for the self-dual Einstein–Maxwell– Higgs model on compact surfaces. Journal of Mathematical Physics, 64(7):071505, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/7/071505/2902638/Mountain-pass-solution-for-the-
self-dual-Einstein.

Lin:2023:GWP

[1467] Hongxia Lin, Heng Zhang, Sen Liu, and Qing Sun. Global well-posedness and decay of the 2D incompressible MHD equations with horizontal magnetic diffusion. Journal of Mathematical Physics, 64(7):071506, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
7/071506/2903534/Global-well-posedness-and-decay-of-the-2D.

Lei:2023:CCV

[1468] Zhoutong Lei and Zhiqiang Shao. Concentration and cavitation in the vanishing pressure limit of solutions to the relativistic Euler equations with the logarithmic equation of state. Journal of Mathematical Physics, 64(7):071507, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/7/071507/2903821/Concentration-and-cavitation-in-
the-vanishing.

Xu:2023:BFL

[1469] Zirui Xu and Qiang Du. Bifurcation and fission in the liquid drop model: a phase-field approach. Journal of Mathematical Physics, 64(7):071508, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/7/071508/2904103/Bifurcation-and-fission-in-the-
liquid-drop-model-A.

Moscolari:2023:UGW

[1470] Massimo Moscolari and Gianluca Panati. Ultra-generalized Wannier bases: Are they relevant to topological transport? Journal of Mathematical Physics, 64(7):071901, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/7/071901/2900786/Ultra-generalized-Wannier-bases-
Are-they-relevant.

Semenoff:2023:BFZ

[1471] Gordon W. Semenoff. Boundary ferromagnetism in zigzag edged graphene. Journal of Mathematical Physics, 64(7):071902, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
071902/2903535/Boundary-ferromagnetism-in-zigzag-edged-graphene.

Bachmann:2023:DAA

[1472] Sven Bachmann, Bruno Nachtergaele, and Siddharth Vadnerkar. Dynamical Abelian anyons with bound states and scattering states. Journal of Mathematical Physics, 64(7):071903, July 28, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/7/071903/2903823/Dynamical-
Abelian-anyons-with-bound-states-and.

Ligthart:2023:IHP

[1473] Laurens T. Ligthart and David Gross. The inflation hierarchy and the polarization hierarchy are complete for the quantum bilocal scenario. Journal of Mathematical Physics, 64(7):072201, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
072201/2901817/The-inflation-hierarchy-and-the-polarization.

Belgiorno:2023:SPK

[1474] F. Belgiorno and S. L. Cacciatori. Spectral properties for the Klein– Gordon Hamiltonian in charged black hole backgrounds. Journal of Mathematical Physics, 64(7):072301, July 28, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/7/072301/2900788/Spectral-
properties-for-the-Klein--Gordon.

Dugan:2023:STH

[1475] William T. Dugan, Loïc Foissy, and Karen Yeats. Sequences of trees and higher-order renormalization group equations. Journal of Mathematical Physics, 64(7):072302, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/7/072302/2902967/Sequences-of-trees-and-
higher-order.

Steininger:2023:CPP

[1476] F. Steininger and P. T. Chru’sciel. The Cauchy problem for the Proca equation in gravitating dielectric media. Journal of Mathematical Physics, 64(7):072501, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/7/072501/2900789/The-Cauchy-problem-
for-the-Proca-equation-in.

Ginoux:2023:ZHB

[1477] Jean-Marc Ginoux and Jaume Llibre. Zero-Hopf bifurcation in the Chua’s circuit. Journal of Mathematical Physics, 64(7):072701, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
072701/2900784/Zero-Hopf-bifurcation-in-the-Chua-s-circuit.

Zhong:2023:LSM

[1478] Wenhu Zhong, Guanggan Chen, and Yuanyuan Zhang. Limit stationary measures of the stochastic magnetohydrodynamic system in a 3D thin domain. Journal of Mathematical Physics, 64(7):072702, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
072702/2900792/Limit-stationary-measures-of-the-stochastic.

Brugues:2023:CSS

[1479] Joaquim Brugués, Sonja Hohloch, Pau Mir, and Eva Miranda. Constructions of b-semitoric systems. Journal of Mathematical Physics, 64(7): 072703, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
7/072703/2903246/Constructions-of-b-semitoric-systems.

McCarney:2023:EIW

[1480] Jordan McCarney. Exact internal waves in the presence of mean currents and rotation. Journal of Mathematical Physics, 64(7):073101, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
073101/2902963/Exact-internal-waves-in-the-presence-of-mean.

Beard:2023:EOT

[1481] K. Beard, A. Stefan, Jr. R.Viator, and A. Welters. Effective operators and their variational principles for discrete electrical network problems. Journal of Mathematical Physics, 64(7):073501, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
073501/2900793/Effective-operators-and-their-variational.

Aktosun:2023:MMS

[1482] Tuncay Aktosun, Ramazan Ercan, and Mehmet Unlu. The Marchenko method to solve the general system of derivative nonlinear Schrödinger equations. Journal of Mathematical Physics, 64(7):073502, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
7/073502/2901405/The-Marchenko-method-to-solve-the-general-
system.

Bailey:2023:IGC

[1483] Michael Bailey and Marco Gualtieri. Integration of generalized complex structures. Journal of Mathematical Physics, 64(7):073503, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/7/
073503/2901947/Integration-of-generalized-complex-structures.

Boggarapu:2023:TFA

[1484] Pradeep Boggarapu, Hatem Mejjaoli, Shyam Swarup Mondal, and P. Jitendra Kumar Senapati. Time-frequency analysis of (k, a)-generalized wavelet transform and applications. Journal of Mathematical Physics, 64(7):073504, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/7/073504/2902664/Time-frequency-analysis-of-k-a-
generalized-wavelet.

Melong:2023:GHV

[1485] Fridolin Melong and Raimar Wulkenhaar. Generalized Heisenberg– Virasoro algebra and matrix models from quantum algebra. Journal of Mathematical Physics, 64(7):073505, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/7/073505/2903076/Generalized-Heisenberg-
Virasoro-algebra-and-matrix.

Raymond:2023:MQC

[1486] Nicolas Raymond and Éric Soccorsi. Magnetic quantum currents in the presence of a Neumann wall. Journal of Mathematical Physics, 64(7): 073506, July 28, 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
7/073506/2903822/Magnetic-quantum-currents-in-the-presence-
of-a.

Seo:2023:EQR

[1487] Yuki Seo. Erratum: The quantum ♮_{α}-Rényi divergence of real order, [J. Math. Phys. 63, 072203 (2022)]. Journal of Mathematical Physics, 64(7):079901, July 28, 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/64/7/079901/2900794/Erratum-
The-quantum-Renyi-divergence-of-real-order. See [1082].

Rong:2023:NSM

[1488] Ting Rong and Fuyi Li. Normalized solutions to the mass supercritical Kirchhoff-type equation with non-trapping potential. Journal of Mathematical Physics, 64(8):081501, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/8/081501/2905675/Normalized-solutions-to-
the-mass-supercritical.

Zhang:2023:CRP

[1489] Minyi Zhang and Changjiang Zhu. Convergence rates to the planar stationary solution to a 2D model of the radiating gas on half space. Journal of Mathematical Physics, 64(8):081502, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081502/2905820/Convergence-rates-to-the-planar-stationary.

Hakkaev:2023:SSP

[1490] Sevdzhan Hakkaev, Milena Stanislavova, and Atanas G. Stefanov. Spectral stability of periodic waves for the Zakharov system. Journal of Mathematical Physics, 64(8):081503, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/8/081503/2906249/Spectral-stability-
of-periodic-waves-for-the.

Liu:2023:PMF

[1491] Lintao Liu, Kaimin Teng, Jie Yang, and Haibo Chen. Properties of minimizers for the fractional Kirchhoff energy functional. Journal of Mathematical Physics, 64(8):081504, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/8/081504/2906532/Properties-of-minimizers-
for-the-fractional.

Sato:2023:NIT

[1492] Naoki Sato and Michio Yamada. Nested invariant tori foliating a vector field and its curl: Toward MHD equilibria and steady Euler flows in toroidal domains without continuous Euclidean isometries. Journal of Mathematical Physics, 64(8):081505, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081505/2906531/Nested-invariant-tori-foliating-a-vector-field-
and.

Liu:2023:GSS

[1493] Yu Liu, Song Meng, Jiayan Wu, and Ting Zhang. Global strong solutions for the multi-dimensional compressible viscoelastic flows with general pressure law. Journal of Mathematical Physics, 64(8):081506, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081506/2906533/Global-strong-solutions-for-the-multi-dimensional.

Fang:2023:TSS

[1494] Beixiang Fang, Piye Sun, and Qin Zhao. Transonic shocks for 2D steady exothermically reacting Euler flows in a finite nozzle. Journal of Mathematical Physics, 64(8):081507, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/081507/2907011/Transonic-
shocks-for-2D-steady-exothermically.

Li:2023:SSR

[1495] Zhi Li and Wenqiang Zhao. Stability of stochastic reaction-diffusion equation under random influences in high regular spaces. Journal of Mathematical Physics, 64(8):081508, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/081508/2907215/Stability-
of-stochastic-reaction-diffusion.

Dinh:2023:LTD

[1496] Van Duong Dinh, Mohamed Majdoub, and Tarek Saanouni. Long time dynamics and blow-up for the focusing inhomogeneous nonlinear Schrödinger equation with spatially growing nonlinearity. Journal of Mathematical Physics, 64(8):081509, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/081509/2907140/Long-time-
dynamics-and-blow-up-for-the-focusing.

Chtioui:2023:CDO

[1497] Taoufik Chtioui, Atef Hajjaji, Sami Mabrouk, and Abdenacer Makhlouf. Cohomologies and deformations of O-operators on Lie triple systems. Journal of Mathematical Physics, 64(8):081701, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081701/2905824/Cohomologies-and-deformations-of-O-operators-
on.

Primc:2023:CRA

[1498] Mirko Primc and Tomislav Sikić. Combinatorial relations among relations for level 2 standard C_{n}(1)-modules. Journal of Mathematical Physics, 64(8):081702, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/081702/2906534/Combinatorial-relations-among-relations-
for-level.

Algethami:2023:QSC

[1499] Dakhilallah Algethami and Andrey Mudrov. Quantum symmetric conjugacy classes of non-exceptional groups. Journal of Mathematical Physics, 64(8):081703, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/081703/2907578/Quantum-symmetric-conjugacy-classes-
of-non.

Poletaeva:2023:LMS

[1500] Elena Poletaeva. On linked modules over the super-Yangian of the superalgebra Q(1). Journal of Mathematical Physics, 64(8):081704, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081704/2908458/On-linked-modules-over-the-super-Yangian-of-
the.

Schulz-Baldes:2023:SLS

[1501] Hermann Schulz-Baldes and Tom Stoiber. Spectral localization for semimetals and callias operators. Journal of Mathematical Physics, 64(8):081901, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/081901/2905837/Spectral-localization-for-semimetals-
and-Callias.

Shapiro:2023:ILD

[1502] Jacob Shapiro. Incomplete localization for disordered chiral strips. Journal of Mathematical Physics, 64(8):081902, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/081902/2906398/Incomplete-
localization-for-disordered-chiral.

Nessi:2023:DSP

[1503] Nicolás Nessi and Lucas Sourrouille. Dynamic and static properties of quantum Hall and harmonic oscillator systems on the noncommutative plane. Journal of Mathematical Physics, 64(8):081903, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
081903/2906701/Dynamic-and-static-properties-of-quantum-Hall-
and.

Sorokanich:2023:BMF

[1504] Stephen Sorokanich. Beyond mean-field: Condensate coupled with pair excitations. Journal of Mathematical Physics, 64(8):081904, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
8/081904/2907736/Beyond-mean-field-Condensate-coupled-with-
pair.

Yang:2023:ESC

[1505] Wei Yang. Exactly solvable complex PT symmetry potential A[(λx) + i tanh(λx)]. Journal of Mathematical Physics, 64(8):082101, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
082101/2905677/Exactly-solvable-complex-PT-symmetry-potential-
A.

Cerjan:2023:SLL

[1506] Alexander Cerjan, Lars Koekenbier, and Hermann Schulz-Baldes. Spectral localizer for line-gapped non-Hermitian systems. Journal of Mathematical Physics, 64(8):082102, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/

aip/jmp/article/64/8/082102/2905836/Spectral-localizer-for-
line-gapped-non-Hermitian.

Anapolitanos:2023:VWI

[1507] Ioannis Anapolitanos, Mariam Badalyan, and Dirk Hundertmark. On the van der Waals interaction between a molecule and a halfinfinite plate. Journal of Mathematical Physics, 64(8):082103, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
082103/2906564/On-the-van-der-Waals-interaction-between-a.

Pathirana:2023:LLN

[1508] Lubashan Pathirana and Jeffrey Schenker. Law of large numbers and central limit theorem for ergodic quantum processes. Journal of Mathematical Physics, 64(8):082201, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/

jmp/article/64/8/082201/2908509/Law-of-large-numbers-and-central-
limit-theorem-for.

Anco:2023:ALR

[1509] Stephen C. Anco and Jordan A. Fazio. Analog of a Laplace–Runge– Lenz vector for particle orbits (time-like geodesics) in Schwarzschild spacetime. Journal of Mathematical Physics, 64(8):082501, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/8/082501/2906622/Analog-of-a-Laplace-Runge--Lenz-vector-
for-particle.

Gasperin:2023:SFN

[1510] Edgar Gasperín and Rafael Pinto. Spin-0 fields and the NP-constants close to spatial infinity in Minkowski spacetime. Journal of Mathematical Physics, 64(8):082502, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/082502/2906567/Spin-0-fields-and-the-NP-constants-
close-to.

Martin:2023:SAH

[1511] Daniel Martin. Staticity of asymptotically hyperbolic minimal mass extensions. Journal of Mathematical Physics, 64(8):082503, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/8/082503/2906706/Staticity-of-asymptotically-hyperbolic-
minimal.

Blitz:2023:TCC

[1512] Samuel Blitz. Toward a classification of conformal hypersurface invariants. Journal of Mathematical Physics, 64(8):082504, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
082504/2907617/Toward-a-classification-of-conformal-hypersurface.

Bohmer:2023:MGU

[1513] Christian G. Böhmer and Erik Jensko. Modified gravity: a unified approach to metric-affine models. Journal of Mathematical Physics, 64(8): 082505, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/
64/8/082505/2908161/Modified-gravity-A-unified-approach-to-
metric.

Wang:2023:DVP

[1514] Yunping Wang and Ercai Chen. Double variational principle for mean dimensions with sub-additive potentials. Journal of Mathematical Physics, 64(8):082701, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/8/082701/2905676/Double-variational-
principle-for-mean-dimensions.

Moor:2023:PHR

[1515] Joris De Moor, Florian Dorsch, and Hermann Schulz-Baldes. Partially hyperbolic random dynamics on Grassmannians. Journal of Mathematical Physics, 64(8):082702, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/082702/2907125/Partially-
hyperbolic-random-dynamics-on.

Pu:2023:DNA

[1516] Zhe Pu and Dingshi Li. Dynamics of the non-autonomous stochastic p-Laplacian parabolic problems on unbounded thin domains. Journal of Mathematical Physics, 64(8):082703, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/64/8/082703/2907795/Dynamics-
of-the-non-autonomous-stochastic-p.

Hohloch:2023:PVD

[1517] S. Hohloch and G. Muarem. Point vortex dynamics on Kähler twistor spaces. Journal of Mathematical Physics, 64(8):082901, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
082901/2905839/Point-vortex-dynamics-on-Kahler-twistor-spaces.

Aydin:2023:BLO

[1518] Cengiz Aydin. From Babylonian lunar observations to Floquet multipliers and Conley–Zehnder indices. Journal of Mathematical Physics, 64(8): 082902, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic). URL https://pubs.aip.org/aip/jmp/article/64/
8/082902/2906705/From-Babylonian-lunar-observations-to-Floquet.

Knauf:2023:MCC

[1519] Andreas Knauf. Mini-course: Classical mechanics and transport. Journal of Mathematical Physics, 64(8):082903, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/8/082903/2908273/Mini-course-Classical-
mechanics-and-transport.

Imamura:2023:NAK

[1520] Takashi Imamura, Matteo Mucciconi, and Tomohiro Sasamoto. New approach to KPZ models through free fermions at positive temperature. Journal of Mathematical Physics, 64(8):083301, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
083301/2908272/New-approach-to-KPZ-models-through-free-fermions.

Yu:2023:PVC

[1521] Jianduo Yu, Siqi Chen, Chuanzhong Li, Mengkun Zhu, and Yang Chen. Painlevé V and confluent Heun equations associated with a perturbed Gaussian unitary ensemble. Journal of Mathematical Physics, 64(8):083501, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/64/8/083501/2905678/Painleve-V-and-
confluent-Heun-equations-associated.

Bertola:2023:HSR

[1522] M. Bertola, J. Harnad, and J. Hurtubise. Hamiltonian structure of rational isomonodromic deformation systems. Journal of Mathematical Physics, 64(8):083502, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/083502/2905679/Hamiltonian-structure-of-rational-
isomonodromic.

Min:2023:HDO

[1523] Chao Min and Yang Chen. Hankel determinant and orthogonal polynomials for a perturbed Gaussian weight: From finite n to large n asymptotics. Journal of Mathematical Physics, 64(8):083503, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
083503/2906702/Hankel-determinant-and-orthogonal-polynomials-
for.

Eom:2023:EPT

[1524] J. Y. Eom, M. Machida, G. Nakamura, G. Nishimura, and C. L. Sun. Expressions of the peak time for time-domain boundary measurements of diffuse light. Journal of Mathematical Physics, 64(8):083504, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/64/8/
083504/2906703/Expressions-of-the-peak-time-for-time-domain.

Tadano:2023:IOE

[1525] Homare Tadano. Improved oscillation estimates and the Hitchin-Thorpe inequality on compact Ricci solitons. Journal of Mathematical Physics, 64(8):083505, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/

article/64/8/083505/2907141/Improved-oscillation-estimates-
and-the-Hitchin.

Nittis:2023:ANE

[1526] Giuseppe De Nittis and Danilo Polo Ojito. About the notion of eigenstates for C^{∗}-algebras and some application in quantum mechanics. Journal of Mathematical Physics, 64(8):083506, August 2023. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://

pubs.aip.org/aip/jmp/article/64/8/083506/2907390/About-the-
notion-of-eigenstates-for-C-algebras-and.

Wang:2023:SLS

[1527] Dan Wang, Mengkun Zhu, and Yang Chen. A singular linear statistic for a perturbed LUE and the Hankel matrices. Journal of Mathematical Physics, 64(8):083507, August 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL

https://pubs.aip.org/aip/jmp/article/64/8/083507/2908163/A-
singular-linear-statistic-for-a-perturbed-LUE.

Solovej:2021:SCH

[1528] Jan Philip Solovej, Rafael Benguria, and Martin Zirnbauer. Special collection in honor of Freeman Dyson. American Institute of Physics

Web site, 2021. URL https://pubs.aip.org/jmp/collection/1381/
Special-collection-in-honor-of-Freeman-Dyson.